A Complete Bibliography of Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous

Technologies (IMWUT)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

10 November 2023 Version 1.09

Title word cross-reference

2 [BGK21, DJY^{+}22]. 3 [ASG^{+}23, AWZK23, BKL^{+}22, BGL18, DCZ^{+}23, GLW^{+}18, HWJ^{+}21, HDFZ22, HHX^{+}20, JWS^{+}23, MQZH23, MLH^{+}23, RWW^{+}22, SLX^{+}18, ZXW^{+}17]. 360 [FSW19]. μ [PLW^{+}22]. n [GSRS20].

-Degree [FSW19]. -Gage [GSRS20]. -Person [DJY^{+}22].

/Check [CCW^{+}20]. /Check-in [CCW^{+}20].

19 [LCR^{+}21, TAHH23].

6-week [CFZ^{+}19]. 60 [SHZ^{+}20].

Abacus [EUHB23]. Absolute [SKS^{+}21]. Abstractions [CND17]. AC [GPW^{+}18]. Academia [DGH^{+}20, GWJ^{+}21]. Academic [KKL21]. Acceleration [GH18]. Accelerometer [CSHN17, XKNM20]. Accelerometers [GLG^{+}19]. Accelerometry [KTH^{+}20]. Accentuation [AOGF^{+}21]. Acceptability [HJL^{+}22]. Access [GL18]. Accessible [BSMP23, KGAH22]. According [LCL20]. Accountability [JSC^{+}18]. Accuracy [BXY^{+}21]. Accurate [ARSN22, ASG^{+}23, ILM^{+}21, LCS^{+}21, PCBK20, WXY^{+}17, ZCA^{+}21]. Achieve [KJP^{+}19]. Achievement [AJV20]. Acoussist [JXZ^{+}20]. Acoustic [AYT21, AWSS22, BLW^{+}23, BMR^{+}20, CPW^{+}21, CCY^{+}23, CYX^{+}19, CNXG22, DWZZ19, GJL^{+}20, GZZH20, GZNL22, HSK^{+}22, IRA^{+}21, JXZ^{+}20, JMC^{+}21, JGZ^{+}21, KDL^{+}21, LZL^{+}22, LLLX22, LLCY21, LYLT21, LT19, LZT23, LYCW20, MLH^{+}23, MSWQ20, PBM^{+}18, STL^{+}19, SXF^{+}22, XZZ21, YZS^{+}20, ZWZ^{+}21b]. Acoustic-based [CYX^{+}19, CNXG22, GJL^{+}20, GZNL22, JGZ^{+}21, XZZ21]. Acoustics [ZXW^{+}17]. Across [CRP21, RH19, CCX^{+}18, CMK22, EWM^{+}20, GZH^{+}21, HFS^{+}19, LGM^{+}22, MH23, SNGK20]. Across-Context [CRP21]. Action [HY22, RLM^{+}19, XZZ21]. Actions [VLR20]. Activation [KSZ^{+}22]. Active [AT19, CCY^{+}23, DMAH18, HAR18, JHL^{+}21, LMS^{+}23, ZXW^{+}17]. Activities [BAT22, DGH^{+}20, HWZZ19, LT19, LZZ^{+}23, SCZ^{+}20, SSR^{+}19, ZZN^{+}20]. Activity [AES^{+}21, AT19, AYT21, ATP^{+}19, AJV20, ACB22, BYW^{+}20, BW20, BGK21, BPA23, CFZ^{+}19, CZP20, CLTS21, CHWZ23, DGS18, FGL18, FTL^{+}19, GwC^{+}23, GMS^{+}23, GDS^{+}20, GF17, GP17, Hal19, HEP21, HEP22, HP20, HNCP22, HAR18, ILH^{+}19, ISY^{+}21, JTM^{+}22, JSGS23, KLM^{+}18, KHZ22, KNP17, KXAH18, KTH^{+}20, KWAP21, LMT19, LH19, LLLX22, LWA23, LGKM20, LKL19, LXC^{+}19, LLG^{+}23, LWC^{+}22, MBB20, MFSK21, MGC^{+}23b, NTKG17, NTT^{+}21, PRC^{+}18, PCYZ18, QCWY19, QWK^{+}19, RH19, SOL19, SGZ^{+}23, SH20, SWLG22, TK17, XKNM20, XLL^{+}21]. Activity-Oriented [ATP^{+}19]. Activity-Stage [GMS^{+}23]. Actors [SGO^{+}23]. Actuate [DH17]. Actuation [CKJH23, FDH21, RPB^{+}18]. Actuator [JCR22]. Actuators [EWM^{+}20]. ad [NZL^{+}20]. AdaEnlight [LLZ^{+}22]. AdaMICA [AY22]. Adaptation [BRYH21, BGK21, HKK^{+}23, IN20, LCW^{+}21, ZSG^{+}21].

Adaptative [WGL^{+}21]. Adaptive [AY22, KKV^{+}22, KPL20, MQZH23, MCCH21, MBB20, QCWY19, SSB^{+}21, WGL^{+}21, FZZ^{+}22, LGM^{+}21]. AdaSpring [LGM^{+}21]. Additional [AKPS18]. Addressing [JW18, PB20]. Adherence [JHK^{+}21, RLY^{+}19, TK17]. Adherence-based [TK17]. ADL [LT19]. Administering [KKZ^{+}18]. Adoption [CMV^{+}18]. Ads [BWUW21]. Adult [KTEM21]. Adults [BFS^{+}17, CBZN19, JHWF22, LCFT19, MLPGP21, MSTT22, WBG^{+}19]. Adversarial [BYW^{+}20, BMR^{+}20, CHWZ23, CLW^{+}23, JHH21, KHZ22, LDSZ19]. Advertisement [NZL^{+}20]. Advertising [NZL^{+}20]. AEROKEY [LYP^{+}22]. AeroTraj [ASG^{+}23]. Aesthetic [HCH^{+}22]. Affective [FDH21, GZC^{+}23, LGL23, LKL22]. after [DDKL19, GKO^{+}23, WTA^{+}20]. Again [CCW^{+}20]. Against [GLW^{+}18, LZZ^{+}20]. Ageing [KTEM21]. Agencies [GWJ^{+}21]. Agent [CSM^{+}20, JKB^{+}18, DSSD23]. Agents [AvdBM^{+}22, CND17, FML^{+}22, LCL20, WMS^{+}21]. Aggregate [FALW18]. Aggregation [LBC^{+}22]. Agnostic [WAL^{+}19]. AI [CKJH23, GJG^{+}18, LCL20, PRS^{+}23]. AI-to-Human [CKJH23]. Aid [LCFT19]. Air [CDL^{+}19, CHZT19, CZT22, CMF^{+}23, EUHB23, GO22, ISSK20, KTM^{+}18, LCC^{+}22, LDC18, LLZ^{+}18, MLX^{+}20, MZT18, PSLS21, WXL^{+}20, SKS^{+}21, MZT18]. Airflow [CZL^{+}22]. AiSee [BMZN20]. Alcohol [MLPGP21, PMGP19]. Alerts [KEI^{+}18]. Algorithm [LGKM20]. Algorithmic [HPWW18]. Alignment [MH23]. All-day [CHC^{+}22]. Alleviating [DH17]. Allocation [RBL^{+}20]. ALWAES [JDZ^{+}21]. Always [AJA^{+}18, TKF^{+}19]. Always-Listening [TKF^{+}19]. Am [AYT21, BWUW21]. Amateur [CLZL18]. Ambient [KOY^{+}19, KSC^{+}17, LYP^{+}22, LZG^{+}21, SvBV^{+}18, TXWZ22, ZPZ^{+}20]. AmbientBreath [LEP21]. Ambulances [JZWL19]. Ambulatory [GDS^{+}20]. American [LHS^{+}23]. AMIR [LMS^{+}23]. among [PCF^{+}23]. Amplitude [LLW^{+}22]. Amplitude-modulated [LLW^{+}22]. Analog [RGL^{+}18]. Analyses [HFX^{+}21]. Analysis [ASTW19, BDPH21, BCK^{+}22, BPA23, CTTMT20, CKLR21, DCMM20, GWJ^{+}21, GZY^{+}21, HHFM17, JKK^{+}17, KVZ^{+}19, KML^{+}21, LAS^{+}19, MDK^{+}22, SH20, LFH18]. Analyzer [HHZ23]. Anatomy [CWP^{+}18]. Anchors [DJL^{+}21]. Angle [HDFZ22]. Angles [GKO18]. Angular [LDW^{+}18]. Animal [KLM^{+}18]. Animo [LEP^{+}19]. Annotation [AT19, LNAH21]. Annotation-Guided [LNAH21]. Annotations [GFK^{+}18]. Annoyance [NZL^{+}20]. Anomalies [FYW^{+}19]. Anomaly [HYH21, LYWZ22]. Answer [JKB^{+}18]. Answering [CKLB22]. Anti [ZLD^{+}21]. Anti-Spoofing [ZLD^{+}21]. Anxiety [SEG^{+}18]. AP [EY20]. AP-independent [EY20]. Apnea [CXT^{+}21]. ApneaDetector [CXT^{+}21]. App [CWH^{+}19, CGH^{+}17, CWP^{+}18, IB23, LSE21, LNAH21, TFL^{+}19, WLZ^{+}19, YXW^{+}20]. Appearance [HBK21]. Appearance-based [HBK21]. Apple [FKI18]. Application [ABL^{+}23]. Applications [BS17, BRS21, Hal19, KBXP23, LL22, LYH^{+}20, LGM^{+}21, PSC^{+}23, WCY^{+}19, WMW^{+}20, ZCA^{+}21]. AppMoD [WBG^{+}19]. Approach [AKB23, CDG^{+}17, DGH^{+}20, GMC18, GCW^{+}18, HWJ^{+}22, KBBN21, LDC18, LLF^{+}21, OGG^{+}18, PTMM^{+}22, RLY^{+}19, SSR^{+}19, WZG^{+}19]. Approaches [HJL^{+}22, MVS^{+}20]. Approaching [KWAP21]. Approximating [RSSN21]. Apps [CCX^{+}18, EZC23, JLD^{+}18, JHWF22, LAH18, LHK^{+}19, OGG^{+}18, RQ20]. Aquilis [KBKH20]. Architecture [JHH21, MCCH21]. Area [LH19]. Areas [ALC^{+}20]. Arm [LWZ^{+}20]. Armband [GMMH21, LLL21]. AROMA [PCYZ18]. Array [GJC^{+}21b, WXW^{+}18]. Arrays [GSS^{+}19, MZST17]. Art [AES^{+}21]. ARticulate [CND22]. Artifacts [GDS^{+}20]. Ask [KKV^{+}22]. ASL [JZG^{+}23, SHZ^{+}20, SPRK23]. AssessBlocks [WTA^{+}20]. Assessing [HEP22, LCX^{+}22, MTH19, SNGK20, WTA^{+}20]. Assessment [BCK^{+}22, CSK^{+}18, CCY^{+}17, DGS18, DBW^{+}22, GLG^{+}19, HMM^{+}20, HP20, KBBN21, KNP17, LSY^{+}18, PLW^{+}22, RHI^{+}23, WXW^{+}18]. Assessments [IAE^{+}19]. Assist [BDM^{+}18]. Assistance [CHC^{+}22]. Assistant [AvdBM^{+}22, AYT21, AZK21, BXY^{+}21, CXR^{+}18, GLS^{+}18, HDW^{+}17, KLT19]. Assistants [BLS^{+}18, BA20, CND22, TKF^{+}19, WAL^{+}19]. Assisted [DZH^{+}22, ELD19, KGL^{+}18, LLL^{+}19, ZLD^{+}21, RLY^{+}19, WXY^{+}19, ZM22]. Assisting [JXZ^{+}20]. Assistive [AKPS18, BMZN20, JMC^{+}21]. Association [HLL17]. Association-Free [HLL17]. Assumptions [BGR^{+}20]. ASSV [DWZZ19]. At-Source [KMC^{+}18]. Ataxia [RHI^{+}23]. Atmospheres [ZAP17]. Atrial [DSR^{+}22]. Attachable [BXY^{+}21]. Attachment [AWZK23]. Attack [FYX^{+}21, GMMH21, SWLZ21]. Attackers [LZZ^{+}20]. Attacks [BMR^{+}20, EBTN22, GLW^{+}18, HWJ^{+}22, LCL17]. Attend [AES^{+}21, ZWZ^{+}21a]. Attention [AKN^{+}19, AHS^{+}18, CZD^{+}21, FSJ^{+}19, GwC^{+}23, GCC^{+}21, HNL22, HYH21, LCG^{+}22, NHL^{+}20]. Attention-Aware [CZD^{+}21]. Attention-based [FSJ^{+}19]. Attention-Guided [NHL^{+}20]. Attentional [LYL^{+}20]. Attitude [BX19]. Au-Id [HWZZ19]. Audience [GDS19, PB20, SGO^{+}23]. Audio [CZL^{+}22, DBBH21, GBLM17, GFK^{+}18, LT19, WYL^{+}21, XYG^{+}20]. Audio-Based [LT19]. Audio-Related [DBBH21]. Auditable [CNS21]. Auditory [KCP^{+}18]. Auditory-verbal [KCP^{+}18]. Audits [FCFW^{+}17]. Augmentation [LGL23]. Augmented [BKL^{+}22, CLZL18, CND22, CNS21, GCM^{+}19, KHZ22, LAX^{+}23, OSG^{+}22, PB20]. Augmented/Mixed/Virtual [CNS21]. Augmenting [DJL^{+}22, FCFW^{+}17, LLPH22, NCM^{+}17]. Auracle [BWT^{+}18]. AuraRing [PWP19]. Authentic [ASV^{+}18]. Authenticating [HSL^{+}20]. Authentication [ATC17, CRS^{+}18, CKHM20, CLD^{+}20, CXZ^{+}21, CNXG22, CYK^{+}23b, GWP^{+}19, GJC^{+}21a, HQY^{+}18, HFS^{+}19, HCYZ18, HWZZ19, HQC^{+}23, HSK^{+}22, HKK^{+}23, JCL^{+}20, JJB22, KTM^{+}18, KPG^{+}20, LKBK19, LYP^{+}22, LFH18, LCR^{+}18, LYCW20, SC22]. Autism [LAX^{+}23]. Auto [CZD^{+}21, RHI^{+}23]. Auto-capture [CZD^{+}21]. Auto-Gait [RHI^{+}23]. Autobiographical [GCP^{+}22]. AutoCompletion [CS22a]. Autoencoder [LS20, MLX^{+}20]. Autoencoder-based [LS20]. Automated [BRS^{+}20, BBH17, CKLR21, CJRG21, CRGR22, CMF^{+}23, CBR23, DBW^{+}22, GLG^{+}19, GWD^{+}17, KYJ^{+}23, LAS^{+}19, LZT23, YPJdB19, ZLB^{+}17]. Automatic [ARD^{+}18, CDA23, CLHW18, HCH^{+}22, HWZZ19, IB23, JDZ^{+}21, KTH^{+}20, NCM^{+}17, RHI^{+}23]. Automatically [DGH^{+}20, WCY^{+}19]. Automating [AGU^{+}22]. Automation [LSY^{+}23, VLR20]. Automotive [HSL^{+}20]. Autonomous [CLZ^{+}21, FRB^{+}21]. Available [JWS^{+}23]. Avalanche [EZC23]. Avatar [LJZ^{+}22]. Aware [ABL^{+}23, CZD^{+}21, CGH^{+}17, DSM17, GMS^{+}23, HFS^{+}19, HYY^{+}23, JDZ^{+}21, JLZ^{+}23, KMC^{+}18, LHW^{+}19, LYH^{+}20, WZF^{+}20, DSSS20, CWH^{+}19, CML22, CYB^{+}23, FTL^{+}19, HY18, KGN^{+}20, LCZ^{+}21, LLZ^{+}22, MFSK21, TLF22, WGL^{+}21, YXW^{+}20]. Awareness [CRGR22, ELD19, GH18, OSG^{+}22, SZJ^{+}17]. Axial [JHK^{+}21].

Baby [HHZ23]. BabyNutri [HHZ23]. Back [AM19, CCW^{+}20]. Backdoor [EBTN22]. Backscatter [KJS20, TZW^{+}19, WRLK22]. Balanced [CLW^{+}23]. Balancing [ATP^{+}19, NZL^{+}20]. Ball [FDL^{+}18]. Ballistocardiogram [HQC^{+}23]. Ballpoint [OWH^{+}20]. Bandaid [RGL^{+}18]. Barometric [MH23]. Barriers [KW23]. Based [AZEL17, AIP^{+}18, ACB22, ATC17, BOH^{+}21, CLZL18, CSM^{+}20, CDL^{+}19, CLD^{+}20, CZP20, DCMM20, DLH^{+}18, DJL^{+}22, DSPR18, EUHB23, FZYZ18, FXC^{+}21, GMMH21, GL18, GWD^{+}17, GZY^{+}21, HBX^{+}17, HHSM18, HNL22, HJL^{+}22, HWJ^{+}22, HHZ23, HSK^{+}22, IAE^{+}19, JCR22, JSF^{+}18, JJB22, JHK^{+}21, KPG^{+}20, LMT19, LKBK19, LCFT19, LT19, LNT^{+}19, LCH^{+}22, LWX^{+}22, MBP^{+}17, MDK^{+}22, MGC^{+}23b, NTO^{+}18, PCYZ18, QCP^{+}18, RDOA22, RBL^{+}20, TZW^{+}19, WXW^{+}18, XFW^{+}20, ASG^{+}23, AZK21, BFS^{+}17, BW20, BGK21, CRS^{+}18, CKHM20, CYX^{+}19, CXZ^{+}21, CLTS21, CNXG22, CLH18, DWHR17, EEY18, EY20, FSJ^{+}19, FWJ^{+}22, GBS^{+}21, GJL^{+}20, GZZH20, GCLL19, GZNL22, GF17, GCC^{+}21, HDW^{+}17, HBK21, IYS22, JSGS23, JMC^{+}21, JGZ^{+}21, KHSW17, KTM^{+}18, KJP^{+}19, KASC21, LLPH22, LS20, LLZ^{+}20, LCW^{+}21, LCC^{+}22, LWA23, LXW^{+}21, LZG^{+}21, LLG^{+}23, LWC^{+}22, LYF^{+}21, MH23, MFSK21, NVP18, PBM^{+}18, PLX^{+}18, RLM^{+}19, SGZ^{+}23, STL^{+}19]. based [SWLG22, TK17, WPH^{+}23, WXY^{+}17, XZZ21, YZ21, ZXY21, XLL^{+}21]. Batch [MBB20, TZW^{+}19]. Batteries [HSL^{+}20, HLS20]. Battery [KBC^{+}20, KGAH22, LHW^{+}19, SBP^{+}22, TKGS17]. Battery-Free [TKGS17, KBC^{+}20, KGAH22]. Battery-less [SBP^{+}22]. Battery-Lifetime-Aware [LHW^{+}19]. Batteryless [FDL^{+}18, GPB^{+}23, KASC21]. Batting [KNP17]. BayesBeat [DSR^{+}22]. Be [ASV^{+}18, CYX^{+}19, JYY21, NCM^{+}17]. Beacon [KKZ^{+}18]. Beamforming [FDL^{+}18]. Beautiful [JHP^{+}17]. Became [LCL20]. Bed [PCBK20]. Bed-Mounted [PCBK20]. Bedrooms [KSC^{+}17]. before [TAHH23]. Behavior [ASV^{+}18, AJA^{+}18, AKB23, CVL20, DYCX21, GNA^{+}19, HFX^{+}21, HW17, HKK^{+}23, IMO^{+}18, JKK^{+}17, KIT^{+}20]. Behavior-irrelevant [DYCX21]. Behavioral [CKHM20, GZY^{+}21]. Behavioral-based [CKHM20]. Behaviors [ASS^{+}19, KPL20, KKL21]. Behaviour [KPG^{+}20, LKZ^{+}21, PSC^{+}17, SWLG22]. Behaviour-Based [KPG^{+}20]. Behavioural [GSRS20]. Behaviours [XL19]. Being [HPWW18, CSK^{+}18, KML^{+}21]. Belgian [AvdBF^{+}19]. Benchmarking [JWCC^{+}20]. Beneficial [KKL21]. Benefits [LKL22]. Better [AJA^{+}18]. Between [GDS19, DCMM20, DWL^{+}22, ISSK20, LFL^{+}23]. Beyond [AES^{+}21, CS22b, JGSC19, LXC^{+}19]. BFree [KBC^{+}20]. Biased [GMC18]. Biases [PLW^{+}22]. Bike [WXL^{+}20]. Bike-share [WXL^{+}20]. BiliScreen [MBP^{+}17]. Biobehavioral [DDKL19]. Biofeedback [STL^{+}19]. Biometric [CXZ^{+}21, HFS^{+}19]. Biometrics [CLD^{+}20, JCL^{+}20]. Biosignals [GCP^{+}22, LEP^{+}19]. Bipolar [HMM^{+}20]. Bit [LH20]. Bit-level [LH20]. Bite [CFF^{+}22, MRCdCRV^{+}21]. BitLight [LH20]. Blind [CHC^{+}22, GAS^{+}18, KIT^{+}20]. Blindness [HSWB21]. BlindType [LYY^{+}17]. Blink [CLY^{+}23, LLWX21]. Blinking [DH17]. BlinkListener [LLWX21]. Block [WTA^{+}20]. Blocks [GCM^{+}19]. Blood [CCJ^{+}17, CH18, FGS^{+}23, GZZ^{+}17, HW17, MNV^{+}19].

Bluetooth [BDPH21]. BMAR [MH23]. Board [LLPH22, PLK21]. Body [ARSN22, AZK^{+}23, BSMP23, CDA23, CJF^{+}23, CBEG17, EWM^{+}20, GJC^{+}21a, GZL^{+}21, KHSW17, LJZ^{+}22, LCR^{+}18, MLH^{+}23, MWJ^{+}21, XKNM20, CLTS21, GNA^{+}19, KHW^{+}23, KTH^{+}20]. Body-worn [XKNM20]. BodyCompass [YYW^{+}20]. Boosting [CKJH23, SGZ^{+}23]. BoostMeUp [CGJC19]. Bootstrapping [HNCP22]. Boundary [CLY^{+}21, LLZ^{+}20]. Bounded [AYT21]. Bowel [CWS^{+}19]. Bracelets [LHC^{+}22]. Bragg [SFR^{+}20]. Brain [LCX^{+}22, OHW^{+}20]. Brainwaves [LP20]. Breaks [DGH^{+}20]. Breakup [TAHH23]. Breathe [PZH^{+}18]. Breathing [CRS^{+}18, CEL^{+}22, GZNL22, HCYZ18, IRA^{+}21, LEP21, PZH^{+}18, STL^{+}19]. Breathing-based [CRS^{+}18]. BreathLive [HCYZ18]. BreathMentor [GZNL22]. BreathTrack [IRA^{+}21]. Breeze [STL^{+}19]. Bringing [MBZ^{+}18, SZJ^{+}17]. Brushing [ASS^{+}21]. Building [CZL^{+}22, DSM17, EY20]. Building- [EY20]. Buildings [BCM^{+}23, FCFW^{+}17]. Built [PM19]. Built-In [PM19]. Bulb [GMY^{+}17]. Bus [FFQ^{+}20, LHW^{+}19]. Buses [PLC^{+}21]. Business [DJN^{+}18]. Butterfly [HQY^{+}18]. Bystander [OSG^{+}22]. Bystanders [AKPS18]. Byte [BDPH21, BS17].

Cables [LYD^{+}23]. CAFI [TLF22].

CAFI-AR [TLF22]. Calibrating [LDC18]. Calibration [CHZT19, GMC18, MZST17]. Calm [KSC^{+}17]. Calorie [BS22]. Camera [ATC17, BDM^{+}18, CML22, DZH^{+}22, GJG^{+}18, HYY^{+}23, LCFT19, LLW^{+}22, LWX^{+}22, TXWZ22]. Camera-Based [LCFT19, LWX^{+}22]. Cameras [ASV^{+}18, ATP^{+}19, CW18, CLL^{+}22, HHX^{+}20, PLK21]. CamRadar [LLW^{+}22]. Can [CYX^{+}19, CGJ^{+}22, LLL21, MMMS^{+}20, NCM^{+}17, RHMZ^{+}21, WMS^{+}21, YLL^{+}20]. Canal [GWP^{+}19]. Cancellation [CCY^{+}23].

Can’t [ASV^{+}18]. CAP [CWH^{+}19]. Capacitive [GPW^{+}18, IY20]. CapHarvester [GPW^{+}18]. CAPTCHA [FCQR20]. Captivates [CRP21]. Capture [ASV^{+}18, NCM^{+}17, QCP^{+}18, CZD^{+}21]. Captured [HHX^{+}20, HWZZ19, IRA^{+}21]. Capturing [CLW^{+}18, CYK^{+}23a, MLMH^{+}18]. Car [CJRG21, LEP21, LKZ^{+}21, MZJ20, PZH^{+}18]. Cardiac [GNA^{+}19, ZPPC^{+}20]. CardiacWave [XLL^{+}21]. Care [ILH^{+}19, JZG^{+}23, KGL^{+}18]. CARIN [BW20]. CarOSense [MZJ20]. Carrier [LBC^{+}22]. Case [AvdBF^{+}19, BSP21, DKZ^{+}22, DDKL19, EBTN22, KKV^{+}22, PRS^{+}23]. Cases [LWYZ21]. Casual [DSPR18]. Cellphone [FZYZ18, TKGS17]. CellSense [FYY^{+}21]. Cellular [FYY^{+}21, LBC^{+}22, SJL^{+}21]. Centered [KKV^{+}22, SSR^{+}19]. Centric [BBBN22]. Chain [WSW^{+}22, WSW^{+}22]. Challenges [BA20, LCR^{+}21]. Change [CWS^{+}19]. Changes [CKL23]. changing [CFF^{+}22]. Channel [FYX^{+}21, GMMH21]. Channels [FZS^{+}23]. CHAR [JSGS23]. Characteristics [CCW^{+}20, HHFM17, LAS^{+}19]. Characterization [CRS^{+}18]. Characterizing [GLX^{+}22, PMGP19]. Charge [MWJ^{+}21]. Charge-Free [MWJ^{+}21]. Charging [WZF^{+}20]. Check [HFX^{+}21, CCW^{+}20]. Check-in [CCW^{+}20, HFX^{+}21]. Checked [BGR^{+}20]. Checkout [DLD^{+}21, FRB^{+}21]. Chemical [JCF^{+}21]. Chest [CXZ^{+}21]. ChestLive [CXZ^{+}21]. Child [CCY^{+}17, KLK^{+}20]. Children [AAM^{+}23, DJL^{+}22, EZC23, GS20, JLZ^{+}23, KSZ^{+}22, KL19, LAX^{+}23, WTA^{+}20]. Chills [HCH^{+}22]. Choice [NJWFM19]. Choices [NNMY23]. Chronic [BLW^{+}23, JHK^{+}21, SKM^{+}19]. Churn [KCuH^{+}19]. Circuit [CLZ^{+}21]. CircuitGIue [LRH^{+}23]. Circuits [CND^{+}20]. Citizen [BSH^{+}17]. City [AvdBM^{+}22, CFLK18, DYL^{+}19, DGZ^{+}21]. City-Wide [DGZ^{+}21]. CityGuard [WZG^{+}19]. Citywide [FSX^{+}18, WZG^{+}19]. class [GSRS20]. Classification [HBH20, ISY^{+}21, ZPPC^{+}20]. Classifying [AKN^{+}19]. Classroom [AKX^{+}19, DBW^{+}22, GRS^{+}22, LMT19]. Climb [DHC^{+}18]. Clinical [GMS^{+}23, GCC^{+}21, RLGN19]. Clippable [CEL^{+}22]. Closures [JLZ^{+}23]. Cloth [BDR17]. Cloud [AM19, CS22b, CWP^{+}18, ZLL^{+}20]. Clouds [ARD^{+}18, CS22a, PSLS21]. Clustering [LWA23]. Clustering-based [LWA23]. cm [ZCA^{+}21]. cm-Accurate [ZCA^{+}21]. CNN [MFSK21, RDOA22]. CNN-based [MFSK21]. Co [AMG^{+}19, LFH18]. Co-analysis [LFH18]. Co-making [AMG^{+}19]. Coach [TAEB17]. Cocaine [GNA^{+}19]. Cochlear [BA20]. COCOA [DXS^{+}22]. Coconut [LAH18]. Code [LNAH21, SLX^{+}18]. Codes [SLX^{+}18]. Coding [HEP21]. Cognition [DSM17]. Cognition-Aware [DSM17]. Cognitive [AVD^{+}17, CSM^{+}20, CRGR22, CGJC19, DSM17, GSRS20, HMM^{+}20, IAE^{+}19, SNGK20]. Cohort [DLH^{+}18, FALW18]. Cohort-Based [DLH^{+}18]. Cohorts [DDP^{+}17]. Cold [TFL^{+}19]. Cold-start [TFL^{+}19]. Collaboration [CWS^{+}19, CCK^{+}18, PRS^{+}23]. Collaborative [FTL^{+}19, GF17, GCM^{+}19, HNL22, JTM^{+}22, ZLB^{+}17]. Collar [KLM^{+}18]. Collect [CLTS21]. Collecting [JLD^{+}18]. Collection [LKL22, PRS^{+}23]. College [MRCdCRV^{+}21, MDK^{+}22]. ColloSSL [JTM^{+}22]. Combining [MFSK21]. Come [CCW^{+}20, CMF^{+}23]. ComFeel [CSQ^{+}20]. Comfort [GKO18]. Comfortable [JSC17]. Command [WAL^{+}19]. Command-Agnostic [WAL^{+}19]. Commands [JGX^{+}22, ZLD^{+}21]. Commercial [HY18, XFW^{+}20].

Commodity [ARSN22, BAT22, BSMP23, CXC^{+}19, FXC^{+}21, GJC^{+}21b, GL18, JJB22, KNL^{+}23, LZT23, LHY^{+}20, RTZ^{+}20, SXF^{+}22, WSW^{+}22, ZWX^{+}20]. Common [DJL^{+}22]. Communicate [CKLR21]. Communicating [FDH21]. Communication [CBR23, GMS^{+}23, HSH^{+}23, HLL17, KJS20, SBP^{+}22, TXWZ22, ZLB^{+}17]. Communities [PCF^{+}23]. Community [GWJ^{+}21, KBXP23]. ComNSense [ARD^{+}18]. Compact [ZCC^{+}20]. Compactor [IB23]. Companion [KXAH18]. Comparative [CKLR21]. Comparing [CCK^{+}18, KTM^{+}18, LCL20]. Comparison [CCX^{+}18]. Complex [JSGS23, PCYZ18]. Complexity [HP20]. Complicated [TAHH23]. Components [LRH^{+}23]. Comprehensive [JZG^{+}23, KEI^{+}18]. Compression [BMG18, LGM^{+}21]. Compressive [FDH21]. Computational [LSW^{+}19, RMG^{+}17, SAD^{+}17]. Computer [DH17, LYC^{+}19, RHI^{+}23, WXY^{+}19]. Computing [AY22, AHS^{+}18, BRYH21, DMdH^{+}22, EBTN22, GZC^{+}23, HMM^{+}20, HWJ^{+}21, KCK^{+}19, KGAH22, LMT19, LGL23, LKL22, LLL^{+}20, XLL^{+}21]. Concept [JSGS23]. Concept-based [JSGS23]. Concepts [CKLR21]. Conceptual [KEI^{+}18]. Concerns [AKPS18]. Concurrency [CQCH21]. Concurrent [FMZN22]. Condition [LLC^{+}20, SKM^{+}19]. Conditional [AT19]. Conditions [CKJH23, JHK^{+}21, ZZN^{+}20]. Conductive [HWJ^{+}21]. Conferencing [DWCY23]. Configurable [LRH^{+}23]. Confirmation [ATP^{+}19]. Conformant [JM18]. Confusion [HWJ^{+}22]. Connected [KFFC18, LMS^{+}23]. Connection [LEP^{+}19]. Connectivity [MBZ^{+}18, OHW^{+}20]. Consent [OSG^{+}22]. Consider [LNT^{+}19]. Consistent [JYY21]. Constellations [LH19]. Constrained [CRS^{+}18, PLC^{+}21, XFW^{+}20]. Constraints [GM19]. Constructing [CZG^{+}21]. Construction [AMG^{+}19, CLHW18, DCZ^{+}23, JSNL23, KMH^{+}21, KHW^{+}23, SRK^{+}19]. Constructions [AMG^{+}19]. Consumption [MRCdCRV^{+}21, PMGP19]. Contact [BRS21, CBT18, HYY^{+}23, LZZ^{+}21, LZZ^{+}23, SZZ^{+}20, TLF22, XLL^{+}21, HQC^{+}23]. Contact-aware [TLF22]. Contact-less [LZZ^{+}21, LZZ^{+}23]. Contactless [CXC^{+}19, HMA21, ALC^{+}20, KRS22, LDW^{+}18, LXC^{+}19, LWX^{+}22]. Contained [BKL^{+}22, KBL18]. ContAuth [CKHM20]. Content [CME^{+}19, FSW19]. Contents [JMvB^{+}19]. Context [BCK^{+}22, BPA23, CTTMT20, CML^{+}20, CWH^{+}19, CGH^{+}17, CRP21, DBY^{+}17, FTL^{+}19, GMS^{+}23, HY18, HYY^{+}23, JLZ^{+}23, KMC^{+}18, KGN^{+}20, LGM^{+}21, LKZ^{+}21, MLPGP21, MFSK21, PRC^{+}18, RLY^{+}19, RLGN19, WGL^{+}21]. Context-adaptive [LGM^{+}21]. Context-assisted [RLY^{+}19]. Context-Aware [CGH^{+}17, GMS^{+}23, HYY^{+}23, JLZ^{+}23, KMC^{+}18, CWH^{+}19, FTL^{+}19, HY18, KGN^{+}20, MFSK21, WGL^{+}21]. Contexts [KGN^{+}20, SNGK20]. Contextual [ASM^{+}18, KBKH20, PLW^{+}22]. Continual [CKHM20]. Continuous [BLW^{+}23, BXY^{+}21, CLC^{+}22, CH18, FGS^{+}23, HY22, HHX^{+}20, HQC^{+}23, HKK^{+}23, KPG^{+}20, LKBK19, LHS^{+}23, ZXW^{+}17]. Continuously [CLT^{+}21, DH17, HW17, LZL^{+}22]. Contrasting [NCM^{+}17]. Contrastive [DXS^{+}22, HEP21]. Control [AZK21, CBEG17, DWHR17, JGSC19, JSC^{+}18, SSB^{+}21, YLL^{+}20]. Controlled [DKZ^{+}22, WCW21]. Controller [BKL^{+}22]. Controlling [CXR^{+}18]. Controls [EBPA22, HMM^{+}20]. ConvBoost [SGZ^{+}23]. Convenient [CH18, DBY^{+}17]. Conversation [BAR^{+}18, LZT23].

Conversational [AvdBM^{+}22, AYT21, CSM^{+}20, FML^{+}22, JKB^{+}18, KCK^{+}19, KXAH18, LCL20]. Conversations [BRS^{+}20]. Conversing [GH18]. Converter [LRH^{+}23]. ConvNets [SGZ^{+}23]. Convolutional [GSS^{+}19, GCC^{+}21, YXW^{+}20]. Cooking [KOY^{+}19, KTEM21]. CoolMoves [AOGF^{+}21]. Cooperative [ZLL^{+}20]. Coordination [PCF^{+}23]. Coping [KGN^{+}20]. Core [BSMP23]. Core-Body [BSMP23]. CORMORANT [HFS^{+}19]. Corneal [GBS^{+}21]. Corpus [AP18]. Correction [JDZ^{+}21]. Cost [HHZ23, JWS^{+}23, JHP^{+}17, LDC18, PCBK20, CLL^{+}22, LCS^{+}21]. Cost-Effective [HHZ23]. COTS [WXW^{+}18]. Count [LKL19]. Counterfeit [WSW^{+}22]. Counterfeit-proof [WSW^{+}22]. Countering [BMR^{+}20]. Counting [EUHB23]. Countries [KVZ^{+}19, MDK^{+}22]. Counts [BS17]. Courses [GRS^{+}22]. Cover [DPW^{+}21]. Coverage [YFX^{+}20]. CovertBand [NTKG17]. COVID [LCR^{+}21, TAHH23]. COVID-19 [LCR^{+}21, TAHH23]. CPU [PM19]. Craving [GNA^{+}19]. Creating [JCF^{+}21]. Creative [PSC^{+}23]. Credibility [LP20]. Credible [WSW^{+}22]. Cricket [KNP17]. Critical [CKLR21, KKZ^{+}18]. Cross [AM19, BGK21, CLL^{+}22, DXS^{+}22, DYL^{+}19, FWJ^{+}22, GHY^{+}19, HJL^{+}22, JXZ^{+}20, LCS^{+}21, LWC^{+}22, QCWY19, RKK18]. Cross-city [DYL^{+}19]. Cross-Cultural [RKK18]. Cross-Dataset [QCWY19]. Cross-Domain [FWJ^{+}22, GHY^{+}19, LWC^{+}22]. Cross-generational [AM19]. Cross-Modal [BGK21, HJL^{+}22]. Cross-target [LCS^{+}21]. CrossGR [LCS^{+}21]. Crowd [ARD^{+}18, GCGD17, HHSM18, HDC18, KMC^{+}18, ZCL21, GJG^{+}18]. Crowd-AI [GJG^{+}18]. Crowd-enabled [HHSM18]. Crowd-Sourcing [ARD^{+}18, KMC^{+}18].

CrowdPickUp [GHvB^{+}17]. CrowdProbe [HDC18]. Crowds [PMGP19]. Crowdsensing [CS19, LLZ^{+}18, PMGP19, WXL^{+}20, ZCL21]. Crowdsourced [GOZ^{+}17, PLC^{+}21]. Crowdsourcing [AvdBF^{+}19, DGZ^{+}21, GAS^{+}18, GHvB^{+}17, HAL^{+}20, IMO^{+}18]. CrowdStory [GOZ^{+}17]. CrowdX [CLHW18]. CSI [BW20]. CSI-based [BW20]. Cue [KTM^{+}18]. Cue-based [KTM^{+}18]. CueAuth [KTM^{+}18]. Cuffless [CCJ^{+}17]. Cultural [KVZ^{+}19, RKK18]. Current [DYL^{+}19, DYCX21, FML^{+}22]. Cursor [DWHR17]. Cursor-based [DWHR17]. Custom [SRK^{+}19]. Customization [LLL^{+}20]. Customized [PRC^{+}21]. Cyber [GLW^{+}18]. Cyber-Physical [GLW^{+}18]. CycleGuard [JMC^{+}21]. Cycling [NJWFM19]. Cyclist [JMC^{+}21]. Cynics [DBBH21].

d [DWCY23, GLJ^{+}21, ASG^{+}23, AWZK23, BKL^{+}22, BGL18, CLL^{+}22, DCZ^{+}23, GLW^{+}18, HWJ^{+}21, HDFZ22, HHX^{+}20, JWS^{+}23, MQZH23, MLH^{+}23, RWW^{+}22, SLX^{+}18, ZXW^{+}17]. d/Deaf [DWCY23, GLJ^{+}21]. D3P [LDK^{+}20]. DAFI [LCW^{+}21]. Daily [BAT22, BPA23, BKM21, HEUH18, LT19, LZZ^{+}23, SSR^{+}19]. DANA [MCCH21]. Dancing [ZXY21]. Dangerous [JHL^{+}21]. DAPPER [GKO^{+}23]. Data [AT19, BSP21, BGK21, BKM21, BSH^{+}17, CLW^{+}18, CLTS21, CLY^{+}21, CND17, DKZ^{+}22, DSR^{+}22, DSSS20, DXS^{+}22, DZH^{+}23, EZC23, FYY^{+}21, FALW18, GDS^{+}20, GOZ^{+}17, GCW^{+}18, GZC^{+}23, IRA^{+}21, IYS22, IMO^{+}18, JLZ^{+}23, JLD^{+}18, KVZ^{+}19, KDL^{+}21, KFFC18, KWAP21, LHW^{+}19, LHS^{+}23, LLC^{+}20, LDSZ19, LSY^{+}18, LDK^{+}20, MCCH21, MGWM21, NJWFM19, RLGN19, RR22, RLCI22, SPRK23, SJL^{+}21, TK17, TFL^{+}19, WZF^{+}20, XKNM20, ZSG^{+}21, dGST21, VJH^{+}18].

Data-Driven [JLZ^{+}23, LHW^{+}19, WZF^{+}20, GZC^{+}23, LDK^{+}20]. Dataset [LGL23, LKL22, OWH^{+}20, QCWY19, WMS^{+}21]. Day [BRA^{+}22, DSM17, JYY21, CHC^{+}22]. Day-to-Day [BRA^{+}22]. DCW [CLZ^{+}21]. De-Identification [CLW^{+}23]. DeActive [HAR18]. Dead [CXC^{+}19, RRS^{+}23, ZM22]. Dead-Zone [CXC^{+}19]. Deaf [BA20, DWCY23, GLJ^{+}21]. Decentralised [BRS21]. Decentralized [FSJ^{+}19, WSW^{+}22]. Decontamination [LCR^{+}21]. Deep [ACB22, CS22a, CRS^{+}18, CDL^{+}19, CS19, CZP20, DCMM20, DJX^{+}23, EY20, FSX^{+}18, GBLM17, GZH^{+}21, GP17, HAR18, HHX^{+}20, HCYZ18, IN20, JZWL19, JSF^{+}18, LS20, LYL^{+}20, LLL^{+}20, LGM^{+}21, MLX^{+}20, MSWQ20, PCYZ18, RRS^{+}23, SC22, SCZ^{+}20, ZLL^{+}20]. DeepPCD [CS22a]. DeepRange [MSWQ20]. Definition [XLL^{+}21]. Deformable [BDR17]. Deformation [WS17]. DeformWear [WS17]. Degree [FSW19]. Delivery [AP18, DGZ^{+}21, DJL^{+}21, DJY^{+}22, GZW^{+}22, JDZ^{+}21, PLC^{+}21]. Demand [DJL^{+}21, DJY^{+}22, LDK^{+}20, GCW^{+}18, GZW^{+}22]. Demands [KPL20]. Demographic [XLX^{+}20]. Demonstration [JHWF22]. Demystifying [WCW21]. Dental [JCL^{+}20]. Dependent [LSSH23]. Deployable [SRK^{+}19]. Deployment [BCM^{+}23, GMY^{+}17, ILH^{+}19]. Deployments [CHZT19]. Depression [DKZ^{+}22, LSY^{+}18, SEG^{+}18]. Depth [PLK21]. Derived [BNW22]. Deriving [HP20]. Dermal [GF17]. Design [ABL^{+}23, AMG^{+}19, BGL18, BWUW21, BWH^{+}22, BCK^{+}22, BCM^{+}23, BDR17, CTTMT20, CEL^{+}22, DPW^{+}21, DBBH21, HWJ^{+}21, HSWB21, JCR22, KKV^{+}22, KJP^{+}19, KGN^{+}20, MGC^{+}23a, PRC^{+}21, PNL^{+}21, Tor23, ZLD^{+}21]. Designing [BNW22, BDM^{+}18, FALW18, KKV^{+}22, KKZ^{+}18, KTEM21, LHY^{+}20, LAX^{+}23]. Designs [FML^{+}22, MVS^{+}20]. Desks [HKSM23]. Detachable [KGL19]. Detailed [LZL^{+}22]. Detect [BLH^{+}17, BAT22, DH17]. Detecting [BFS^{+}17, BWT^{+}18, CML^{+}20, CXT^{+}21, CKL23, CCY^{+}17, CBT18, DMAH18, GH18, HBH20, IRA^{+}21, KME^{+}19, MKK^{+}21, SEG^{+}18, ZZN^{+}20]. Detection [BAR^{+}18, BRS^{+}20, BPA23, CRGR22, DSR^{+}22, DZH^{+}22, GBS^{+}21, GDS^{+}20, HCH^{+}22, HCYZ18, ILM^{+}21, JGSC19, JHL^{+}21, KDL^{+}21, LZZ^{+}20, LYLT21, LZT23, LLW^{+}22, LYWZ22, RDOA22, RHMZ^{+}21, SOL19, SWLZ21, STL^{+}19, YPJdB19]. Determinants [JHK^{+}21]. Determination [GWD^{+}17]. Determining [LDS^{+}18]. Developer [BRYH21]. Developers [EZC23]. Developing [CMV^{+}18, EZC23, LAH18]. Development [Hal19, KCK^{+}19, RQ20]. Developmental [YPJdB19]. Device [BSP21, BMZN20, CKLB22, CDG^{+}17, CYX^{+}19, CEL^{+}22, DH17, FS18, GLS^{+}18, HLS20, HHFM17, HW17, IN20, KKZ^{+}18, LRH^{+}21, LILN19, LYP^{+}22, LCW^{+}21, QCP^{+}18, WYL^{+}21, CMK22, RR22]. Device-Free [CYX^{+}19, CDG^{+}17, GLS^{+}18, LCW^{+}21]. Devices [ARSN22, ATC17, BDPH21, BMR^{+}20, BX19, CS22b, CRS^{+}18, CXZ^{+}21, CML22, CLY^{+}23, CND17, EBPA22, FXC^{+}21, FZS^{+}23, GBLM17, HLL17, HWJ^{+}21, HFS^{+}19, HAL^{+}20, HBK21, KBKH20, LSSH23, LFH18, LLF^{+}21, LJK^{+}21, LLZ^{+}22, LHY^{+}20, MH23, MWJ^{+}21, RKK18, RDOA22, RSSN21, WS17, YZ21, YLMR20, ZLL^{+}20, ZWZ^{+}21b]. DeXAR [ACB22]. Diabetes [BSP21]. Diagnosing [WCY^{+}19]. Diagnosis [AGU^{+}22]. Diagnostic [PNL^{+}21]. Diaphragmatic [GZNL22]. Diaries [CWS^{+}19]. Diary [BRA^{+}22]. Didn’t [BA20]. Diet [NNMY23]. Difference [ISSK20]. Differences [RKK18]. Different [GRS^{+}22]. Differentiable [LL22].

Differential [KBXP23]. Differentiating [MMG^{+}19]. Digital [BML17, CBZN19, GZY^{+}21, LCC^{+}22, WMW^{+}20]. Digitally [DJL^{+}22]. Dimension [DLD^{+}21, MCCH21]. Dimension-Adaptive [MCCH21]. DIO [AMG^{+}19]. Direct [CLZ^{+}21, MQZH23]. Direct-Circuit-Writing [CLZ^{+}21]. Directed [SKM^{+}19]. Disabilities [MW22]. Disasters [WTA^{+}20]. Discovering [AJA^{+}18, ASM^{+}18]. Discriminate [AES^{+}21]. Discrimination [LLZ^{+}20]. Discriminative [LWC^{+}22]. Disease [BLW^{+}23, LAS^{+}19]. Disengagement [PL23]. Disentangled [JSNL23, SWLG22]. Disorder [LAX^{+}23]. Disorders [MBP^{+}17]. DisPad [CJF^{+}23]. Displacement [CJF^{+}23]. Display [BDR17, CME^{+}19, HSH^{+}23, KCFI19, LHS^{+}21]. Displays [GO22, KTM^{+}18, KSC^{+}17, MOD21, PTK18, RZ21, SFR^{+}20]. Dissecting [CWP^{+}18]. DistFL [LCZ^{+}21]. Distilling [LZZ^{+}23]. Distinct [SvBV^{+}18]. Distraction [KGN^{+}20]. Distributed [FDL^{+}18, ZSG^{+}21]. Distribution [HXH^{+}18, LCZ^{+}21]. Distribution-aware [LCZ^{+}21]. Diverse [ZPPC^{+}20]. Diversity [DSSD23]. DLP [LH20]. Do [KML^{+}21]. Does [CGH^{+}17, LP20, LKL19, PTK18]. Dog [GZC^{+}23]. Doing [AYT21]. Domain [BCK^{+}22, BGK21, DZZ^{+}21, FWJ^{+}22, JSNL23, LCW^{+}21, MBB20, GHY^{+}19, LWC^{+}22]. Don’t [LLPH22]. Door [DMAH18]. Doppler [BGK21, SCZ^{+}20]. Doppler-based [BGK21]. Dot [BGR^{+}20]. Douleur [JCF^{+}21]. Downlink [KSCS19]. DrawingPresence [MTH19]. DRG [LHY^{+}22]. DRG-Keyboard [LHY^{+}22]. Drift [DYL^{+}19]. Drink [AZK21]. Drinking [BFS^{+}17, MLPGP21]. Drinks [JMvB^{+}19, PMGP19]. Drip [LXW^{+}21]. Driven [ARD^{+}18, AYM^{+}22, GLJ^{+}21, JLZ^{+}23, KCK^{+}19, LHW^{+}19, WZF^{+}20, WAL^{+}19, GZC^{+}23, LDK^{+}20]. Driver [BW20, BCK^{+}22, JNB^{+}20, LKZ^{+}21].

Drivers [BBH17, HSL^{+}20, KCP^{+}18, KML^{+}21, WMS^{+}21]. DriverSonar [JHL^{+}21]. Driving [CS19, CJRG21, JHL^{+}21]. Drone [AZEL17, ASG^{+}23, KDL^{+}21, LCC^{+}22, RPB^{+}18, YLL^{+}20]. Drone-based [ASG^{+}23]. DronePrint [KDL^{+}21]. Drones [DZH^{+}22]. Drop [CRMG22]. DropMonitor [LXW^{+}21]. Drowsiness [RMG^{+}17, SAD^{+}17]. Drug [GNA^{+}19]. Drug-Seeking [GNA^{+}19]. Dry [DH17]. Dual [DSSD23, DWHR17, ILM^{+}21, LYQ^{+}21, LHY^{+}22, LCG^{+}22, LHY^{+}20]. Dual-Attention [LCG^{+}22]. Dual-layer [DSSD23]. Dual-radio [ILM^{+}21]. Dual-Sided [DWHR17]. DualBlink [DH17]. DualRing [LYQ^{+}21]. Duco [CLZ^{+}21]. Duet [VJH^{+}18]. During [HBX^{+}17, SGO^{+}23, CH18, DGS18, HSH^{+}23, JLZ^{+}23, TAHH23]. Dwell [IYS22]. Dyadic [KLK^{+}20]. Dynamic [GMS^{+}23, GCW^{+}18, JZWL19, LLG^{+}23, RBL^{+}20]. Dynamics [CYK^{+}23a, GH18, XL19]. Dysphagia [AZK21].

Ear [BWT^{+}18, GWP^{+}19, GJC^{+}21a,

JGX^{+}22, KSC^{+}21, SXF^{+}22]. Ear-mounted [BWT^{+}18]. Ear-worn [KSC^{+}21]. Earable [CCY^{+}23, LZL^{+}22]. Earables [VBS^{+}21]. EarAcE [CCY^{+}23]. EarBit [BLH^{+}17]. Earbuds [GZH^{+}21]. EarCommand [JGX^{+}22]. Earcons [BBH17]. EarEcho [GWP^{+}19]. EarIO [LZL^{+}22]. Early [JYY21, RDOA22]. EarlyScreen [KSZ^{+}22]. Earmonitor [SXF^{+}22]. Earphones [JGZ^{+}21, SXF^{+}22, CGJ^{+}22]. Ears [MJ19]. eat2pic [NNMY23]. Eating [BLH^{+}17, BWT^{+}18, CFF^{+}22, CBT18, CWS^{+}19, NNMY23, ZZN^{+}20]. Eating-Painting [NNMY23]. Eavesdropping [CYK^{+}23a]. ECG [PCBK20, WXW^{+}18]. Echo [GWP^{+}19, BGR^{+}20]. EchoSpot [LLCY21]. EchoWhisper [GJL^{+}20]. Ecological [CSK^{+}18, PLW^{+}22]. Economy [CMV^{+}18, WXL^{+}20]. EDEN [KBBN21]. Edge [DZH^{+}22, EBTN22, GLC^{+}21, JCL^{+}20, LLL^{+}20, ZLL^{+}20]. Edge-Case [EBTN22]. Editorial [AKS^{+}17]. EduSense [AKX^{+}19]. Effect [AJV20, BSH^{+}17, CQCH21, CSHN17, LYWZ22, SvBV^{+}18]. Effective [CTTMT20, HP20, HHZ23, RZ21, WGL^{+}21]. Effectiveness [DLH^{+}18, NZL^{+}20]. Effects [ASV^{+}18, CJRG21, CRGR22, CMF^{+}23, GRS^{+}22, HKK^{+}23, SvBH^{+}19]. Efficacy [LH19]. efficiency [GLC^{+}21]. Efficient [ASC^{+}22, CKLB22, CLTS21, GwC^{+}23, ILM^{+}21, NTT^{+}21, PLC^{+}21, PRS^{+}23, RR22, KPK^{+}18, RDOA22]. Effort [HAY^{+}17]. EFRing [CLYZ22]. Eggly [LAX^{+}23]. Egocentric [MLH^{+}23]. Eight [MDK^{+}22]. Elderly [KGL^{+}18]. Electric [CLYZ22, GPW^{+}18, LHW^{+}19, LDK^{+}20, MGY^{+}23, WZF^{+}20]. Electrical [KAH18, NTT^{+}21]. Electro [GF17]. Electrocardiogram [CLC^{+}22]. Electrodermal [DGS18, GDS^{+}20]. Electromagnetic [LYP^{+}22, LLW^{+}22, PWP19]. Electromyography [FFT^{+}18]. Electronic [LRH^{+}23]. Electronics [LRH^{+}21]. EMA [KME^{+}19, PLW^{+}22]. Emanations [LLW^{+}22]. Embedded [GBLM17, LLW^{+}22, PTMM^{+}22, SFR^{+}20]. Embedding [CPC^{+}18, CWH^{+}19, JWS^{+}23, LCG^{+}22, SJL^{+}21, XLX^{+}20]. Embeddings [LT19]. Emergency [KEI^{+}18]. EMG [GMMH21]. Emotion [BCK^{+}22, HPWW18, LKZ^{+}21, PRS^{+}23]. Emotional [CSK^{+}18, DCMM20, DGS18, FML^{+}22, GSRS20, GCP^{+}22]. Emotions [CGJC19, FDH21]. Empathetic [KLK^{+}20, LKZ^{+}21]. Empirical [BRA^{+}22, WMW^{+}20]. Empowering [BBBN22, CCY^{+}23]. EMS [HDW^{+}17]. EMS-based [HDW^{+}17]. Enable [GH22, LL22]. Enabled [CLW^{+}23, DLD^{+}21, JZWL19, LLZ^{+}18, HHSM18]. Enabling [CS22a, CW18, CLYZ22, GMY^{+}17, JGSC19, KBC^{+}20, LYQ^{+}21, LHY^{+}22, LGY^{+}20, LCH^{+}22, MZT18, PLK21, RPF^{+}18, ZWX^{+}20]. Encoding [LH20]. Encounters [CLHW18]. Encourage [MGC^{+}23b]. Encouraging [CYK^{+}19]. End [CND17, LLF^{+}21, PRC^{+}21, RQ20, ZLL^{+}20, ZWZ^{+}21a]. End-to-End [LLF^{+}21, PRC^{+}21, ZWZ^{+}21a]. End-User [CND17, RQ20]. Endowing [FZZ^{+}22]. Energy [AWZK23, ASC^{+}22, FDL^{+}18, GPW^{+}18, ILM^{+}21, JSC17, KASC21, KPK^{+}18, LLZ^{+}22, PLC^{+}21, PCF^{+}23, PSC^{+}23, RDOA22]. Energy-aware [LLZ^{+}22]. Energy-Ball [FDL^{+}18]. Energy-Constrained [PLC^{+}21]. Energy-Efficient [ASC^{+}22, ILM^{+}21, KPK^{+}18, RDOA22]. Enforcing [KBBN21]. Engagement [AJA^{+}18, DGS18, DBW^{+}22, GSRS20, GRS^{+}22, HSH^{+}23, HKK^{+}18, KCuH^{+}19]. EngageMon [HKK^{+}18]. Engaging [KXAH18, MGWM21]. Engine [KOY^{+}19]. Enhance [AHB17, JCF^{+}21]. Enhanced [HHSM18, OWH^{+}20, KHG^{+}19, XLX^{+}20]. Enhancement [DWL^{+}22, FYY^{+}21, JWK^{+}21, LLZ^{+}22, ZWZ^{+}21b]. Enhancing [CLHW18, CS19, DJL^{+}22, OSG^{+}22, VS19]. Enriching [LFL^{+}23]. Ensemble [FSX^{+}18, HY18]. Ensembles [GP17]. Entangled [YHW^{+}18]. Entity [LCL20]. Entropy [DSSS20]. Entry [GYL^{+}20, LLL^{+}19, LHFZ22, LYY^{+}17]. Environment [CYB^{+}23, Hal19, HQY^{+}18, LCH^{+}22, MKK^{+}21, RTZ^{+}20, SHZ^{+}20, ZAP17]. Environment-aware [CYB^{+}23]. Environment-Independent [HQY^{+}18, LCH^{+}22, SHZ^{+}20]. Environments [AHS^{+}18, ACB22, BLH^{+}17, CYB^{+}23, CMK22, GCGD17, JM18, KJS20, KOY^{+}19, KPG^{+}20, KPK^{+}18, LMS^{+}23, RRS^{+}23, WXY^{+}19]. EOG [RLM^{+}19].

EOG-based [RLM^{+}19]. Epidemic [HFX^{+}21]. Epidermal [DHC^{+}18, KHW^{+}23]. Episodes [BFS^{+}17, BLH^{+}17, BWT^{+}18, CBT18, JM18]. equipped [BMR^{+}20]. Error [AGU^{+}22, BSH^{+}17, HJL^{+}22, JHWF22]. Errors [AYM^{+}22]. ESPRESSO [DSSS20]. Estimate [AVD^{+}17, BS21, BS22, SSB^{+}21]. Estimating [HDFZ22, VJH^{+}18]. Estimation [CDG^{+}17, CLY^{+}21, GKO^{+}23, LNT^{+}19, MLH^{+}23, RWW^{+}22, YZ21]. Euphoria [GNA^{+}19]. Evaluating [BDM^{+}18, ELD19, HJL^{+}22, LGM^{+}22, LHY^{+}20]. Evaluation [AMG^{+}19, CBZN19, CEL^{+}22, CBEG17, CBR23, JWCC^{+}20, KYJ^{+}23, LEP21, PNL^{+}21, RRS^{+}23, RHMZ^{+}21]. Evening [RLY^{+}19]. Event [GOZ^{+}17, HYH21]. Events [DMAH18]. Every [BDPH21, BS17, JYY21]. Everyday [AWZK23, CLZ^{+}21, DCMM20, JMvB^{+}19, KLK^{+}20, KFFC18, MWJ^{+}21, OSG^{+}22, PTMM^{+}22]. Evidence [JHK^{+}21]. Evolution [OGG^{+}18]. evolutionary [LGM^{+}21]. Evolving [JSC^{+}18]. Examination [JHP^{+}17]. Examining [MLPGP21, SKM^{+}19]. Example [CLW^{+}23]. Exercise [KNL^{+}23]. Exercises [RPF^{+}18]. Exergy [PSC^{+}23]. Existing [LYD^{+}23]. Exit [RDOA22]. Expanding [LDK^{+}20]. Expectations [FDH21, TKF^{+}19]. Experience [GRS^{+}22, JCF^{+}21, MTH19, NTO^{+}18, RQ20]. Experiences [CSK^{+}18, CYK^{+}19, GCM^{+}19]. Experiencing [KAH18]. Experiments [DDP^{+}17]. Explain [KYJ^{+}23]. Explainable [ACB22, JSGS23]. Explanations [KYJ^{+}23]. Exploiting [LYWZ22]. Exploration [CYK^{+}19, GFK^{+}18, KCuH^{+}19]. Exploratory [KML^{+}21]. Exploring [AVD^{+}17, AZEL17, BWUW21, BBH17, BML17, CYK^{+}19, CJRG21, DPW^{+}21, FCFW^{+}17, GJL^{+}20, GKO18, HKSM23, HCH^{+}22, HSWB21, IAE^{+}19, JYY21, JWK^{+}21, KJKL19, KMK^{+}19, LH19, LCC^{+}22, LFL^{+}23, LKZ^{+}21, WTA^{+}20, ZCN^{+}20]. ExpressEar [VBS^{+}21]. Expressions [CLT^{+}21, CGJ^{+}22, GJC^{+}21b, VBS^{+}21]. Expressive [LYQ^{+}21]. Extending [IY20, YFX^{+}20]. External [CBR23]. Externalities [JHP^{+}17]. Extracting [CPC^{+}18, DSM17, YHW^{+}18]. Extraction [ASC^{+}22, KTH^{+}20, LS20]. Eye [AKN^{+}19, AIP^{+}18, CLY^{+}23, LSSH23, LLZ^{+}18, LLWX21, MFSK21]. Eye-Movement [LSSH23]. Eyeglass [CRP21, RMG^{+}17, SAD^{+}17]. Eyes [CCY^{+}17, DH17, LYY^{+}17]. Eyes-Free [LYY^{+}17]. EyeSpyVR [AIP^{+}18]. Eyewear [CZD^{+}21, DWHR17, RLM^{+}19, XZZ21].

FabHandWear [PRC^{+}21]. Fabric [CJF^{+}23]. Fabricating [KBL18]. Fabrication [PRC^{+}21, SRK^{+}19, SFR^{+}20]. Fabrics [LSW^{+}19]. Face [CRB^{+}21, KSC^{+}21, KLT19, LZT23, SWLZ21]. Face-To-Face [LZT23]. FaceBit [CRB^{+}21]. FaceSense [KSC^{+}21]. Facial [CLT^{+}21, CYK^{+}23a, CGJ^{+}22, GJC^{+}21b, LZL^{+}22, LYF^{+}21, RLM^{+}19, SNGK20, VBS^{+}21, XZZ21]. Facilitating [LCR^{+}21]. Facilitators [KW23]. Factor [ATC17, CYK^{+}23b, CNXG22]. Factors [CTTMT20]. Factory [QWK^{+}19, XKNM20]. Fair [LWA23]. FairCharge [WZF^{+}20]. Fairness [WZF^{+}20]. Fairness-Aware [WZF^{+}20]. Fall [LYLT21]. Families [BGR^{+}20]. Family [LFL^{+}23]. FarmChat [JKB^{+}18]. Farmer [JKB^{+}18]. FarSight [NVP18]. Fast [ASG^{+}23, JSNL23, LDK^{+}20, RR22, WGL^{+}21, WRLK22, ZSG^{+}21]. Fast-Moving [WRLK22]. Fatigue [BOH^{+}21, RMG^{+}17, SAD^{+}17]. Favorable [CKJH23]. Feasibility [BOH^{+}21, SS22].

Feature [KLM^{+}18]. Features [IAE^{+}19, JWK^{+}21, LSSH23, LS20, SNGK20, SC22, WTA^{+}20, YZS^{+}20]. Federated [CMK22, EBTN22, FRS^{+}20, GLC^{+}21, KBBN21, LWA23, LCZ^{+}21]. Feedback [CFZ^{+}19, CEL^{+}22, HPWW18, MFSK21, MGC^{+}23b, RPB^{+}18, XYG^{+}20]. Feel [CGJ^{+}22, HPWW18]. Feeling [HSH^{+}23]. Fever [BSMP23]. FeverPhone [BSMP23]. FewShotBP [FGS^{+}23]. FG [LZZ^{+}21]. FG-LiquID [LZZ^{+}21]. Fi [CML22, FWJ^{+}22, GZZ^{+}21, HDC18, LLZ^{+}20, LCS^{+}21]. Fiber [KHG^{+}19, LYD^{+}23, SFR^{+}20]. Fiber-enhanced [KHG^{+}19]. Fiber-Optic [LYD^{+}23]. Fibrillation [DSR^{+}22]. Field [CLYZ22, GPW^{+}18, LHS^{+}21, MGC^{+}23b, PTK18, SSB^{+}21, CHZT19, HQC^{+}23]. Finding [LHS^{+}21, PRC^{+}18]. Fine [CXR^{+}18, GOZ^{+}17, JHH21, JHL^{+}21, LLLX22, LZZ^{+}21, LLC^{+}20, LYD^{+}23, MLX^{+}20, VBS^{+}21, WXL^{+}20]. Fine-Grain [CXR^{+}18]. Fine-Grained [GOZ^{+}17, JHH21, JHL^{+}21, LYD^{+}23, MLX^{+}20, VBS^{+}21, LLLX22, LZZ^{+}21, LLC^{+}20, WXL^{+}20]. Finger [BDM^{+}18, CLYZ22, EUHB23, HDFZ22, LLL^{+}19, PWP19, SZZ^{+}20, ZXW^{+}17]. Finger-Counting [EUHB23]. Finger-Mounted [SZZ^{+}20]. Finger-Worn [BDM^{+}18]. Fingernail [IY20]. Fingerprint [HDFZ22, LHFZ22, TFL^{+}19, WXY^{+}17]. Fingerprinting [FZS^{+}23]. FingerReader2.0 [BDM^{+}18]. Fingers [CLD^{+}20, GKO18]. Fingertip [LHY^{+}22]. FingerTrak [HHX^{+}20]. Fire [WZG^{+}19]. Firmware [CWP^{+}18]. First [FCQR20]. Fitbit [CWP^{+}18, LP20]. Fitness [BNW22, CWP^{+}18, GLS^{+}18, RR22]. Five [KVZ^{+}19]. FLAME [CMK22]. Fleets [LHW^{+}19, WZF^{+}20, XFW^{+}20]. Flexible [AZS^{+}18, BRYH21, CJF^{+}23, FZZ^{+}22, GPB^{+}23]. Flicker [KKZ^{+}18]. Floor [CZG^{+}21, EEY18, EY20, FMZN22].

Floorplan [CLHW18]. Flows [NTO^{+}18]. FlowSense [CZL^{+}22]. Flu [DGLQ19]. Fluctuations [DSM17, MTH19]. Fluid [HEUH18]. FluidMeter [HEUH18]. FluSense [ALC^{+}20]. Fly [CMF^{+}23]. FMT [LCFT19]. FocalPoint [MQZH23]. Focus [WXY^{+}19]. Focused [BRYH21, JCR22]. Folded [LGM^{+}22]. Food [BS21, BS22, CWS^{+}19, HHZ23, JWCC^{+}20, MRCdCRV^{+}21]. Foot [ASTW19, HDW^{+}17]. Foot-Mounted [ASTW19]. FootNotes [GFK^{+}18]. Footprint [TFL^{+}19]. Footprints [WMW^{+}20]. Footstep [HW21]. FootStriker [HDW^{+}17]. Foraging [MOD21]. Force [GPB^{+}23, LLL^{+}19]. ForceSticker [GPB^{+}23]. Forearm [FFT^{+}18]. Forecasting [AKB23, LS20, OGG^{+}18, WZG^{+}19]. Forensics [LCC^{+}22]. Forma [KHSW17]. Fortifying [CXZ^{+}21]. FORTNIoT [CVL20]. Foster [HKSM23]. Foundations [JWCC^{+}20]. Four [BWH^{+}22]. Frames [GGJ^{+}22]. Framework [AYM^{+}22, CKHM20, FRS^{+}20, HNL22, JHH21, KME^{+}19, LAS^{+}19, LYL^{+}20, LLL^{+}20, SEG^{+}18, SJL^{+}21, WGL^{+}21]. Free [CYX^{+}19, GZL^{+}21, GKO^{+}23, HLL17, LYY^{+}17, MWJ^{+}21, SRR^{+}20, TKGS17, ZZN^{+}20, ZWZ^{+}21a, CZX^{+}22, CDG^{+}17, GLS^{+}18, HNL22, KBC^{+}20, KGAH22, LCW^{+}21]. Free-Hand [SRR^{+}20]. Free-Living [ZZN^{+}20]. Free-style [ZWZ^{+}21a]. Freehand [TLF22]. Frequency [BOH^{+}21, KKZ^{+}18, WYL^{+}21]. Fresnel [LLG^{+}23]. Friendly [EZC23, LAH18]. Frisson [HCH^{+}22]. Fruit [LJK^{+}21]. Fruits [LJK^{+}21]. Frustration [HKK^{+}23]. Frying [KOY^{+}19]. Full [CLT^{+}21, DLD^{+}21]. Full-Dimension [DLD^{+}21]. Fully [RGL^{+}18]. Fully-Analog [RGL^{+}18]. Function [KJP^{+}19]. Functional [OHW^{+}20, PRC^{+}21]. Functioning [AKB23, HMM^{+}20]. Fusing [DCZ^{+}23]. Fusion [AZK^{+}23, CML22, DZH^{+}22, EEY18, GZH^{+}21, GOZ^{+}17, KSZ^{+}22, LYL^{+}20, WMS^{+}21]. Fusion-based [EEY18]. Future [FML^{+}22, KEI^{+}18, LCR^{+}21].

Gage [GSRS20]. Gain [WXY^{+}17]. Gait [ASTW19, CLL^{+}22, RHI^{+}23]. Game [IAE^{+}19]. Game-Based [IAE^{+}19]. Games [HKK^{+}18, LAX^{+}23]. Gamified [STL^{+}19]. Garages [CLH18]. Garment [AWSS22, FDH21]. Gated [LYH^{+}20]. Gauging [HEUH18]. Gaze [CCY^{+}17, HBK21, IYS22, KTM^{+}18]. GC [HNL22]. GC-Loc [HNL22]. General [BCM^{+}23, GLG^{+}19, LH19]. General-Purpose [BCM^{+}23, LH19]. Generalizability [LWYZ21]. Generalizable [LWC^{+}22]. Generalization [MDK^{+}22]. Generalizing [SNGK20]. Generate [MGY^{+}23]. Generated [LCL17]. Generating [DZH^{+}23]. Generation [CS22b, GOZ^{+}17, HYS^{+}23, LKBK19, LNAH21, SPRK23]. generational [AM19]. Generative [CHWZ23]. Generic [ZLD^{+}21]. Genuine [SLX^{+}18]. Geo [GFK^{+}18]. Geo-referenced [GFK^{+}18]. Geographic [JHP^{+}17]. Geographical [KCuH^{+}19]. GeoLifecycle [KCuH^{+}19]. Geomagnetic [NHL^{+}20]. Geometry [CLD^{+}20]. Geophones [PCBK20]. Geryon [DZH^{+}22]. Gesture [AWSS22, CLD^{+}20, FWJ^{+}22, GZZ^{+}21, GLX^{+}22, JGZ^{+}21, LCS^{+}21, LHY^{+}22, LCL17, LWZ^{+}20, LCH^{+}22, PSLS21, SHZ^{+}20]. Gestures [EUHB23, KTM^{+}18, LGM^{+}22, LHY^{+}20, PLX^{+}18, EUHB23]. Get [YLL^{+}20]. Getting [MT19]. GHz [SHZ^{+}20]. GIobalFusion [LYL^{+}20]. Glabella [HW17]. Glasses [FZZ^{+}22, MLMH^{+}18, MFSK21]. GlassMessaging [JGZ^{+}23]. Glaze [KSCS19]. Global [LYL^{+}20]. GlucoMine [BSP21]. GlucoScreen [WPH^{+}23]. Glucose [GZZ^{+}17, WPH^{+}23]. Go [RHMZ^{+}21, RZ21]. Goal [AJV20, SKM^{+}19]. Goal-Directed [SKM^{+}19]. GoalKeeper [KJKL19]. Goals [AJV20]. Good [BA20, CKP^{+}20]. Google [AYT21]. GoPose [RWW^{+}22]. Gossip [BBBN22]. Government [CLY^{+}21]. GPS [NJWFM19, ZM22]. GPS-assisted [ZM22]. Gradient [WXY^{+}17]. Grain [CXR^{+}18]. Grained [GOZ^{+}17, JHH21, JHL^{+}21, LYD^{+}23, MLX^{+}20, VBS^{+}21, LLLX22, LZZ^{+}21, LLC^{+}20, WXL^{+}20]. Grammar [ARD^{+}18]. Grammar-Driven [ARD^{+}18]. Grandchildren [LFL^{+}23]. Grandparents [LFL^{+}23]. Graph [CWH^{+}19, CLH18, GZY^{+}21, HNL22, LCG^{+}22, LYH^{+}20, YXW^{+}20]. Graph-Based [GZY^{+}21, CLH18]. Graphics [LHS^{+}21]. Grasped [FFT^{+}18]. Grating [SFR^{+}20]. Grocery [BMZN20]. Group [GRS^{+}22, GH18, HSH^{+}23]. Group-wise [GRS^{+}22]. Groups [GH18, PSC^{+}17]. Guarantee [HHSM18]. Guard [CLC^{+}22]. Guardians [DBBH21]. Guessing [LCL17]. Guidance [CEL^{+}22, JHWF22, LHS^{+}21]. Guided [LNAH21, NHL^{+}20, PZH^{+}18]. Guiding [KIT^{+}20]. Gym [RPF^{+}18].

Hall [GGJ^{+}22, CSHN17]. Hall-Effect [CSHN17]. Hand [CYK^{+}23b, FFT^{+}18, HHX^{+}20, JJB22, KBL18, LYQ^{+}21, LLL21, LHY^{+}20, PRC^{+}21, SRR^{+}20, LLL21]. Hand-to-Hand [LHY^{+}20]. Handcrafted [SC22]. Handed [LLL^{+}19, YLL^{+}20, EBPA22]. Handheld [CS22b, LYY^{+}17]. Handling [AYM^{+}22, KKL21, LYH^{+}20]. Handwriting [BXY^{+}21, DJL^{+}22, FZZ^{+}22, OWH^{+}20, YZS^{+}20, ZWZ^{+}21a]. Handwriting-Assistant [BXY^{+}21]. Handwritten [DWZZ19]. Hanging [CLZ^{+}21]. Happiness [BKM21]. Haptic [CQCH21, DPB21, FMP^{+}23, CEL^{+}22]. Hard [BA20, DWCY23, GLJ^{+}21, SZJ^{+}17]. Hard-of-hearing [SZJ^{+}17]. Harnessing [TK17]. Harvester [GPW^{+}18]. Harvesting [AWZK23, MGC^{+}23a, PSC^{+}23]. Hawkes [OGG^{+}18]. HCI [BWH^{+}22]. Head [GO22, HYH21, LHS^{+}21, RZ21, YZ21]. Head-Mounted [RZ21]. Headphone [CHC^{+}22]. Headsets [AIP^{+}18, LLL^{+}19]. Heal [GZC^{+}23]. Health [AJA^{+}18, CPK^{+}19, JLZ^{+}23, JHK^{+}21, LS20, LCX^{+}22, LDS^{+}18, LLC^{+}20, MGC^{+}23b, PL23, RLGN19, RLCI22]. Healthier [HKSM23, NNMY23]. HealthWalks [LLC^{+}20]. Healthy [CWS^{+}19, HMM^{+}20]. Hear [CPW^{+}21, ZWZ^{+}21b]. Hearing [BA20, DWCY23, GLJ^{+}21, JGX^{+}22, SZJ^{+}17]. Heart [CLC^{+}22, HCYZ18, WXW^{+}18, XLL^{+}21, HQC^{+}23]. HeartQuake [PCBK20]. HeartSteps [LGKM20]. Heat [AVD^{+}17]. Hello [CKP^{+}20]. Help [MSTT22, WBG^{+}19]. Helping [MT19, WBG^{+}19]. HERMAS [SJL^{+}21]. Heterogeneity [LNT^{+}19]. Heterogeneous [CWH^{+}19, DSSS20, GKO^{+}23, LRH^{+}23, LSY^{+}18, LWYZ21]. Heuristic [BRYH21]. HeyTeddy [KCK^{+}19]. Hidden [LLW^{+}22]. Hierarchical [CHWZ23, KNP17, LWA23]. High [XLL^{+}21]. High-Definition [XLL^{+}21]. Higher [MMG^{+}19]. Highly [BDR17, CRGR22, KYJ^{+}23]. HMGAN [CHWZ23]. Holding [YPJdB19]. Holistic [KNL^{+}23]. Home [ASM^{+}18, ACB22, BMR^{+}20, CND17, CVL20, HAY^{+}17, HAL^{+}20, JSC^{+}18, KTEM21, LWZ^{+}20, LSY^{+}23, RTZ^{+}20, SCZ^{+}20]. Homes [HNCP22, VJH^{+}18]. Hometown [DYL^{+}19]. Honeysuckle [LNAH21]. Hop [MZST17]. Hospital [ALC^{+}20]. Hotness [HY18]. Hotspot [ELD19]. House [SCZ^{+}20]. Human [AES^{+}21, AZEL17, AT19, ARSN22, AZK^{+}23, BSH^{+}17, CLTS21, CML22, CHWZ23, CKJH23, DZH^{+}23, DCZ^{+}23, FSX^{+}18, FSJ^{+}19, FYY^{+}21, FRS^{+}20, GwC^{+}23, HEUH18, HEP21, HEP22, HP20, HNCP22, HBH20, HWZZ19, JTM^{+}22, JCR22, JSGS23, JSF^{+}18, KHZ22, KTH^{+}20, LGL23, LWA23, LZZ^{+}23, LSW^{+}19, LLG^{+}23, LYC^{+}19, LWYZ21, MBB20, MFSK21, PCYZ18, RH19, RWW^{+}22, RBL^{+}20, SOL19, SH20, SJL^{+}21, SWLG22, YFX^{+}20, PRS^{+}23]. Human-Computer [LYC^{+}19]. Human-Drone [AZEL17]. Hybrid [FSW19]. Hygiene [ASS^{+}19]. Hyper [AvdBM^{+}22]. Hyperactivity [AAM^{+}23]. HyRise [EEY18].

I3 [LYC^{+}19]. ICT [CHZT19]. Id [HWZZ19]. IDE [LAH18]. Ideas [FML^{+}22]. ID’em [CPC^{+}18]. Identification [CDG^{+}17, CLY^{+}21, CML22, CLW^{+}23, DYCX21, FGL18, FFQ^{+}20, FXC^{+}21, HWZZ19, HW21, KDL^{+}21, LCC^{+}22, RPB^{+}18, SYYL22, BRS21, CLL^{+}22, KBBN21]. Identifier [LZZ^{+}21]. Identifying [ATQ^{+}21, ASS^{+}21, CWS^{+}19, FMZN22]. Identity [VJH^{+}18, FXC^{+}21]. IDrone [RPB^{+}18]. If [DBY^{+}17, LCL20]. Ignorance [DWCY23]. iLid [RMG^{+}17, SAD^{+}17]. Illness [ALC^{+}20]. ILLOC [GGJ^{+}22]. Illumination [KSC^{+}17]. Image [DCMM20, DCZ^{+}23, HDFZ22, HBH20]. Images [LLW^{+}22]. Imaging [AVD^{+}17, AKN^{+}19, RSSN21]. IMar [HY22]. Immersive [RPF^{+}18]. iMon [HBK21]. Impact [CFZ^{+}19, ELD19, KVZ^{+}19]. Impacts [FYW^{+}19]. Impaired [AKPS18, BMZN20, LGY^{+}20]. Impairment [XYG^{+}20]. Impairments [BDM^{+}18, JXZ^{+}20, KLT19]. Implants [BA20]. Implementation [BDR17]. Implications [YPJdB19]. Improve [CVL20, DPW^{+}21, RLY^{+}19]. Improvements [KEI^{+}18]. Improving [BSP21, CGJC19, ZLB^{+}17]. IMU [BGK21, GMMH21, JWK^{+}21, KWAP21, LHS^{+}23, LYQ^{+}21, LHY^{+}22, LGY^{+}20, MFSK21, OWH^{+}20, SPRK23, SZZ^{+}20]. IMU-Enhanced [OWH^{+}20]. IMUTube [KTH^{+}20]. In-Air [GO22]. In-App [LNAH21]. In-Car [PZH^{+}18]. in-class [GSRS20]. In-ear [SXF^{+}22]. In-field [CHZT19]. In-Hall [GGJ^{+}22]. In-Home [HAY^{+}17]. In-packet [SYYL22]. In-Situ [BRA^{+}22, LSY^{+}23, RLCI22]. In-Store [FRB^{+}21]. In-the-wild [KVZ^{+}19, SPRK23, BOH^{+}21, ASS^{+}19, KME^{+}19]. In-Vehicle [JCR22, KML^{+}21, KCP^{+}18, KPL20, ZCC^{+}20]. INAGT [WMS^{+}21]. Inattentional [HSWB21]. Inaudible [LYLT21]. Incentivized [AJA^{+}18]. Inclusive [DWCY23]. Incomplete [VJH^{+}18]. Inconsistency [LJZ^{+}22]. Incremental [MBB20]. Independent [GZZ^{+}21, GJC^{+}21a, HQY^{+}18, KLM^{+}18, LCH^{+}22, LYCW20, SHZ^{+}20, EY20, SZJ^{+}17]. Index [CDL^{+}19, CLYZ22]. Indicators [ATQ^{+}21]. Individual [BRS21, GRS^{+}22, LLC^{+}20, PB20]. Individualized [CWS^{+}19, LSSH23]. Indoor [ARD^{+}18, CS22a, CS22b, CLHW18, CYB^{+}23, DJL^{+}21, DJY^{+}22, GAS^{+}18, GZW^{+}22, HNL22, HXH^{+}18, JM18, KJS20, KASC21, KPK^{+}18, LCW^{+}21, MMMS^{+}20, NHL^{+}20, PBM^{+}18, VS19, ZCA^{+}21, ZM22]. Induced [MGY^{+}23]. Inducing [CKJH23]. Inductive [CPC^{+}18]. Industrial [LYWZ22]. Inertial [ASS^{+}19, ASS^{+}21, AZK^{+}23, BRS^{+}20, BXY^{+}21, GZH^{+}21, VS19]. Infant [YPJdB19]. Infants [GLG^{+}19]. Infarction [RDOA22]. Inference [CSHN17, IN20, JLZ^{+}23, LKZ^{+}21, LYC^{+}19, MLX^{+}20, MDK^{+}22, XLX^{+}20, ZLL^{+}20]. Inferring [ASS^{+}19, JLD^{+}18, MRCdCRV^{+}21, MOD21, SSW^{+}17]. Influence [KPL20]. Influences [HPWW18]. Influenza [ALC^{+}20]. Influenza-Like [ALC^{+}20]. InfoPrint [JWS^{+}23]. Informatics [DDP^{+}17, ECF^{+}20, FALW18, KW23, MW22]. Information [AvdBM^{+}22, AT19, AKPS18, AP18, AYM^{+}22, CPC^{+}18, CLW^{+}18, HWJ^{+}23, JWS^{+}23, KCFI19, KL19, LYL^{+}20, MJ19, NTKG17, QWK^{+}19].

InformationSense [BDR17]. Infrared [JMvB^{+}19, KGL^{+}18]. Infrastructure [BCM^{+}23, GAS^{+}18, HNL22, MZJ20]. Infrastructure-free [HNL22]. Infusion [LXW^{+}21]. Inherent [KBXP23]. Inkjet [CND^{+}20]. Inkjet-Printed [CND^{+}20]. Input [AAM^{+}23, CYX^{+}19, DSPR18, EBPA22, FZZ^{+}22, GKO18, HJL^{+}22, SRK^{+}19, WS17, YZS^{+}20]. InSight [JNB^{+}20]. Insights [TK17, TAHH23]. Insomnia [HAY^{+}17]. Inspector [HAL^{+}20]. Instability [MSL^{+}19]. Installation [GAS^{+}18]. Instance [KSZ^{+}22]. Instant [KKL21]. Instantaneous [LDW^{+}18]. Instruction [QWK^{+}19]. Instructional [MMMS^{+}20]. Intake [HEUH18]. Integrated [Hal19]. Integrating [ILH^{+}19, SC22]. Integrity [ASM^{+}18, KBKH20, MMMS^{+}20]. Intelligent [CZT22, CND22, DPW^{+}21, GF17, LYC^{+}19, NZL^{+}20]. Intelligibility [JSC^{+}18]. Intelligible [CVL20]. Intensity [HXH^{+}18]. Intent [IYS22]. Interact [KML^{+}21]. Interacting [CLY^{+}23]. Interaction [AZEL17, CYX^{+}19, CLYZ22, CJRG21, DWL^{+}22, FMA^{+}19, GZC^{+}23, HYS^{+}23, IY20, JCR22, KJKL19, KLK^{+}20, LLL^{+}19, LYQ^{+}21, LMS^{+}23, MGC^{+}23a, MTF^{+}18, PRS^{+}23, QCP^{+}18, SvBV^{+}18, SvBH^{+}19, SS22, TLF22]. Interactions [AYT21, AWZK23, CKP^{+}20, CCY^{+}17, CND22, EUHB23, KCP^{+}18, LYC^{+}19]. Interactive [AIP^{+}18, ABL^{+}23, GMY^{+}17, JWS^{+}23, JHWF22, LRH^{+}21, LH20, NNMY23, RPF^{+}18, SFR^{+}20]. Interactivity [BGL18]. Interconnecting [LRH^{+}23]. Interest [CYK^{+}19, DYL^{+}19, LYC^{+}19, MOD21, TFL^{+}19, GLC^{+}21, WXY^{+}19]. Interface [DWHR17, FKI18, GJL^{+}20, KMH^{+}21, RGL^{+}18, SKS^{+}21, TK17, VLR20]. Interfaces [HSWB21, HYS^{+}23, IB23, KHW^{+}23, WXZ^{+}20, YLL^{+}20]. Interference [AZK^{+}23, BW20]. Interferometry [MFSK21]. Interior [SCZ^{+}20]. Intermittent [AY22, KGAH22, LILN19, VJH^{+}18]. Intermittently [BRYH21, IN20, LILN19]. Intermittently-Powered [IN20]. Internal [LCR^{+}18]. Internet [ABL^{+}23, ASM^{+}18, BMG18, FDL^{+}18, FMA^{+}19, GM19, KPK^{+}18, RPF^{+}18]. Internet-of-Things [KPK^{+}18]. Interpersonal [SGO^{+}23]. Interpretation [CDA23, LS20, PNL^{+}21]. Interpreter [JZG^{+}23]. Interpreting [LKL19, RLGN19]. Interruptibility [KPL20]. Interrupting [KCP^{+}18]. Interruption [GF17]. Intersections [CS19]. Intervention [AJA^{+}18, CPK^{+}19, LEP21, LJZ^{+}22, YPJdB19]. Interventions [KKV^{+}22, KML^{+}21, KMK^{+}19, MKK^{+}21, PZH^{+}18, PL23]. Interview [MGWM21]. Intimate [KFFC18]. Intra [MWJ^{+}21]. Intra-Body [MWJ^{+}21]. Intrusive [FXC^{+}21, GZZ^{+}17, LYWZ22]. Intuitive [SKS^{+}21]. Invasive [PCBK20, HDC18]. Inverse [MOD21]. Investigating [CQCH21, CSM^{+}20, CMF^{+}23, DLH^{+}18, DPB21, FMP^{+}23, HJL^{+}22, RLCI22, TKF^{+}19]. Invisible [GCGD17, LCL20]. Involvement [LKBK19]. IoT [KPK^{+}18, FZS^{+}23, HLL17, HAL^{+}20, JSC^{+}18, KSCS19, KPG^{+}20, LHK^{+}19, LYH^{+}20, WXL^{+}20]. IR [DJX^{+}23]. IR-UWB [DJX^{+}23]. Iris [HWJ^{+}23, SWLZ21]. IriTrack [SWLZ21]. irrelevant [DYCX21]. Irritable [CWS^{+}19]. ISACS [FRB^{+}21]. Isolated [MBZ^{+}18]. iSpray [CZT22]. Issue [DWCY23].

Jacdac [DMdH^{+}22]. Janus [ILM^{+}21]. Jaundice [MBP^{+}17]. Jawbone [CBT18]. JITAIs [KKV^{+}22]. Job [MMG^{+}19, SSR^{+}19]. Joint [CJF^{+}23, DZZ^{+}21, LSW^{+}19, LSY^{+}18]. Joint-Motion [CJF^{+}23]. Joints [ARSN22]. Jump [GWD^{+}17]. Jumping [GWD^{+}17]. Just [CPK^{+}19, GS20, KKV^{+}22, LEP21, LDS^{+}18, PZH^{+}18, PL23, PLX^{+}18].

Just-In-Time [CPK^{+}19, KKV^{+}22, LDS^{+}18, PL23, LEP21, PLX^{+}18].

Keep [LGY^{+}20]. Keeping [HKK^{+}23]. Keratoconus [GBS^{+}21]. Key [LKBK19]. Keyboard [GSS^{+}19, PRS^{+}23, LHY^{+}22]. Keyless [MZJ20]. Keylogging [GMMH21]. Keys [MGY^{+}23]. Keystroke [HKK^{+}23]. Keyword [AO19, WYL^{+}21]. Kinetic [AWZK23]. Kit [KMH^{+}21]. Kitchen [KOY^{+}19, KTEM21]. Knitted [MVS^{+}20]. Knocker [GCLL19]. Know [JJB22, LCG^{+}22]. Knowledge [DGH^{+}20, RH19].

Label [GKO^{+}23]. Label-Free [GKO^{+}23]. Labeled [HAL^{+}20]. Lacrosse [JWK^{+}21]. Landmarks [AvdBM^{+}22]. Language [CXR^{+}18, HHIF19, JGZ^{+}21, MZW^{+}18, PLK21, RR22, LHS^{+}23]. Laptops [DYCX21]. Large [BDR17, CLZ^{+}21, EUHB23, FFQ^{+}20, HHFM17, JWCC^{+}20, LT19, LKL19, SJL^{+}21, WLZ^{+}19, WZF^{+}20, ZPPC^{+}20]. Large-Scale [CLZ^{+}21, FFQ^{+}20, HHFM17, LT19, LKL19, WZF^{+}20, JWCC^{+}20, SJL^{+}21]. LASense [LLLX22]. Laser [MFSK21, SCZ^{+}20]. Last [PLC^{+}21]. Last-Mile [PLC^{+}21]. Latency [WRLK22]. Latent [FZS^{+}23]. LAUREATE [LGL23]. LAX [JWK^{+}21]. LAX-Score [JWK^{+}21]. Layer [GH22, HQY^{+}18, DSSD23]. LBSNs [KCuH^{+}19]. Lead [GWJ^{+}21]. Leading [CCW^{+}20]. Leakage [BS21, BS22, DYCX21, JGSC19, MJ19, NTKG17]. LeakDoctor [WCY^{+}19]. Leaks [GMMH21, WCY^{+}19]. Learn [BGR^{+}20, NTT^{+}21]. Learned [DDP^{+}17, LCR^{+}21]. Learner [FWJ^{+}22]. Learners [GP17]. Learning [AT19, BBBN22, CS22a, CQCH21, CRS^{+}18, CKHM20, CDL^{+}19, CS19, CZP20, CMK22, DKZ^{+}22, DCMM20, DXS^{+}22, DYL^{+}19, DGZ^{+}21, DZZ^{+}21, DJX^{+}23, DPB21, EY20, EBTN22, FSX^{+}18, FMP^{+}23, FRS^{+}20, FWJ^{+}22, FML^{+}22, HHIF19, HAR18, HHX^{+}20, ISY^{+}21, JTM^{+}22, JZWL19, JHH21, JSNL23, KHZ22, KBBN21, LILN19, LWA23, LGKM20, LDC18, LNT^{+}19, LDSZ19, LYL^{+}20, LLL^{+}20, LCZ^{+}21, LCH^{+}22, LSY^{+}18, MSWQ20, OWH^{+}20, PCYZ18, QCWY19, RZ21, SOL19, SEG^{+}18, SH20, SSB^{+}21, SWLG22, SCZ^{+}20, TFL^{+}19, WZG^{+}19, WMS^{+}21, XL19, YZS^{+}20, GLC^{+}21]. Learning-based [EY20]. Learning-Enabled [JZWL19]. Lectures [DGS18]. LemurDx [AAM^{+}23]. Length [GWD^{+}17]. Less [NCM^{+}17, LZZ^{+}21, LZZ^{+}23, SBP^{+}22]. Lessons [DDP^{+}17, LCR^{+}21]. Letters [TAHH23]. Level [MRCdCRV^{+}21, RTZ^{+}20, BXY^{+}21, LXW^{+}21, LH20]. Leveraging [AT19, AZS^{+}18, BAT22, CLY^{+}21, CYK^{+}23a, DYCX21, DBY^{+}17, GCW^{+}18, HYY^{+}23, HW21, KKV^{+}22, KBXP23, LLW^{+}22, LYY^{+}17, LYCW20, SRR^{+}20, VSY^{+}20]. LiDAR [ASG^{+}23, CRMG22]. Life [KFFC18]. Lifeloggers [PSC^{+}17]. Lifetime [LHW^{+}19]. Light [CLH18, GMY^{+}17, HXH^{+}18, HWJ^{+}23, JNB^{+}20, KASC21, LH20, MTF^{+}18, TXWZ22, ZPZ^{+}20, LLZ^{+}22, MJ19]. Light-weight [CLH18]. Lighting [SSB^{+}21]. Lightitude [HXH^{+}18]. Lights [MJ19]. Lightweight [CWS^{+}19, KPK^{+}18, LEP^{+}19]. Like [CGJ^{+}22, FALW18, GS20, ALC^{+}20]. Limit [AKPS18, GL18]. Limited [NCM^{+}17]. Limits [LLLX22, LZZ^{+}21, XYX21]. Line [LKBK19]. Lines [GPW^{+}18]. Linguistics [DZZ^{+}21]. Liquid [CRMG22, LZZ^{+}21, RTZ^{+}20, SYYL22, LZZ^{+}21]. LiSee [CHC^{+}22]. Listen [CLD^{+}20, LLWX21]. Listening [DBBH21, TKF^{+}19]. Literature [ECF^{+}20, JCR22]. Live [CYK^{+}23a, FSW19, KLK^{+}20, SGO^{+}23]. Live-view [KLK^{+}20]. Liveness [HCYZ18, LZZ^{+}20]. Liver [MBP^{+}17].

Living [BAT22, LT19, ZZN^{+}20]. Load [AVD^{+}17, CSM^{+}20, CRGR22, ELD19]. Loc [DJY^{+}22, HNL22]. Local [AvdBM^{+}22]. Localization [CPW^{+}21, CLH18, DJL^{+}21, DJY^{+}22, EEY18, GAS^{+}18, GL18, GGJ^{+}22, HNL22, JSNL23, KPK^{+}18, LCW^{+}21, NHL^{+}20, TZW^{+}19, WXY^{+}19, WXY^{+}17, ZCA^{+}21]. Location [CFLK18, CGH^{+}17, GHY^{+}19, HHSM18, JDZ^{+}21, KBBN21, RKK18, SZJ^{+}17, TFL^{+}19, WRLK22, ZCL21]. Location-Aware [JDZ^{+}21]. Location-independent [SZJ^{+}17]. Locations [LLCY21]. Lockout [KJKL19]. Logger [JWCC^{+}20]. Logging [PSC^{+}17]. Long [BLS^{+}18, GS20, HHFM17, HKK^{+}23, SBP^{+}22, TK17, XYX21, ZCN^{+}20]. Long-Range [SBP^{+}22, ZCN^{+}20]. Long-Term [BLS^{+}18, GS20, HHFM17, HKK^{+}23, TK17]. Longitudinal [JSC^{+}18, JKK^{+}17, JHK^{+}21, KML^{+}21]. Look [PTK18]. LoRa [CZX^{+}22, XYX21, ZCN^{+}20]. LoRaWAN [GGJ^{+}22]. Loss [MNV^{+}19]. Lost [RRS^{+}23]. Love [TAHH23]. Love/ [TAHH23]. Low [AO19, CLL^{+}22, CHC^{+}22, GBLM17, HKK^{+}23, IB23, JNB^{+}20, JWS^{+}23, LCS^{+}21, LZL^{+}22, LDC18, LLZ^{+}22, PCBK20, RDOA22, RHMZ^{+}21, RMG^{+}17, SAD^{+}17, WRLK22, ZCA^{+}21]. Low-Cost [JWS^{+}23, LDC18, PCBK20, CLL^{+}22, LCS^{+}21]. Low-Latency [WRLK22]. Low-Light [JNB^{+}20, LLZ^{+}22]. Low-Power [ZCA^{+}21, LZL^{+}22, RDOA22, RMG^{+}17, SAD^{+}17]. Low-Resolution [RHMZ^{+}21]. Low-resource [GBLM17]. Low-Vision [IB23, CHC^{+}22]. Lower [MMG^{+}19]. LSTM [GP17]. LSVP [RZ21]. Luminance [SSB^{+}21]. LumNet [SSB^{+}21]. LungTrack [CXC^{+}19].

M [NZL^{+}20]. MAAT [LHK^{+}19]. MAC [FYW^{+}19]. Machine [LILN19, LNT^{+}19, OWH^{+}20, RPF^{+}18, WZG^{+}19]. Macronutrients [HHZ23]. Mag [IB23]. Magnetic [AZK^{+}23, CSHN17, KNL^{+}23]. Mago [CSHN17]. MAIL [NHL^{+}20]. Maintenance [GAS^{+}18]. MakeCode [KGAH22]. Makes [LYKH18]. Making [NNMY23, PB20, ZPPC^{+}20, AMG^{+}19]. Malice [DWCY23]. Manage [WBG^{+}19]. Management [AHS^{+}18, BSP21, CGH^{+}17, FCFW^{+}17, GF17, IMO^{+}18, KGN^{+}20, KGL^{+}18, SKM^{+}19, WSW^{+}22]. Manager [HYS^{+}23]. Maneuver [CRGR22]. Manipulation [MQZH23]. Manual [ADY^{+}20]. Map [JSNL23]. Mapping [ARD^{+}18, ECF^{+}20, PBM^{+}18]. Markets [WCW21]. Mask [ATP^{+}19]. Masks [CRB^{+}21, LCR^{+}21]. Mat [LLPH22]. Matching [EBPA22, KSC^{+}17, RDOA22]. Materials [CPC^{+}18, JWS^{+}23]. Math [EUHB23]. Math-Based [EUHB23]. Matter [CSQ^{+}20]. Matters [BDPH21]. Me [AZEL17, CMF^{+}23, FALW18, GS20, NZL^{+}20, PSC^{+}17]. Mealtime [CLRH19]. Meaningful [LYKH18]. Meaningless [LYKH18]. Means [CYK^{+}19]. Measure [GDS19, KKZ^{+}18, LP20, SGO^{+}23, Tor23]. Measurement [AAM^{+}23, CRP21, FGS^{+}23, LDW^{+}18, LDC18, Tor23]. Measurements [CKKM20]. Measures [RMG^{+}17, SAD^{+}17]. Measuring [FYW^{+}19, MTF^{+}18, SvBH^{+}19]. Mechanism [HYH21]. Mechanisms [AWZK23, KJKL19]. Media [AHB17, GWJ^{+}21, PMGP19]. Mediated [ZAP17]. Medication [KGL^{+}18]. Meditation [HBX^{+}17]. Meets [RPF^{+}18, WXL^{+}20]. Members [GDS19, PB20, SGO^{+}23]. Memory [BRS21, CBZN19, CSM^{+}20, GCP^{+}22, LGL23, LCFT19, LYY^{+}17, NCM^{+}17]. MemX [CZD^{+}21]. Men [WMW^{+}20]. Menstrual [MNV^{+}19]. MenstruLoss [MNV^{+}19]. Mental [AM19, CND17, FDH21, LCL20]. Mesh [DCZ^{+}23, DCZ^{+}23]. Message [KKL21].

Messaging [JGZ^{+}23]. Meta [KW23].

Meta-Synthesis [KW23]. Metal [AZK^{+}23]. MetaTP [ZSG^{+}21]. Method [AWSS22, HLL17, HHIF19, KOY^{+}19, MTH19, MGWM21, PCYZ18, SKS^{+}21]. Methods [SRK^{+}19]. METIER [CZP20]. Metrics [BNW22]. mHealth [ASS^{+}19, AJA^{+}18, KJP^{+}19, KMK^{+}19, MKK^{+}21]. MI [AZK^{+}23, DCZ^{+}23]. MI-Mesh [DCZ^{+}23]. MI-Poser [AZK^{+}23]. Mic [CYK^{+}23a]. micro [KME^{+}19]. micro-Stress [KME^{+}19]. MicroCam [HYY^{+}23]. Microcontrollers [LL22]. Microgesture [CLYZ22]. Microinteraction [PLW^{+}22]. Microphone [AZS^{+}18, BMR^{+}20, GJC^{+}21b]. Microphone-equipped [BMR^{+}20]. Microscope [HYY^{+}23]. Microtask [GCGD17]. Microwave [BS21, BS22]. Mid [EUHB23, KTM^{+}18, PSLS21, SKS^{+}21]. Mid-Air [EUHB23, KTM^{+}18, PSLS21, SKS^{+}21]. Midas [DZH^{+}23]. Mile [PLC^{+}21]. Millimeter [ARSN22, BXY^{+}21, CS22b, DCZ^{+}23, LZZ^{+}21, LZZ^{+}23, LXW^{+}21, LWZ^{+}20, PSLS21, RSSN21, SHZ^{+}20]. Millimeter-level [BXY^{+}21, LXW^{+}21]. Millimeter-Wave [ARSN22, CS22b, PSLS21, RSSN21, LZZ^{+}21, LZZ^{+}23, SHZ^{+}20]. MilliPCD [CS22b]. MindfulWatch [HBX^{+}17]. Miniature [HHX^{+}20]. Miniaturised [KGL^{+}18]. Miniaturized [JMvB^{+}19]. Minimize [AT19]. Minority [WMW^{+}20]. Mirror [KLK^{+}20]. MiShape [ARSN22]. Missing [LYH^{+}20]. Mites [BCM^{+}23]. Mitigating [BSH^{+}17, HWJ^{+}22]. Mitigation [AZK^{+}23, EBTN22]. Mixed [CKL23, CNS21, GSS^{+}19, dGST21]. Mixup [LWC^{+}22]. ML [IYS22]. ML-based [IYS22]. mmASL [SHZ^{+}20]. mmStress [LZZ^{+}23]. MMTSA [GwC^{+}23]. MmWave [CYB^{+}23, CLL^{+}22, DZH^{+}22, DZH^{+}23, LCH^{+}22, XLL^{+}21]. mmWave-based [XLL^{+}21]. Mobile [AvdBF^{+}19, ATQ^{+}21, ASTW19, ATC17, BAR^{+}18, BS17, BRS21, BX19, CTTMT20, CMV^{+}18, CML^{+}20, CS19, CML22, CPK^{+}19, CNS21, DSM17, DGLQ19, EBPA22, FSW19, FCQR20, FS18, GBLM17, GSS^{+}19, GKO^{+}23, GMC18, HLS20, HHFM17, HFS^{+}19, HYS^{+}23, HKK^{+}18, HBK21, JZWL19, JLD^{+}18, JJB22, JWCC^{+}20, KMC^{+}18, KVZ^{+}19, KNL^{+}23, KBKH20, LSSH23, LLL^{+}19, LKL22, LDS^{+}18, LHK^{+}19, LDSZ19, LGM^{+}21, LCZ^{+}21, LLF^{+}21, LLZ^{+}22, LWX^{+}22, LAX^{+}23, MLX^{+}20, MZST17, MDK^{+}22, MT19, MSTT22, MMG^{+}19, MGC^{+}23b, OHW^{+}20, OGG^{+}18, PNL^{+}21, PLK21, PL23, PLX^{+}18, QCP^{+}18, RKK18, RSSN21, SvBV^{+}18, SvBH^{+}19, WBG^{+}19, WCY^{+}19, WXL^{+}20, YFX^{+}20, YLMR20, ZWZ^{+}21b, dGST21]. Mobile-First [FCQR20]. Mobilephone [LLZ^{+}18]. Mobilephone-Enabled [LLZ^{+}18]. MobilePhys [LWX^{+}22]. Mobileportation [YLMR20]. Mobility [CMF^{+}23, DJN^{+}18, FSX^{+}18, FSJ^{+}19, FYY^{+}21, FRS^{+}20, JM18, JSF^{+}18, LLC^{+}20, LWYZ21, RBL^{+}20, SJL^{+}21, XLX^{+}20, YFX^{+}20]. MobiVQA [CKLB22]. Modal [BGK21, CHWZ23, GwC^{+}23, GOZ^{+}17, HJL^{+}22, HFS^{+}19, HKK^{+}18, VLR20]. Modalities [RH19]. Modality [DXS^{+}22]. Mode [CSHN17]. Model [ASS^{+}19, CZP20, CHWZ23, CPK^{+}19, HYS^{+}23, JSGS23, LLL^{+}20, LGM^{+}21, LLG^{+}23, MLX^{+}20, WGL^{+}21, YZ21]. Model-based [YZ21]. ModElec [HWJ^{+}21]. Modeling [AKB23, DDKL19, FZYZ18, JSF^{+}18, KVZ^{+}19, LJZ^{+}22, LSY^{+}18, OGG^{+}18, WLZ^{+}19, YFX^{+}20]. Modelling [BRS21, CDL^{+}19]. Models [AM19, CRS^{+}18, CND17, FDH21, KVZ^{+}19, LCL20, LWYZ21, MDK^{+}22]. Moderation [AJV20]. Modification [AP18]. modulated [LLW^{+}22]. Module [ZLB^{+}17]. Modus [GCGD17]. Moisture [CZX^{+}22, DJX^{+}23]. Mom [MT19]. Moment [BAR^{+}18, CZD^{+}21, CGJC19, MSTT22]. Momentary [CSK^{+}18, DBW^{+}22, PLW^{+}22]. Moments [CKP^{+}20, KCP^{+}18]. Money [CMV^{+}18]. Monitor [GZNL22]. Monitoring [BS21, BS22, BLW^{+}23, BSMP23, CLC^{+}22, CCJ^{+}17, CH18, CXC^{+}19, CZL^{+}22, GLW^{+}18, GZZ^{+}17, HMA21, HBX^{+}17, HDC18, HAY^{+}17, ISSK20, JNB^{+}20, KOY^{+}19, KRS22, KNL^{+}23, LAAE^{+}19, LXW^{+}21, LLZ^{+}18, LZG^{+}21, LYD^{+}23, MZT18, MBP^{+}17, MNV^{+}19, PCBK20, SBP^{+}22, WXL^{+}20, YYW^{+}20, ZCC^{+}20]. Mood [LS20, MDK^{+}22, MSL^{+}19, Tor23]. mORAL [ASS^{+}19]. Morphology [CFLK18]. Mothers [KME^{+}19]. Motifs [XKNM20]. Motion [AOGF^{+}21, BAT22, BX19, CLC^{+}22, CXZ^{+}21, CJF^{+}23, CJRG21, EBPA22, FSW19, GZZH20, GZH^{+}21, IAE^{+}19, LFH18, LSW^{+}19, MH23, QCP^{+}18, RPB^{+}18, SXF^{+}22, WYL^{+}21]. Motion-Based [QCP^{+}18, MH23]. Motion-resilient [SXF^{+}22]. Motions [HWZZ19]. Motivations [MT19]. Motivators [GM19]. Motor [KJP^{+}19, MW22, NTT^{+}21]. Motorized [CJRG21]. Mounted [ASTW19, DJX^{+}23, GO22, PCBK20, RZ21, SZZ^{+}20, BWT^{+}18, CLT^{+}21]. Movement [CBEG17, GLG^{+}19, GZL^{+}21, LSSH23, LJZ^{+}22, MFSK21, ZPPC^{+}20]. Movements [CBT18, LZL^{+}22]. Mover [LWYZ21]. Moving [BX19, LGM^{+}22, WRLK22]. MSLife [GZY^{+}21]. mTeeth [ASS^{+}21]. mTransSee [LCH^{+}22]. Much [KL19, LDS^{+}18]. Multi [BYW^{+}20, BCK^{+}22, CTTMT20, CDL^{+}19, CZP20, CYB^{+}23, CHWZ23, CMK22, CPK^{+}19, CCK^{+}18, DKZ^{+}22, DSSD23, DGH^{+}20, EEY18, GwC^{+}23, GBLM17, GOZ^{+}17, GCW^{+}18, HY18, HY22, HFS^{+}19, HYH21, HKK^{+}18, KSZ^{+}22, LEP21, LSY^{+}18, MZST17, NHL^{+}20, PCYZ18, SOL19, SH20, VLR20, YHW^{+}18, ZWX^{+}20, ZZN^{+}20].

Multi-agent [DSSD23]. Multi-device [CMK22]. Multi-Domain [BCK^{+}22]. Multi-Head [HYH21]. Multi-Hop [MZST17]. Multi-Modal [CHWZ23, GwC^{+}23, GOZ^{+}17, HFS^{+}19, HKK^{+}18, VLR20]. Multi-Person [YHW^{+}18, CYB^{+}23, ZWX^{+}20]. Multi-perspective [CTTMT20]. Multi-Scale [NHL^{+}20, KSZ^{+}22]. Multi-Sensor [DGH^{+}20, EEY18, ZZN^{+}20]. Multi-sensory [LEP21]. Multi-source [GCW^{+}18]. Multi-Stage [CPK^{+}19]. Multi-Task [CDL^{+}19, CZP20, DKZ^{+}22, PCYZ18, SH20, GBLM17, LSY^{+}18, SOL19]. Multi-Touch [CCK^{+}18]. Multi-user [HY22]. Multi-view [BYW^{+}20, HY18]. MultiCell [FZYZ18]. Multicore [AY22]. Multimodal [AYM^{+}22, CDG^{+}17, GMS^{+}23, LMS^{+}23, ZPPC^{+}20, ZAP17]. Multipath [WXY^{+}19]. Multipath-prevalent [WXY^{+}19]. Multiple [AKB23, BOH^{+}21, CPW^{+}21, FMZN22, FZYZ18, FYW^{+}19, GZY^{+}21, LRH^{+}23, VSY^{+}20, XKNM20]. MultiSense [ZWX^{+}20]. Multisensor [LYL^{+}20, SSR^{+}19, WMS^{+}21]. Multitasking [KPL20]. Multivariate [MCCH21, OGG^{+}18]. Muscle [DSPR18, KAH18, LYY^{+}17, NTT^{+}21]. Muscle-Based [DSPR18]. MuscleIO [DSPR18]. Museum [CCK^{+}18]. Music [AP18, HCH^{+}22, LHC^{+}22, NTKG17]. Mutual [AT19]. My [BA20, CGH^{+}17, CND17, GMMH21, JLD^{+}18, MT19, SLX^{+}18]. Myocardial [RDOA22]. Myself [ASV^{+}18].

N95 [LCR^{+}21]. Nanogenerator [AZS^{+}18]. Naptics [CH18]. Narrative [MGC^{+}23b]. Narrative-Based [MGC^{+}23b]. Nation [LKL19]. Natural [CXR^{+}18, CCY^{+}17, MKK^{+}21, RR22, WTA^{+}20]. Navigating [EZC23]. Navigation [CLZL18, GAS^{+}18]. NCE [RLM^{+}19]. Near [AZEL17, HQC^{+}23, JMvB^{+}19, KGL^{+}18].

Near-field [HQC^{+}23]. Near-Infrared [JMvB^{+}19, KGL^{+}18]. Nearby [KIT^{+}20]. Neck [CLT^{+}21]. Neck-mounted [CLT^{+}21]. NeckFace [CLT^{+}21]. Necklace [ZZN^{+}20]. NeckSense [ZZN^{+}20]. Need [CLTS21, CGH^{+}17, LNT^{+}19]. Needs [JSC^{+}18, KL19, OSG^{+}22]. Neglect [KKL21]. Network [CHWZ23, FYY^{+}21, GwC^{+}23, GMS^{+}23, GBLM17, HY18, HAL^{+}20, JLD^{+}18, LCG^{+}22, LL22, LDSZ19, LYH^{+}20, LMS^{+}23, XLX^{+}20, YXW^{+}20]. Networks [BYW^{+}20, FZYZ18, GSS^{+}19, GCC^{+}21, LBC^{+}22]. Neural [GBLM17, GSS^{+}19, HY18, JHH21, KSZ^{+}22, LL22, LYH^{+}20, MCCH21, WYL^{+}21]. NeuralGait [LCX^{+}22]. Neurocognitive [LSSH23]. Neurocognitive-State-Dependent [LSSH23]. Neurofeedback [LAX^{+}23]. News [GWJ^{+}21]. NF [HQC^{+}23]. NF-Heart [HQC^{+}23]. no [AAM^{+}23]. Noise [CCY^{+}23, CDG^{+}17, KBXP23, LKBK19, MLMH^{+}18, SvBV^{+}18]. Noise-Robust [MLMH^{+}18]. Noisy [DSR^{+}22]. Nomadic [YLMR20]. Non [CBT18, FXC^{+}21, GZZ^{+}17, HDC18, HQC^{+}23, JM18, LHS^{+}21, LYWZ22, PCBK20, PLW^{+}22, XLL^{+}21]. Non-Conformant [JM18]. Non-Contact [CBT18, XLL^{+}21, HQC^{+}23]. Non-Intrusive [FXC^{+}21, GZZ^{+}17, LYWZ22]. Non-Invasive [PCBK20, HDC18]. Non-registered [LHS^{+}21]. Non-response [PLW^{+}22]. Nonlinear [LYWZ22]. Nonsense [BSH^{+}17]. Nonvisual [GFK^{+}18]. Normalization [MBB20]. Norms [ASM^{+}18]. NOSE [KOY^{+}19]. Noticeability [LJZ^{+}22]. Notices [LNAH21]. Notification [CTTMT20, KKL21]. Notifications [DSPR18, JW18]. Novel [KOY^{+}19, MVS^{+}20, PTMM^{+}22]. Nudge [NNMY23]. Nursing [ILH^{+}19]. Nutrients [BS22]. Nutrition [JWCC^{+}20].

o [GMMH21]. Obfuscation [KMC^{+}18]. Object [DZH^{+}22, DZZ^{+}21, GCLL19, LCFT19, SKS^{+}21]. Objective [AAM^{+}23, HP20]. Objects [CHC^{+}22, DCMM20, FFT^{+}18, FKI18, LGM^{+}22, MQZH23, MWJ^{+}21, TLF22]. Observation [PTK18]. Observations [DGLQ19]. Obstructive [BLW^{+}23]. Occupancy [MZJ20, RHMZ^{+}21]. Occupants [FMZN22]. Occupied [KSCS19]. Odometry [VS19]. Odor [KOY^{+}19]. ODSearch [RR22]. OFDMA [TZW^{+}19]. Off [AIP^{+}18, PLX^{+}18]. Off-Screen-based [PLX^{+}18]. Off-the-Shelf [AIP^{+}18]. Office [CTTMT20]. Offline [BWUW21, MH23, XL19]. offs [BDR17, RHMZ^{+}21]. Oh [CND17]. OHMDs [JGZ^{+}23]. Ok [AYT21]. Older [CBZN19, JHWF22, KTEM21, LCFT19, MT19, MSTT22, WBG^{+}19]. On-board [PLK21]. On-Body [CDA23, CJF^{+}23, CLTS21, GNA^{+}19, KHW^{+}23, KTH^{+}20]. On-Demand [DJL^{+}21, DJY^{+}22, GZW^{+}22]. On-Device [CKLB22, IN20, LILN19, WYL^{+}21, RR22]. On-road [KYJ^{+}23]. On-Skin [KBL18, KMH^{+}21]. On-the-go [RZ21]. One [BRA^{+}22, CND22, EBPA22, FWJ^{+}22, LLL^{+}19, MRCdCRV^{+}21, YLL^{+}20]. One-Handed [LLL^{+}19, YLL^{+}20, EBPA22]. One-Shot [CND22, FWJ^{+}22]. Ones [FML^{+}22]. OnHW [OWH^{+}20]. Online [BWUW21, CCX^{+}18, FSX^{+}18, FTL^{+}19, GLW^{+}18, JDZ^{+}21, KDL^{+}21, LT19, OWH^{+}20, WBG^{+}19, WMW^{+}20, XL19, ZLB^{+}17]. Only [IYS22]. Ontology [BPA23]. Open [CLY^{+}21, KDL^{+}21, LKL22, NJWFM19]. Open-set [KDL^{+}21]. Operandi [GCGD17]. Operation [LGY^{+}20]. Operations [LYWZ22]. Opioid [GCC^{+}21]. OpiTrack [GCC^{+}21]. Opportune [CKP^{+}20, KCP^{+}18]. Opportunistic [CLHW18]. Opportunities [LFL^{+}23, SKM^{+}19]. Opportunity [CML^{+}20]. Optic [LYD^{+}23]. Optical [RZ21, SFR^{+}20]. Optimal [KJP^{+}19]. Optimized [AO19, LLL^{+}19]. Optimizes [CFF^{+}22]. Optimizing [IB23, LGKM20]. Optimum [LHS^{+}21]. OptiStructures [SFR^{+}20]. OptoSense [ZPZ^{+}20]. Oral [ASS^{+}19, PB20, TAEB17]. Orbiting [CBEG17]. Orbits [EBPA22]. Orchestration [LMT19]. Organizational [SSR^{+}19]. Orientation [FCQR20, KLM^{+}18, YZ21]. Oriented [ATP^{+}19]. Outdoor [JDZ^{+}21]. Outliers [BKM21]. Output [DSPR18, SRK^{+}19]. Oven [BS21, BS22]. Overlaying [KSCS19]. Overthere [SKS^{+}21]. Ownership [LJZ^{+}22].

P2 [DJY^{+}22]. P2-Loc [DJY^{+}22]. PACE [CPW^{+}21]. packet [SYYL22]. PackquID [SYYL22]. PACT [SBP^{+}22]. Pain [JCF^{+}21, WXY^{+}17]. Painting [FKI18, NNMY23]. Pairing [LKBK19]. Pancreatic [DDKL19, MBP^{+}17]. Pandemic [LCR^{+}21, TAHH23]. Panel [CYX^{+}19]. Pantomime [PSLS21]. Paper [DJL^{+}22]. Paper-Based [DJL^{+}22]. Papers [DJL^{+}22]. Paradox [JW18]. Parcel [PLC^{+}21]. Parent [AAM^{+}23, CLRH19, FML^{+}22, KLK^{+}20]. Parent-child [KLK^{+}20]. Parents [GS20, KL19]. PARK [LAS^{+}19]. Parking [CLH18]. Parkinson [LAS^{+}19]. ParkLoc [CLH18]. PARROT [ABL^{+}23]. Partial [DGLQ19, KHZ22]. Participatory [BSH^{+}17]. Partition [WGL^{+}21]. Passenger [LLPH22]. Passengers [BW20]. Passerby [IMO^{+}18]. Passive [AAM^{+}23, BKL^{+}22, CPW^{+}21, DPB21, DDKL19, FMP^{+}23, HQY^{+}18, HWJ^{+}23, KME^{+}19, MSL^{+}19, RGL^{+}18]. Passphrase [GJC^{+}21a, LYCW20].

Passphrase-Independent [GJC^{+}21a, LYCW20]. Passwords [GMMH21, LCL17]. Path [CMF^{+}23, XYG^{+}20]. Patient [CWS^{+}19]. Patient-Provider [CWS^{+}19]. Patients [AZK21, BOH^{+}21, HMM^{+}20]. Patterned [GH22]. Patterns [BPA23, LKL19, SWLG22]. Pavement [SS22]. Paving [XYG^{+}20]. PCB [KBL18]. PCIAS [LDW^{+}18]. Pedestrian [HW21, ZM22]. Pedestrians [KIT^{+}20]. Peeled [CND^{+}20]. Peeled-and-Transferred [CND^{+}20]. Peer [DBY^{+}17]. Peer-to-Peer [DBY^{+}17]. Pen [DJL^{+}22]. Pen- [DJL^{+}22]. Pencil [FYX^{+}21]. Pens [OWH^{+}20]. People [AKB23, BDM^{+}18, FALW18, GLJ^{+}21, JXZ^{+}20, KHSW17, KLT19, LGY^{+}20, MW22, SZJ^{+}17, VSY^{+}20, XYG^{+}20]. PEPPER [BBBN22]. Perceive [BBH17]. Perceived [BOH^{+}21, LKL22]. Perception [HPWW18]. Perceptions [DBBH21, RKK18]. Performance [CRS^{+}18, CGJC19, DSM17, GKO^{+}23, JWK^{+}21, KKL21, RRS^{+}23, RHMZ^{+}21, SNGK20, SGO^{+}23, SSR^{+}19]. Performances [HCH^{+}22]. Performers [MMG^{+}19]. Perils [FMA^{+}19]. Perinatal [GLG^{+}19]. Persistent [GCM^{+}19]. Person [CDG^{+}17, DJY^{+}22, FXC^{+}21, SSW^{+}17, SSR^{+}19, YHW^{+}18, CYB^{+}23, KLK^{+}20, ZWX^{+}20]. Person- [DJY^{+}22]. Person-Centered [SSR^{+}19]. Person-to-person [SSW^{+}17]. Personal [AvdBM^{+}22, AP18, CTTMT20, DDP^{+}17, ECF^{+}20, FALW18, KW23, MZT18, MGWM21, MW22, PB20, RLGN19, RLCI22]. Personalidad [KVZ^{+}19]. Personalised [HHSM18]. Personality [KVZ^{+}19]. Personalizable [CEL^{+}22]. Personalization [GKO^{+}23, GLJ^{+}21, MBB20, MDK^{+}22]. Personalized [CZD^{+}21, FSJ^{+}19, FTL^{+}19, FGS^{+}23, FZZ^{+}22, GLS^{+}18, JW18, LS20, LHC^{+}22, LWA23, LGKM20, LWX^{+}22]. Personas [SSR^{+}19]. Persons [CPW^{+}21]. perspective [CTTMT20]. Perspectives [CLRH19]. Pervasive [CME^{+}19, DZH^{+}23, MOD21]. Phase [LDC18]. Phases [IRA^{+}21, STL^{+}19]. Phenotyping [GZY^{+}21]. Phone [CYK^{+}23b, GSS^{+}19, KGN^{+}20, LNT^{+}19, LGY^{+}20]. Phone-Based [LNT^{+}19]. Phones [DGLQ19, JJB22]. Photo [AHB17, AM19]. Photo-Sharing [AHB17]. Photoplethysmography [DSR^{+}22]. Photos [LFL^{+}23, ZCL21]. Phyjama [KHG^{+}19]. Physical [ADY^{+}20, AJV20, AKB23, DMdH^{+}22, FZS^{+}23, FKI18, GLW^{+}18, GYL^{+}20, HQY^{+}18, HWJ^{+}21, KCK^{+}19, KXAH18, LCL20, LGKM20, MGC^{+}23b, PRC^{+}18, SZZ^{+}20, TK17]. Physical-Layer [HQY^{+}18]. Physiological [BRS^{+}20, CRP21, GDS19, GF17, KHG^{+}19, LWX^{+}22, SNGK20]. Piano [DPB21, FMP^{+}23]. Pickup [GHvB^{+}17]. Picture [NCM^{+}17]. Pictures [AM19]. Pipeline [PRC^{+}21]. Place [KTEM21]. Placement [NZL^{+}20]. Places [HSWB21]. Plan [CZG^{+}21]. Planning [ASG^{+}23, CWS^{+}19, CRGR22, NTO^{+}18]. Plans [PRC^{+}18]. Platform [BRYH21, CCY^{+}23, CRP21, CRB^{+}21, ALC^{+}20, MGC^{+}23b, QCP^{+}18, WPJ^{+}19]. Platforms [JDZ^{+}21, PLK21]. Play [WTA^{+}20, DMdH^{+}22]. Playing [NTT^{+}21]. Plotter [CLZ^{+}21]. Plug [DMdH^{+}22]. Plug-and-play [DMdH^{+}22]. Plugin [LAH18]. PMC [LLL^{+}20]. PMF [FRS^{+}20]. Pneumatic [CEL^{+}22, SRK^{+}19]. Pneumatic-haptic [CEL^{+}22]. Pocket [LGY^{+}20]. PocketCare [DGLQ19]. PocketThumb [DWHR17]. POI [DYL^{+}19, HFX^{+}21]. Point [ARD^{+}18, CS22a, CS22b, CYK^{+}19, GL18, GLC^{+}21, JZG^{+}23, PSLS21, SCZ^{+}20]. Point-of-Care [JZG^{+}23]. Point-of-Interest [CYK^{+}19, GLC^{+}21].

Pointing [SKS^{+}21]. POIs [CCX^{+}18]. Polarization [VSY^{+}20]. Policing [ELD19]. Pollution [CZT22, MLX^{+}20, MZT18]. Popularity [OGG^{+}18]. Population [CLY^{+}21, FZYZ18, LNT^{+}19, WLZ^{+}19, ZPPC^{+}20]. Popup [MT19]. Portable [KKZ^{+}18, JSC17]. Pose [AZK^{+}23, HHX^{+}20, MLH^{+}23, RWW^{+}22]. Poser [AZK^{+}23]. PoseSonic [MLH^{+}23]. Position [GZZ^{+}21, LHS^{+}21, VJH^{+}18]. Position-Independent [GZZ^{+}21]. Positioning [DLD^{+}21, GZW^{+}22, HXH^{+}18, HWJ^{+}23, MMMS^{+}20]. Positive [CFZ^{+}19]. Possibilities [BML17]. Postal [AvdBF^{+}19]. Posture [FZZ^{+}22, YYW^{+}20]. Posture-adaptive [FZZ^{+}22]. Potential [SRR^{+}20]. Potentials [MGY^{+}23]. Power [AO19, FDL^{+}18, GPW^{+}18, LKBK19, MWJ^{+}21, ZCA^{+}21, LZL^{+}22, RDOA22, RMG^{+}17, SAD^{+}17]. Powered [BRYH21, ISSK20, IN20, LILN19, MGC^{+}23a, ZPZ^{+}20, AZS^{+}18]. Powerline [MGY^{+}23]. PPG [KRS22]. PPGface [CGJ^{+}22]. Practical [AKX^{+}19, SLX^{+}18]. Practicalities [CNS21]. Practically [HWJ^{+}22]. Practice [ZXY21]. Practices [ADY^{+}20, GM19]. Pragmatists [DBBH21]. Pre [LYD^{+}23]. Pre-Existing [LYD^{+}23]. Precise [LDW^{+}18, LLZ^{+}20, PWP19]. Prediabetes [WPH^{+}23]. Predict [DDKL19, PM19]. Predicting [CMV^{+}18, DKZ^{+}22, FSX^{+}18, GSRS20, JM18, KSZ^{+}22, KCP^{+}18, LCG^{+}22, MSTT22, OHW^{+}20]. Prediction [CWH^{+}19, CRGR22, EY20, FSJ^{+}19, FTL^{+}19, FSW19, FRS^{+}20, GCW^{+}18, HY18, HYH21, IYS22, JSF^{+}18, KASC21, KCuH^{+}19, LLPH22, LDK^{+}20, MSL^{+}19, RBL^{+}20, ZSG^{+}21]. Predictions [CVL20]. Predictive [HEP21]. PREFER [GLC^{+}21]. Preferences [RKK18, TKF^{+}19]. Prefetching [BS17]. Pregnant [KME^{+}19]. Preschool [KSZ^{+}22]. Presence [MTH19]. Present [JYY21]. Presentation [SWLZ21]. Presentations [PB20, TAEB17].

Presenters [DWCY23, GDS19]. Preservation [KBXP23, GLC^{+}21]. preserving [DZH^{+}23, FRS^{+}20, GHY^{+}19, LLL^{+}20]. Pressure [CCJ^{+}17, CH18, FGS^{+}23, HW17, KCFI19]. prevalent [WXY^{+}19]. Price [GCW^{+}18]. Prime [CND17]. Printable [BGL18]. Printed [AWZK23, CND^{+}20]. Printer [GLW^{+}18]. Printers [JWS^{+}23]. Printing [HWJ^{+}21, SLX^{+}18]. Prints [JWS^{+}23]. PrinType [LHFZ22]. PrISM [AYM^{+}22]. PrISM-Tracker [AYM^{+}22]. Privacy [AKPS18, ATP^{+}19, ABL^{+}23, ASM^{+}18, CFLK18, CLW^{+}23, CGH^{+}17, DZH^{+}23, DBBH21, EZC23, FFQ^{+}20, FRS^{+}20, GHY^{+}19, HHSM18, JW18, KBXP23, KMC^{+}18, KBBN21, KBKH20, LBC^{+}22, LKL22, LAH18, LNAH21, LDSZ19, LLL^{+}20, MT19, OSG^{+}22, PSC^{+}17, PLX^{+}18, WCY^{+}19, ZCL21, dGST21, GLC^{+}21]. Privacy-Aware [ABL^{+}23]. Privacy-Enhanced [HHSM18]. Privacy-Enhancing [OSG^{+}22]. Privacy-Friendly [EZC23, LAH18]. privacy-preservation [GLC^{+}21]. Privacy-preserving [DZH^{+}23, FRS^{+}20, GHY^{+}19, LLL^{+}20]. PrivacyShield [PLX^{+}18]. PrivateBus [FFQ^{+}20]. Proactive [CKP^{+}20, KCP^{+}18, MSTT22]. Proactively [CKJH23]. Probe [BWUW21, HDC18, MGC^{+}23a]. Probing [FML^{+}22, JMvB^{+}19]. Procedure [AYM^{+}22]. Process [GLW^{+}18, OGG^{+}18, QWK^{+}19, WCW21]. Processes [MT19]. Processing [DSSS20, HHSM18]. ProCMotive [MBZ^{+}18]. Productivity [CSQ^{+}20]. Profiling [BWUW21]. Program [JYY21]. Programmability [MBZ^{+}18]. Programmable [GH22]. Programming [KGAH22, LSY^{+}23]. Progress [CFZ^{+}19]. Projections [LH20]. Prompto [CSM^{+}20].

Prompts [CSM^{+}20]. proof [WSW^{+}22]. ProspecFit [CBZN19]. Prospective [CBZN19]. Protect [LZZ^{+}20]. Protection [FFQ^{+}20, KBKH20]. Prototyping [HWJ^{+}21, KBC^{+}20, KMH^{+}21, KHW^{+}23, LRH^{+}21]. Provider [CWS^{+}19]. Provides [CHC^{+}22]. Providing [MSTT22]. Provisioning [PLX^{+}18]. Proxemics [MMMS^{+}20, MTF^{+}18]. Proxies [VLR20]. ProxiFit [KNL^{+}23]. Proximity [DGLQ19, HLL17, ILM^{+}21, KNL^{+}23, SSW^{+}17]. Proxy [HYS^{+}23]. Pruning [LL22]. Psychopathology [KSZ^{+}22]. PTSD [SRR^{+}20]. Public [CW18, HSWB21, KIT^{+}20, KML^{+}21, PTK18, RBL^{+}20]. Pull [CFZ^{+}19]. PulmoListener [BLW^{+}23]. Pulmonary [BLW^{+}23]. Pulse [CCJ^{+}17]. Purpose [BCM^{+}23, LH19, LCG^{+}22, NJWFM19]. Purposes [JLD^{+}18]. Push [CFZ^{+}19]. Pushing [GL18, LLLX22, LZZ^{+}21, XYX21]. Pyjamas [KHG^{+}19]. Python [KBC^{+}20].

QR [SLX^{+}18]. Quadmetric [LLL^{+}19]. Qualitative [LP20]. Quality [BSH^{+}17, CDL^{+}19, CHZT19, CJRG21, GLX^{+}22, HHSM18, IMO^{+}18, KNP17, LDC18, LLZ^{+}18, WXL^{+}20]. Quantifiable [GCW^{+}18]. Quantifying [DCMM20, JWK^{+}21]. Quantitative [GWJ^{+}21]. Queries [JKB^{+}18, RLCI22]. Question [CKLB22]. QuickTalk [HLL17]. Quite [KPG^{+}20]. QwertyRing [GYL^{+}20].

Radar [CZG^{+}21, DZH^{+}22, DZH^{+}23, DJX^{+}23, KRS22, PSLS21]. Radars [CLL^{+}22, CYB^{+}23]. Radiation [LYP^{+}22, MGY^{+}23]. Radio [HAY^{+}17, JSNL23, ILM^{+}21]. RAMT [BX19]. Random [KBXP23]. Randomized [DKZ^{+}22]. Range [SBP^{+}22, XYX21, ZCN^{+}20]. Ranging [JMC^{+}21, MSWQ20, NVP18]. Rapid [CDG^{+}17, IB23, PNL^{+}21]. Rataplan [VLR20]. Rate [CFF^{+}22, LAAE^{+}19, LXW^{+}21, WXW^{+}18]. Ray [KASC21, DZZ^{+}21]. rConverse [BAR^{+}18]. Re [BRS21, CLL^{+}22, CCW^{+}20, KBBN21]. Re-check-in [CCW^{+}20]. Re-identification [BRS21, CLL^{+}22, KBBN21]. Reach [CHC^{+}22]. Readerless [WPH^{+}23]. Readily [JWS^{+}23]. Readily-Available [JWS^{+}23]. Reading [DBW^{+}22, ISY^{+}21]. Readmission [DDKL19]. Ready [SZZ^{+}20]. Real [AM19, BWUW21, BX19, CYK^{+}19, CML22, DZH^{+}22, GMS^{+}23, GJG^{+}18, HBX^{+}17, KWAP21, LAAE^{+}19, LWZ^{+}20, MBB20, MVS^{+}20, PLK21, STL^{+}19, ZLL^{+}20]. Real-Time [BWUW21, HBX^{+}17, MBB20, MVS^{+}20, BX19, CYK^{+}19, CML22, DZH^{+}22, GMS^{+}23, LWZ^{+}20, PLK21, STL^{+}19, ZLL^{+}20]. Real-World [KWAP21, MVS^{+}20]. Reality [AOGF^{+}21, BKL^{+}22, CLZL18, CKL23, CND22, CNS21, CJRG21, GCM^{+}19, GCP^{+}22, JCF^{+}21, LAX^{+}23, OSG^{+}22, PB20, RPF^{+}18, dGST21]. RealityReplay [CKL23]. Really [CGH^{+}17]. Realtime [AP18]. Recall [NCM^{+}17, RLY^{+}19]. Receptivity [CSM^{+}20, CPK^{+}19, KMK^{+}19, MKK^{+}21]. Reckoning [RRS^{+}23, ZM22]. Recognition [AES^{+}21, AT19, AYT21, AWSS22, ACB22, BYW^{+}20, BW20, BGK21, CYX^{+}19, CZP20, CHWZ23, DZZ^{+}21, FWJ^{+}22, GZZ^{+}21, GLX^{+}22, GwC^{+}23, GMS^{+}23, GCLL19, GP17, Hal19, HEP21, HEP22, HY22, HJL^{+}22, HP20, HNCP22, HAR18, ILH^{+}19, JTM^{+}22, JSGS23, KLM^{+}18, KHZ22, KNP17, KLT19, KTH^{+}20, KWAP21, LCS^{+}21, LWA23, LHS^{+}23, LT19, LXC^{+}19, LWZ^{+}20, LCH^{+}22, LHFZ22, LMS^{+}23, LLG^{+}23, LWC^{+}22, LYF^{+}21, MZW^{+}18, MBB20, MFSK21, OWH^{+}20, PSLS21, PCYZ18, QCWY19, QWK^{+}19, RH19, SHZ^{+}20, SGZ^{+}23, SWLG22, WXY^{+}19, WPJ^{+}19, XKNM20, XZZ21, ZWZ^{+}21a]. Recognition-Based [HJL^{+}22]. Recognize [DGH^{+}20, GCP^{+}22, YZS^{+}20]. Recognizer [GLJ^{+}21, JGZ^{+}21]. Recognizes [SLX^{+}18]. Recognizing [FFT^{+}18, SCZ^{+}20]. Recommendation [DYL^{+}19, GHY^{+}19, LHC^{+}22, TFL^{+}19, WZF^{+}20, GLC^{+}21]. Recommendations [CBR23, DLH^{+}18, DBY^{+}17, DBBH21, RQ20]. Recommender [BBBN22]. Reconfigurable [KHW^{+}23]. Reconstructing [BXY^{+}21, HYS^{+}23, LSW^{+}19]. Reconstruction [ASG^{+}23, HHZ23, WYL^{+}21]. Records [HFX^{+}21, ILH^{+}19]. Recovery [FYY^{+}21, JYY21]. Recurrent [AO19, GSS^{+}19]. Recursive [FMZN22]. Redeployment [JZWL19]. Reducing [CZT22, NTT^{+}21]. referenced [GFK^{+}18]. Refinement [MH23]. Reflected [LZG^{+}21]. Reflection [BWH^{+}22, FALW18, KLK^{+}20, KXAH18]. Reflective [BNW22]. Refractive [AGU^{+}22]. Region [WXY^{+}19]. Region-of-interest [WXY^{+}19]. registered [LHS^{+}21]. Registration [SKS^{+}21]. Regular [IRA^{+}21]. Regulating [CGJC19, KJKL19]. Regulation [CFF^{+}22]. REHASH [BRYH21]. Reinforcement [DGZ^{+}21, JZWL19, LGKM20]. Relacks [KJS20]. Related [AJA^{+}18, DBBH21, MLX^{+}20]. Relationship [LNT^{+}19, TAHH23, WMW^{+}20]. Relationships [DCMM20]. Relative [DLD^{+}21]. Relatives [MT19]. Reliable [DSR^{+}22, GLG^{+}19, KJS20, LCC^{+}22]. Remember [BML17]. Reminder [BML17]. Remote [CBEG17, LZZ^{+}20, QCP^{+}18]. Remotion [QCP^{+}18]. Rendering [XYG^{+}20]. Repair [WAL^{+}19]. Replay [HW21, QCP^{+}18]. Replaying [CKL23]. Replicable [PSC^{+}17]. Report [CSK^{+}18, PRS^{+}23, RLY^{+}19]. Reports [MRCdCRV^{+}21]. Representation [FMZN22, KNP17, LDSZ19, SH20, YXW^{+}20]. Representations [GBLM17, SC22]. Requirements [RQ20]. Research [LGL23, MW22]. Resilience [ATQ^{+}21]. Resilient [VLR20, SXF^{+}22]. Resistance [HW21]. Resistant [GJC^{+}21a]. Resolution [LLF^{+}21, RHMZ^{+}21]. Resonance [LCR^{+}18]. Resource [CRS^{+}18, RR22, RBL^{+}20, XFW^{+}20, GBLM17]. Resource-Constrained [CRS^{+}18, XFW^{+}20]. Resources [BWH^{+}22]. Respiration [BAR^{+}18, CXC^{+}19, HBX^{+}17, LXC^{+}19, LZG^{+}21, YHW^{+}18, ZWX^{+}20]. Respiratory [GNA^{+}19, LAAE^{+}19]. Response [CYK^{+}23b, HSK^{+}22, PLW^{+}22]. Retail [DJN^{+}18, FRB^{+}21]. Retention [DPB21]. Rethinking [KEI^{+}18]. Retinoscopy [AGU^{+}22]. RetroSphere [BKL^{+}22]. Reveal [BKM21, MMMS^{+}20, WMW^{+}20]. Revealing [XL19]. ReVibe [RLY^{+}19]. Review [JCR22, MW22]. Reviewing [FML^{+}22]. Revisitation [CCX^{+}18, CCW^{+}20]. Revisiting [BWH^{+}22]. Revive [HFX^{+}21]. RF [CKKM20, CLL^{+}22, CYK^{+}23a, DSSD23, DZZ^{+}21, FXC^{+}21, GZL^{+}21, LCC^{+}22, LXC^{+}19, RGL^{+}18, SYYL22, VJH^{+}18, WXW^{+}18, WXY^{+}19, WSW^{+}22, YHW^{+}18, ZCC^{+}20]. RF-based [LCC^{+}22]. RF-Chain [WSW^{+}22]. RF-Data [VJH^{+}18]. RF-ECG [WXW^{+}18]. RF-Focus [WXY^{+}19]. RF-Identity [FXC^{+}21]. RF-Mic [CYK^{+}23a]. RF-ray [DZZ^{+}21]. RFaceID [LYF^{+}21]. RFCam [CML22]. RFID [CYK^{+}23a, DLD^{+}21, FXC^{+}21, GMY^{+}17, HQY^{+}18, HWZZ19, LXW^{+}21, LYF^{+}21, WXW^{+}18, WXY^{+}19, WPJ^{+}19]. RFID-based [LXW^{+}21, LYF^{+}21]. RFID-Enabled [DLD^{+}21]. RFIDs [CXC^{+}19, FGL18, WSW^{+}22]. RFTemp [BS21]. RGB [CLL^{+}22]. RGB-D [CLL^{+}22]. Rhythms [DDKL19]. Rich [CLW^{+}18].

Ride [GCW^{+}18]. Ride-on-demand [GCW^{+}18]. Right [MSTT22]. Ring [BRA^{+}22, CLYZ22, GYL^{+}20, KJP^{+}19, LGY^{+}20, YLL^{+}20]. Ring-sized [YLL^{+}20]. Rings [LHS^{+}23, LYQ^{+}21, LHY^{+}22]. RISC [XFW^{+}20]. Risk [DDKL19, HFS^{+}19, KBBN21, RHI^{+}23, WZG^{+}19]. Risk-Aware [HFS^{+}19]. Risks [LBC^{+}22, LKL22, dGST21]. Road [ZSG^{+}21, KYJ^{+}23]. Roads [KML^{+}21]. RoboCOP [TAEB17]. Robot [HBH20, LYWZ22]. Robotic [GZC^{+}23, TAEB17]. Robots [DHC^{+}18]. Robust [CPC^{+}18, CDG^{+}17, CJF^{+}23, DSSD23, DZH^{+}22, EEY18, EBTN22, GLX^{+}22, GZH^{+}21, KLM^{+}18, LWA23, MLMH^{+}18, RPB^{+}18, XKNM20]. Robustness [CKJH23]. ROI [JSF^{+}18]. ROI-Based [JSF^{+}18]. Role [ADY^{+}20, DJN^{+}18, FCFW^{+}17, GCGD17, SSR^{+}19]. Room [DBW^{+}22, HSH^{+}23, SRK^{+}19, SFR^{+}20]. Room-Scale [SRK^{+}19, SFR^{+}20]. Route [NJWFM19]. Routing [JHP^{+}17]. RoVaR [DSSD23]. Rule [BRA^{+}22]. Run [HYS^{+}23]. Run-time [HYS^{+}23]. Running [HDW^{+}17]. Runtime [LGM^{+}21]. Runtime-evolutionary [LGM^{+}21].

s [PRC^{+}18]. S3 [FYX^{+}21]. Safeguarding [GLW^{+}18, LCR^{+}21]. Safety [CS19, JMC^{+}21, MSTT22]. Sampling [RQ20]. SAR [RSSN21]. Satisfaction [AJV20, PM19]. SATURN [AZS^{+}18]. Savings [JSC17]. Scalability [CBR23]. Scalable [CLZ^{+}21, EY20, IB23, SBP^{+}22, ZCA^{+}21]. Scale [AKX^{+}19, CLZ^{+}21, FFQ^{+}20, HHFM17, HAL^{+}20, LT19, LKL19, NHL^{+}20, SRK^{+}19, SFR^{+}20, WZF^{+}20, JWCC^{+}20, KSZ^{+}22, SJL^{+}21]. Scaling [AvdBF^{+}19, HAR18]. SCAN [MZST17]. Scanners [JMvB^{+}19]. Scenarios [LWZ^{+}20, LCZ^{+}21]. Scene [CRGR22].

Schedule [PLC^{+}21]. Scheduling [CME^{+}19, LHW^{+}19, XFW^{+}20]. Scheme [XLL^{+}21]. School [CCK^{+}18, JLZ^{+}23]. Science [BSH^{+}17, YPJdB19]. Scleral [MBP^{+}17]. Sclerosis [AKB23, BOH^{+}21, GZY^{+}21]. Score [JWK^{+}21]. Scratch [HNCP22]. ScratchThat [WAL^{+}19]. ScraTouch [IY20]. Screen [DJL^{+}22, PLX^{+}18, TXWZ22]. Screen-to-Camera [TXWZ22]. Screening [GLG^{+}19, SRR^{+}20, WPH^{+}23]. Screens [LGM^{+}22]. Scrolling [LYC^{+}19]. Search [JHH21, RR22]. Seat [CJRG21, CJRG21]. Seating [GRS^{+}22]. Second [CNXG22, HHIF19, KLK^{+}20]. Second-factor [CNXG22]. Second-person [KLK^{+}20]. Secret [LKBK19]. Secure [LYP^{+}22]. Securing [SS22]. Security [FMA^{+}19, MT19, WBG^{+}19]. See [HSWB21, LJK^{+}21, NZL^{+}20]. Seeking [GNA^{+}19]. Segment [GwC^{+}23]. SegmentatiOn [DSSS20]. SeismoWatch [CCJ^{+}17]. Selecting [MQZH23]. Selection [IYS22, KLM^{+}18, LSE21]. Selective [BS17]. Self [ADY^{+}20, AZS^{+}18, BKL^{+}22, DDP^{+}17, DLD^{+}21, HEP22, ISY^{+}21, JTM^{+}22, JYY21, JHK^{+}21, KBL18, KKZ^{+}18, LSE21, MRCdCRV^{+}21, PRS^{+}23, RLY^{+}19, SOL19, SKM^{+}19, WXZ^{+}20, ZPZ^{+}20]. Self-Administering [KKZ^{+}18]. Self-Checkout [DLD^{+}21]. Self-Contained [BKL^{+}22, KBL18]. Self-Experiments [DDP^{+}17]. Self-Powered [ZPZ^{+}20, AZS^{+}18]. Self-report [PRS^{+}23, RLY^{+}19]. Self-Reports [MRCdCRV^{+}21]. Self-Supervised [HEP22, JTM^{+}22, SOL19, ISY^{+}21]. Self-Tracking [ADY^{+}20, JYY21, JHK^{+}21, LSE21, SKM^{+}19]. Semantic [HYS^{+}23, LWC^{+}22, XLX^{+}20, YXW^{+}20]. Semantic-aware [YXW^{+}20]. Semantic-Discriminative [LWC^{+}22]. Semantic-enhanced [XLX^{+}20]. Semi [ZLB^{+}17]. Semi-Automated [ZLB^{+}17]. SenCAPTCHA [FCQR20]. Sensation [JCF^{+}21]. Sensations [FMP^{+}23]. Sense [CKKM20, LJZ^{+}22, ZPPC^{+}20]. SenseCollect [CLTS21]. Sensemaking [GF17, RLGN19]. Senses [CSQ^{+}20]. Sensing [ATQ^{+}21, AIP^{+}18, AKX^{+}19, AWSS22, AAM^{+}23, BCM^{+}23, BSMP23, BSH^{+}17, CCY^{+}23, CPC^{+}18, CLW^{+}18, CZX^{+}22, CLYZ22, CZL^{+}22, CKJH23, DSSD23, DZH^{+}23, DJX^{+}23, DMAH18, DDKL19, FMZN22, FFT^{+}18, GZZ^{+}21, GLX^{+}22, GBLM17, GZL^{+}21, GKO^{+}23, GMC18, GNA^{+}19, GJG^{+}18, HW17, HYY^{+}23, HKK^{+}18, KSC^{+}21, KBXP23, KOY^{+}19, KHG^{+}19, KNL^{+}23, KME^{+}19, LH19, LLZ^{+}20, LZL^{+}22, LLLX22, LLCY21, LYLT21, LZZ^{+}21, LZT23, LZZ^{+}23, LLC^{+}20, LXW^{+}21, LWZ^{+}20, LCH^{+}22, LWX^{+}22, LSY^{+}18, LYC^{+}19, LYCW20, MZJ20, MLX^{+}20, MLH^{+}23, MVS^{+}20, MLPGP21, MRCdCRV^{+}21, MDK^{+}22, MMG^{+}19, MSL^{+}19, OHW^{+}20, PTMM^{+}22, RH19, RTZ^{+}20, RMG^{+}17, RLM^{+}19, SZZ^{+}20, SAD^{+}17, SXF^{+}22, VBS^{+}21, XFW^{+}20, XYX21, YFX^{+}20, YPJdB19, ZWX^{+}20, ZPPC^{+}20, ZCN^{+}20, ZPZ^{+}20, ZWZ^{+}21b, ZCC^{+}20, ZSG^{+}21]. Sensing-Based [MDK^{+}22]. Sensitive [IN20]. Sensor [AZK^{+}23, ACB22, BAR^{+}18, BWT^{+}18, BKM21, CLC^{+}22, CZX^{+}22, CSHN17, CLTS21, CHZT19, CBT18, DSSS20, DXS^{+}22, DGH^{+}20, EEY18, FZZ^{+}22, FCFW^{+}17, GZH^{+}21, GF17, GMC18, HPWW18, ISSK20, JCR22, KBXP23, KLM^{+}18, KHZ22, KJP^{+}19, KBC^{+}20, LH19, LLW^{+}22, LWC^{+}22, MZST17, MCCH21, MFSK21, MNV^{+}19, SGZ^{+}23, WYL^{+}21, XKNM20, ZZN^{+}20]. Sensor-Based [ACB22, CLTS21, KJP^{+}19, LWC^{+}22, SGZ^{+}23]. Sensor-free [CZX^{+}22]. Sensor-Orientation-Independent [KLM^{+}18]. Sensors [AES^{+}21, ASS^{+}19, ASS^{+}21, ASTW19, AYM^{+}22, BFS^{+}17, BRS^{+}20, BLH^{+}17, BXY^{+}21, CDA23, CML^{+}20, CZP20, CJF^{+}23, DHC^{+}18, DGS18, FYX^{+}21, FCQR20, GDS19, GWD^{+}17, GPB^{+}23, KASC21, LDC18, LYH^{+}20, LYD^{+}23, MLMH^{+}18, MVS^{+}20, PCYZ18, PM19, QWK^{+}19, RGL^{+}18, RHMZ^{+}21, SGO^{+}23, SFR^{+}20]. Sensory [CQCH21, LEP21]. Sensurfaces [PTMM^{+}22]. Sequences [NHL^{+}20]. Sequential [HWZZ19]. SeRaNDiP [KBXP23]. Series [BMG18, DSSS20]. Service [AvdBF^{+}19, DBY^{+}17]. Services [DLD^{+}21, GCW^{+}18, HHSM18, HYS^{+}23]. Serving [AvdBM^{+}22]. Set [EUHB23, KDL^{+}21]. Sets [KHZ22]. Settings [CTTMT20, GMS^{+}23, PM19]. Sexual [WMW^{+}20]. Sexual-Minority [WMW^{+}20]. Shadow [HBH20]. ShadowSense [HBH20]. Shallow [ISSK20]. ShaPe [DSSS20, CFF^{+}22, KHSW17]. Shape-changing [CFF^{+}22]. Share [AKPS18, WXL^{+}20]. Shared [AMG^{+}19, GBLM17, HSWB21]. Sharing [AHB17, CXR^{+}18, CYK^{+}19, GM19, HCH^{+}22, LEP^{+}19, LDK^{+}20, PSC^{+}17, WXL^{+}20]. ShaZam [MWJ^{+}21]. Shear [KCFI19]. Shelf [AIP^{+}18]. Shield [ZCL21]. Shoppers [BMZN20]. Shopping [BDM^{+}18]. Shot [CND22, FWJ^{+}22]. Shots [KNP17]. Should [CKLR21]. Shower [KFFC18]. Side [FYX^{+}21, FZS^{+}23, GMMH21]. Side-Channel [FYX^{+}21, GMMH21]. Side-Channels [FZS^{+}23]. Sided [DWHR17]. SiFi [GL18]. Sign [GZL^{+}21, JGZ^{+}21, LHS^{+}23, MZW^{+}18, PLK21, ZCC^{+}20]. Signal [MH23]. Signaling [SJL^{+}21]. Signals [CDA23, CQCH21, CZX^{+}22, DWZZ19, GLX^{+}22, GNA^{+}19, HMA21, HNL22, HY22, HAY^{+}17, LLLX22, LXC^{+}19, LZG^{+}21, LYCW20, SHZ^{+}20, SSW^{+}17, SYYL22, VSY^{+}20, WYL^{+}21, YHW^{+}18, YYW^{+}20, ZWZ^{+}21b]. Signature [DWZZ19]. Signatures [KDL^{+}21]. SignFi [MZW^{+}18]. SignRing [LHS^{+}23]. Silent [GJL^{+}20, JGX^{+}22]. Silently [CLC^{+}22]. Silhouettes [ARSN22, HHX^{+}20]. Silver [CND^{+}20]. Simple [GCW^{+}18, PCYZ18, SKS^{+}21]. Simplify [PSC^{+}23]. Simulator [LEP21]. Simultaneously [VSY^{+}20]. Single [CLYZ22, CCK^{+}18, GL18, KNL^{+}23, SCZ^{+}20]. Single-Touch [CCK^{+}18]. Situ [BRA^{+}22, CBZN19, CKL23, LSY^{+}23, RLCI22]. Situated [KTM^{+}18]. Situation [CRGR22]. Situational [ELD19]. Situations [CKLR21]. Size [CFF^{+}22, GH18]. sized [YLL^{+}20]. Sketches [SRR^{+}20]. Ski [GWD^{+}17]. SkillFence [HWJ^{+}22]. Skills [NTT^{+}21, WCW21, ZLB^{+}17]. Skin [DHC^{+}18, KBL18, KMH^{+}21, MGY^{+}23]. SkinKit [KMH^{+}21]. SkinLink [KHW^{+}23]. SkinWire [KBL18]. Sleep [CH18, CLW^{+}18, CXT^{+}21, DLH^{+}18, HAY^{+}17, LP20, LXC^{+}19, YYW^{+}20, ZPPC^{+}20]. Sleep-tracking [LP20]. SleepGuard [CLW^{+}18]. Slow [PZH^{+}18]. Small [MQZH23]. Smart [ASM^{+}18, ACB22, BLS^{+}18, BMR^{+}20, BA20, CKP^{+}20, CZD^{+}21, CXZ^{+}21, CLYZ22, CLY^{+}23, CRP21, CND17, CND22, CVL20, CRB^{+}21, DPW^{+}21, DWHR17, GS20, GZH^{+}21, HKSM23, HNCP22, HAL^{+}20, JSC17, JGSC19, JSC^{+}18, LZZ^{+}20, LFH18, LHC^{+}22, LL22, LWZ^{+}20, LJK^{+}21, LSY^{+}23, MJ19, MFSK21, RTZ^{+}20, VJH^{+}18, XZZ21, YZ21]. Smart-Eyewear [DWHR17]. Smart-Home [ACB22]. SmartASL [JZG^{+}23]. Smartglasses [MLH^{+}23]. SmartKC [GBS^{+}21]. SmartLOC [DJL^{+}21]. Smartphone [AIP^{+}18, AZK21, BOH^{+}21, BKM21, CCX^{+}18, CRMG22, DJL^{+}21, DMAH18, GBS^{+}21, GJL^{+}20, GZH^{+}21, HYY^{+}23, IRA^{+}21, IB23, JMC^{+}21, JHWF22, JHK^{+}21, KJKL19, LDW^{+}18, LCX^{+}22, LGY^{+}20, LLWX21, LZG^{+}21, LYKH18, MLPGP21, MRCdCRV^{+}21, NVP18, PBM^{+}18, PRS^{+}23, RQ20, STL^{+}19, SLX^{+}18, SC22, WPH^{+}23, BFS^{+}17, MBP^{+}17]. Smartphone-Based [AIP^{+}18, BOH^{+}21, JHK^{+}21, AZK21, GBS^{+}21, JMC^{+}21, LZG^{+}21, NVP18, PBM^{+}18, STL^{+}19, WPH^{+}23, MBP^{+}17, BFS^{+}17]. Smartphones [BSMP23, CNXG22, CGH^{+}17, DCMM20, GCLL19, GZZ^{+}17, JNB^{+}20, JCL^{+}20, LYC^{+}19, LYCW20, PM19, SZJ^{+}17, TAHH23, Tor23]. Smartring [BRA^{+}22]. Smartwatch [CLW^{+}18, CGJC19, HBX^{+}17, JKK^{+}17, KGL19, LFH18, LZT23, LEP^{+}19, SPRK23]. Smartwatch-Based [HBX^{+}17]. Smartwatches [BAT22, CXT^{+}21, GKO18, HEUH18, HSK^{+}22, LAAE^{+}19, RLCI22, ZWZ^{+}21a]. SmileAuth [JCL^{+}20]. Smoke [CZG^{+}21]. SmokeMon [ASC^{+}22]. Smoking [ASC^{+}22, CML^{+}20]. SmokingOpp [CML^{+}20]. Smooth [XYG^{+}20]. Snooping [GSS^{+}19]. Social [AHB17, CTTMT20, CCY^{+}17, FML^{+}22, GWJ^{+}21, GH18, HBH20, LEP^{+}19, MLPGP21, PMGP19, SEG^{+}18, WBG^{+}19, WMW^{+}20]. Soft [GH22]. Software [LRH^{+}23]. Soil [CZX^{+}22, DJX^{+}23, ISSK20]. Soil-Monitoring [ISSK20]. Solar [PCF^{+}23]. Someone [DBBH21]. Sonar [JHL^{+}21, LZZ^{+}20]. Songs [DPB21, FMP^{+}23]. SonicASL [JGZ^{+}21]. SonicFace [GJC^{+}21b]. Sound [BAT22, BA20, DMAH18, GJC^{+}21a, GLJ^{+}21, HCYZ18, LXC^{+}19, SZJ^{+}17]. Sound-Activity [LXC^{+}19]. Sounds [HW21]. SoundSignaling [AP18]. SoundTrak [ZXW^{+}17]. Source [KMC^{+}18, GCW^{+}18]. Sourcing [ARD^{+}18, KMC^{+}18]. Space [BGL18, BCK^{+}22, CCX^{+}18, IB23, JCR22, PBM^{+}18]. Spaces [CND22, KIT^{+}20]. SpaceX [IB23]. Spacing [EWM^{+}20]. SpARklingPaper [DJL^{+}22]. Sparse [FMZN22, LH19, PSLS21]. Spatial [BBH17, QCWY19, WXY^{+}17, dGST21].

Spatial-Temporal [QCWY19]. Spatio [HYH21, WLZ^{+}19, YXW^{+}20]. Spatio-Temporal [HYH21, WLZ^{+}19, YXW^{+}20]. Spatiotemporal [CQCH21]. Speaker [BLS^{+}18, LCL20]. Speakers [CKP^{+}20, GS20, LZZ^{+}20]. Spear [ZCL21]. Spectral [HHZ23, HWJ^{+}23]. Spectroscopy [JMvB^{+}19, KGL^{+}18]. Spectrum [KSCS19, LAX^{+}23]. Speech [CYK^{+}23a, DWL^{+}22, GJL^{+}20, JGX^{+}22, MLMH^{+}18, WPJ^{+}19, WAL^{+}19, ZWZ^{+}21b, MLMH^{+}18]. Speed [LDW^{+}18]. Spell [ZWZ^{+}21a]. SplitSR [LLF^{+}21]. Spondyloarthritis [JHK^{+}21]. Spoofing [GJC^{+}21a, ZLD^{+}21]. Spoofing-Resistant [GJC^{+}21a]. Spoon [CFF^{+}22]. Spot [PRC^{+}18]. Spotting [AO19, LLCY21]. Spraying [CZT22]. Sprintz [BMG18]. SquiggleMilli [RSSN21]. SSpoon [CFF^{+}22]. Stage [CPK^{+}19, GMS^{+}23, ZPPC^{+}20]. Standard [GGJ^{+}22]. Standing [HKSM23]. start [TFL^{+}19]. State [AES^{+}21, AYM^{+}22, HEP22, JNB^{+}20, KMK^{+}19, LSSH23]. State-of-Receptivity [KMK^{+}19]. State-of-the-Art [AES^{+}21]. States [DCMM20]. Stationary [YFX^{+}20]. Statistical [GZY^{+}21]. Status [MTH19, WMW^{+}20]. Steady [SZZ^{+}20]. Step [LKL19]. Stick [GPW^{+}18]. Stick-on [GPW^{+}18]. Still [PTK18]. Stimulant [JCF^{+}21]. Stimulation [KAH18, LEP21, NTT^{+}21]. Stimuli [AHB17]. Stochastic [JHH21]. Stock [ECF^{+}20]. Storage [AM19]. Store [FRB^{+}21]. Storyline [GOZ^{+}17]. StoryTeller [EY20]. Storytelling [LFL^{+}23]. Strategies [KGN^{+}20]. Stray [GPW^{+}18]. Streak [LKL19]. Stream [LLZ^{+}22]. Streaming [FSW19, ZWZ^{+}21a]. Streets [JXZ^{+}20]. Strength [GZZH20]. Strength-based [GZZH20]. Stress [ATQ^{+}21, HMA21, KME^{+}19, LS20, LZZ^{+}23, SvBH^{+}19, WTA^{+}20, KME^{+}19].

Stress-Resilience [ATQ^{+}21]. Stressful [BRS^{+}20]. Strike [HDW^{+}17]. Strip [WPH^{+}23]. Stroke [GLG^{+}19, KJP^{+}19]. Strokes [BXY^{+}21]. Stroking [FMP^{+}23]. Structure [XKNM20]. Structures [SRK^{+}19, SFR^{+}20]. Student [CSK^{+}18, DGS18, DBW^{+}22, GRS^{+}22]. Students [MRCdCRV^{+}21, MDK^{+}22]. Study [AvdBF^{+}19, BRA^{+}22, BOH^{+}21, CFZ^{+}19, DKZ^{+}22, DDKL19, GS20, ILH^{+}19, JSC^{+}18, JKK^{+}17, JWCC^{+}20, KKV^{+}22, KML^{+}21, KEI^{+}18, KTEM21, LP20, PRS^{+}23, PSC^{+}17, WMW^{+}20, YLL^{+}20]. Studying [KFFC18]. style [ZWZ^{+}21a]. Stylistic [AP18]. Stylus [FYX^{+}21]. Substitution [FS18]. Substrates [CND^{+}20]. Subtle [CYK^{+}23a, LYQ^{+}21, LHY^{+}22, PLX^{+}18, SCZ^{+}20, YLL^{+}20]. Sucrose [JMvB^{+}19]. SugarMate [GZZ^{+}17]. Suggests [KVZ^{+}19]. SUME [XLX^{+}20]. SunBox [TXWZ22]. Super [LLF^{+}21]. Super-Resolution [LLF^{+}21]. Supervised [HEP22, JTM^{+}22, SOL19, SEG^{+}18, SH20, GZW^{+}22, ISY^{+}21]. Supply [WSW^{+}22]. Support [BWH^{+}22, BMZN20, GAS^{+}18, HYS^{+}23, JHWF22, MSTT22, PRC^{+}18, SKM^{+}19, ZXY21]. Supporting [EZC23, HHIF19, KWAP21, LMT19, LGL23, LSE21, PCF^{+}23, PLX^{+}18, WAL^{+}19]. Supports [TK17]. Surface [HYY^{+}23, LH20, SCZ^{+}20]. Surfaces [ASS^{+}21, CLZ^{+}21, GH22, GYL^{+}20, IY20, PTMM^{+}22, ZPZ^{+}20]. Surgery [DDKL19, WGL^{+}21]. Surrogate [BOH^{+}21]. Surrounding [CHC^{+}22]. Surveillance [ALC^{+}20]. Survey [AHS^{+}18, LRH^{+}21]. Survival [DJN^{+}18]. Survive [HFX^{+}21]. Survivors [KJP^{+}19]. Sustainable [FCFW^{+}17]. Sustained [MGC^{+}23b]. Sustaining [WXZ^{+}20]. Sweet [PRC^{+}18]. Swing [JHH21]. SwingNet [JHH21]. SwipePass [CNXG22]. Swivel [CJRG21]. SwiVR [CJRG21]. SwiVR-Car-Seat [CJRG21]. Symmetric [MGY^{+}23].

Symptoms [DGLQ19, GZY^{+}21]. Synapse [JHWF22]. Synchronization [MH23]. Synchronized [ZXY21]. Synchrony [CBEG17, GDS19, SGO^{+}23]. SyncUp [ZXY21]. Syndrome [CWS^{+}19, DH17]. Syndromic [ALC^{+}20]. Synthesis [KW23]. Synthetic [SPRK23]. System [CLZL18, CTTMT20, CZD^{+}21, CWP^{+}18, CNS21, DGZ^{+}21, DJY^{+}22, EEY18, FRB^{+}21, GZNL22, HBX^{+}17, HQC^{+}23, HBK21, ILH^{+}19, JZWL19, JDZ^{+}21, JLZ^{+}23, KSC^{+}21, KJP^{+}19, KGN^{+}20, KXAH18, KPK^{+}18, LILN19, LCC^{+}22, LLZ^{+}18, LYD^{+}23, MFSK21, NNMY23, NVP18, NTO^{+}18, PNL^{+}21, PLX^{+}18, WZF^{+}20, WXL^{+}20]. Systematic [JCR22, JWCC^{+}20, RHMZ^{+}21]. Systems [AHS^{+}18, BBBN22, BML17, CZL^{+}22, CME^{+}19, DSM17, FYW^{+}19, FFQ^{+}20, FALW18, GMY^{+}17, HP20, HNCP22, HWJ^{+}22, IN20, KW23, LDK^{+}20, MLX^{+}20, SBP^{+}22, YFX^{+}20].

Table [CYX^{+}19]. Tablet [DJL^{+}22]. Tabletop [CCK^{+}18, FKI18]. Tactile [FMP^{+}23, GZC^{+}23, KCFI19]. Tag [WXW^{+}18, WXY^{+}19, ZCA^{+}21]. TagFree [FGL18]. Tags [KLM^{+}18, MTF^{+}18, SBP^{+}22]. Taking [ECF^{+}20]. Tale [TAHH23]. Talk [CW18, CKP^{+}20, WMS^{+}21]. Tangible [FKI18, LFH18]. TAO [BPA23]. Tape [CND^{+}20]. Tapping [BOH^{+}21, FMP^{+}23]. target [LCS^{+}21]. Targeted [BWUW21]. Task [CDL^{+}19, CZP20, DKZ^{+}22, ELD19, GHvB^{+}17, HP20, KPL20, LHS^{+}21, PCYZ18, RHI^{+}23, SH20, XFW^{+}20, GBLM17, LSY^{+}18, SOL19]. Task-Load [ELD19]. Tasks [DZH^{+}23, KCP^{+}18]. Tattoo [WPJ^{+}19]. Taxi [WZF^{+}20]. Taxonomy [LRH^{+}21]. Teacher [MMMS^{+}20]. Teaching [CKKM20]. Team [GMS^{+}23, JWK^{+}21]. Technical [BCK^{+}22, KEI^{+}18]. Technique [HJL^{+}22, IY20]. Techniques [RRS^{+}23].

Technologies [AKPS18, BWH^{+}22, GWJ^{+}21, JYY21, KTEM21, LP20]. Technology [ADY^{+}20, CLRH19, ELD19, HMM^{+}20, OSG^{+}22]. Teeth [ASS^{+}21]. Telecommunication [LYD^{+}23]. TelecomTM [LYD^{+}23]. Telepresence [YLMR20]. Temperature [BS21, BSMP23, HSH^{+}23, ISSK20]. Template [RDOA22]. Temporal [BPA23, CKL23, GwC^{+}23, GCC^{+}21, HYH21, LS20, MTH19, QCWY19, WLZ^{+}19, XKNM20, YXW^{+}20]. Term [BLS^{+}18, GS20, HHFM17, HKK^{+}23, TK17]. Test [KCK^{+}19, PNL^{+}21, WPH^{+}23]. Test-Driven [KCK^{+}19]. Testing [CRMG22, QCP^{+}18]. Text [FZZ^{+}22, GYL^{+}20, LLL^{+}19, LHFZ22, LYY^{+}17]. Theatre [SGO^{+}23]. Their [AKPS18, TAHH23]. Them [BRA^{+}22]. There [CKP^{+}20, DBW^{+}22]. Thermal [AVD^{+}17, AKN^{+}19, AHB17, ASC^{+}22, HSH^{+}23, HHX^{+}20, RHMZ^{+}21]. Thermometers [HLS20]. Thermostats [JGSC19]. Thin [AZS^{+}18, GPB^{+}23]. Things [ABL^{+}23, LHK^{+}19, ASM^{+}18, BMG18, FDL^{+}18, FMA^{+}19, GM19, KPK^{+}18, RPF^{+}18]. Third [LLZ^{+}18]. Third-Eye [LLZ^{+}18]. Three [FMP^{+}23, LGM^{+}22]. Through-wall [LLZ^{+}20, ZCN^{+}20]. Through-Wrist [HSK^{+}22]. Thumb [CLYZ22, LLL^{+}19, LYY^{+}17]. Thumb-to-Finger [LLL^{+}19]. Thumb-to-Index-Finger [CLYZ22]. ThumbAir [GO22]. Time [BWUW21, BMG18, CCJ^{+}17, CKP^{+}20, CPK^{+}19, DSM17, FYW^{+}19, GL18, HBX^{+}17, IN20, KKV^{+}22, LDS^{+}18, MBB20, MVS^{+}20, PL23, DSSS20, BX19, CYK^{+}19, CML22, DZH^{+}22, GMS^{+}23, HYS^{+}23, LEP21, LWZ^{+}20, PLK21, PLX^{+}18, STL^{+}19, ZLL^{+}20]. Time-Based [GL18]. Time-of-Day [DSM17]. Time-Sensitive [IN20]. timE-Series [DSSS20]. Timing [LDS^{+}18].

Tiny [WS17]. Toccata [LMT19]. Today [KPG^{+}20]. Together [XL19]. Told [HPWW18]. Too [CSQ^{+}20, KL19, LDS^{+}18]. Tool [ABL^{+}23, HWJ^{+}21, JXZ^{+}20, JMC^{+}21]. Toolkit [AMG^{+}19, DSM17, PSC^{+}23, Hal19]. Toolkits [FCFW^{+}17, LRH^{+}21]. Tools [CSK^{+}18, RLGN19]. Topographer [GBS^{+}21]. Topography [ASC^{+}22]. Topology [LYH^{+}20]. Topology-Aware [LYH^{+}20]. Total [GCP^{+}22]. Touch [AZEL17, CLD^{+}20, CCK^{+}18, DWHR17, FZZ^{+}22, GKO18, GZC^{+}23, HBH20, IY20, IAE^{+}19, KSC^{+}21, KTM^{+}18, LGM^{+}22, MGY^{+}23, PTMM^{+}22, PLX^{+}18, SZZ^{+}20, WXZ^{+}20]. Touch-and-Heal [GZC^{+}23]. Touch-Based [AZEL17]. TouchKey [MGY^{+}23]. Touchpad [LYY^{+}17]. Tourgether [CYK^{+}19]. Tourist [CYK^{+}19]. Toy [WTA^{+}20]. TraceMatch [CBEG17]. Traces [NJWFM19]. Tracing [BRS21, KASC21]. Tracing-based [KASC21]. Track [DH17, LLL21, VSY^{+}20, KHSW17]. Tracker [GCC^{+}21, RR22, AYM^{+}22]. Trackers [BNW22, TK17]. Tracking [AKN^{+}19, ADY^{+}20, AYM^{+}22, AZK^{+}23, BKL^{+}22, BX19, CFZ^{+}19, CLT^{+}21, CJF^{+}23, CYB^{+}23, CBT18, CWP^{+}18, DSSD23, DGLQ19, FSW19, GJC^{+}21b, GZZH20, GZH^{+}21, HHX^{+}20, HBK21, JHH21, JYY21, JHK^{+}21, KHSW17, LSSH23, LSE21, LCFT19, LZL^{+}22, MMMS^{+}20, PWP19, SBP^{+}22, SKM^{+}19, SWLZ21, WRLK22, ZXW^{+}17, ZCA^{+}21, LP20]. Tract [LYCW20]. Trade [BDR17, RHMZ^{+}21]. Trade-offs [BDR17, RHMZ^{+}21]. Traditional [CS22b]. Traffic [BDPH21, HAL^{+}20, JLD^{+}18, LKZ^{+}21, LMS^{+}23, LYD^{+}23, ZSG^{+}21]. Training [CKKM20, CBZN19, CSM^{+}20, DZH^{+}23, DJL^{+}22, KWAP21, LAX^{+}23, STL^{+}19, ZLB^{+}17]. Trajectory [ASG^{+}23]. Transaction [DBY^{+}17]. Transfer [CHZT19, DYL^{+}19, FDL^{+}18, KCFI19, LCH^{+}22, MWJ^{+}21, QCWY19, RH19]. Transferred [CND^{+}20]. Transit [CCJ^{+}17]. Transition [AYM^{+}22]. Translation [PLK21]. Transmissions [KSCS19]. Transparency [BWUW21, CNS21]. Transport [CSHN17]. Transportation [FYW^{+}19]. Travel [FYW^{+}19, LCG^{+}22, NTO^{+}18]. Treatment [LDS^{+}18]. Tremolo [NTT^{+}21]. Trends [AJA^{+}18]. Trial [JHWF22, MGC^{+}23b]. Trial-and-Error [JHWF22]. Trials [DKZ^{+}22]. Triboelectric [AZS^{+}18]. Triggering [HCH^{+}22]. Trimodal [YLL^{+}20]. Trip [LCG^{+}22, NJWFM19]. Trust [CRGR22, DPW^{+}21]. Trustworthy [HHSM18]. Turning [AvdBM^{+}22, LH20]. TurnsMap [CS19]. TwinkleTwinkle [CLY^{+}23]. TwistIn [LFH18]. Two [ATC17, CYK^{+}23b, DDP^{+}17, LDC18]. Two-Factor [CYK^{+}23b]. Two-Phase [LDC18]. Type [FS18]. Types [AKN^{+}19, SvBV^{+}18]. Typing [GO22, LHY^{+}22].

UAV [DJX^{+}23, PLC^{+}21]. UAV-Mounted [DJX^{+}23]. UbiEar [SZJ^{+}17]. UbiquiTouch [WXZ^{+}20]. Ubiquitous [AHS^{+}18, EEY18, EBTN22, FGS^{+}23, GMY^{+}17, HFS^{+}19, HYS^{+}23, JSC17, JGZ^{+}23, JHH21, LRH^{+}21, LHC^{+}22, LYD^{+}23, WXZ^{+}20, ZPZ^{+}20]. ULoc [ZCA^{+}21]. Ultra [AO19, CZG^{+}21, MZJ20, WRLK22]. Ultra-Low-Power [AO19]. Ultra-Wideband [MZJ20]. Ultrasound [DWL^{+}22]. UltraSpeech [DWL^{+}22]. Unaided [GMC18]. Unannotated [IRA^{+}21]. Uncertainty [AYM^{+}22, CML22, RKK18]. Uncertainty-aware [CML22]. Unclonable [SLX^{+}18]. Unconstrained [AAM^{+}23, BLH^{+}17]. Uncontrolled [JXZ^{+}20]. Undergoing [HFX^{+}21]. Underground [ISSK20]. Underlying [BWUW21]. Understand [SSR^{+}19]. Understanding [ADY^{+}20, AM19, BRA^{+}22, BLS^{+}18, BA20, CMV^{+}18, CLRH19, CCW^{+}20, CVL20, EWM^{+}20, GM19, KIT^{+}20, KGN^{+}20, KL19, LSE21, LKL22, LSY^{+}23, MT19, NJWFM19, OSG^{+}22, PL23, PRC^{+}18]. Underwater [RRS^{+}23]. Uneven [HXH^{+}18]. Unevenly [ZSG^{+}21]. Unevenly-Distributed [ZSG^{+}21]. Unfamiliar [CND22]. Unfolding [WYL^{+}21]. Uniqueness [CFLK18]. Unit [AO19]. Units [RLM^{+}19]. Unlabeled [GZW^{+}22]. Unlimited [NCM^{+}17]. Unlock [CYK^{+}23b]. Unmodified [CKJH23, IY20]. Unobtrusive [ASC^{+}22, BCK^{+}22, DGS18, GDS19, GWD^{+}17, HSH^{+}23, HSWB21, HW17, HW21, LEP21, RLM^{+}19]. Unobtrusively [AVD^{+}17, CGJC19]. Unravelling [dGST21]. Unsupervised [QWK^{+}19, XKNM20]. Uplink [GGJ^{+}22]. upon [KLK^{+}20]. Upper [MLH^{+}23, RLM^{+}19, XZZ21]. Urban [AvdBM^{+}22, CCX^{+}18, CFLK18, CDL^{+}19, CCW^{+}20, CLY^{+}21, CZT22, CMF^{+}23, DJN^{+}18, FZYZ18, GCW^{+}18, HYH21, JSF^{+}18, LDC18, WXL^{+}20, XL19, XFW^{+}20, XLX^{+}20]. Usability [DPW^{+}21, QCP^{+}18]. Usable [EUHB23, LYP^{+}22]. Usage [AVD^{+}17, BRA^{+}22, CWH^{+}19, HHFM17, WLZ^{+}19, YXW^{+}20]. Use [BSP21, BGR^{+}20, BLS^{+}18, CMV^{+}18, CCY^{+}17, GS20, GCC^{+}21, JHWF22, KJKL19, LYKH18, LWYZ21, TK17]. Used [LHS^{+}21]. User [AOGF^{+}21, AYM^{+}22, AM19, BWUW21, BBBN22, CKHM20, CLD^{+}20, CZP20, CNXG22, CND17, CVL20, DYCX21, FSW19, FDH21, GWJ^{+}21, GLJ^{+}21, GMC18, HWZZ19, HYS^{+}23, HQC^{+}23, HKK^{+}23, JCL^{+}20, JCF^{+}21, KKV^{+}22, KGN^{+}20, KCuH^{+}19, LCL20, LJZ^{+}22, LCL17, LYCW20, MGC^{+}23a, MOD21, NZL^{+}20, PM19, RQ20, TKF^{+}19, TAHH23, TFL^{+}19, VJH^{+}18, VLR20, WLZ^{+}19, XLX^{+}20, HY22, LLL21].

User-Avatar [LJZ^{+}22]. User-Centered [KKV^{+}22]. User-Centric [BBBN22]. User-Content [FSW19]. User-Driven [AYM^{+}22, GLJ^{+}21]. User-Generated [LCL17]. User-Powered [MGC^{+}23a]. Users [AKPS18, BA20, CHC^{+}22, GJL^{+}20, IB23, KKV^{+}22, KXAH18, NNMY23]. Using [AES^{+}21, ASG^{+}23, AIP^{+}18, AHB17, ASS^{+}19, ASS^{+}21, AWZK23, ASC^{+}22, AWSS22, ASTW19, ASM^{+}18, AYM^{+}22, AAM^{+}23, AZK^{+}23, BFS^{+}17, BAR^{+}18, BRS^{+}20, BLH^{+}17, BCK^{+}22, BGK21, BA20, BPA23, BSMP23, CDA23, CCJ^{+}17, CRMG22, CLW^{+}18, CZX^{+}22, CML^{+}20, CSHN17, CDG^{+}17, CZP20, CZG^{+}21, CLYZ22, CLZ^{+}21, CZL^{+}22, CKL23, CWS^{+}19, CND22, CJRG21, DCMM20, DGS18, DWZZ19, DMAH18, DGLQ19, DJL^{+}22, FMZN22, FZZ^{+}22, FMP^{+}23, FSW19, FCQR20, FZS^{+}23, GWP^{+}19, GLG^{+}19, GJC^{+}21b, GDS19, GH18, GL18, GYL^{+}20, GPW^{+}18, GNA^{+}19, GLS^{+}18, GZY^{+}21, GCP^{+}22, HMA21, HEUH18, HEP22, HY18, HSL^{+}20, HWJ^{+}21, HNL22, HXH^{+}18, HBH20, HWJ^{+}23, HWZZ19, IY20, IYS22, JGZ^{+}23, JNB^{+}20, JMvB^{+}19, JCL^{+}20, JHL^{+}21, JWS^{+}23, JMC^{+}21, JGZ^{+}21, JZG^{+}23, KRS22, KNL^{+}23, KGL^{+}18, KBKH20, LSSH23, LZZ^{+}20, LEP21, LYP^{+}22, LDW^{+}18, LCS^{+}21, LLLX22, LCX^{+}22, LHS^{+}23, LAAE^{+}19, LCR^{+}18, LXC^{+}19, LLWX21, LZG^{+}21, LYD^{+}23, MZW^{+}18, MLMH^{+}18]. Using [MBB20, MRCdCRV^{+}21, MMG^{+}19, NJWFM19, NTT^{+}21, NHL^{+}20, OHW^{+}20, PLC^{+}21, PCBK20, PB20, PCYZ18, PM19, PRS^{+}23, QWK^{+}19, RZ21, RTZ^{+}20, RWW^{+}22, SBP^{+}22, SHZ^{+}20, SSW^{+}17, SYYL22, SNGK20, SWLZ21, SH20, SCZ^{+}20, SXF^{+}22, SGO^{+}23, VJH^{+}18, WZG^{+}19, WMS^{+}21, XKNM20, YPJdB19, ZPPC^{+}20, ZXW^{+}17, ZWZ^{+}21b, ZWZ^{+}21a, GF17, GP17, HW17, HAY^{+}17, KNP17, NTKG17, WXY^{+}17]. Usual [AWSS22]. Utility [ATP^{+}19, CLW^{+}23]. Utilize [PM19]. UVLens [CLY^{+}21]. UWB [ASTW19, DJX^{+}23, ZCA^{+}21]. UWB-Tag [ZCA^{+}21].

V [MLMH^{+}18]. V-Speech [MLMH^{+}18]. V2iFi [ZCC^{+}20]. Valves [GH22]. Variability [WXW^{+}18]. Variable [DBY^{+}17]. Varying [OSG^{+}22]. Vehicle [BX19, CLZL18, CJRG21, JCR22, JHP^{+}17, KML^{+}21, LDK^{+}20, NVP18, SS22, ZCC^{+}20, KCP^{+}18, KPL20]. Vehicle-Pavement [SS22]. Vehicles [BBH17, CKLR21, CRGR22, CBR23, KYJ^{+}23, MBZ^{+}18]. Vehicular [CLH18]. Velcro [FZZ^{+}22]. Velocity [GWD^{+}17]. Ventilation [CZL^{+}22]. verbal [KCP^{+}18]. Verification [DWZZ19, ILH^{+}19, LWYZ21]. VeriMask [LCR^{+}21]. Versatile [CCY^{+}23, CND^{+}20]. Vertical [CLZ^{+}21, SSB^{+}21]. Vetting [WCW21]. via [BXY^{+}21, CS22b, CCY^{+}23, CYK^{+}23a, DZH^{+}22, DYL^{+}19, DMAH18, EBPA22, FFT^{+}18, FYY^{+}21, FRS^{+}20, GBLM17, GLC^{+}21, HFX^{+}21, HDFZ22, HQC^{+}23, JHH21, KHG^{+}19, LLPH22, LFH18, LLZ^{+}20, LCW^{+}21, LLCY21, LYLT21, LZZ^{+}23, LCG^{+}22, LLC^{+}20, LWZ^{+}20, LKZ^{+}21, LCH^{+}22, LHFZ22, LSY^{+}18, LWYZ21, MJ19, MSWQ20, MWJ^{+}21, QCWY19, SLX^{+}18, WYL^{+}21, YXW^{+}20, ZCC^{+}20, ZSG^{+}21]. VibPath [CYK^{+}23b]. Vibration [CYK^{+}23b, FMZN22, FMP^{+}23, KCFI19, MLMH^{+}18]. Vibrations [JJB22, SCZ^{+}20]. Vibroacoustic [GCLL19]. Vibroacoustic-based [GCLL19]. VibroMap [EWM^{+}20]. Vibrometry [SCZ^{+}20]. VibroSense [SCZ^{+}20]. Vibrotactile [EWM^{+}20, XYG^{+}20]. Video [DWCY23, FSW19, KTEM21, KTH^{+}20, LLZ^{+}22, LMS^{+}23, RZ21]. Videos [DZH^{+}23, LT19, RHI^{+}23, SPRK23]. view [BYW^{+}20, HY18, KLK^{+}20]. Viewport [FSW19]. Village [CLY^{+}21]. Virtual [AOGF^{+}21, CXR^{+}18, CNS21, CJRG21, GCP^{+}22, JCF^{+}21, KTH^{+}20, KWAP21, LHS^{+}23, MQZH23, RPF^{+}18, XYG^{+}20]. ViscoCam [AZK21]. Viscoelastic [LLPH22]. Viscosity [AZK21]. Visible [HWJ^{+}23]. ViSig [CDA23]. Vision [CS22b, CLL^{+}22, DH17, IB23, LLPH22, RHI^{+}23, WXY^{+}19, ZXY21, CHC^{+}22]. Vision-assisted [WXY^{+}19]. Vision-based [LLPH22, ZXY21]. Vision-RF [CLL^{+}22]. Vision2Sensor [RH19]. Visit [HFX^{+}21]. Visits [CCK^{+}18]. Visual [ATP^{+}19, BDM^{+}18, CKLB22, CDA23, DSSD23, JXZ^{+}20, KLT19, LHS^{+}21, MGC^{+}23b, PSC^{+}17, SSB^{+}21, XYG^{+}20]. Visualization [CKLR21]. Visualizations [CRGR22, CMF^{+}23]. Visualizing [RKK18]. Visually [AKPS18, BMZN20, LGY^{+}20]. Vital [GZL^{+}21, ZCC^{+}20]. Vocabulary [HHIF19]. VocaBura [HHIF19]. Vocal [LCR^{+}18, LYCW20]. VocalLock [LYCW20]. Voice [CXZ^{+}21, CYK^{+}23a, CLW^{+}23, GJC^{+}21a, HWJ^{+}22, KPL20, LCR^{+}18, TKF^{+}19, WCW21, WAL^{+}19, ZLD^{+}21]. Voice-Based [HWJ^{+}22, CXZ^{+}21]. Voice-controlled [WCW21]. Voice-Driven [WAL^{+}19]. VoiceCloak [CLW^{+}23]. VoltKey [LKBK19]. VPS [KCFI19]. VR [AIP^{+}18, MTH19]. VREcall [GCP^{+}22]. vs [CCX^{+}18, CFZ^{+}19, KVZ^{+}19]. Vulnerable [CWP^{+}18].

W [MZT18, RLM^{+}19]. W-Air [MZT18]. Waiting [ALC^{+}20]. Walk [LKL19]. Walking [CPW^{+}21, HHIF19, KIT^{+}20]. wall [LLZ^{+}20, ZCN^{+}20]. Want [JJB22]. Warm [FDH21]. was [MT19]. Watching [CGJ^{+}22, GLW^{+}18, LBC^{+}22]. Water [CZT22]. Wave [ARSN22, CS22b, DCZ^{+}23, LWZ^{+}20, PSLS21, RSSN21, LZZ^{+}21, LZZ^{+}23, SHZ^{+}20]. Waveform [CLC^{+}22]. Waves [HCH^{+}22]. Way [JSC17]. Ways [CLTS21]. Weak [GZW^{+}22]. Weak-supervised [GZW^{+}22]. Weakly [SEG^{+}18, SH20]. Wearable [AES^{+}21, ASV^{+}18, ATP^{+}19, ASC^{+}22, AYM^{+}22, ATC17, BRS^{+}20, BDPH21, BSP21, BLH^{+}17, BDM^{+}18, BMZN20, CCJ^{+}17, CZP20, CHWZ23, CBT18, DKZ^{+}22, DH17, DHC^{+}18, DWHR17, GWJ^{+}21, GWP^{+}19, GLG^{+}19, GDS19, GH18, GWD^{+}17, GCC^{+}21, HMM^{+}20, HW17, KBXP23, KGL19, LCFT19, LCR^{+}18, LLL21, MWJ^{+}21, MTF^{+}18, PCYZ18, QWK^{+}19, RGL^{+}18, RDOA22, SWLG22, SGO^{+}23, WS17, YPJdB19]. Wearable-based [GCC^{+}21, SWLG22]. Wearables [CLT^{+}21, GWJ^{+}21, GP17, GZY^{+}21, Hal19, HEP22, JZG^{+}23, LS20, MZT18, PRC^{+}21, SH20]. WearBreathing [LAAE^{+}19]. Wearing [JKK^{+}17]. Weather [MLX^{+}20]. Weather-Related [MLX^{+}20]. Webcam [DPW^{+}21]. Websites [CCX^{+}18]. week [CFZ^{+}19]. Weekly [LKL19]. Weight [KNL^{+}23, LLPH22, CLH18]. Well [CSK^{+}18, KML^{+}21]. Well-being [CSK^{+}18, KML^{+}21]. WePos [GZW^{+}22]. Wheels [LCG^{+}22]. Which [Tor23]. While [HHIF19, BDM^{+}18]. Whiteboard [CCK^{+}18]. Who [BWUW21, DWCY23, GLJ^{+}21, HFX^{+}21]. WhoIsZuki [MGC^{+}23b]. Wi [CML22, FWJ^{+}22, GZZ^{+}21, HDC18, LLZ^{+}20, LCS^{+}21, LJK^{+}21]. Wi-Fi [CML22, FWJ^{+}22, GZZ^{+}21, HDC18, LLZ^{+}20, LCS^{+}21]. Wi-Fruit [LJK^{+}21]. Wi-Learner [FWJ^{+}22]. WiBorder [LLZ^{+}20]. Wide [DGZ^{+}21, LH19]. Wide-Area [LH19]. Wideband [CZG^{+}21, MZJ20]. Widgets [CBEG17, MGC^{+}23a]. WiFi [FFQ^{+}20, GLX^{+}22, GL18, GLS^{+}18, GZW^{+}22, HY22, LCW^{+}21, LZG^{+}21, LLG^{+}23, MZW^{+}18, RTZ^{+}20, RWW^{+}22, SSW^{+}17, VS19, VSY^{+}20, WXY^{+}17, ZWX^{+}20]. WiFi-based [LCW^{+}21, LLG^{+}23, WXY^{+}17]. Wild [ASV^{+}18, BOH^{+}21, BLW^{+}23, CFZ^{+}19, DDKL19, GSRS20, GHvB^{+}17, GZY^{+}21, PTK18, ASS^{+}19, KVZ^{+}19, KME^{+}19, SPRK23]. Will [CCW^{+}20, HFX^{+}21, WMW^{+}20]. Willingness [AKPS18]. Wind [PSC^{+}23]. WiNE [BS22]. WiPhone [LZG^{+}21]. Wireless [BW20, FDL^{+}18, GPB^{+}23, HMA21, JSNL23, KEI^{+}18, LYP^{+}22, RGL^{+}18, WPJ^{+}19, XYX21, YYW^{+}20, ZLD^{+}21]. Wisdom [ZCL21]. wise [GRS^{+}22]. WiStress [HMA21]. Without [WXY^{+}17, CKKM20]. Work [DGH^{+}20, GCGD17, JYY21, ZAP17]. Worker [IMO^{+}18]. Workers [DGH^{+}20, GCGD17]. Workforce [AvdBF^{+}19]. Workplace [HKSM23, MMG^{+}19]. World [GJG^{+}18, KWAP21, LAAE^{+}19, MVS^{+}20]. Worn [ASS^{+}21, BDM^{+}18, CLC^{+}22, LHS^{+}21, LHY^{+}20, ASS^{+}19, KSC^{+}21, XKNM20]. WR [LLL21]. WR-Hand [LLL21]. Wrist [ASS^{+}19, ASS^{+}21, BAT22, CLC^{+}22, HHX^{+}20, HSK^{+}22, LHY^{+}20]. Wrist-Worn [ASS^{+}21, CLC^{+}22, LHY^{+}20, ASS^{+}19]. WristAcoustic [HSK^{+}22]. Write [ZWZ^{+}21a]. Writing [CLZ^{+}21]. Wuhan [HFX^{+}21].

X [JSGS23]. X-CHAR [JSGS23]. Xnavi [NTO^{+}18].

Years [GWJ^{+}21]. Young

[BFS^{+}17, KL19, MLPGP21, WTA^{+}20]. Yourself [KPG^{+}20]. YouTube [LBC^{+}22].

Zero [CXC^{+}19, FMA^{+}19, HAY^{+}17, LKBK19]. Zero-Effort [HAY^{+}17]. Zero-Interaction [FMA^{+}19]. Zero-Involvement [LKBK19]. Zone [CXC^{+}19, LLG^{+}23]. Zygarde [IN20].

References

Arakawa:2023:LUU

[AAM^{+}23] Riku Arakawa, Karan Ahuja, Kristie Mak, Gwendolyn Thompson, Sam Shaaban, Oliver Lindhiem, and Mayank Goel. LemurDx: Using unconstrained passive sensing for an objective measurement of hyperactivity in children with no parent input. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):46:1– 46:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596244.

Alhirabi:2023:PIP

[ABL^{+}23] Nada Alhirabi, Stephanie Beaumont, Jose Tomas Llanos, Dulani Meedeniya, Omer Rana, and Charith Perera. PARROT: Interactive privacy-aware Internet of things application design tool. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):1:1– 1:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580880.

Arrotta:2022:DDE

[ACB22] Luca Arrotta, Gabriele Civitarese, and Claudio Bettini. DeXAR: Deep explainable sensor-based activity recognition in smarthome environments. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1): 1:1–1:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517224.

Abtahi:2020:UPP

[ADY^{+}20] Parastoo Abtahi, Victoria Ding, Anna C. Yang, Tommy Bruzzese, Alyssa B. Romanos, Elizabeth L. Murnane, Sean Follmer, and James A. Landay. Understanding physical practices and the role of technology in manual self-tracking. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):115:1– 115:24, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3432236.

Abedin:2021:ADB

[AES^{+}21] Alireza Abedin, Mahsa Ehsanpour, Qinfeng Shi, Hamid Rezatofighi, and Damith C. Ranasinghe. Attend and discriminate: Beyond the state-of-the-art for human activity recognition using wearable sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1): 1:1–1:22, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3448083.

Aggarwal:2022:TAR

[AGU^{+}22] Aditya Aggarwal, Siddhartha Gairola, Uddeshya Upadhyay, Akshay P. Vasishta, Diwakar Rao, Aditya Goyal, Kaushik Murali, Nipun Kwatra, and Mohit Jain. Towards automating retinoscopy for refractive error diagnosis. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):97:1– 97:??, September 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3550283.

Akazue:2017:UTS

[AHB17] Moses Akazue, Martin Halvey, and Lynne Baillie. Using thermal stimuli to enhance photo-sharing in social media. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090050.

Anderson:2018:SAM

[AHS^{+}18] Christoph Anderson, Isabel Hübener, Ann-Kathrin Seipp, Sandra Ohly, Klaus David, and Veljko Pejovic. A survey of attention management systems in ubiquitous computing environments. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–27, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3214261.

Ahuja:2018:EIE

[AIP^{+}18] Karan Ahuja, Rahul Islam, Varun Parashar, Kuntal Dey, Chris Harrison, and Mayank Goel. EyeSpyVR: Interactive eye sensing using off-the-shelf, smartphonebased VR headsets. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–10, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3214260.

Alshurafa:2018:MAB

[AJA^{+}18] Nabil Alshurafa, Jayalakshmi Jain, Rawan Alharbi, Gleb Iakovlev, Bonnie Spring, and Angela Pfammatter. Is more always better?: Discovering incentivized mHealth intervention engagement related to health behavior trends. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 26, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287031.

Alqahtani:2020:EGM

[AJV20] Deemah Alqahtani, Caroline Jay, and Markel Vigo. The effect of goal moderation on the achievement and satisfaction of physical activity goals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):116:1– 116:18, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3432209.

Antar:2023:BMA

[AKB23] Anindya Das Antar, Anna Kratz, and Nikola Banovic. Behavior modeling approach for forecasting physical functioning of people with multiple sclerosis. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):7:1– 7:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3580887.

Abdelrahman:2019:CAT

[AKN^{+}19] Yomna Abdelrahman, Anam Ahmad Khan, Joshua Newn, Eduardo Velloso, Sherine Ashraf Safwat, James Bailey, Andreas Bulling, Frank Vetere, and Albrecht Schmidt. Classifying attention types with thermal imaging and eye tracking. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 27, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3351227.

Ahmed:2018:LPC

[AKPS18] Tousif Ahmed, Apu Kapadia, Venkatesh Potluri, and Manohar Swaminathan. Up to a limit?: Privacy concerns of bystanders and their willingness to share additional information with visually impaired users of assistive technologies. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1– 27, September 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3264899.

Abowd:2017:E

[AKS^{+}17] Gregory D. Abowd, Vassilis Kostakos, Silvia Santini, James Scott, and Koji Yatani. Editorial. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(1):1, March 2017. CODEN ???? ISSN 2474-9567. URL http:

//dl.acm.org/citation.
cfm?id=3075960.

Ahuja:2019:EPC

[AKX^{+}19] Karan Ahuja, Dohyun Kim, Franceska Xhakaj, Virag Varga, Anne Xie, Stanley Zhang, Jay Eric Townsend, Chris Harrison, Amy Ogan, and Yuvraj Agarwal. EduSense: Practical classroom sensing at scale. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1–26, September 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351229.

Hossain:2020:FCS

[ALC^{+}20] Forsad Al Hossain, Andrew A. Lover, George A. Corey, Nicholas G. Reich, and Tauhidur Rahman. FluSense: a contactless syndromic surveillance platform for influenza-like illness in hospital waiting areas. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1): 1:1–1:28, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3381014.

Axtell:2019:BRP

[AM19] Benett Axtell and Cosmin Munteanu. Back to real pictures: a crossgenerational understanding of users’ mental models of photo cloud storage. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 24, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351232.

Arora:2019:DED

[AMG^{+}19] Jatin Arora, Kartik Mathur, Manvi Goel, Piyush Kumar, Abhijeet Mishra, and Aman Parnami. Design and evaluation of DIO construction toolkit for co-making shared constructions. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 25, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369833.

Amoh:2019:ORU

[AO19] Justice Amoh and Kofi M.

Odame. An optimized recurrent unit for ultra-low-power keyword spotting. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–17, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328907.

Ahuja:2021:CUM

[AOGF^{+}21] Karan Ahuja, Eyal Ofek, Mar Gonzalez-Franco, Christian Holz, and Andrew D. Wilson. CoolMoves: User motion accentuation in virtual reality. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):52:1– 52:23, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463499.

Ananthabhotla:2018:SRS

[AP18] Ishwarya Ananthabhotla and Joseph A. Paradiso. SoundSignaling: Realtime, stylistic modification of a personal music corpus for information delivery. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 23, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3287032.

Abdelaal:2018:CGD

[ARD^{+}18] Mohamed Abdelaal, Daniel Reichelt, Frank Dürr, Kurt Rothermel, Lavinia Runceanu, Susanne Becker, and Dieter Fritsch. ComNSense: Grammar-driven crowdsourcing of point clouds for automatic indoor mapping. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–26, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3191733.

Adhikari:2022:MAH

[ARSN22] Aakriti Adhikari, Hem Regmi, Sanjib Sur, and Srihari Nelakuditi. MiShape: Accurate human silhouettes and body joints from commodity millimeterwave devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):96:1– 96:??, September 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3550300.

Alharbi:2022:SUE

[ASC^{+}22] Rawan Alharbi, Soroush Shahi, Stefany Cruz, Lingfeng Li, Sougata Sen, Mahdi Pedram, Christopher Romano, Josiah Hester, Aggelos K. Katsaggelos, and Nabil Alshurafa. Smoke-Mon: Unobtrusive extraction of smoking topography using wearable energyefficient thermal. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):155:1– 155:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3569460.

Ahmad:2023:ATP

[ASG^{+}23] Fawad Ahmad, Christina Suyong Shin, Rajrup Ghosh, John D’Ambrosio, Eugene Chai, Karthikeyan Sundaresan, and Ramesh Govindan. AeroTraj: Trajectory planning for fast, and accurate 3D reconstruction using a drone-based LiDAR. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):83:1– 83:??, September 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3610911.

Apthorpe:2018:DSH

[ASM^{+}18] Noah Apthorpe, Yan Shvartzshnaider, Arunesh Mathur, Dillon Reisman, and Nick Feamster. Discovering smart home Internet of Things privacy norms using contextual integrity. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–23, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3214262.

Akther:2019:MMM

[ASS^{+}19] Sayma Akther, Nazir Saleheen, Shahin Alan Samiei, Vivek Shetty, Emre Ertin, and Santosh Kumar. mORAL: an mHealth model for inferring oral hygiene behaviors in-the-wild using wristworn inertial sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–25, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3314388.

Akther:2021:MIB

[ASS^{+}21] Sayma Akther, Nazir Saleheen, Mithun Saha, Vivek Shetty, and Santosh Kumar. mTeeth: Identifying brushing teeth surfaces using wrist-worn inertial sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):

53:1–53:25, June 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463494.

Anderson:2019:MGA

[ASTW19] Boyd Anderson, Mingqian Shi, Vincent Y. F. Tan, and Ye Wang. Mobile gait analysis using foot-mounted UWB sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1–22, September 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351231.

Alharbi:2018:CME

[ASV^{+}18] Rawan Alharbi, Tammy Stump, Nilofar Vafaie, Angela Pfammatter, Bonnie Spring, and Nabil Alshurafa. I can’t be myself: Effects of wearable cameras on the capture of authentic behavior in the wild. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1– 40, September 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264900.

Adaimi:2019:LAL

[AT19] Rebecca Adaimi and Edison Thomaz. Leveraging active learning and conditional mutual information to minimize data annotation in human activity recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 23, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351228.

Azimpourkivi:2017:CBT

[ATC17]  Mozhgan  Azimpourkivi,

Umut Topkara, and Bogdan Carbunar. Camera based two factor authentication through mobile and wearable devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1–37, September 2017. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3131904.

Alharbi:2019:MMB

[ATP^{+}19] Rawan Alharbi, Mariam Tolba, Lucia C. Petito, Josiah Hester, and Nabil Alshurafa. To mask or not to mask?: Balancing privacy with visual confirmation utility in activityoriented wearable cameras. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 29, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351230.

Adler:2021:IMS

[ATQ^{+}21] Daniel A. Adler, Vincent W.-S. Tseng, Gengmo Qi, Joseph Scarpa, Srijan Sen, and Tanzeem Choudhury. Identifying mobile sensing indicators of stress-resilience. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):51:1– 51:32, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463528.

Abdelrahman:2017:CHE

[AVD^{+}17] Yomna Abdelrahman, Eduardo Velloso, Tilman Dingler, Albrecht Schmidt, and Frank Vetere. Cognitive heat: Exploring the usage of thermal imaging to unobtrusively estimate cognitive load. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1–20, September 2017. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3130898.

Acer:2019:SCM

[AvdBF^{+}19] Utku Günay Acer, Marc van den Broeck, Claudio Forlivesi, Florian Heller, and Fahim Kawsar. Scaling crowdsourcing with mobile workforce: a case study with Belgian postal service. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–32, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3328906.

Acer:2022:CPA

[AvdBM^{+}22] Utku Günay Acer, Marc van den Broeck, Chulhong Min, Mallesham Dasari, and Fahim Kawsar. The city as a personal assistant: Turning urban landmarks into conversational agents for serving hyper local information. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2): 40:1–40:??, July 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3534573.

Amesaka:2022:GRM

[AWSS22] Takashi Amesaka, Hiroki Watanabe, Masanori Sugimoto, and Buntarou Shizuki. Gesture recognition method using acoustic sensing on usual garment. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):41:1– 41:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534579.

AlArabi:2023:HEE

[AWZK23] Abul Al Arabi, Xue Wang, Yang Zhang, and Jeeeun Kim. E3D: Harvesting energy from everyday kinetic interactions using 3D printed attachment mechanisms. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):84:1– 84:??, September 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3610897.

Akhunov:2022:AAM

[AY22] Khakim Akhunov and Kasim Sinan Yildirim. AdaMICA: Adaptive multicore intermittent computing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):98:1– 98:??, September 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3550304.

Arakawa:2022:PTF

[AYM^{+}22] Riku Arakawa, Hiromu Yakura, Vimal Mollyn, Suzanne Nie, Emma Russell, Dustin P. DeMeo, Haarika A. Reddy, Alexander K. Maytin, Bryan T. Carroll, Jill Fain Lehman, and Mayank Goel. PrISM-Tracker: a framework for multimodal procedure tracking using wearable sensors and state transition information with userdriven handling of errors and uncertainty. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):156:1– 156:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3569504.

Adaimi:2021:OGW

[AYT21] Rebecca Adaimi, Howard Yong, and Edison Thomaz. Ok Google, what am I doing?: Acoustic activity recognition bounded by conversational assistant interactions. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):2:1– 2:24, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3448090.

Abtahi:2017:DNM

[AZEL17] Parastoo Abtahi, David Y.

Zhao, Jane L. E., and James A. Landay. Drone near me: Exploring touchbased human-drone interaction. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1–8, September 2017. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3130899.

An:2021:VSB

[AZK21] Kecheng An, Qian Zhang, and Elaine Kwong. ViscoCam: Smartphone-based drink viscosity control assistant for dysphagia patients. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1): 3:1–3:25, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3448109.

Arakawa:2023:MPH

[AZK^{+}23] Riku Arakawa, Bing Zhou, Gurunandan Krishnan, Mayank Goel, and Shree K. Nayar. MI-Poser: Human body pose tracking using magnetic and inertial sensor fusion with metal interference mitigation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):85:1– 85:??, September 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3610891.

Arora:2018:STF

[AZS^{+}18] Nivedita Arora, Steven L. Zhang, Fereshteh Shahmiri, Diego Osorio, Yi-Cheng Wang, Mohit Gupta, Zhengjun Wang, Thad Starner, Zhong Lin Wang, and Gregory D. Abowd. SATURN: a thin and flexible self-powered microphone leveraging triboelectric nanogenerator. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–28, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214263.

Blair:2020:IDS

[BA20] Johnna Blair and Saeed Abdullah. It didn’t sound good with my cochlear implants: Understanding the challenges of using smart assistants for deaf and hard of hearing users. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):118:1– 118:27, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3432194.

Bari:2018:RMM

[BAR^{+}18] Rummana Bari, Roy J.

Adams, Md. Mahbubur Rahman, Megan Battles Parsons, Eugene H. Buder, and Santosh Kumar. rConverse: Moment by moment conversation detection using a mobile respiration sensor. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–27, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191734.

Bhattacharya:2022:LSW

[BAT22] Sarnab Bhattacharya, Rebecca Adaimi, and Edison Thomaz. Leveraging sound and wrist motion to detect activities of daily living with commodity smartwatches. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2): 42:1–42:??, July 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534582.

Belal:2022:PEU

[BBBN22] Yacine Belal, Aurélien Bellet, Sonia Ben Mokhtar, and Vlad Nitu. PEPPER: Empowering user-centric recommender systems over gossip learning. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):101:1– 101:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550302.

Beattie:2017:EHD

[BBH17] David Beattie, Lynne Baillie, and Martin Halvey. Exploring how drivers perceive spatial earcons in automated vehicles. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1–24, September 2017. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3130901.

Bethge:2022:TDS

[BCK^{+}22] David Bethge, Luis Falconeri Coelho, Thomas Kosch, Satiyabooshan Murugaboopathy, Ulrich von Zadow, Albrecht Schmidt, and Tobias Grosse-Puppendahl. Technical design space analysis for unobtrusive driver emotion assessment using multi-domain context. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):159:1– 159:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3569466.

Boovaraghavan:2023:MDD

[BCM^{+}23] Sudershan Boovaraghavan, Chen Chen, Anurag Maravi, Mike Czapik, Yang Zhang, Chris Harrison, and Yuvraj Agarwal. Mites: Design and deployment of a general-purpose sensing infrastructure for buildings. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1): 2:1–2:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580865.

Boldu:2018:FDE

[BDM^{+}18] Roger Boldu, Alexandru Dancu, Denys J. C. Matthies, Thisum Buddhika, Shamane Siriwardhana, and Suranga Nanayakkara. Finger-Reader2.0: Designing and evaluating a wearable fingerworn camera to assist people with visual impairments while shopping. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1–19, September 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3264904.

Barman:2021:EBM

[BDPH21] Ludovic Barman, Alexandre Dumur, Apostolos Pyrgelis, and Jean-Pierre Hubaux. Every byte matters: Traffic analysis of Bluetooth wearable devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):54:1– 54:45, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3463512.

Butscher:2017:ITO

[BDR17] Simon Butscher, Maximilian Dürr, and Harald Reiterer. InformationSense: Trade-offs for the design and the implementation of a large highly deformable cloth display. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

                                        //dl.acm.org/citation.
                                        cfm?id=3090053.

Bae:2017:DDE

[BFS^{+}17] Sangwon Bae, Denzil Ferreira, Brian Suffoletto, Juan C. Puyana, Ryan Kurtz, Tammy Chung, and Anind K. Dey. Detecting drinking episodes in young adults using Smartphonebased sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090051.

Bhalla:2021:IOC

[BGK21] Sejal Bhalla, Mayank Goel, and Rushil Khurana. IMU2Doppler: Cross-modal domain adaptation for Doppler-based activity recognition using IMU data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):145:1– 145:20, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3494994.

Ballagas:2018:DSP

[BGL18] Rafael Ballagas, Sarthak Ghosh, and James Landay. The design space of 3D printable interactivity. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–21, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214264.

Beneteau:2020:ACH

[BGR^{+}20] Erin Beneteau, Yini Guan, Olivia K. Richards, Mingrui Ray Zhang, Julie A. Kientz, Jason Yip, and Alexis Hiniker. Assumptions checked: How families learn about and use the Echo Dot. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1): 3:1–3:23, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3380993.

Balaji:2022:RSC

[BKL^{+}22] Ananta Narayanan Balaji, Clayton Kimber, David Li, Shengzhi Wu, Ruofei Du, and David Kim. Retro-Sphere: Self-contained passive 3D controller tracking for augmented reality. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):157:1– 157:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3569479.

Buda:2021:OSS

[BKM21] Teodora Sandra Buda, Mohammed Khwaja, and Aleksandar Matic. Outliers in smartphone sensor data reveal outliers in daily happiness. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):5:1– 5:19, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3448095.

Bedri:2017:EUW

[BLH^{+}17] Abdelkareem Bedri, Richard Li, Malcolm Haynes, Raj Prateek Kosaraju, Ishaan Grover, Temiloluwa Prioleau, Min Yan Beh, Mayank Goel, Thad Starner, and Gregory Abowd. EarBit: Using wearable sensors to detect eating episodes in unconstrained environments. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1–20, September 2017. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3130902.

Bentley:2018:ULT

[BLS^{+}18] Frank Bentley, Chris Luvogt, Max Silverman, Rushani Wirasinghe, Brooke White, and Danielle Lottridge. Understanding the longterm use of smart speaker assistants. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1–24, September 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3264901.

Bhalla:2023:PCA

[BLW^{+}23] Sejal Bhalla, Salaar Liaqat, Robert Wu, Andrea S. Gershon, Eyal de Lara, and Alex Mariakakis. PulmoListener: Continuous acoustic monitoring of chronic obstructive pulmonary disease in the wild. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):86:1– 86:??, September 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3610889.

Blalock:2018:STS

[BMG18] Davis Blalock, Samuel Madden, and John Guttag. Sprintz: Time series compression for the Internet of Things. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1–23, September 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3264903.

Brewer:2017:HRW

[BML17] R. N. Brewer, M. R. Morris, and S. E. Lindley. How to remember what to remember: Exploring possibilities for digital reminder systems.

Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1– 20, September 2017. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3130903.

Bhattacharya:2020:CAA

[BMR^{+}20] Sourav Bhattacharya, Dionysis Manousakas, Alberto Gil C. P. Ramos, Stylianos I. Venieris, Nicholas D. Lane, and Cecilia Mascolo. Countering acoustic adversarial attacks in microphoneequipped smart home devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):73:1–73:24, June 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3397332.

Boldu:2020:AAW

[BMZN20] Roger Boldu, Denys J. C.

Matthies, Haimo Zhang, and Suranga Nanayakkara. AiSee: an assistive wearable device to support visually impaired grocery shoppers. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4): 119:1–119:25, December 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3432196.

Bentvelzen:2022:DRD

[BNW22] Marit Bentvelzen, Jasmin Niess, and Paweł W. Woźniak. Designing reflective derived metrics for fitness trackers. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):158:1– 158:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3569475.

Barrios:2021:SBT

[BOH^{+}21] Liliana Barrios, Pietro Oldrati, Marc Hilty, David Lindlbauer, Christian Holz, and Andreas Lutterotti. Smartphone-based tapping frequency as a surrogate for perceived fatigue: an in-the-wild feasibility study in multiple sclerosis patients. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):89:1– 89:30, September 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3478098.

Boovaraghavan:2023:TCD

[BPA23] Sudershan Boovaraghavan, Prasoon Patidar, and Yuvraj Agarwal. TAO: Context detection from daily activity patterns using temporal analysis and ontology. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3): 87:1–87:??, September 2023. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3610896.

Bardot:2022:ORR

[BRA^{+}22] Sandra Bardot, Bradley Rey, Lucas Audette, Kevin Fan, Da-Yuan Huang, Jun Li, Wei Li, and Pourang Irani. One ring to rule them all: an empirical understanding of day-to-day smartring usage through insitu diary study. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):100:1– 100:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550315.

Bari:2020:ADS

[BRS^{+}20] Rummana Bari, Md. Mahbubur Rahman, Nazir Saleheen, Megan Battles Parsons, Eugene H. Buder, and Santosh Kumar. Automated detection of stressful conversations using wearable physiological and inertial sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):117:1– 117:23, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3432210.

Bedogni:2021:MMI

[BRS21] Luca Bedogni, Shakila Khan Rumi, and Flora D. Salim. Modelling memory for individual re-identification in decentralised mobile contact tracing applications. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1): 4:1–4:21, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3448088.

Bakar:2021:RFD

[BRYH21] Abu Bakar, Alexander G. Ross, Kasim Sinan Yildirim, and Josiah Hester. RE-HASH: a flexible, developer focused, heuristic adaptation platform for intermittently powered computing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3): 87:1–87:42, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478077.

Baumann:2017:EBC

[BS17] Paul Baumann and Silvia Santini. Every byte counts: Selective prefetching for mobile applications. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090052.

Banerjee:2021:RMM

[BS21] Avishek Banerjee and Kannan Srinivasan. RFTemp: Monitoring microwave oven leakage to estimate food temperature. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):144:1– 144:25, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3494967.

Banerjee:2022:WMM

[BS22] Avishek Banerjee and Kannan Srinivasan. WiNE: Monitoring microwave oven leakage to estimate food nutrients and calorie. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):99:1– 99:??, September 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3550313.

Budde:2017:PSP

[BSH^{+}17] Matthias Budde, Andrea Schankin, Julien Hoffmann, Marcel Danz, Till Riedel, and Michael Beigl. Participatory sensing or participatory nonsense?: Mitigating the effect of human error on data quality in citizen science. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1– 23, September 2017. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3131900.

Breda:2023:FAC

[BSMP23] Joseph Breda, Mastafa Springston, Alex Mariakakis, and Shwetak Patel. FeverPhone: Accessible core-body temperature sensing for fever monitoring using commodity smartphones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):3:1– 3:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3580850.

Bartolome:2021:GCI

[BSP21] Abigail Bartolome, Sahaj Shah, and Temiloluwa Prioleau. GlucoMine: a case for improving the use of wearable device data in diabetes management. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):90:1– 90:24, September 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3478109.

Bai:2020:CWC

[BW20] Yunhao Bai and Xiaorui Wang. CARIN: Wireless CSI-based driver activity recognition under the interference of passengers. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1): 2:1–2:28, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3380992.

Bentvelzen:2022:RRH

[BWH^{+}22] Marit Bentvelzen, Paweł W.

Wo’zniak, Pia S. F. Herbes, Evropi Stefanidi, and Jasmin Niess. Revisiting reflection in HCI: Four design resources for technologies that support reflection. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1): 2:1–2:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3517233.

Bi:2018:ADE

[BWT^{+}18] Shengjie Bi, Tao Wang, Nicole Tobias, Josephine Nordrum, Shang Wang, George Halvorsen, Sougata Sen, Ronald Peterson, Kofi Odame, Kelly Caine, Ryan Halter, Jacob Sorber, and David Kotz. Auracle: Detecting eating episodes with an ear-mounted sensor. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1– 27, September 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3264902.

Barbosa:2021:WDP

[BWUW21] Natã M. Barbosa, Gang Wang, Blase Ur, and Yang Wang. Who am I?: a design probe exploring real-time transparency about online and offline user profiling underlying targeted Ads. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):88:1– 88:32, September 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3478122.

Bi:2019:RRT

[BX19] Chongguang Bi and Guoliang Xing. RAMT: Realtime attitude and motion tracking for mobile devices in moving vehicle. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–21, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328909.

Bu:2021:HAR

[BXY^{+}21] Yanling Bu, Lei Xie, Yafeng Yin, Chuyu Wang, Jingyi Ning, Jiannong Cao, and Sanglu Lu. Handwritingassistant: Reconstructing continuous strokes with millimeter-level accuracy via attachable inertial sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):146:1–146:25, December 2021. CODEN ???? ISSN 2474-9567 (electronic).

URL https://dl.acm.org/
doi/10.1145/3494956.

Bai:2020:AMV

[BYW^{+}20] Lei Bai, Lina Yao, Xianzhi Wang, Salil S. Kanhere, Bin Guo, and Zhiwen Yu. Adversarial multiview networks for activity recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):42:1– 42:22, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397323.

Clarke:2017:RCB

[CBEG17] Christopher Clarke, Alessio Bellino, Augusto Esteves, and Hans Gellersen. Remote control by body movement in synchrony with orbiting widgets: an evaluation of TraceMatch. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1– 22, September 2017. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3130910.

Colley:2023:SEC

[CBR23] Mark Colley, Julian Britten, and Enrico Rukzio. Scalability in external communication of automated vehicles: Evaluation and recommendations. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):51:1– 51:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596248.

Chun:2018:DEE

[CBT18] Keum San Chun, Sarnab Bhattacharya, and Edison Thomaz. Detecting eating episodes by tracking jawbone movements with a noncontact wearable sensor. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–21, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191736.

Chan:2019:PSE

[CBZN19] Samantha W. T. Chan, Thisum Buddhika, Haimo Zhang, and Suranga Nanayakkara. ProspecFit: In situ evaluation of digital prospective memory training for older adults. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1–20, September 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351235.

Carek:2017:SWC

[CCJ^{+}17] Andrew M. Carek, Jordan Conant, Anirudh Joshi, Hyolim Kang, and Omer T. Inan. SeismoWatch: Wearable cuffless blood pressure monitoring using pulse transit time. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1–16, September 2017. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3130905.

Clayphan:2018:CST

[CCK^{+}18] Andrew Clayphan, Anthony Collins, Judy Kay, Nathan Slawitschka, and Jenny Horder. Comparing a single-touch whiteboard and a multi-touch tabletop for collaboration in school museum visits. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–23, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191738.

Chen:2020:WYC

[CCW^{+}20] Zhilong Chen, Hancheng Cao, Huangdong Wang, Fengli Xu, Vassilis Kostakos, and Yong Li. Will you come back /check-in again?: Understanding characteristics leading to urban revisitation and re-check-in. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):76:1– 76:27, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3411812.

Cao:2018:RUS

[CCX^{+}18] Hancheng Cao, Zhilong Chen, Fengli Xu, Yong Li, and Vassilis Kostakos. Revisitation in urban space vs. online: a comparison across POIs, Websites, and smartphone apps. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 24, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287034.

Chong:2017:DGT

[CCY^{+}17] Eunji Chong, Katha Chanda, Zhefan Ye, Audrey Southerland, Nataniel Ruiz, Rebecca M. Jones, Agata Rozga, and James M. Rehg. Detecting gaze towards eyes in natural social interactions and its use in child assessment. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1–20, September 2017. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3131902.

Cao:2023:EEV

[CCY^{+}23] Yetong Cao, Chao Cai, Anbo Yu, Fan Li, and Jun Luo. EarAcE: Empowering versatile acoustic sensing via earable active noise cancellation platform. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):47:1– 47:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596242.

Cao:2023:VAI

[CDA23] Yifeng Cao, Ashutosh Dhekne, and Mostafa Ammar. ViSig: Automatic interpretation of visual body signals using on-body sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):4:1– 4:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580797.

Chen:2017:RMD

[CDG^{+}17] Yuanying Chen, Wei Dong, Yi Gao, Xue Liu, and Tao Gu. Rapid: a multimodal and device-free approach using noise estimation for robust person identification. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1– 27, September 2017. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3130906.

Chen:2019:DMT

[CDL^{+}19] Ling Chen, Yifang Ding, Dandan Lyu, Xiaoze Liu, and Hanyu Long. Deep multi-task learning based urban air quality index modelling. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1):1–17, March 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314389.

Choi:2022:DEC

[CEL^{+}22] Kyung Yun Choi, Neska El-Haouij, Jinmo Lee, Rosalind W. Picard, and Hiroshi Ishii. Design and evaluation of a clippable and personalizable pneumatichaptic feedback device for breathing guidance. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1): 6:1–6:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517234.

Chen:2022:SSC

[CFF^{+}22] Yang Chen, Katherine Fennedy, Anna Fogel, Shengdong Zhao, Chao Zhang, Lijuan Liu, and Chingchiuan Yen. SSpoon: a shapechanging spoon that optimizes bite size for eating rate regulation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):105:1– 105:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550312.

Cao:2018:UCU

[CFLK18] Hancheng Cao, Jie Feng, Yong Li, and Vassilis Kostakos. Uniqueness in the city: Urban morphology and location privacy. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–20, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214265.

Cauchard:2019:PIP

[CFZ^{+}19] Jessica R. Cauchard, Jeremy Frey, Octavia Zahrt, Krister Johnson, Alia Crum, and James A. Landay. The positive impact of push vs pull progress feedback: a 6-week activity tracking study in the wild. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 23, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351234.

Chitkara:2017:DAR

[CGH^{+}17] Saksham Chitkara, Nishad Gothoskar, Suhas Harish, Jason I. Hong, and Yuvraj Agarwal. Does this app really need my location?: Context-aware privacy management for smartphones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1– 22, September 2017. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3132029.

Choi:2022:PLW

[CGJ^{+}22] Seokmin Choi, Yang Gao, Yincheng Jin, Se jun Kim, Jiyang Li, Wenyao Xu, and Zhanpeng Jin. PPGface: Like what you are watching? Earphones can “Feel” your facial expressions. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):48:1– 48:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534597.

Costa:2019:BIC

[CGJC19] Jean Costa, François Guimbretière, Malte F. Jung, and Tanzeem Choudhury. BoostMeUp: Improving cognitive performance in the moment by unobtrusively regulating emotions with a smartwatch. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–23, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328911.

Carek:2018:NCC

[CH18] Andrew Carek and Christian Holz. Naptics: Convenient and continuous blood pressure monitoring during sleep. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1–22, September 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264906.

Chen:2022:LHP

[CHC^{+}22] Kaixin Chen, Yongzhi Huang, Yicong Chen, Haobin Zhong, Lihua Lin, Lu Wang, and Kaishun Wu. LiSee: a headphone that provides all-day assistance for blind and low-vision users to reach surrounding objects. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):104:1– 104:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550282.

Chen:2023:HHM

[CHWZ23] Ling Chen, Rong Hu, Menghan Wu, and Xin Zhou. HMGAN: a hierarchical multi-modal generative adversarial network model for wearable human activity recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):88:1– 88:??, September 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3610909.

Cheng:2019:IFC

[CHZT19] Yun Cheng, Xiaoxi He, Zimu Zhou, and Lothar Thiele. ICT: In-field calibration transfer for air quality sensor deployments. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–19, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314393.

Chen:2023:DFB

[CJF^{+}23] Xiaowei Chen, Xiao Jiang, Jiawei Fang, Shihui Guo, Juncong Lin, Minghong Liao, Guoliang Luo, and Hongbo Fu. DisPad: Flexible on-body displacement of fabric sensors for robust joint-motion tracking. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1): 5:1–5:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580832.

Colley:2021:SCS

[CJRG21] Mark Colley, Pascal Jansen, Enrico Rukzio, and Jan Gugenheimer. SwiVR-Car-Seat: Exploring vehicle motion effects on interaction quality in virtual reality automated driving using a motorized swivel seat. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):150:1– 150:26, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3494968.

Chauhan:2020:CCL

[CKHM20] Jagmohan Chauhan, Young D. Kwon, Pan Hui, and Cecilia Mascolo. ContAuth: Continual learning framework for behavioral-based user authentication. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):122:1– 122:23, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3432203.

Cho:2023:AHA

[CKJH23] Sungjae Cho, Yoonsu Kim, Jaewoong Jang, and Inseok Hwang. AI-to-human actuation: Boosting unmodified AI’s robustness by proactively inducing favorable human sensing conditions. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1): 6:1–6:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580812.

Cai:2020:TRS

[CKKM20] Hong Cai, Belal Korany, Chitra R. Karanam, and Yasamin Mostofi. Teaching RF to sense without RF training measurements. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):120:1– 120:22, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3432224.

Cho:2023:RDR

[CKL23] Hyunsung Cho, Matthew L.

Komar, and David Lindlbauer. RealityReplay: Detecting and replaying temporal changes in situ using mixed reality. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):90:1– 90:??, September 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3610888.

Cao:2022:MED

[CKLB22] Qingqing Cao, Prerna Khanna, Nicholas D. Lane, and Aruna Balasubramanian. MobiVQA: Efficient ondevice visual question answering. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):44:1–44:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3534619.

Colley:2021:HSA

[CKLR21] Mark Colley, Svenja Krauss, Mirjam Lanzer, and Enrico Rukzio. How should automated vehicles communicate critical situations?: a comparative analysis of visualization concepts. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):94:1– 94:23, September 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3478111.

Cha:2020:HTN

[CKP^{+}20] Narae Cha, Auk Kim, Cheul Young Park, Soowon Kang, Mingyu Park, Jae-Gil Lee, Sangsu Lee, and Uichin Lee. Hello there! Is now a good time to talk?: Opportune moments for proactive interactions with smart speakers. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):74:1– 74:28, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3411810.

Cao:2022:GYH

[CLC^{+}22] Yetong Cao, Fan Li, Huijie Chen, Xiaochen liu, Li Zhang, and Yu Wang. Guard your heart silently: Continuous electrocardiogram waveform monitoring with wrist-worn motion sensor. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):103:1– 103:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550307.

Chen:2020:LYF

[CLD^{+}20] Huijie Chen, Fan Li, Wan Du, Song Yang, Matthew Conn, and Yu Wang. Listen to your fingers: User authentication based on geometry biometrics of touch gesture. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):75:1– 75:23, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3411809.

Cherian:2018:PLW

[CLH18] Jim Cherian, Jun Luo, and Shen-Shyang Ho. Park-Loc: Light-weight graphbased vehicular localization in parking garages. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1– 23, September 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3264909.

Chen:2018:CEA

[CLHW18] Huijie Chen, Fan Li, Xiaojun Hei, and Yu Wang. CrowdX: Enhancing automatic construction of indoor floorplan with opportunistic encounters. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 21, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287037.

Cao:2022:CVR

[CLL^{+}22] Dongjiang Cao, Ruofeng Liu, Hao Li, Shuai Wang, Wenchao Jiang, and Chris Xiaoxuan Lu. Cross vision-RF gait re-identification with low-cost RGB-D cameras and mmWave radars. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):102:1– 102:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550325.

Chen:2019:UPP

[CLRH19] Ying-Yu Chen, Ziyue Li, Daniela Rosner, and Alexis Hiniker. Understanding parents’ perspectives on mealtime technology. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–19, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314392.

Chen:2021:NCT

[CLT^{+}21] Tuochao Chen, Yaxuan Li, Songyun Tao, Hyunchul Lim, Mose Sakashita, Ruidong Zhang, Francois Guimbretiere, and Cheng Zhang. NeckFace: Continuously tracking full facial expressions on neck-mounted wearables. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):58:1– 58:31, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463511.

Chen:2021:SWN

[CLTS21] Wenqiang Chen, Shupei Lin, Elizabeth Thompson, and John Stankovic. SenseCollect: We need efficient ways to collect on-body sensorbased human activity data! Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3): 91:1–91:27, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478119.

Chang:2018:SCR

[CLW^{+}18] Liqiong Chang, Jiaqi Lu, Ju Wang, Xiaojiang Chen, Dingyi Fang, Zhanyong Tang, Petteri Nurmi, and Zheng Wang. SleepGuard: Capturing rich sleep information using smartwatch sensing data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1–34, September 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264908.

Chen:2023:VAE

[CLW^{+}23] Meng Chen, Li Lu, Junhao Wang, Jiadi Yu, Yingying Chen, Zhibo Wang, Zhongjie Ba, Feng Lin, and Kui Ren. VoiceCloak: Adversarial example enabled voice De-identification with balanced privacy and utility. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2): 48:1–48:??, June 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596266.

Chen:2021:UUV

[CLY^{+}21] Longbiao Chen, Chenhui Lu, Fangxu Yuan, Zhihan Jiang, Leye Wang, Daqing Zhang, Ruixiang Luo, Xiaoliang Fan, and Cheng Wang. UVLens: Urban village boundary identification and population estimation leveraging open government data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):57:1– 57:26, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3463495.

Cheng:2023:TIY

[CLY^{+}23] Haiming Cheng, Wei Lou, Yanni Yang, Yi pu Chen, and Xinyu Zhang. TwinkleTwinkle: Interacting with your smart devices by eye blink. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):50:1– 50:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3596238.

Chen:2022:EET

[CLYZ22] Taizhou Chen, Tianpei Li, Xingyu Yang, and Kening Zhu. EFRing: Enabling thumb-to-index-finger microgesture interaction through electric field sensing using single smart ring. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):161:1– 161:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3569478.

Cheng:2021:DAL

[CLZ^{+}21] Tingyu Cheng, Bu Li, Yang Zhang, Yunzhi Li, Charles Ramey, Eui Min Jung, Yepu Cui, Sai Ganesh Swaminathan, Youngwook Do, Manos Tentzeris, Gregory D. Abowd, and Hyun-Joo Oh. Duco: Autonomous large-scale direct-circuit-writing (DCW) on vertical everyday surfaces using a scalable hanging plotter. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3): 92:1–92:25, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478118.

Cao:2018:AAR

[CLZL18] Chu Cao, Zhenjiang Li, Pengfei Zhou, and Mo Li. Amateur: Augmented reality based vehicle navigation system. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1–24, December 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287033.

Clinch:2019:SCP

[CME^{+}19] Sarah Clinch, Mateusz Mikusz, Ivan Elhart, Nigel Davies, and Marc Langheinrich. Scheduling content in pervasive display systems. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 37, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3369826.

Colley:2023:CFM

[CMF^{+}23] Mark Colley, Luca-Maxim Meinhardt, Alexander Fassbender, Michael Rietzler, and Enrico Rukzio. Come fly with me: Investigating the effects of path visualizations in automated urban air mobility. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):52:1– 52:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3596249.

Cho:2022:FFL

[CMK22] Hyunsung Cho, Akhil Mathur, and Fahim Kawsar. FLAME: Federated learning across multi-device environments. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3): 107:1–107:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3550289.

Chatterjee:2020:SDS

[CML^{+}20] Soujanya Chatterjee, Alexander Moreno, Steven Lloyd Lizotte, Sayma Akther, Emre Ertin, Christopher P. Fagundes, Cho Lam, James M. Rehg, Neng Wan, David W. Wetter, and Santosh Kumar. SmokingOpp: Detecting the smoking ‘opportunity’ context using mobile sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):4:1– 4:26, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3380987.

Chen:2022:RUA

[CML22] Hongkai Chen, Sirajum Munir, and Shan Lin. RF-Cam: Uncertainty-aware fusion of camera and Wi-Fi for real-time human identification with mobile devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2): 47:1–47:??, July 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534588.

Centellegher:2018:MMU

[CMV^{+}18] Simone Centellegher, Giovanna Miritello, Daniel Villatoro, Devyani Parameshwar, Bruno Lepri, and Nuria Oliver. Mobile money: Understanding and predicting its adoption and use in a developing economy. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 18, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287035.

Clark:2017:DDA

[CND17] Meghan Clark, Mark W. Newman, and Prabal Dutta. Devices and data and agents, oh my: How smart home abstractions prime end-user mental models. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1– 26, September 2017. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3132031.

Cheng:2020:STI

[CND^{+}20] Tingyu Cheng, Koya Narumi, Youngwook Do, Yang Zhang, Tung D. Ta, Takuya Sasatani, Eric Markvicka, Yoshihiro Kawahara, Lining Yao, Gregory D. Abowd, and HyunJoo Oh. Silver Tape: Inkjet-printed circuits peeled-and-transferred on versatile substrates. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1): 6:1–6:17, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3381013.

Clark:2022:AOS

[CND22] Meghan Clark, Mark W.

Newman, and Prabal Dutta. ARticulate: One-shot interactions with intelligent assistants in unfamiliar smart spaces using augmented reality. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):7:1– 7:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517235.

Cloete:2021:AAM

[CNS21] Richard Cloete, Chris Norval, and Jatinder Singh. Auditable augmented/mixed/ virtual reality: The practicalities of mobile system transparency. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):149:1– 149:24, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3495001.

Chen:2022:SAB

[CNXG22] Yongliang Chen, Tao Ni, Weitao Xu, and Tao Gu. SwipePass: Acoustic-based second-factor user authentication for smartphones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):106:1– 106:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3550292.

Chadalavada:2018:IIS

[CPC^{+}18] Perumal Varun Chadalavada, Goutham Palaniappan, Vimal Kumar Chandran, Khai Truong, and Daniel Wigdor. ID’em: Inductive sensing for embedding and extracting information in robust materials. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1–28, September 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3264907.

Choi:2019:MSR

[CPK^{+}19] Woohyeok Choi, Sangkeun Park, Duyeon Kim, Youn kyung Lim, and Uichin Lee. Multi-stage receptivity model for mobile justin-time health intervention. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1– 26, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3328910.

Cai:2021:WHY

[CPW^{+}21] Chao Cai, Henglin Pu, Peng Wang, Zhe Chen, and Jun Luo. We hear your PACE: Passive acoustic localization of multiple walking persons. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2): 55:1–55:24, June 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463510.

Carson:2021:IES

[CQCH21] Iain Carson, Aaron Quigley, Loraine Clarke, and Uta Hinrichs. Investigating the effect of sensory concurrency on learning haptic spatiotemporal signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1): 6:1–6:30, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3448102.

Curtiss:2021:FSF

[CRB^{+}21] Alexander Curtiss, Blaine Rothrock, Abu Bakar, Nivedita Arora, Jason Huang, Zachary Englhardt, Aaron-Patrick Empedrado, Chixiang Wang, Saad Ahmed, Yang Zhang, Nabil Alshurafa, and Josiah Hester. FaceBit: Smart face masks platform. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):151:1– 151:44, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3494991.

Colley:2022:ESD

[CRGR22] Mark Colley, Max Rädler, Jonas Glimmann, and Enrico Rukzio. Effects of scene detection, scene prediction, and maneuver planning visualizations on trust, situation awareness, and cognitive load in highly automated vehicles. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):49:1– 49:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534609.

Chan:2022:TDL

[CRMG22] Justin Chan, Ananditha Raghunath, Kelly E. Michaelsen, and Shyamnath Gollakota. Testing a drop of liquid using smartphone LiDAR. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1): 3:1–3:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517256.

Chwalek:2021:CSE

[CRP21] Patrick Chwalek, David Ramsay, and Joseph A. Paradiso. Captivates: a smart eyeglass platform for across-context physiological measurement. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):93:1– 93:32, September 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3478079.

Chauhan:2018:PCD

[CRS^{+}18] Jagmohan Chauhan, Jathushan Rajasegaran, Suranga Seneviratne, Archan Misra, Aruna Seneviratne, and Youngki Lee. Performance characterization of deep learning models for breathing-based authentication on resourceconstrained devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 24, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3287036.

Chen:2019:TED

[CS19] Dongyao Chen and Kang G. Shin. TurnsMap: Enhancing driving safety at intersections with mobile crowdsensing and deep learning. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 22, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3351236.

Cai:2022:DEA

[CS22a] Pingping Cai and Sanjib Sur. DeepPCD: Enabling AutoCompletion of indoor point clouds with deep learning. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):43:1– 43:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3534611.

Cai:2022:MBT

[CS22b] Pingping Cai and Sanjib Sur. MilliPCD: Beyond traditional vision indoor point cloud generation via handheld millimeterwave devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):160:1– 160:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3569497.

Chen:2017:MMT

[CSHN17] Ke-Yu Chen, Rahul C.

Shah, Jonathan Huang, and Lama Nachman. Mago: Mode of transport inference using the Hall-effect magnetic sensor and accelerometer. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090054.

Chan:2018:SEE

[CSK^{+}18] Larry Chan, Vedant Das Swain, Christina Kelley, Kaya de Barbaro, Gregory D. Abowd, and Lauren Wilcox. Students’ experiences with ecological momentary assessment tools to report on emotional well-being. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–20, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3191735.

Chan:2020:PIR

[CSM^{+}20] Samantha W. T. Chan, Shardul Sapkota, Rebecca Mathews, Haimo Zhang, and Suranga Nanayakkara. Prompto: Investigating receptivity to prompts based on cognitive load from memory training conversational agent. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):121:1– 121:23, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3432190.

Constantinides:2020:CPM

[CSQ^{+}20] Marios Constantinides, Sanja Sćepanović, Daniele Quercia, Hongwei Li, Ugo Sassi, and Michael Eggleston. ComFeel: Productivity is a matter of the senses too. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):123:1– 123:21, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3432234.

Cavdar:2020:MPA

[CTTMT20] Seyma Kucukozer Cavdar, Tugba Taskaya-Temizel, Mirco Musolesi, and Peter Tino. A multi-perspective analysis of social context and personal factors in Office settings for the design of an effective mobile notification system. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):15:1– 15:38, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3381000.

Coppers:2020:FIP

[CVL20] Sven Coppers, Davy Vanacken, and Kris Luyten. FORT-NIoT: Intelligible predictions to improve user understanding of smart home behavior. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):124:1– 124:24, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3432225.

Cao:2018:EPC

[CW18] Siyuan Cao and He Wang.

Enabling public cameras to talk to the public. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–20, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3214266.

Chen:2019:CCA

[CWH^{+}19] Xinlei Chen, Yu Wang, Jiayou He, Shijia Pan, Yong Li, and Pei Zhang. CAP: Context-aware app usage prediction with heterogeneous graph embedding. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–25, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3314391.

Classen:2018:AVF

[CWP^{+}18] Jiska Classen, Daniel Wegemer, Paul Patras, Tom Spink, and Matthias Hollick. Anatomy of a vulnerable fitness tracking system: Dissecting the Fitbit cloud, app, and firmware. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–24, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3191737.

Chung:2019:IPI

[CWS^{+}19] Chia-Fang Chung, Qiaosi Wang, Jessica Schroeder, Allison Cole, Jasmine Zia, James Fogarty, and Sean A.

Munson. Identifying and planning for individualized change: Patientprovider collaboration using lightweight food diaries in healthy eating and irritable bowel syndrome. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–27, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314394.

Chen:2019:LTC

[CXC^{+}19] Lili Chen, Jie Xiong, Xiaojiang Chen, Sunghoon Ivan Lee, Daqing Zhang, Tao Yan, and Dingyi Fang. LungTrack: Towards contactless and zero dead-zone respiration monitoring with commodity RFIDs. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 22, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351237.

Campagna:2018:CFG

[CXR^{+}18] Giovanni Campagna, Silei Xu, Rakesh Ramesh, Michael Fischer, and Monica S. Lam. Controlling fine-grain sharing in natural language with a virtual assistant. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1– 28, September 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3264905.

Chen:2021:ADS

[CXT^{+}21] Xianda Chen, Yifei Xiao, Yeming Tang, Julio Fernandez-Mendoza, and Guohong Cao. ApneaDetector: Detecting sleep apnea with smartwatches. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):59:1– 59:22, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3463514.

Chen:2021:CFV

[CXZ^{+}21] Yanjiao Chen, Meng Xue, Jian Zhang, Qianyun Guan, Zhiyuan Wang, Qian Zhang, and Wei Wang. ChestLive: Fortifying voice-based authentication with chest motion biometric on smart devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):148:1–148:25, December 2021. CODEN ???? ISSN 2474-9567 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3494962.

Chen:2023:EAM

[CYB^{+}23] Weiyan Chen, Hongliu Yang, Xiaoyang Bi, Rong Zheng, Fusang Zhang, Peng Bao, Zhaoxin Chang, Xujun Ma, and Daqing Zhang. Environment-aware multiperson tracking in indoor environments with MmWave radars. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):89:1– 89:??, September 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3610902.

Chang:2019:TET

[CYK^{+}19] Yung-Ju Chang, Chu-Yuan Yang, Ying-Hsuan Kuo, Wen-Hao Cheng, Chun-Liang Yang, Fang-Yu Lin, I-Hui Yeh, Chih-Kuan Hsieh, Ching-Yu Hsieh, and Yu-Shuen Wang. Tourgether: Exploring tourists’ real-time sharing of experiences as a means of encouraging pointof-interest exploration. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 25, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369832.

Chen:2023:RML

[CYK^{+}23a] Yunzhong Chen, Jiadi Yu, Linghe Kong, Hao Kong, Yanmin Zhu, and Yi-Chao Chen. RF-Mic: Live voice eavesdropping via capturing subtle facial speech dynamics leveraging RFID. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):49:1– 49:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3596259.

Choi:2023:VTF

[CYK^{+}23b] Seokmin Choi, Junghwan Yim, Se Jun Kim, Yincheng Jin, Di Wu, and Zhanpeng Jin. VibPath: Two-factor authentication with your hand’s vibration response to unlock your phone. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):91:1– 91:??, September 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3610894.

Chen:2019:YTC

[CYX^{+}19] Mingshi Chen, Panlong Yang, Jie Xiong, Maotian Zhang, Youngki Lee, Chaocan Xiang, and Chang Tian. Your table can be an input panel: Acousticbased device-free interaction recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–21, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314390.

Chang:2021:MAA

[CZD^{+}21] Yuhu Chang, Yingying Zhao, Mingzhi Dong, Yujiang Wang, Yutian Lu, Qin Lv, Robert P. Dick, Tun Lu, Ning Gu, and Li Shang. MemX: an attention-aware smart eyewear system for personalized moment auto-capture. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2): 56:1–56:23, June 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463509.

Chen:2021:CFP

[CZG^{+}21] Weiyan Chen, Fusang Zhang, Tao Gu, Kexing Zhou, Zixuan Huo, and Daqing Zhang. Constructing floor plan through smoke using ultra wideband radar. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):147:1– 147:29, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3494977.

Chhaglani:2022:FMA

[CZL^{+}22] Bhawana Chhaglani, Camellia Zakaria, Adam Lechowicz, Jeremy Gummeson, and Prashant Shenoy. FlowSense: Monitoring airflow in building ventilation systems using audio sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1): 5:1–5:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3517258.

Chen:2020:MDM

[CZP20] Ling Chen, Yi Zhang, and Liangying Peng. METIER: a deep multi-task learning based activity and user recognition model using wearable sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):5:1– 5:18, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3381012.

Cheng:2022:IRU

[CZT22] Yun Cheng, Zimu Zhou, and Lothar Thiele. iSpray: Reducing urban air pollution with intelligent water spraying. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):4:1– 4:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517227.

Chang:2022:SFS

[CZX^{+}22] Zhaoxin Chang, Fusang Zhang, Jie Xiong, Junqi Ma, Beihong Jin, and Daqing Zhang. Sensor-free soil moisture sensing using LoRa signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):45:1– 45:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534608.

Dunbar:2021:SLA

[DBBH21] Julia C. Dunbar, Emily Bascom, Ashley Boone, and Alexis Hiniker. Is someone listening?: Audiorelated privacy perceptions and design recommendations from guardians, pragmatists, and cynics. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):98:1– 98:23, September 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3478091.

DiSalvo:2022:RRA

[DBW^{+}22] Betsy DiSalvo, Dheeraj Bandaru, Qiaosi Wang, Hong Li, and Thomas Plötz. Reading the room: Automated, momentary assessment of student engagement in the classroom: Are we there yet? Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):112:1– 112:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3550328.

Doryab:2017:ICL

[DBY^{+}17] Afsaneh Doryab, Victoria Bellotti, Alaaeddine Yousfi, Shuobi Wu, John M. Carroll, and Anind K. Dey. If it’s convenient: Leveraging context in peer-to-peer variable service transaction recommendations. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1–28, September 2017. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3130913.

Darvariu:2020:QRB

[DCMM20] Victor-Alexandru Darvariu, Laura Convertino, Abhinav Mehrotra, and Mirco Musolesi. Quantifying the relationships between everyday objects and emotional states through deep learning based image analysis using smartphones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1): 7:1–7:21, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3380997.

Ding:2023:MMH

[DCZ^{+}23] Han Ding, Zhenbin Chen, Cui Zhao, Fei Wang, Ge Wang, Wei Xi, and Jizhong Zhao. MI-Mesh: 3D human mesh construction by fusing image and millimeter wave. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):10:1– 10:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580861.

Doryab:2019:MBR

[DDKL19] Afsaneh Doryab, Anind K.

Dey, Grace Kao, and Carissa Low. Modeling biobehavioral rhythms with passive sensing in the wild: a case study to predict readmission risk after pancreatic surgery.

Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–21, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3314395.

Daskalova:2017:LLT

[DDP^{+}17] Nediyana Daskalova, Karthik Desingh, Alexandra Papoutsaki, Diane Schulze, Han Sha, and Jeff Huang. Lessons learned from two cohorts of personal informatics self-experiments. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1– 22, September 2017. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3130911.

DiLascio:2020:MSA

[DGH^{+}20] Elena Di Lascio, Shkurta Gashi, Juan Sebastian Hidalgo, Beatrice Nale, Maike E. Debus, and Silvia Santini. A multi-sensor approach to automatically recognize breaks and work activities of knowledge workers in academia. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):78:1– 78:20, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3411821.

Dong:2019:PTF

[DGLQ19] Wen Dong, Tong Guan, Bruno Lepri, and Chunming Qiao. PocketCare: Tracking the flu with mobile phones using partial observations of proximity and symptoms. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1– 23, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3328912.

DiLascio:2018:UAS

[DGS18] Elena Di Lascio, Shkurta Gashi, and Silvia Santini. Unobtrusive assessment of students’ emotional engagement during lectures using electrodermal activity sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1– 21, September 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264913.

deGuzman:2021:USP

[dGST21] Jaybie Agullo de Guzman, Aruna Seneviratne, and Kanchana Thilakarathna. Unravelling spatial privacy risks of mobile mixed reality data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):14:1– 14:26, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3448103.

Ding:2021:CWC

[DGZ^{+}21] Yi Ding, Baoshen Guo, Lin Zheng, Mingming Lu, Desheng Zhang, Shuai Wang, Sang Hyuk Son, and Tian He. A city-wide crowdsourcing delivery system with reinforcement learning. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):97:1– 97:22, September 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3478117.

Dementyev:2017:DWD

[DH17] Artem Dementyev and Christian Holz. DualBlink: A wearable device to continuously detect, track, and actuate blinking for alleviating dry eyes and computer vision syndrome. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(1):1:1–1:19, March 2017. CODEN ???? ISSN 2474-9567. URL http:

//dl.acm.org/citation.
cfm?id=3053330.

Dementyev:2018:ERW

[DHC^{+}18] Artem Dementyev, Javier Hernandez, Inrak Choi, Sean Follmer, and Joseph Paradiso. Epidermal robots: Wearable sensors that climb on the skin. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1–22, September 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264912.

Ding:2021:SIL

[DJL^{+}21] Yi Ding, Dongzhe Jiang, Yunhuai Liu, Desheng Zhang, and Tian He. Smart-LOC: Indoor localization with smartphone anchors for on-demand delivery. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):153:1– 153:24, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3494972.

Drey:2022:SEC

[DJL^{+}22] Tobias Drey, Jessica Janek, Josef Lang, Dietmar Puschmann, Michael Rietzler, and Enrico Rukzio. SpARklingPaper: Enhancing common penand paper-based handwriting training for children by digitally augmenting papers using a tablet screen. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):113:1– 113:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550337.

DSilva:2018:RUM

[DJN^{+}18] Krittika D’Silva, Kasthuri Jayarajah, Anastasios Noulas, Cecilia Mascolo, and Archan Misra. The role of urban mobility in retail business survival. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1–22, September 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264910.

Ding:2023:SMS

[DJX^{+}23] Rong Ding, Haiming Jin, Dong Xiang, Xiaocheng Wang, Yongkui Zhang, Dingman Shen, Lu Su, Wentian Hao, Mingyuan Tao, Xinbing Wang, and Chenghu Zhou. Soil moisture sensing with UAVmounted IR-UWB radar and deep learning. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):11:1– 11:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580867.

Ding:2022:PLP

[DJY^{+}22] Yi Ding, Dongzhe Jiang, Yu Yang, Yunhuai Liu, Tian He, and Desheng Zhang. P2-Loc: a person-2-person indoor localization system in on-demand delivery. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1): 9:1–9:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517238.

Dai:2022:MTL

[DKZ^{+}22] Ruixuan Dai, Thomas Kannampallil, Jingwen Zhang, Nan Lv, Jun Ma, and Chenyang Lu. Multi-task learning for randomized controlled trials: a case study on predicting depression with wearable data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):50:1– 50:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534591.

Duan:2021:FDR

[DLD^{+}21] Chunhui Duan, Jiajun Liu, Xuan Ding, Zhenhua Li, and Yunhao Liu. Fulldimension relative positioning for RFID-enabled selfcheckout services. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):7:1– 7:23, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3448094.

Daskalova:2018:IEC

[DLH^{+}18] Nediyana Daskalova, Bongshin Lee, Jeff Huang, Chester Ni, and Jessica Lundin. Investigating the effectiveness of cohortbased sleep recommendations. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1–19, September 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3264911.

Dissanayake:2018:DDE

[DMAH18] Thilina Dissanayake, Takuya Maekawa, Daichi Amagata, and Takahiro Hara. Detecting door events using a smartphone via active sound sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1–26, December 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287038.

Devine:2022:PPP

[DMdH^{+}22] James Devine, Michal Moskal, Peli de Halleux, Thomas Ball, Steve Hodges, Gabriele D’Amone, David Gakure, Joe Finney, Lorraine Underwood, Kobi Hartley, Paul Kos, and Matt Oppenheim. Plug-and-play physical computing with Jacdac. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):110:1– 110:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550317.

Donchev:2021:IRP

[DPB21] Rumen Donchev, Erik Pescara, and Michael Beigl. Investigating retention in passive haptic learning of piano songs. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):60:1– 60:14, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463513.

Do:2021:SWC

[DPW^{+}21] Youngwook Do, Jung Wook Park, Yuxi Wu, Avinandan Basu, Dingtian Zhang, Gregory D. Abowd, and Sauvik Das. Smart webcam cover: Exploring the design of an intelligent webcam cover to improve usability and trust. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):154:1– 154:21, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3494983.

Dingler:2017:BCA

[DSM17] Tilman Dingler, Albrecht Schmidt, and Tonja Machulla. Building cognition-aware systems: a mobile toolkit for extracting time-of-day fluctuations of cognitive performance. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1–15, September 2017. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3132025.

Duente:2018:MMB

[DSPR18] Tim Duente, Justin Schulte, Max Pfeiffer, and Michael Rohs. MuscleIO: Musclebased input and output for casual notifications. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–21, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214267.

Das:2022:BRA

[DSR^{+}22] Sarkar Snigdha Sarathi Das, Subangkar Karmaker Shanto, Masum Rahman, Md Saiful Islam, Atif Hasan Rahman, Mohammad M. Masud, and Mohammed Eunus Ali. BayesBeat: Reliable atrial fibrillation detection from noisy photoplethysmography data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1): 8:1–8:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517247.

Dasari:2023:RRM

[DSSD23] Mallesham Dasari, Ramanujan K. Sheshadri, Karthikeyan Sundaresan, and Samir R. Das. RoVaR: Robust multi-agent tracking through dual-layer diversity in visual and RF sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1): 8:1–8:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3580854.

Deldari:2020:EES

[DSSS20] Shohreh Deldari, Daniel V. Smith, Amin Sadri, and Flora Salim. ESPRESSO: Entropy and ShaPe awaRe timE-Series SegmentatiOn for processing heterogeneous sensor data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):77:1– 77:24, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3411832.

Davis:2023:IMI

[DWCY23] Josh Urban Davis, Hongwei Wang, Parmit K. Chilana, and Xing-Dong Yang. “It’s not an issue of malice, but of ignorance”: Towards inclusive video conferencing for presenters who are d/deaf or hard of hearing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):92:1– 92:??, September 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3610901.

Dobbelstein:2017:PWD

[DWHR17] David Dobbelstein, Christian Winkler, Gabriel Haas, and Enrico Rukzio. PocketThumb: a wearable dual-sided touch interface for cursor-based control of smart-eyewear. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090055.

Ding:2022:USE

[DWL^{+}22] Han Ding, Yizhan Wang, Hao Li, Cui Zhao, Ge Wang, Wei Xi, and Jizhong Zhao. UltraSpeech: Speech enhancement by interaction between ultrasound and speech. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):111:1–111:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic).

URL https://dl.acm.org/
doi/10.1145/3550303.

Ding:2019:AHS

[DWZZ19] Feng Ding, Dong Wang, Qian Zhang, and Run Zhao. ASSV: Handwritten signature verification using acoustic signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1–22, September 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3351238.

Deldari:2022:CCM

[DXS^{+}22] Shohreh Deldari, Hao Xue, Aaqib Saeed, Daniel V. Smith, and Flora D. Salim. COCOA: Cross modality contrastive learning for sensor data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):108:1– 108:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3550316.

Ding:2021:LIB

[DYCX21] Dian Ding, Lanqing Yang, Yi-Chao Chen, and Guangtao Xue. Leakage or identification: Behaviorirrelevant user identification leveraging leakage current on laptops. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):152:1– 152:23, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3494984.

Ding:2019:LHC

[DYL^{+}19] Jingtao Ding, Guanghui Yu, Yong Li, Depeng Jin, and Hui Gao. Learning from hometown and current city:

Cross-city POI recommendation via interest drift and transfer learning. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1–28, December 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369822.

Deng:2022:GEA

[DZH^{+}22] Kaikai Deng, Dong Zhao, Qiaoyue Han, Shuyue Wang, Zihan Zhang, Anfu Zhou, and Huadong Ma. Geryon: Edge assisted real-time and robust object detection on drones via mmWave radar and camera fusion. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):109:1– 109:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550298.

Deng:2023:MGM

[DZH^{+}23] Kaikai Deng, Dong Zhao, Qiaoyue Han, Zihan Zhang, Shuyue Wang, Anfu Zhou, and Huadong Ma. Midas: Generating mmWave radar data from videos for training pervasive and privacy-preserving human sensing tasks. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):9:1– 9:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3580872.

Ding:2021:RRJ

[DZZ^{+}21] Han Ding, Linwei Zhai, Cui Zhao, Songjiang Hou, Ge Wang, Wei Xi, Jizhong Zhao, and Yihong Gong. RF-ray: Joint RF and linguistics domain learning for object recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):96:1– 96:24, September 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3478115.

Esteves:2022:OHI

[EBPA22] Augusto Esteves, Elizabeth Bouquet, Ken Pfeuffer, and Florian Alt. One-handed input for mobile devices via motion matching and orbits controls. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):51:1– 51:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3534624.

Elhattab:2022:RFL

[EBTN22] Fatima Elhattab, Sara Bouchenak, Rania Talbi, and Vlad Nitu. Robust federated learning for ubiquitous computing through mitigation of edge-case backdoor attacks. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):162:1– 162:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3569492.

Epstein:2020:MTS

[ECF^{+}20] Daniel A. Epstein, Clara Caldeira, Mayara Costa Figueiredo, Xi Lu, Lucas M. Silva, Lucretia Williams, Jong Ho Lee, Qingyang Li, Simran Ahuja, Qiuer Chen, Payam Dowlatyari, Craig Hilby, Sazeda Sultana, Elizabeth V. Eikey, and Yunan Chen. Mapping and taking stock of the personal informatics literature. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):126:1– 126:38, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3432231.

Elbakly:2018:HRU

[EEY18] Rizanne Elbakly, Moustafa Elhamshary, and Moustafa Youssef. HyRise: a robust and ubiquitous multisensor fusion-based floor localization system. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1– 23, September 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264914.

Engelbrecht:2019:EIT

[ELD19] Hendrik Engelbrecht, Stephan G. Lukosch, and Dragos Datcu. Evaluating the impact of technology assisted hotspot policing on situational awareness and task-load. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–18, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314396.

Ehtesham-Ul-Haque:2023:AGL

[EUHB23]  Md  Ehtesham-Ul-Haque

and Syed Masum Billah. Abacus Gestures: a large set of math-based usable finger-counting gestures for mid-air interactions. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):93:1– 93:??, September 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3610898.

Elsayed:2020:VUS

[EWM^{+}20] Hesham Elsayed, Martin Weigel, Florian Müller, Martin Schmitz, Karola Marky, Sebastian Günther, Jan Riemann, and Max Mühlhäuser. VibroMap: Understanding the spacing of vibrotactile actuators across the body. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):125:1– 125:16, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3432189.

Elbakly:2020:SSB

[EY20] Rizanne Elbakly and Moustafa Youssef. The StoryTeller: Scalable building- and APindependent deep learningbased floor prediction. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1): 8:1–8:20, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3380979.

Ekambaranathan:2023:NDA

[EZC23] Anirudh Ekambaranathan, Jun Zhao, and George Chalhoub. Navigating the data avalanche: Towards supporting developers in developing privacy-friendly children’s apps. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):53:1– 53:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596267.

Feustel:2018:PLM

[FALW18] Clayton Feustel, Shyamak Aggarwal, Bongshin Lee, and Lauren Wilcox. People like me: Designing for reflection on aggregate cohort data in personal informatics systems. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1– 21, September 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264917.

Finnigan:2017:AAE

[FCFW^{+}17] S. Mitchell Finnigan, A. K. Clear, G. Farr-Wharton, K. Ladha, and R. Comber. Augmenting audits: Exploring the role of sensor toolkits in sustainable buildings management. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090075.

Feng:2020:SMF

[FCQR20] Yunhe Feng, Qing Cao, Hairong Qi, and Scott Ruoti. SenCAPTCHA: a mobile-first CAPTCHA using orientation sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):43:1– 43:26, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397312.

Foo:2021:UEM

[FDH21] Esther W. Foo, Lucy E.

Dunne, and Brad Holschuh. User expectations and mental models for communicating emotions through compressive & warm affective garment actuation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):10:1– 10:25, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3448097.

Fan:2018:EBW

[FDL^{+}18] Xiaoran Fan, Han Ding, Sugang Li, Michael Sanzari, Yanyong Zhang, Wade Trappe, Zhu Han, and Richard E. Howard. Energyball: Wireless power transfer for batteryless Internet of Things through distributed beamforming. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–22, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214268.

Fang:2020:PPI

[FFQ^{+}20] Zhihan Fang, Boyang Fu, Zhou Qin, Fan Zhang, and Desheng Zhang. Private-Bus: Privacy identification and protection in large-scale bus WiFi systems. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1): 9:1–9:23, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3380990.

Fan:2018:WYH

[FFT^{+}18] Junjun Fan, Xiangmin Fan, Feng Tian, Yang Li, Zitao Liu, Wei Sun, and Hongan Wang. What is that in your hand?: Recognizing grasped objects via forearm electromyography sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 24, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287039.

Fan:2018:TAI

[FGL18] Xiaoyi Fan, Wei Gong, and Jiangchuan Liu. TagFree activity identification with RFIDs. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1):1–23, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191739.

Fan:2023:FTP

[FGS^{+}23] Feiyi Fan, Yang Gu, Jianfei Shen, Fan Dong, and Yiqiang Chen. FewShotBP: Towards personalized ubiquitous continuous blood pressure measurement. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):94:1– 94:??, September 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3610918.

Fujinami:2018:PAA

[FKI18] Kaori Fujinami, Mami Kosaka, and Bipin Indurkhya. Painting an Apple with an Apple: a tangible tabletop interface for painting with physical objects.

Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 22, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287040.

Fomichev:2019:PZI

[FMA^{+}19] Mikhail Fomichev, Max Maass, Lars Almon, Alejandro Molina, and Matthias Hollick. Perils of zerointeraction security in the Internet of Things. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–38, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314397.

Fu:2022:SEL

[FML^{+}22] Yue Fu, Rebecca Michelson, Yifan Lin, Lynn K. Nguyen, Tala June Tayebi, and Alexis Hiniker. Social emotional learning with conversational agents: Reviewing current designs and probing parents’ ideas for future ones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):52:1– 52:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534622.

Fang:2023:IPH

[FMP^{+}23] Likun Fang, Timo Müller, Erik Pescara, Nikola Fischer, Yiran Huang, and Michael Beigl. Investigating passive haptic learning of piano songs using three tactile sensations of vibration, stroking and tapping. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):95:1– 95:??, September 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3610899.

Fagert:2022:RSR

[FMZN22] Jonathon Fagert, Mostafa Mirshekari, Pei Zhang, and Hae Young Noh. Recursive sparse representation for identifying multiple concurrent occupants using floor vibration sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):10:1– 10:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517229.

Falcao:2021:ISA

[FRB^{+}21] João Diogo Falcão, Carlos Ruiz, Adeola Bannis, Hae Young Noh, and Pei Zhang. ISACS: In-store autonomous checkout system for retail. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):99:1– 99:26, September 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3478086.

Feng:2020:PPP

[FRS^{+}20] Jie Feng, Can Rong, Funing Sun, Diansheng Guo, and Yong Li. PMF: a privacypreserving human mobility prediction framework via federated learning. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):10:1– 10:21, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3381006.

Finley:2018:MDT

[FS18] Benjamin Finley and Tapio Soikkeli. Mobile device type substitution. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–20, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191740.

Fan:2019:DAB

[FSJ^{+}19] Zipei Fan, Xuan Song, Renhe Jiang, Quanjun Chen, and Ryosuke Shibasaki. Decentralized attentionbased personalized human mobility prediction. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 26, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369830.

Feng:2019:VPL

[FSW19] Xianglong Feng, Viswanathan Swaminathan, and Sheng Wei. Viewport prediction for live 360-degree mobile video streaming using user-content hybrid motion tracking. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–22, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328914.

Fan:2018:ODE

[FSX^{+}18] Zipei Fan, Xuan Song, Tianqi Xia, Renhe Jiang, Ryosuke Shibasaki, and Ritsu Sakuramachi. Online deep ensemble learning for predicting citywide human mobility. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1–21, September 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264915.

Fan:2019:PCA

[FTL^{+}19] Yali Fan, Zhen Tu, Yong Li, Xiang Chen, Hui Gao, Lin Zhang, Li Su, and Depeng Jin. Personalized contextaware collaborative online activity prediction. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 28, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369829.

Feng:2022:WLT

[FWJ^{+}22] Chao Feng, Nan Wang, Yicheng Jiang, Xia Zheng, Kang Li, Zheng Wang, and Xiaojiang Chen. Wi-Learner: Towards oneshot learning for crossdomain Wi-Fi based gesture recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):114:1– 114:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550318.

Feng:2021:RIN

[FXC^{+}21] Chao Feng, Jie Xiong, Liqiong Chang, Fuwei Wang, Ju Wang, and Dingyi Fang. RF-Identity: Nonintrusive person identification based on commodity RFID devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):9:1– 9:23, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3448101.

Fang:2019:MMI

[FYW^{+}19] Zhihan Fang, Yu Yang, Shuai Wang, Boyang Fu, Zixing Song, Fan Zhang, and Desheng Zhang. MAC: Measuring the impacts of anomalies on travel time of multiple transportation systems. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–24, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328913.

Farrukh:2021:SSC

[FYX^{+}21] Habiba Farrukh, Tinghan Yang, Hanwen Xu, Yuxuan Yin, He Wang, and Z. Berkay Celik. S3: Sidechannel attack on stylus pencil through sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1): 8:1–8:25, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3448085.

Fang:2021:CHM

[FYY^{+}21] Zhihan Fang, Yu Yang, Guang Yang, Yikuan Xian, Fan Zhang, and Desheng Zhang. CellSense: Human mobility recovery via cellular network data enhancement. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):100:1– 100:22, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478087.

Feng:2023:FID

[FZS^{+}23] Justin Feng, Tianyi Zhao, Shamik Sarkar, Dominic Konrad, Timothy Jacques, Danijela Cabric, and Nader Sehatbakhsh. Fingerprinting IoT devices using latent physical side-channels. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2): 54:1–54:??, June 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596247.

Fang:2018:MUP

[FZYZ18] Zhihan Fang, Fan Zhang, Ling Yin, and Desheng Zhang. MultiCell: Urban population modeling based on multiple cellphone networks. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1–25, September 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264916.

Fang:2022:HVE

[FZZ^{+}22] Fengyi Fang, Hongwei Zhang, Lishuang Zhan, Shihui Guo, Minying Zhang, Juncong Lin, Yipeng Qin, and Hongbo Fu. Handwriting Velcro: Endowing AR glasses with personalized and posture-adaptive text input using flexible touch sensor. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):163:1– 163:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3569461.

Gleason:2018:CIM

[GAS^{+}18] Cole Gleason, Dragan Ahmetovic, Saiph Savage, Carlos Toxtli, Carl Posthuma, Chieko Asakawa, Kris M. Kitani, and Jeffrey P. Bigham. Crowdsourcing the installation and maintenance of indoor localization infrastructure to support blind navigation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–25, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191741.

Georgiev:2017:LRM

[GBLM17] Petko Georgiev, Sourav Bhattacharya, Nicholas D. Lane, and Cecilia Mascolo. Low-resource multitask audio sensing for mobile and embedded devices via shared deep neural network representations. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1– 19, September 2017. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3131895.

Gairola:2021:SSB

[GBS^{+}21] Siddhartha Gairola, Murtuza Bohra, Nadeem Shaheer, Navya Jayaprakash, Pallavi Joshi, Anand Balasubramaniam, Kaushik Murali, Nipun Kwatra, and Mohit Jain. SmartKC: Smartphone-based corneal topographer for keratoconus detection. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):155:1– 155:27, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3494982.

Gullapalli:2021:OWB

[GCC^{+}21] Bhanu Teja Gullapalli, Stephanie Carreiro, Brittany P. Chapman, Deepak Ganesan, Jan Sjoquist, and Tauhidur Rahman. Opi-Track: a wearable-based clinical opioid use tracker with temporal convolutional attention networks. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):102:1– 102:29, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478107.

Gadiraju:2017:MOC

[GCGD17] Ujwal Gadiraju, Alessandro Checco, Neha Gupta, and Gianluca Demartini. Modus operandi of crowd workers: The invisible role of microtask work environments. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1– 29, September 2017. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3130914.

Gong:2019:KVB

[GCLL19] Taesik Gong, Hyunsung Cho, Bowon Lee, and Sung-Ju Lee. Knocker: Vibroacoustic-based object recognition with smartphones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1–21, September 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3351240.

Guo:2019:BCP

[GCM^{+}19] Anhong Guo, Ilter Canberk, Hannah Murphy, Andrés Monroy-Hernández, and Rajan Vaish. Blocks: Collaborative and persistent augmented reality experiences. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 24, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3351241.

Gupta:2022:TVU

[GCP^{+}22] Kunal Gupta, Sam W. T. Chan, Yun Suen Pai, Nicholas Strachan, John Su, Alexander Sumich, Suranga Nanayakkara, and Mark Billinghurst. Total VREcall: Using biosignals to recognize emotional autobiographical memory in virtual reality. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2): 55:1–55:??, July 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534615.

Guo:2018:SQA

[GCW^{+}18] Suiming Guo, Chao Chen, Jingyuan Wang, Yaxiao Liu, Ke Xu, Daqing Zhang, and Dah Ming Chiu. A simple but quantifiable approach to dynamic price prediction in ride-on-demand services leveraging multisource urban data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1– 24, September 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264922.

Gashi:2019:UUW

[GDS19] Shkurta Gashi, Elena Di Lascio, and Silvia Santini. Using unobtrusive wearable sensors to measure the physiological synchrony between presenters and audience members. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–19, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3314400.

Gashi:2020:DAA

[GDS^{+}20] Shkurta Gashi, Elena Di Lascio, Bianca Stancu, Vedant Das Swain, Varun Mishra, Martin Gjoreski, and Silvia Santini. Detection of artifacts in ambulatory electrodermal activity data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):44:1– 44:31, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3397316.

Goyal:2017:IIM

[GF17] Nitesh Goyal and Susan R. Fussell. Intelligent interruption management using electro dermal activity based physiological sensor for collaborative sensemaking. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1– 21, September 2017. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3130917.

Gleason:2018:FGR

[GFK^{+}18] Cole Gleason, Alexander J.

Fiannaca, Melanie Kneisel, Edward Cutrell, and Meredith Ringel Morris. Foot-Notes: Geo-referenced audio annotations for nonvisual exploration. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1– 24, September 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264919.

Guo:2022:IHL

[GGJ^{+}22] Dongfang Guo, Chaojie Gu, Linshan Jiang, Wenjie Luo, and Rui Tan. IL-LOC: In-hall localization with standard LoRaWAN uplink frames. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):13:1– 13:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517245.

Gedik:2018:DCG

[GH18] Ekin Gedik and Hayley Hung. Detecting conversing groups using social dynamics from wearable acceleration: Group size awareness.

Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 24, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3287041.

Gonzalez:2022:LLP

[GH22] Jesse T. Gonzalez and Scott E. Hudson. Layer by layer, patterned valves enable programmable soft surfaces. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):12:1–12:??, March 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3517251.

Goncalves:2017:CCT

[GHvB^{+}17] Jorge Goncalves, Simo Hosio, Niels van Berkel, Furqan Ahmed, and Vassilis Kostakos. CrowdPickUp: Crowdsourcing task pickup in the wild. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1–22, September 2017. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3130916.

Gao:2019:PPC

[GHY^{+}19] Chen Gao, Chao Huang, Yue Yu, Huandong Wang, Yong Li, and Depeng Jin. Privacy-preserving cross-domain location recommendation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–21, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314398.

Gao:2021:VES

[GJC^{+}21a] Yang Gao, Yincheng Jin, Jagmohan Chauhan, Seokmin Choi, Jiyang Li, and Zhanpeng Jin. Voice in ear: Spoofing-resistant and passphrase-independent body sound authentication. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1): 12:1–12:25, March 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3448113.

Gao:2021:STF

[GJC^{+}21b] Yang Gao, Yincheng Jin, Seokmin Choi, Jiyang Li, Junjie Pan, Lin Shu, Chi Zhou, and Zhanpeng Jin. SonicFace: Tracking facial expressions using a commodity microphone array. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4): 156:1–156:33, December 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3494988.

Guo:2018:CAC

[GJG^{+}18] Anhong Guo, Anuraag Jain, Shomiron Ghose, Gierad Laput, Chris Harrison, and Jeffrey P. Bigham. Crowd-AI camera sensing in the real world. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1–20, September 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3264921.

Gao:2020:EEA

[GJL^{+}20] Yang Gao, Yincheng Jin, Jiyang Li, Seokmin Choi, and Zhanpeng Jin. EchoWhisper: Exploring an acousticbased silent speech interface for smartphone users. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):80:1– 80:27, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3411830.

Gil:2018:FAE

[GKO18] Hyunjae Gil, Hongmin Kim, and Ian Oakley. Fingers and angles: Exploring the comfort of touch input on smartwatches. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 21, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287042.

Gong:2023:DLF

[GKO^{+}23] Taesik Gong, Yewon Kim, Adiba Orzikulova, Yunxin Liu, Sung Ju Hwang, Jinwoo Shin, and Sung-Ju Lee. DAPPER: Label-free performance estimation after personalization for heterogeneous mobile sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):55:1– 55:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596256.

Gong:2018:SPL

[GL18] Wei Gong and Jiangchuan Liu. SiFi: Pushing the limit of time-based WiFi localization using a single commodity access point. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–21, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3191742.

Guo:2021:PPI

[GLC^{+}21] Yeting Guo, Fang Liu, Zhiping Cai, Hui Zeng, Li Chen, Tongqing Zhou, and Nong Xiao. PREFER: Point-of-interest REcommendation with efficiency and privacypreservation via federated edge leaRning. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):13:1– 13:25, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3448099.

Gao:2019:TRA

[GLG^{+}19] Yan Gao, Yang Long, Yu Guan, Anna Basu, Jessica Baggaley, and Thomas Ploetz. Towards reliable, automated general movement assessment for perinatal stroke screening in infants using wearable accelerometers. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–22, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3314399.

Goodman:2021:TUD

[GLJ^{+}21] Steven M. Goodman, Ping Liu, Dhruv Jain, Emma J. McDonnell, Jon E. Froehlich, and Leah Findlater. Toward user-driven sound recognizer personalization with people who are d/deaf or hard of hearing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):63:1– 63:23, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463501.

Guo:2018:DFP

[GLS^{+}18] Xiaonan Guo, Jian Liu, Cong Shi, Hongbo Liu, Yingying Chen, and Mooi Choo Chuah. Device-free personalized fitness assistant using WiFi. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1–23, December 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287043.

Gao:2018:WSY

[GLW^{+}18] Yang Gao, Borui Li, Wei Wang, Wenyao Xu, Chi Zhou, and Zhanpeng Jin. Watching and safeguarding your 3D printer: Online process monitoring against cyber-physical attacks. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1– 27, September 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264918.

Gao:2022:TRG

[GLX^{+}22] Ruiyang Gao, Wenwei Li, Yaxiong Xie, Enze Yi, Leye Wang, Dan Wu, and Daqing Zhang. Towards robust gesture recognition by characterizing the sensing quality of WiFi signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):11:1– 11:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517241.

Garg:2019:UMC

[GM19] Radhika Garg and Christopher Moreno. Understanding motivators, constraints, and practices of sharing Internet of Things. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–21, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328915.

Grammenos:2018:YSY

[GMC18] Andreas Grammenos, Cecilia Mascolo, and Jon Crowcroft. You are sensing, but are you biased?: a user unaided sensor calibration approach for mobile sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–26, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191743.

Gazzari:2021:MAL

[GMMH21] Matthias Gazzari, Annemarie Mattmann, Max Maass, and Matthias Hollick. My(o) armband leaks passwords: an EMG and IMU based keylogging sidechannel attack. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):157:1– 157:24, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3494986.

Gao:2023:RTC

[GMS^{+}23] Chenyang Gao, Ivan Marsic, Aleksandra Sarcevic, Waverly Gestrich-Thompson, and Randall S. Burd. Realtime context-aware multimodal network for activity and activity-stage recognition from team communication in dynamic clinical settings. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):12:1– 12:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580798.

Gummeson:2017:RLB

[GMY^{+}17] Jeremy Gummeson, James Mccann, Chouchang (JACK) Yang, Damith Ranasinghe, Scott Hudson, and Alanson Sample. RFID light bulb: Enabling ubiquitous deployment of interactive RFID systems. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090077.

Gullapalli:2019:BSC

[GNA^{+}19] Bhanu Teja Gullapalli, Annamalai Natarajan, Gustavo A. Angarita, Robert T. Malison, Deepak Ganesan, and Tauhidur Rahman. Onbody sensing of cocaine craving, euphoria and drugseeking behavior using cardiac and respiratory signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1– 31, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3328917.

Gil:2022:TAT

[GO22] Hyunjae Gil and Ian Oakley.

ThumbAir: In-air typing for head mounted displays. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):164:1– 164:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3569474.

Guo:2017:CFG

[GOZ^{+}17] Bin Guo, Yi Ouyang, Cheng Zhang, Jiafan Zhang, Zhiwen Yu, Di Wu, and Yu Wang. CrowdStory: Fine-grained event storyline generation by fusion of multi-modal crowdsourced data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1–19, September 2017. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3130920.

Guan:2017:EDL

[GP17] Yu Guan and Thomas Plötz.

Ensembles of deep LSTM learners for activity recognition using wearables. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090076.

Gupta:2023:FWB

[GPB^{+}23] Agrim Gupta, Daegue Park, Shayaun Bashar, Cedric Girerd, Nagarjun Bhat, Siddhi Mundhra, Tania K. Morimoto, and Dinesh Bharadia. ForceSticker: Wireless, batteryless, thin & flexible force sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):13:1– 13:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580793.

Gulati:2018:CSC

[GPW^{+}18] Manoj Gulati, Farshid Salemi Parizi, Eric Whitmire, Sidhant Gupta, Shobha Sundar Ram, Amarjeet Singh, and Shwetak N. Patel. CapHarvester: a stick-on capacitive energy harvester using stray electric field from AC power lines. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1–20, September 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264920.

Gao:2022:IGW

[GRS^{+}22] Nan Gao, Mohammad Saiedur Rahaman, Wei Shao, Kaixin Ji, and Flora D. Salim. Individual and group-wise classroom seating experience: Effects on student engagement in different courses. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3): 115:1–115:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550335.

Garg:2020:JLM

[GS20] Radhika Garg and Subhasree Sengupta. He is just like me: a study of the long-term use of smart speakers by parents and children. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):11:1– 11:24, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3381002.

Gao:2020:GPC

[GSRS20] Nan Gao, Wei Shao, Mohammad Saiedur Rahaman, and Flora D. Salim. n-Gage: Predicting in-class emotional, behavioural and cognitive engagement in the wild. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):79:1– 79:26, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3411813.

Giallanza:2019:KSM

[GSS^{+}19] Tyler Giallanza, Travis Siems, Elena Smith, Erik Gabrielsen, Ian Johnson, Mitchell A. Thornton, and Eric C. Larson. Keyboard snooping from mobile phone arrays with mixed convolutional and recurrent neural networks. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–22, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328916.

Gao:2023:MMM

[GwC^{+}23] Ziqi Gao, Yuntao wang, Jianguo Chen, Junliang Xing, Shwetak Patel, Xin Liu, and Yuanchun Shi. MMTSA: Multi-modal temporal segment attention network for efficient human activity recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):96:1– 96:??, September 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3610872.

Groh:2017:ASV

[GWD^{+}17] Benjamin H. Groh, Frank Warschun, Martin Deininger, Thomas Kautz, Christine Martindale, and Bjoern M. Eskofier. Automated ski velocity and jump length determination in ski jumping based on unobtrusive and wearable sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1– 17, September 2017. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3130918.

Gan:2021:YWQ

[GWJ^{+}21] Yanglei Gan, Tianyi Wang, Alireza Javaheri, Elaheh Momeni-Ortner, Milad Dehghani, Mehdi Hosseinzadeh, and Reza Rawassizadeh. 11 years with wearables: Quantitative analysis of social media, academia, news agencies, and lead user community from 2009–2020 on wearable technologies. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1): 11:1–11:26, March 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3448096.

Gao:2019:EUE

[GWP^{+}19] Yang Gao, Wei Wang, Vir V. Phoha, Wei Sun, and Zhanpeng Jin. EarEcho: Using ear canal echo for wearable authentication. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 24, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3351239.

Gu:2020:QTE

[GYL^{+}20] Yizheng Gu, Chun Yu, Zhipeng Li, Zhaoheng Li, Xiaoying Wei, and Yuanchun Shi. QwertyRing: Text entry on physical surfaces using a ring. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):128:1– 128:29, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3432204.

Guo:2023:THD

[GZC^{+}23] Shihui Guo, Lishuang Zhan, Yancheng Cao, Chen Zheng, Guyue Zhou, and Jiangtao Gong. Touch-and-heal: Data-driven affective computing in tactile interaction with robotic dog. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):56:1– 56:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596258.

Gong:2021:RIM

[GZH^{+}21] Jian Gong, Xinyu Zhang, Yuanjun Huang, Ju Ren, and Yaoxue Zhang. Robust inertial motion tracking through deep sensor fusion across smart earbuds and smartphone. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):62:1– 62:26, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463517.

Gong:2021:RVS

[GZL^{+}21] Jian Gong, Xinyu Zhang, Kaixin Lin, Ju Ren, Yaoxue Zhang, and Wenxun Qiu. RF vital sign sensing under free body movement. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):101:1– 101:22, September 2021.

CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478090.

Gong:2022:BAB

[GZNL22] Yanbin Gong, Qian Zhang, Bobby H. P. NG, and Wei Li. BreathMentor: Acoustic-based diaphragmatic breathing monitor system. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):53:1–53:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3534595.

Guo:2022:WWS

[GZW^{+}22] Baoshen Guo, Weijian Zuo, Shuai Wang, Wenjun Lyu, Zhiqing Hong, Yi Ding, Tian He, and Desheng Zhang. WePos: Weaksupervised indoor positioning with unlabeled WiFi for on-demand delivery. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):54:1– 54:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3534574.

Guo:2021:MDB

[GZY^{+}21] Gabriel Guo, Hanbin Zhang, Liuyi Yao, Huining Li, Chenhan Xu, Zhengxiong Li, and Wenyao Xu. MSLife: Digital behavioral phenotyping of multiple sclerosis symptoms in the wild using wearables and graph-based statistical analysis. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):158:1– 158:35, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3494970.

Gu:2017:SNI

[GZZ^{+}17] Weixi Gu, Yuxun Zhou, Zimu Zhou, Xi Liu, Han Zou, Pei Zhang, Costas J. Spanos, and Lin Zhang. SugarMate: Non-intrusive blood glucose monitoring with smartphones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1– 27, September 2017. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3130919.

Gao:2021:TPI

[GZZ^{+}21] Ruiyang Gao, Mi Zhang, Jie Zhang, Yang Li, Enze Yi, Dan Wu, Leye Wang, and Daqing Zhang. Towards position-independent sensing for gesture recognition with Wi-Fi. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):61:1– 61:28, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3463504.

Ge:2020:ASB

[GZZH20] Linfei Ge, Qian Zhang, Jin Zhang, and Qianyi Huang. Acoustic strength-based motion tracking. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):127:1– 127:19, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3432215.

Haladjian:2019:WDT

[Hal19] Juan Haladjian. The Wearables Development Toolkit: an integrated development environment for activity recognition applications. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 26, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3369813.

Huang:2020:IIC

[HAL^{+}20] Danny Yuxing Huang, Noah Apthorpe, Frank Li, Gunes Acar, and Nick Feamster.

IoT inspector: Crowdsourcing labeled network traffic from smart home devices at scale. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):46:1– 46:21, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397333.

Hossain:2018:DSA

[HAR18] H. M. Sajjad Hossain, MD Abdullah Al Haiz Khan, and Nirmalya Roy. DeActive: Scaling activity recognition with active deep learning. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–23, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214269.

Hsu:2017:ZEH

[HAY^{+}17] Chen-Yu Hsu, Aayush Ahuja, Shichao Yue, Rumen Hristov, Zachary Kabelac, and Dina Katabi. Zero-effort in-home sleep and insomnia monitoring using radio signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1– 18, September 2017. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3130924.

Hu:2020:SDH

[HBH20] Yuhan Hu, Sara Maria Bejarano, and Guy Hoffman. ShadowSense: Detecting human touch in a social robot using shadow image classification. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):132:1– 132:24, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3432202.

Huynh:2021:IAB

[HBK21] Sinh Huynh, Rajesh Krishna Balan, and JeongGil Ko. iMon: Appearancebased gaze tracking system on mobile devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):161:1– 161:26, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3494999.

Hao:2017:MSB

[HBX^{+}17] Tian Hao, Chongguang Bi, Guoliang Xing, Roxane Chan, and Linlin Tu. MindfulWatch: a smartwatchbased system for real-time respiration monitoring during meditation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1–19, September 2017. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3130922.

He:2022:FWE

[HCH^{+}22] Yan He, George Chernyshov, Jiawen Han, Dingding Zheng, Ragnar Thomsen, Danny Hynds, Muyu Liu, Yuehui Yang, Yulan Ju, Yun Suen Pai, Kouta Minamizawa, Kai Kunze, and Jamie A. Ward. Frisson waves: Exploring automatic detection, triggering and sharing of aesthetic chills in music performances. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):118:1– 118:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550324.

Huang:2018:BLD

[HCYZ18] Chenyu Huang, Huangxun Chen, Lin Yang, and Qian Zhang. BreathLive: Liveness detection for heart sound authentication with deep breathing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–25, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3191744.

Hong:2018:CNI

[HDC18] Hande Hong, Girisha Durrel De Silva, and Mun Choon Chan. CrowdProbe: Noninvasive crowd monitoring with Wi-Fi probe. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1– 23, September 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3264925.

He:2022:EFA

[HDFZ22] Ke He, Yongjie Duan, Jianjiang Feng, and Jie Zhou. Estimating 3D finger angle via fingerprint image. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):14:1– 14:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3517243.

Hassan:2017:FEB

[HDW^{+}17] Mahmoud Hassan, Florian Daiber, Frederik Wiehr, Felix Kosmalla, and Antonio Krüger. FootStriker: An EMS-based foot strike assistant for running. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(1):2:1–2:18, March 2017. CODEN ???? ISSN 2474-9567. URL http:

//dl.acm.org/citation.
cfm?id=3053332.

Haresamudram:2021:CPC

[HEP21] Harish Haresamudram, Irfan Essa, and Thomas Plötz. Contrastive predictive coding for human activity recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):65:1– 65:26, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463506.

Haresamudram:2022:ASS

[HEP22] Harish Haresamudram, Irfan Essa, and Thomas Plötz. Assessing the state of selfsupervised human activity recognition using wearables. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3): 116:1–116:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550299.

Hamatani:2018:FGH

[HEUH18] Takashi Hamatani, Moustafa Elhamshary, Akira Uchiyama, and Teruo Higashino. FluidMeter: Gauging the human daily fluid intake using smartwatches. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1– 25, September 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3264923.

Hintze:2019:CUR

[HFS^{+}19] Daniel Hintze, Matthias Füller, Sebastian Scholz, Rainhard D. Findling, Muhammad Muaaz, Philipp Kapfer, Eckhard Koch, and René Mayrhofer. CORMORANT: Ubiquitous risk-aware multimodal biometric authentication across mobile devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 23, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3351243.

Han:2021:WWS

[HFX^{+}21] Zhenyu Han, Haohao Fu, Fengli Xu, Zhen Tu, Yang Yu, Pan Hui, and Yong Li. Who will survive and revive undergoing the epidemic: Analyses about POI visit behavior in Wuhan via check-in records. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):64:1– 64:20, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463525.

Hintze:2017:LSL

[HHFM17] Daniel Hintze, Philipp Hintze, Rainhard D. Findling, and René Mayrhofer. A large-scale, long-term analysis of mobile device usage characteristics. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090078.

Hautasaari:2019:VMS

[HHIF19] Ari Hautasaari, Takeo Hamada, Kuntaro Ishiyama, and Shogo Fukushima. VocaBura: a method for supporting second language vocabulary learning while walking. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1–23, December 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3369824.

Hashem:2018:CEP

[HHSM18] Tanzima Hashem, Rubaba Hasan, Flora Salim, and Mehnaz Tabassum Mahin. Crowd-enabled processing of trustworthy, privacyenhanced and personalised location based services with quality guarantee. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 25, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3287045.

Hu:2020:FCH

[HHX^{+}20] Fang Hu, Peng He, Songlin Xu, Yin Li, and Cheng Zhang. FingerTrak: Continuous 3D hand pose tracking by deep learning hand silhouettes captured by miniature thermal cameras on wrist. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):71:1– 71:24, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3397306.

Hu:2023:BCE

[HHZ23] Haiyan Hu, Qianyi Huang, and Qian Zhang. BabyNutri: a cost-effective baby food macronutrients analyzer based on spectral reconstruction. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):15:1– 15:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580858.

Henderson:2022:ICM

[HJL^{+}22] Jay Henderson, Tanya R.

Jonker, Edward Lank, Daniel Wigdor, and Ben Lafreniere. Investigating cross-modal approaches for evaluating error acceptability of a recognition-based input technique. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):15:1– 15:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517262.

Huynh:2018:EMM

[HKK^{+}18] Sinh Huynh, Seungmin Kim, JeongGil Ko, Rajesh Krishna Balan, and Youngki Lee. EngageMon: Multi-modal engagement sensing for mobile games. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1):

1–27, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3191745.

Huh:2023:LTE

[HKK^{+}23] Jun Ho Huh, Sungsu Kwag, Iljoo Kim, Alexandr Popov, Younghan Park, Geumhwan Cho, Juwon Lee, Hyoungshick Kim, and Choong-Hoon Lee. On the longterm effects of continuous keystroke authentication: Keeping user frustration low through behavior adaptation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):58:1– 58:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3596236.

Haliburton:2023:ESS

[HKSM23] Luke Haliburton, Saba Kheirinejad, Albrecht Schmidt, and Sven Mayer. Exploring smart standing desks to foster a healthier workplace. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2): 57:1–57:??, June 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3596260.

Ham:2017:QAF

[HLL17] Seongmin Ham, Jihyung Lee, and Kyunghan Lee. QuickTalk: an associationfree communication method for IoT devices in proximity. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1– 18, September 2017. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3130921.

He:2020:MDB

[HLS20] Liang He, Youngmoon Lee, and Kang G. Shin. Mobile device batteries as thermometers. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):12:1– 12:21, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3381015.

Ha:2021:WCS

[HMA21] Unsoo Ha, Sohrab Madani, and Fadel Adib. WiStress: Contactless stress monitoring using wireless signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3): 103:1–103:37, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478121.

Hafiz:2020:WCT

[HMM^{+}20] Pegah Hafiz, Kamilla Woznica Miskowiak, Alban Maxhuni, Lars Vedel Kessing, and Jakob Eyvind Bardram. Wearable computing technology for assessment of cognitive functioning of bipolar patients and healthy controls. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):129:1– 129:22, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3432219.

Hiremath:2022:BHA

[HNCP22] Shruthi K. Hiremath, Yasutaka Nishimura, Sonia Chernova, and Thomas Plötz. Bootstrapping human activity recognition systems for smart homes from scratch. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3): 119:1–119:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3550294.

He:2022:GLG

[HNL22] Tao He, Qun Niu, and Ning Liu. GC-Loc: a graph attention based framework for collaborative indoor localization using infrastructurefree signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):165:1– 165:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3569495.

Hiremath:2020:DEH

[HP20] Shruthi K. Hiremath and Thomas Plötz. Deriving effective human activity recognition systems through objective task complexity assessment. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):131:1– 131:24, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3432227.

Hollis:2018:BTH

[HPWW18] Victoria Hollis, Alon Pekurovsky, Eunika Wu, and Steve Whittaker. On being told how we feel: How algorithmic sensor feedback influences emotion perception. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1– 31, September 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3264924.

Huang:2023:NHN

[HQC^{+}23] Yandao Huang, Minghui Qiu, Lin Chen, Zhencan Peng, Qian Zhang, and Kaishun Wu. NF-Heart: a near-field non-contact continuous user authentication system via ballistocardiogram. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):16:1– 16:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3580851.

Han:2018:BEI

[HQY^{+}18] Jinsong Han, Chen Qian, Yuqin Yang, Ge Wang, Han Ding, Xin Li, and Kui Ren. Butterfly: Environmentindependent physical-layer authentication for passive RFID. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1–21, December 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3287044.

Haliburton:2023:FTR

[HSH^{+}23] Luke Haliburton, Svenja Yvonne Schött, Linda Hirsch, Robin Welsch, and Albrecht Schmidt.

Feeling the temperature of the room: Unobtrusive thermal display of engagement during group communication. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):14:1– 14:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580820.

Huh:2022:WTW

[HSK^{+}22] Jun Ho Huh, Hyejin Shin, HongMin Kim, Eunyong Cheon, Youngeun Song, Choong-Hoon Lee, and Ian Oakley. WristAcoustic: Through-wrist acoustic response based authentication for smartwatches. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):167:1– 167:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3569473.

He:2020:ADU

[HSL^{+}20] Liang He, Yuanchao Shu, Youngmoon Lee, Dongyao Chen, and Kang G. Shin. Authenticating drivers using automotive batteries. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):130:1– 130:27, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3432198.

Hirsch:2021:SSE

[HSWB21] Linda Hirsch, Christina Schneegass, Robin Welsch, and Andreas Butz. To see or not to see: Exploring inattentional blindness for the design of unobtrusive interfaces in shared public places. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1): 15:1–15:25, March 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3448123.

Holz:2017:GCS

[HW17] Christian Holz and Edward J. Wang. Glabella: Continuously sensing blood pressure behavior using an unobtrusive wearable device. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1–23, September 2017. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3132024.

Huang:2021:UPI

[HW21] Long Huang and Chen Wang. Unobtrusive pedestrian identification by leveraging footstep sounds with replay resistance. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):160:1– 160:19, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3494963.

He:2021:MDT

[HWJ^{+}21] Liang He, Jarrid A. Wittkopf, Ji Won Jun, Kris Erickson, and Rafael Tico Ballagas. ModElec: a design tool for prototyping physical computing devices using conductive 3D printing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4): 159:1–159:20, December 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3495000.

Hooda:2022:SSA

[HWJ^{+}22] Ashish Hooda, Matthew Wallace, Kushal Jhunjhunwalla, Earlence Fernandes, and Kassem Fawaz. SkillFence: a systems approach to practically mitigating voice-based confusion attacks. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):16:1– 16:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3517232.

Hu:2023:IPV

[HWJ^{+}23] Jiawei Hu, Yanxiang Wang, Hong Jia, Wen Hu, Mahbub Hassan, Brano Kusy, Ashraf Uddin, and Moustafa Youssef. Iris: Passive visible light positioning using light spectral information. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):97:1– 97:??, September 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3610913.

Huang:2019:AIA

[HWZZ19] Anna Huang, Dong Wang, Run Zhao, and Qian Zhang. Au-Id: Automatic user identification and authentication through the motions captured from sequential human activities using RFID. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1– 26, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3328919.

Hu:2018:LIP

[HXH^{+}18] Yiqing Hu, Yan Xiong, Wenchao Huang, Xiang-Yang Li, Panlong Yang, Yanan Zhang, and Xufei Mao. Lightitude: Indoor positioning using uneven light intensity distribution. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–25, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214270.

He:2018:MVC

[HY18] Zhiyuan He and Su Yang.

Multi-view commercial hotness prediction using contextaware neural network ensemble. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1–19, December 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3287046.

He:2022:IMU

[HY22] Jing He and Wei Yang.

IMar: Multi-user continuous action recognition with WiFi signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):117:1– 117:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550311.

Huang:2021:MHS

[HYH21] Huiqun Huang, Xi Yang, and Suining He. Multihead spatio-temporal attention mechanism for urban anomaly event prediction. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3): 104:1–104:21, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478099.

Huang:2023:IPM

[HYS^{+}23] Tian Huang, Chun Yu, Weinan Shi, Bowen Wang, David Yang, Yihao Zhu, Zhaoheng Li, and Yuanchun Shi. Interaction proxy manager: Semantic model generation and run-time support for reconstructing ubiquitous user interfaces of mobile services. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):99:1– 99:??, September 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3610929.

Hu:2023:MLS

[HYY^{+}23] Yongquan Hu, Hui-Shyong Yeo, Mingyue Yuan, Haoran Fan, Don Samitha Elvitigala, Wen Hu, and Aaron Quigley. Micro-Cam: Leveraging smartphone microscope camera for context-aware contact surface sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):98:1– 98:??, September 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3610921.

Intarasirisawat:2019:ETM

[IAE^{+}19] Jittrapol Intarasirisawat, Chee Siang Ang, Christos Efstratiou, Luke William Feidhlim Dickens, and Rupert Page. Exploring the touch and motion features in game-based cognitive assessments. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 25, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351245.

Islam:2023:SMA

[IB23] Md Touhidul Islam and Syed Masum Billah. SpaceX Mag: an automatic, scalable, and rapid space compactor for optimizing smartphone app interfaces for low-vision users. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):59:1– 59:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3596253.

Inoue:2019:IAR

[ILH^{+}19] Sozo Inoue, Paula Lago, Tahera Hossain, Tittaya Mairittha, and Nattaya Mairittha. Integrating activity recognition and nursing care records: The system, deployment, and a verification study. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1–24, September 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3351244.

Istomin:2021:JDR

[ILM^{+}21] Timofei Istomin, Elia Leoni, Davide Molteni, Amy L. Murphy, Gian Pietro Picco, and Maurizio Griva. Janus: Dual-radio accurate and energy-efficient proximity detection. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):162:1– 162:33, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3494978.

Iwamoto:2018:PCW

[IMO^{+}18] Eiichi Iwamoto, Masaki Matsubara, Chihiro Ota, Satoshi Nakamura, Tsutomu Terada, Hiroyuki Kitagawa, and Atsuyuki Morishima. Passerby crowdsourcing: Workers’ behavior and data quality management. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1–20, December 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287047.

Islam:2020:ZTS

[IN20] Bashima Islam and Shahriar Nirjon. Zygarde: Timesensitive on-device deep inference and adaptation on intermittently-powered systems. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):82:1–82:29, September 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3411808.

Islam:2021:BDR

[IRA^{+}21] Bashima Islam, Md Mahbubur Rahman, Tousif Ahmed, Mohsin Yusuf Ahmed, Md Mehedi Hasan, Viswam Nathan, Korosh Vatanparvar, Ebrahim Nemati, Jilong Kuang, and Jun Alex Gao. Breath-Track: Detecting regular breathing phases from unannotated acoustic data captured by a smartphone. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):124:1– 124:22, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478123.

Ikeda:2020:SMS

[ISSK20] Natsuki Ikeda, Ryo Shigeta, Junichiro Shiomi, and Yoshihiro Kawahara. Soilmonitoring sensor powered by temperature difference between air and shallow underground soil. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):13:1– 13:22, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3380995.

Islam:2021:SSL

[ISY^{+}21] Md. Rabiul Islam, Shuji Sakamoto, Yoshihiro Yamada, Andrew W. Vargo, Motoi Iwata, Masakazu Iwamura, and Koichi Kise. Selfsupervised learning for reading activity classification. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3): 105:1–105:22, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478088.

Ikematsu:2020:SEI

[IY20] Kaori Ikematsu and Shota Yamanaka. ScraTouch: Extending interaction technique using fingernail on unmodified capacitive touch surfaces. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):81:1–81:19, September 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3411831.

Isomoto:2022:DSM

[IYS22] Toshiya Isomoto, Shota Yamanaka, and Buntarou Shizuki. Dwell selection with ML-based intent prediction using only gaze data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3): 120:1–120:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550301.

Jiang:2021:DCP

[JCF^{+}21] Chutian Jiang, Yanjun Chen, Mingming Fan, Liuping Wang, Luyao Shen, Nianlong Li, Wei Sun, Yu Zhang, Feng Tian, and Teng Han. Douleur: Creating pain sensation with chemical stimulant to enhance user experience in virtual reality. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):66:1– 66:26, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463527.

Jiang:2020:SUD

[JCL^{+}20] Hongbo Jiang, Hangcheng Cao, Daibo Liu, Jie Xiong, and Zhichao Cao. SmileAuth: Using dental edge biometrics for user authentication on smartphones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):84:1– 84:24, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3411806.

Jansen:2022:DSH

[JCR22] Pascal Jansen, Mark Colley, and Enrico Rukzio. A design space for human sensor and actuator focused in-vehicle interaction based on a systematic literature review. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2): 56:1–56:??, July 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534617.

Jiang:2021:AAO

[JDZ^{+}21] Dongzhe Jiang, Yi Ding, Hao Zhang, Yunhuai Liu, Tian He, Yu Yang, and Desheng Zhang. AL-WAES: an automatic outdoor location-aware correction system for online delivery platforms. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):107:1– 107:24, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478081.

Jain:2019:BCE

[JGSC19] Milan Jain, Mridula Gupta, Amarjeet Singh, and Vikas Chandan. Beyond control: Enabling smart thermostats for leakage detection. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–21, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314401.

Jin:2022:EHY

[JGX^{+}22] Yincheng Jin, Yang Gao, Xuhai Xu, Seokmin Choi, Jiyang Li, Feng Liu, Zhengxiong Li, and Zhanpeng Jin. EarCommand: “Hearing” your silent speech commands in ear. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):57:1– 57:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534613.

Jin:2021:SAB

[JGZ^{+}21] Yincheng Jin, Yang Gao, Yanjun Zhu, Wei Wang, Jiyang Li, Seokmin Choi, Zhangyu Li, Jagmohan Chauhan, Anind K. Dey, and Zhanpeng Jin. SonicASL: an acoustic-based sign language gesture recognizer using earphones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):67:1– 67:30, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463519.

Janaka:2023:GTU

[JGZ^{+}23] Nuwan Janaka, Jie Gao, Lin Zhu, Shengdong Zhao, Lan Lyu, Peisen Xu, Maximilian Nabokow, Silang Wang, and Yanch Ong. GlassMessaging: Towards ubiquitous messaging using OHMDs.

Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3): 100:1–100:??, September 2023. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3610931.

Jia:2021:SUF

[JHH21] Hong Jia, Jiawei Hu, and Wen Hu. SwingNet: Ubiquitous fine-grained swing tracking framework via stochastic neural architecture search and adversarial learning. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):106:1– 106:21, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478082.

Jones:2021:DLA

[JHK^{+}21] Simon L. Jones, William Hue, Ryan M. Kelly, Rosemarie Barnett, Violet Henderson, and Raj Sengupta. Determinants of longitudinal adherence in smartphone-based selftracking for chronic health conditions: Evidence from axial spondyloarthritis. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):16:1– 16:24, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3448093.

Jiang:2021:DFG

[JHL^{+}21] Hongbo Jiang, Jingyang Hu, Daibo Liu, Jie Xiong, and Mingjie Cai. Driver-Sonar: Fine-grained dangerous driving detection using active sonar. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):108:1– 108:22, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478084.

Johnson:2017:BWC

[JHP^{+}17] I. Johnson, J. Henderson, C. Perry, J. Schöning, and B. Hecht. Beautiful … but at what cost?: An examination of externalities in geographic vehicle routing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

                                        //dl.acm.org/citation.
                                        cfm?id=3090080.

Jin:2022:SIG

[JHWF22] Xiaofu Jin, Xiaozhu Hu, Xiaoying Wei, and Mingming Fan. Synapse: Interactive guidance by demonstration with trial-and-error support for older adults to use smartphone apps. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):121:1– 121:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550321.

Jiokeng:2022:WKY

[JJB22] Kevin Jiokeng, Gentian Jakllari, and André-Luc Beylot. I want to know your hand: Authentication on commodity mobile phones based on your hand’s vibrations. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):58:1– 58:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534575.

Jain:2018:FCA

[JKB^{+}18] Mohit Jain, Pratyush Kumar, Ishita Bhansali, Q. Vera Liao, Khai Truong, and Shwetak Patel. FarmChat: a conversational agent to answer Farmer queries. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 22, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3287048.

Jeong:2017:SWB

[JKK^{+}17] Hayeon Jeong, Heepyung Kim, Rihun Kim, Uichin Lee, and Yong Jeong. Smartwatch wearing behavior analysis: a longitudinal study. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1–31, September 2017. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3131892.

Jin:2018:WTC

[JLD^{+}18] Haojian Jin, Minyi Liu, Kevan Dodhia, Yuanchun Li, Gaurav Srivastava, Matthew Fredrikson, Yuvraj Agarwal, and Jason I. Hong. Why are they collecting my data?: Inferring the purposes of network traffic in mobile apps. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 27, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3287051.

Jiang:2023:DDC

[JLZ^{+}23] Zhihan Jiang, Lin Lin, Xinchen Zhang, Jianduo Luan, Running Zhao, Longbiao Chen, James Lam, Ka-Man Yip, Hung-Kwan So, Wilfred H. S. Wong, Patrick Ip, and Edith C. H. Ngai. A data-driven context-aware health inference system for children during school closures. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):18:1– 18:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580800.

Jayarajah:2018:PEN

[JM18] Kasthuri Jayarajah and Archan Misra. Predicting episodes of non-conformant mobility in indoor environments. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1–24, December 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287050.

Jin:2021:CSB

[JMC^{+}21] Wenqiang Jin, Srinivasan Murali, Youngtak Cho, Huadi Zhu, Tianhao Li, Rachael Thompson Panik, Anika Rimu, Shuchisnigdha Deb, Kari Watkins, Xu Yuan, and Ming Li. CycleGuard: a smartphone-based assistive tool for cyclist safety using acoustic ranging. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):163:1– 163:30, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3494992.

Jiang:2019:PSC

[JMvB^{+}19] Weiwei Jiang, Gabriele Marini, Niels van Berkel, Zhanna Sarsenbayeva, Zheyu Tan, Chu Luo, Xin He, Tilman Dingler, Jorge Goncalves, Yoshihiro Kawahara, and Vassilis Kostakos. Probing sucrose contents in everyday drinks using miniaturized near-infrared spectroscopy scanners. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 25, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3369834.

Janveja:2020:IMS

[JNB^{+}20] Ishani Janveja, Akshay Nambi, Shruthi Bannur, Sanchit Gupta, and Venkat Padmanabhan. InSight: Monitoring the state of the driver in low-light using smartphones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):83:1– 83:29, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3411819.

Jain:2017:PUS

[JSC17] Milan Jain, Amarjeet Singh, and Vikas Chandan. Portable+: A ubiquitous and smart way towards comfortable energy savings. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090079.

Jakobi:2018:ENI

[JSC^{+}18] Timo Jakobi, Gunnar Stevens, Nico Castelli, Corinna Ogonowski, Florian Schaub, Nils Vindice, Dave Randall, Peter Tolmie, and Volker Wulf. Evolving needs in IoT control and accountability: a longitudinal study on smart home intelligibility. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 28, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287049.

Jiang:2018:DRB

[JSF^{+}18] Renhe Jiang, Xuan Song, Zipei Fan, Tianqi Xia, Quanjun Chen, Qi Chen, and Ryosuke Shibasaki. Deep ROI-based modeling for urban human mobility prediction. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–29, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3191746.

Jeyakumar:2023:XCC

[JSGS23] Jeya Vikranth Jeyakumar, Ankur Sarker, Luis Antonio Garcia, and Mani Srivastava. X-CHAR: a concept-based explainable complex human activity recognition model. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):17:1– 17:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3580804.

Jiang:2023:FRM

[JSNL23] Weina Jiang, Lin Shi, Qun Niu, and Ning Liu. Fast radio map construction with domain disentangled learning for wireless localization. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3): 101:1–101:??, September 2023. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3610922.

Jain:2022:CCS

[JTM^{+}22] Yash Jain, Chi Ian Tang, Chulhong Min, Fahim Kawsar, and Akhil Mathur. ColloSSL: Collaborative selfsupervised learning for human activity recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1): 17:1–17:??, March 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3517246.

Jackson:2018:APP

[JW18] Corey Brian Jackson and Yang Wang. Addressing the privacy paradox through personalized privacy notifications. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–25, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214271.

Jung:2020:FSE

[JWCC^{+}20] Jisu Jung, Lyndal Wellard-Cole, Colin Cai, Irena Koprinska, Kalina Yacef, Margaret Allman-Farinelli, and Judy Kay. Foundations for systematic evaluation and benchmarking of a mobile food logger in a large-scale nutrition study. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):47:1– 47:25, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397327.

Jung:2021:LSQ

[JWK^{+}21] Woosub Jung, Amanda Watson, Scott Kuehn, Erik Korem, Ken Koltermann, Minglong Sun, Shuangquan Wang, Zhenming Liu, and Gang Zhou. LAX-Score: Quantifying team performance in lacrosse and exploring IMU features towards performance enhancement. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):109:1– 109:28, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478076.

Jiang:2023:IEI

[JWS^{+}23] Weiwei Jiang, Chaofan Wang, Zhanna Sarsenbayeva, Andrew Irlitti, Jing Wei, Jarrod Knibbe, Tilman Dingler, Jorge Goncalves, and Vassilis Kostakos. InfoPrint: Embedding interactive information in 3D prints using low-cost readily-available printers and materials. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):102:1– 102:??, September 2023. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3610933.

Jin:2020:AAA

[JXZ^{+}20] Wenqiang Jin, Mingyan Xiao, Huadi Zhu, Shuchisnigdha Deb, Chen Kan, and Ming Li. Acoussist: an acoustic assisting tool for people with visual impairments to cross uncontrolled streets. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):133:1– 133:30, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3432216.

Jones:2021:CWP

[JYY21] Jasmine Jones, Ye Yuan, and Svetlana Yarosh. Be consistent, work the program, be present every day: Exploring technologies for self-tracking in early recovery. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):164:1– 164:26, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3494955.

Jin:2023:SPC

[JZG^{+}23] Yincheng Jin, Shibo Zhang, Yang Gao, Xuhai Xu, Seokmin Choi, Zhengxiong Li, Henry J. Adler, and Zhanpeng Jin. SmartASL: “Point-of-Care” comprehensive ASL interpreter using wearables. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):60:1– 60:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3596255.

Ji:2019:DRL

[JZWL19] Shenggong Ji, Yu Zheng, Zhaoyuan Wang, and Tianrui Li. A deep reinforcement learning-enabled dynamic redeployment system for mobile ambulances. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–20, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314402.

Knibbe:2018:EEM

[KAH18] J. Knibbe, A. Alsmith, and K. Hornbæk. Experiencing electrical muscle stimulation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1–14, September 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264928.

Kim:2021:RTB

[KASC21] Daeyong Kim, Junick Ahn, Jun Shin, and Hojung Cha. Ray tracing-based light energy prediction for indoor batteryless sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):17:1– 17:27, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3448086.

Khalfoun:2021:EEL

[KBBN21] Besma Khalfoun, Sonia Ben Mokhtar, Sara Bouchenak, and Vlad Nitu. EDEN: Enforcing location privacy through re-identification risk assessment: a federated learning approach. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):68:1– 68:25, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3463502.

Kortbeek:2020:BEB

[KBC^{+}20] Vito Kortbeek, Abu Bakar, Stefany Cruz, Kasim Sinan Yildirim, Przemysław Pawełczak, and Josiah Hester. BFree: Enabling battery-free sensor prototyping with Python. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4): 135:1–135:39, December 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3432191.

Kumar:2020:AUC

[KBKH20] Abhishek Kumar, Tristan Braud, Young D. Kwon, and Pan Hui. Aquilis: Using contextual integrity for privacy protection on mobile devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):137:1– 137:28, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3432205.

Kao:2018:SFS

[KBL18] Hsin-Liu Cindy Kao, Abdelkareem Bedri, and Kent Lyons. SkinWire: Fabricating a self-contained on-skin PCB for the hand. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1– 23, September 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264926.

Kalupahana:2023:SLI

[KBXP23] Ayanga Imesha Kumari Kalupahana, Ananta Narayanan Balaji, Xiaokui Xiao, and Li-Shiuan Peh. SeRaNDiP: Leveraging inherent sensor random noise for differential privacy preservation in wearable community sensing applications. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):61:1– 61:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596252.

Kim:2019:VTD

[KCFI19] Lawrence H. Kim, Pablo Castillo, Sean Follmer, and Ali Israr. VPS tactile display: Tactile information transfer of vibration, pressure, and shear. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–17, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3328922.

Kim:2019:HCT

[KCK^{+}19] Yoonji Kim, Youngkyung Choi, Daye Kang, Minkyeong Lee, Tek-Jin Nam, and Andrea Bianchi. HeyTeddy: Conversational test-driven development for physical computing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1–21, December 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3369838.

Kim:2018:IDI

[KCP^{+}18] Auk Kim, Woohyeok Choi, Jungmi Park, Kyeyoon Kim, and Uichin Lee. Interrupting drivers for interactions: Predicting opportune moments for in-vehicle proactive auditory-verbal tasks. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 28, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3287053.

Kwon:2019:GUE

[KCuH^{+}19] Young D. Kwon, Dimitris Chatzopoulos, Ehsan ul Haq, Raymond Chi-Wing Wong, and Pan Hui. GeoLifecycle: User engagement of geographical exploration and churn prediction in LBSNs. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 29, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351250.

Kolamunna:2021:DAS

[KDL^{+}21] Harini Kolamunna, Thilini Dahanayaka, Junye Li, Suranga Seneviratne, Kanchana Thilakaratne, Albert Y. Zomaya, and Aruna Seneviratne. DronePrint: Acoustic signatures for open-set drone detection and identification with online data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):20:1– 20:31, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3448115.

Kumar:2018:RFW

[KEI^{+}18] Sumeet Kumar, Hakan Erdogmus, Bob Iannucci, Martin Griss, and João Diogo Falcão. Rethinking the future of wireless emergency alerts: a comprehensive study of technical and conceptual improvements. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–33, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3214274.

Kwon:2018:CSS

[KFFC18] Hyosun Kwon, Joel E. Fischer, Martin Flintham, and James Colley. The connected shower: Studying intimate data in everyday life. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 22, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3287054.

Kraemer:2022:BFM

[KGAH22] Christopher Kraemer, Amy Guo, Saad Ahmed, and Josiah Hester. Batteryfree MakeCode: Accessible programming for intermittent computing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):18:1– 18:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3517236.

Klakegg:2018:AMM

[KGL^{+}18] Simon Klakegg, Jorge Goncalves, Chu Luo, Aku Visuri, Alexey Popov, Niels van Berkel, Zhanna Sarsenbayeva, Vassilis Kostakos, Simo Hosio, Scott Savage, Alexander Bykov, Igor Meglinski, and Denzil Ferreira. Assisted medication management in elderly care using miniaturised nearinfrared spectroscopy. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–24, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214272.

Khurana:2019:DSM

[KGL19] Rushil Khurana, Mayank Goel, and Kent Lyons. Detachable smartwatch: More than a wearable. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–14, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328921.

Kim:2020:UUC

[KGN^{+}20] Inyeop Kim, Hwarang Goh, Nematjon Narziev, Youngtae Noh, and Uichin Lee. Understanding user contexts and coping strategies for context-aware phone distraction management system design. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):134:1– 134:33, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3432213.

Kiaghadi:2019:PPS

[KHG^{+}19] Ali Kiaghadi, Seyedeh Zohreh Homayounfar, Jeremy Gummeson, Trisha Andrew, and Deepak Ganesan. Phyjama: Physiological sensing via fiber-enhanced pyjamas. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 29, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351247.

Kalyanaraman:2017:FTT

[KHSW17]  Avinash  Kalyanaraman,

Dezhi Hong, Elahe Soltanaghaei, and Kamin Whitehouse. Forma Track: Tracking people based on body shape. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1– 21, September 2017. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3130926.

Ku:2023:SBC

[KHW^{+}23] Pin-Sung Ku, Kunpeng Huang, Nancy Wang, Boaz Ng, Alicia Chu, and Hsin-Liu Cindy Kao. Skin-Link: On-body construction and prototyping of reconfigurable epidermal interfaces. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2): 62:1–62:??, June 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596241.

Kang:2022:AAL

[KHZ22] Hua Kang, Qianyi Huang, and Qian Zhang. Augmented adversarial learning for human activity recognition with partial sensor sets. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3): 122:1–122:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550285.

Kayukawa:2020:GBP

[KIT^{+}20] Seita Kayukawa, Tatsuya Ishihara, Hironobu Takagi, Shigeo Morishima, and Chieko Asakawa. Guiding blind pedestrians in public spaces by understanding walking behavior of nearby pedestrians. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):85:1– 85:22, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3411825.

Kim:2019:GEI

[KJKL19] Jaejeung Kim, Hayoung Jung, Minsam Ko, and Uichin Lee. GoalKeeper: Exploring interaction lockout mechanisms for regulating smartphone use. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–29, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314403.

Kim:2019:TDR

[KJP^{+}19] Yoojung Kim, Hee-Tae Jung, Joonwoo Park, Yangsoo Kim, Nathan Ramasarma, Paolo Bonato, Eun Kyoung Choe, and Sunghoon Ivan Lee. Towards the design of a ring sensor-based mHealth system to achieve optimal motor function in stroke survivors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1–26, December 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369817.

Katanbaf:2020:RRB

[KJS20] Mohamad Katanbaf, Vivek Jain, and Joshua R. Smith. Relacks: Reliable backscatter communication in indoor environments. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):48:1– 48:24, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397314.

Kim:2021:BNI

[KKL21] Minhyung Kim, Inyeop Kim, and Uichin Lee. Beneficial neglect: Instant message notification handling behaviors and academic performance. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):18:1– 18:26, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3448089.

Kabir:2022:AUC

[KKV^{+}22] Kazi Sinthia Kabir, Stacey A.

Kenfield, Erin L. Van Blarigan, June M. Chan, and Jason Wiese. Ask the users: a case study of leveraging user-centered design for designing just-in-time adaptive interventions (JITAIs). Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2): 59:1–59:??, July 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534612.

Karkar:2018:BDP

[KKZ^{+}18] Ravi Karkar, Rafal Kocielnik, Xiaoyi Zhang, Jasmine Zia, George N. Ioannou, Sean A. Munson, and James Fogarty. Beacon: Designing a portable device for self-administering a measure of critical flicker frequency. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1– 27, September 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3264927.

Kuzminykh:2019:HMT

[KL19] Anastasia Kuzminykh and Edward Lank. How much is too much?: Understanding the information needs of parents of young children. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1– 21, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3328923.

Kim:2020:DME

[KLK^{+}20] Wonjung Kim, Seungchul Lee, Seonghoon Kim, Sungbin Jo, Chungkuk Yoo, Inseok Hwang, Seungwoo Kang, and Junehwa Song. Dyadic mirror: Everyday second-person live-view for empathetic reflection upon parent-child interaction. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3): 86:1–86:29, September 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3411815.

Kamminga:2018:RSO

[KLM^{+}18] Jacob W. Kamminga, Duc V.

Le, Jan Pieter Meijers, Helena Bisby, Nirvana Meratnia, and Paul J. M. Havinga. Robust sensor-orientation-independent feature selection for animal activity recognition on collar tags. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–27, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191747.

Kianpisheh:2019:FRA

[KLT19] Mohammad Kianpisheh, Franklin Mingzhe Li, and Khai N. Truong. Face recognition assistant for people with visual impairments. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 24, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3351248.

Kandappu:2018:OSP

[KMC^{+}18] Thivya Kandappu, Archan Misra, Shih-Fen Cheng, Randy Tandriansyah, and Hoong Chuin Lau. Obfuscation at-source: Privacy in context-aware mobile crowd-sourcing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–24, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3191748.

King:2019:MSE

[KME^{+}19] Zachary D. King, Judith Moskowitz, Begum Egilmez, Shibo Zhang, Lida Zhang, Michael Bass, John Rogers, Roozbeh Ghaffari, Laurie Wakschlag, and Nabil Alshurafa. micro-Stress EMA: a passive sensing framework for detecting in-the-wild stress in pregnant mothers. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 22, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351249.

Ku:2021:SCK

[KMH^{+}21] Pin-Sung Ku, Md. Tahmidul Islam Molla, Kunpeng Huang, Priya Kattappurath, Krithik Ranjan, and Hsin-Liu Cindy Kao. SkinKit: Construction kit for on-skin interface prototyping. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):165:1– 165:23, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3494989.

Kunzler:2019:ESR

[KMK^{+}19] Florian Künzler, Varun Mishra, Jan-Niklas Kramer, David Kotz, Elgar Fleisch, and Tobias Kowatsch. Exploring the state-of-receptivity for mHealth interventions. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 27, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3369805.

Koch:2021:WDD

[KML^{+}21] Kevin Koch, Varun Mishra, Shu Liu, Thomas Berger, Elgar Fleisch, David Kotz, and Felix Wortmann. When do drivers interact with invehicle well-being interventions?: an exploratory analysis of a longitudinal study on public roads. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):19:1– 19:30, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3448116.

Kim:2023:PPM

[KNL^{+}23] Jiha Kim, Younho Nam, Jungeun Lee, Young-Joo Suh, and Inseok Hwang. ProxiFit: Proximity magnetic sensing using a single commodity mobile toward holistic weight exercise monitoring. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):105:1– 105:??, September 2023. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3610920.

Khan:2017:ARQ

[KNP17] Aftab Khan, James Nicholson, and Thomas Plötz. Activity recognition for quality assessment of batting shots in cricket using a hierarchical representation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):1– 31, September 2017. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3130927.

Khaloo:2019:NNO

[KOY^{+}19] Pooya Khaloo, Brandon Oubre, Jeremy Yang, Tauhidur Rahman, and Sunghoon Ivan Lee. NOSE: a novel odor sensing engine for ambient monitoring of the frying cooking method in kitchen environments. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–25, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328920.

Krasovec:2020:QYT

[KPG^{+}20] Andraz Krasovec, Daniel Pellarini, Dimitrios Geneiatakis, Gianmarco Baldini, and Veljko Pejović. Not quite yourself today: Behaviourbased continuous authentication in IoT environments.

Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4): 136:1–136:29, December 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3432206.

Kwak:2018:EEL

[KPK^{+}18] Myeongcheol Kwak, Youngmong Park, Junyoung Kim, Jinyoung Han, and Taekyoung Kwon. An energyefficient and lightweight indoor localization system for Internet-of-Things (IoT) environments. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–28, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191749.

Kim:2020:IVM

[KPL20] Auk Kim, Jung-Mi Park, and Uichin Lee. Interruptibility for in-vehicle multitasking: Influence of voice task demands and adaptive behaviors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):14:1– 14:22, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3381009.

Khan:2022:CMP

[KRS22] Usman Mahmood Khan, Luca Rigazio, and Muhammad Shahzad. Contactless monitoring of PPG using radar. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):123:1– 123:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550330.

Kucera:2017:TCD

[KSC^{+}17] Jan Kucera, James Scott, Nicholas Chen, Patrick Olivier, and Steve Hodges. Towards calm displays: Matching ambient illumination in bedrooms. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090081.

Kakaraparthi:2021:FSF

[KSC^{+}21] Vimal Kakaraparthi, Qijia Shao, Charles J. Carver, Tien Pham, Nam Bui, Phuc Nguyen, Xia Zhou, and Tam Vu. FaceSense: Sensing face touch with an earworn system. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):110:1– 110:27, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478129.

Kapetanovic:2019:GOO

[KSCS19] Zerina Kapetanovic, Ali Saffari, Ranveer Chandra, and Joshua R. Smith. Glaze: Overlaying occupied spectrum with downlink IoT transmissions. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1–21, December 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369825.

Kalanadhabhatta:2022:EMS

[KSZ^{+}22] Manasa Kalanadhabhatta, Adrelys Mateo Santana, Zhongyang Zhang, Deepak Ganesan, Adam S. Grabell, and Tauhidur Rahman. EarlyScreen: Multiscale instance fusion for predicting neural activation and psychopathology in preschool children. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):60:1– 60:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534583.

Kuoppamaki:2021:DKT

[KTEM21] Sanna Kuoppamäki, Sylvaine Tuncer, Sara Eriksson, and Donald McMillan. Designing kitchen technologies for ageing in place: a video study of older adults’ cooking at home. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):69:1– 69:19, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463516.

Kwon:2020:IAE

[KTH^{+}20] Hyeokhyen Kwon, Catherine Tong, Harish Haresamudram, Yan Gao, Gregory D. Abowd, Nicholas D. Lane, and Thomas Plötz. IMU-Tube: Automatic extraction of virtual on-body accelerometry from video for human activity recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3): 87:1–87:29, September 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3411841.

Khamis:2018:CCT

[KTM^{+}18] Mohamed Khamis, Ludwig Trotter, Ville Mäkelä, Emanuel von Zezschwitz, Jens Le, Andreas Bulling, and Florian Alt. CueAuth: Comparing touch, mid-air gestures, and gaze for cuebased authentication on situated displays. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1–22, December 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287052.

Khwaja:2019:MPV

[KVZ^{+}19] Mohammed Khwaja, Sumer S. Vaid, Sara Zannone, Gabriella M. Harari, A. Aldo Faisal, and Aleksandar Matic. Modeling personality vs. modeling personalidad: In-the-wild mobile data analysis in five countries suggests cultural impact on personality models. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 24, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351246.

Kabir:2023:MSB

[KW23] Kazi Sinthia Kabir and Jason Wiese. A meta-synthesis of the barriers and facilitators for personal informatics systems. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):103:1– 103:??, September 2023. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3610893.

Kwon:2021:ARW

[KWAP21] Hyeokhyen Kwon, Bingyao Wang, Gregory D. Abowd, and Thomas Plötz. Approaching the real-world: Supporting activity recognition training with virtual IMU data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):111:1– 111:32, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478096.

Kocielnik:2018:RCC

[KXAH18] Rafal Kocielnik, Lillian Xiao, Daniel Avrahami, and Gary Hsieh. Reflection companion: a conversational system for engaging users in reflection on physical activity. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–26, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214273.

Kim:2023:WWE

[KYJ^{+}23] Gwangbin Kim, Dohyeon Yeo, Taewoo Jo, Daniela Rus, and SeungJun Kim. What and when to explain?: On-road evaluation of explanations in highly automated vehicles. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):104:1– 104:??, September 2023. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3610886.

Liaqat:2019:WRW

[LAAE^{+}19] Daniyal Liaqat, Mohamed Abdalla, Pegah Abed-Esfahani, Moshe Gabel, Tatiana Son, Robert Wu, Andrea Gershon, Frank Rudzicz, and Eyal De Lara. WearBreathing: Real world respiratory rate monitoring using smartwatches. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–22, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3328927.

Li:2018:CIP

[LAH18] Tianshi Li, Yuvraj Agarwal, and Jason I. Hong. Coconut: an IDE plugin for developing privacy-friendly apps. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 35, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287056.

Langevin:2019:PFA

[LAS^{+}19] Raina Langevin, Mohammad Rafayet Ali, Taylan Sen, Christopher Snyder, Taylor Myers, E. Ray Dorsey, and Mohammed Ehsan Hoque. The PARK framework for automated analysis of Parkinson’s disease characteristics. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–22, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328925.

Lyu:2023:EDM

[LAX^{+}23] Yue Lyu, Pengcheng An, Yage Xiao, Zibo Zhang, Huan Zhang, Keiko Katsuragawa, and Jian Zhao. Eggly: Designing mobile augmented reality neurofeedback training games for children with autism spectrum disorder. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):67:1– 67:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3596251.

Lakshmanan:2022:PRW

[LBC^{+}22] Nitya Lakshmanan, Abdelhak Bentaleb, Byoungjun Choi, Roger Zimmermann, Jun Han, and Min Suk Kang. On privacy risks of watching YouTube over cellular networks with carrier aggregation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):19:1– 19:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3517261.

Li:2022:RDF

[LCC^{+}22] Zhengxiong Li, Baicheng Chen, Xingyu Chen, Chenhan Xu, Yuyang Chen, Feng Lin, Changzhi Li, Karthik Dantu, Kui Ren, and Wenyao Xu. Reliable digital forensics in the air: Exploring an RF-based drone identification system. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2): 63:1–63:??, July 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3534598.

Li:2019:FWC

[LCFT19] Franklin Mingzhe Li, Di Laura Chen, Mingming Fan, and Khai N. Truong. FMT: a wearable camera-based object tracking memory aid for older adults. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1–25, September 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351253.

Liao:2022:WKW

[LCG^{+}22] Chengwu Liao, Chao Chen, Suiming Guo, Zhu Wang, Yaxiao Liu, Ke Xu, and Daqing Zhang. Wheels know why you travel: Predicting trip purpose via a dual-attention graph embedding network. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):22:1– 22:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517239.

Liu:2022:MEE

[LCH^{+}22] Haipeng Liu, Kening Cui, Kaiyuan Hu, Yuheng Wang, Anfu Zhou, Liang Liu, and Huadong Ma. mTransSee: Enabling environment-independent mmWave sensing based gesture recognition via transfer learning. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):23:1– 23:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517231.

Liu:2017:GAU

[LCL17] Can Liu, Gradeigh D. Clark, and Janne Lindqvist. Guessing attacks on usergenerated gesture passwords. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(1):3:1–3:24, March 2017. CODEN ???? ISSN 2474-9567. URL http:

//dl.acm.org/citation.
cfm?id=3053331.

Lee:2020:WIC

[LCL20] Sunok Lee, Minji Cho, and Sangsu Lee. What if conversational agents became invisible?: Comparing users’ mental models according to physical entity of AI speaker. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):88:1– 88:24, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3411840.

Liu:2018:VRU

[LCR^{+}18] Rui Liu, Cory Cornelius, Reza Rawassizadeh, Ronald Peterson, and David Kotz. Vocal resonance: Using internal body voice for wearable authentication. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–23, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191751.

Long:2021:VFD

[LCR^{+}21] Yan Long, Alexander Curtiss, Sara Rampazzi, Josiah Hester, and Kevin Fu. VeriMask: Facilitating decontamination of N95 masks in the COVID-19 pandemic: Challenges, lessons learned, and safeguarding the future. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3): 119:1–119:29, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478105.

Li:2021:CAL

[LCS^{+}21] Xinyi Li, Liqiong Chang, Fangfang Song, Ju Wang, Xiaojiang Chen, Zhanyong Tang, and Zheng Wang. CrossGR: Accurate and lowcost cross-target gesture recognition using Wi-Fi. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1): 21:1–21:23, March 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3448100.

Li:2021:DWB

[LCW^{+}21] Hang Li, Xi Chen, Ju Wang, Di Wu, and Xue Liu. DAFI: WiFi-based device-free indoor localization via domain adaptation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):167:1– 167:21, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3494954.

Li:2022:NAB

[LCX^{+}22] Huining Li, Huan Chen, Chenhan Xu, Zhengxiong Li, Hanbin Zhang, Xiaoye Qian, Dongmei Li, Ming chun Huang, and Wenyao Xu. NeuralGait: Assessing brain health using your smartphone. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):169:1– 169:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3569476.

Liu:2021:DDA

[LCZ^{+}21] Bingyan Liu, Yifeng Cai, Ziqi Zhang, Yuanchun Li, Leye Wang, Ding Li, Yao Guo, and Xiangqun Chen. DistFL: Distribution-aware federated learning for mobile scenarios. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):168:1– 168:26, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3494966.

Lin:2018:CLC

[LDC18] Yuxiang Lin, Wei Dong, and Yuan Chen. Calibrating low-cost sensors by a two-phase learning approach for urban air quality measurement. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–18, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191750.

Luo:2020:DDD

[LDK^{+}20] Man Luo, Bowen Du, Konstantin Klemmer, Hongming Zhu, Hakan Ferhatosmanoglu, and Hongkai Wen. D3P: Data-driven demand prediction for fast expanding electric vehicle sharing systems. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):21:1– 21:21, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3381005.

Liao:2018:JTT

[LDS^{+}18] Peng Liao, Walter Dempsey, Hillol Sarker, Syed Monowar Hossain, Mustafa al’Absi, Predrag Klasnja, and Susan Murphy. Just-in-time but not too much: Determining treatment timing in mobile health. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 21, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287057.

Liu:2019:PAN

[LDSZ19] Sicong Liu, Junzhao Du, Anshumali Shrivastava, and Lin Zhong. Privacy adversarial network: Representation learning for mobile data privacy. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 18, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369816.

Li:2018:PPC

[LDW^{+}18] Zeshui Li, Haipeng Dai, Wei Wang, Alex X. Liu, and Guihai Chen. PCIAS: Precise and contactless measurement of instantaneous angular speed using a smartphone. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 24, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287055.

Liu:2019:ASB

[LEP^{+}19] Fannie Liu, Mario Esparza, Maria Pavlovskaia, Geoff Kaufman, Laura Dabbish, and Andrés Monroy-Hernández. Animo: Sharing biosignals on a smartwatch for lightweight social connection. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–19, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314405.

Lee:2021:AUJ

[LEP21] Jinmo Lee, Neska Elhaouij, and Rosalind Picard. AmbientBreath: Unobtrusive just-in-time breathing intervention using multi-sensory stimulation and its evaluation in a car simulator. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2): 71:1–71:30, June 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463493.

Leung:2018:TTA

[LFH18] Ho-Man Colman Leung, Chi-Wing Fu, and Pheng-Ann Heng. TwistIn: Tangible authentication of smart devices via motion coanalysis with a smartwatch. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1– 24, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3214275.

Li:2023:EOA

[LFL^{+}23] Zisu Li, Li Feng, Chen Liang, Yuru Huang, and Mingming Fan. Exploring the opportunities of AR for enriching storytelling with family photos between grandparents and grandchildren. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):108:1– 108:??, September 2023. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3610903.

Liao:2020:PHR

[LGKM20] Peng Liao, Kristjan Greenewald, Predrag Klasnja, and Susan Murphy. Personalized HeartSteps: a reinforcement learning algorithm for optimizing physical activity. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):18:1– 18:22, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3381007.

Laporte:2023:LDS

[LGL23] Matias Laporte, Martin Gjoreski, and Marc Langheinrich. LAUREATE: a dataset for supporting research in affective computing and human memory augmentation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3): 106:1–106:??, September 2023. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3610892.

Liu:2021:ACA

[LGM^{+}21] Sicong Liu, Bin Guo, Ke Ma, Zhiwen Yu, and Junzhao Du. AdaSpring: Context-adaptive and runtimeevolutionary deep model compression for mobile applications. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):24:1– 24:22, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3448125.

Li:2022:ETT

[LGM^{+}22] Dengyun Li, Xin Ge, Qingzhou Ma, Brinda Mehra, Jie Liu, Teng Han, and Can Liu. Evaluating three touch gestures for moving objects across folded screens. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):124:1– 124:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3550309.

Liu:2020:KPY

[LGY^{+}20] Guanhong Liu, Yizheng Gu, Yiwen Yin, Chun Yu, Yuntao Wang, Haipeng Mi, and Yuanchun Shi. Keep the phone in your pocket: Enabling smartphone operation with an IMU ring for visually impaired people. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):58:1– 58:23, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397308.

Laput:2019:EES

[LH19] Gierad Laput and Chris Harrison. Exploring the efficacy of sparse, generalpurpose sensor constellations for wide-area activity sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–19, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328926.

Liu:2020:BTD

[LH20] Song Liu and Tian He. Bit-Light: Turning DLP projections into an interactive surface through bitlevel light encoding. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):141:1– 141:23, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3432228.

Li:2022:TUP

[LHC^{+}22] Jiayu Li, Zhiyu He, Yumeng Cui, Chenyang Wang, Chong Chen, Chun Yu, Min Zhang, Yiqun Liu, and Shaoping Ma. Towards ubiquitous personalized music recommendation with smart bracelets. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(3):125:1– 125:??, September 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3550333.

Liu:2022:PTE

[LHFZ22] Zongjian Liu, Jieling He, Jianjiang Feng, and Jie Zhou. PrinType: Text entry via fingerprint recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4): 174:1–174:??, December 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3569491.

Lindquist:2019:MMA

[LHK^{+}19] Wyatt Lindquist, Sumi Helal, Ahmed Khaled, Gerald Kotonya, and Jaejoon Lee. MAAT: Mobile apps as things in the IoT. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 22, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369823.

Lin:2021:TFO

[LHS^{+}21] Georgianna Lin, Malcolm Haynes, Sarthak Srinivas, Pramod Kotipalli, and Thad Starner. Towards finding the optimum position in the visual field for a head worn display used for task guidance with nonregistered graphics. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):22:1– 22:26, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3448091.

Li:2023:SCA

[LHS^{+}23] Jiyang Li, Lin Huang, Siddharth Shah, Sean J. Jones, Yincheng Jin, Dingran Wang, Adam Russell, Seokmin Choi, Yang Gao, Junsong Yuan, and Zhanpeng Jin. SignRing: Continuous American Sign Language recognition using IMU rings and virtual IMU data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):107:1– 107:??, September 2023.

CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3610881.

Li:2019:DDB

[LHW^{+}19] Songyuan Li, Shibo He, Shuai Wang, Tian He, and Jiming Chen. Data-driven battery-lifetime-aware scheduling for electric bus fleets. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 22, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3369810.

Lu:2020:DEH

[LHY^{+}20] Yiqin Lu, Bingjian Huang, Chun Yu, Guahong Liu, and Yuanchun Shi. Designing and evaluating hand-to-hand gestures with dual commodity wristworn devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):20:1– 20:27, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3380984.

Liang:2022:DKE

[LHY^{+}22] Chen Liang, Chi Hsia, Chun Yu, Yukang Yan, Yuntao Wang, and Yuanchun Shi. DRG-Keyboard: Enabling subtle gesture typing on the fingertip with dual IMU rings. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):170:1– 170:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3569463.

Lee:2019:ILD

[LILN19] Seulki Lee, Bashima Islam, Yubo Luo, and Shahriar Nirjon. Intermittent learning: On-device machine learning on intermittently powered system. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1–30, December 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369837.

Liu:2021:WFS

[LJK^{+}21] Yutong Liu, Landu Jiang, Linghe Kong, Qiao Xiang, Xue Liu, and Guihai Chen. Wi-Fruit: See through fruits with smart devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):169:1– 169:29, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3494971.

Li:2022:MNU

[LJZ^{+}22] Zhipeng Li, Yu Jiang, Yihao Zhu, Ruijia Chen, Ruolin Wang, Yuntao Wang, Yukang Yan, and Yuanchun Shi. Modeling the noticeability of user-avatar movement inconsistency for sense of body ownership intervention. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):64:1– 64:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3534590.

Lee:2019:VCS

[LKBK19] Kyuin Lee, Neil Klingensmith, Suman Banerjee, and Younghyun Kim. VoltKey: Continuous secret key generation based on power line noise for zero-involvement pairing and authentication. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 26, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3351251.

Lim:2019:HDN

[LKL19] Brian Y. Lim, Judy Kay, and Weilong Liu. How does a nation walk?: Interpreting large-scale step count activity with weekly streak patterns. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–46, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328928.

Lee:2022:UPR

[LKL22] Hyunsoo Lee, Soowon Kang, and Uichin Lee. Understanding privacy risks and perceived benefits in open dataset collection for mobile affective computing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):61:1– 61:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534623.

Liu:2021:ECE

[LKZ^{+}21] Shu Liu, Kevin Koch, Zimu Zhou, Simon Föll, Xiaoxi He, Tina Menke, Elgar Fleisch, and Felix Wortmann. The empathetic car: Exploring emotion inference via driver behaviour and traffic context. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):117:1– 117:34, September 2021.

CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478078.

Liberis:2022:DNN

[LL22] Edgar Liberis and Nicholas D. Lane. Differentiable neural network pruning to enable smart applications on microcontrollers. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):171:1– 171:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3569468.

Lin:2020:HSF

[LLC^{+}20] Zongyu Lin, Shiqing Lyu, Hancheng Cao, Fengli Xu, Yuqiong Wei, Hanan Samet, and Yong Li. HealthWalks: Sensing fine-grained individual health condition via mobility data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):138:1– 138:26, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3432229.

Lian:2021:ESY

[LLCY21] Jie Lian, Jiadong Lou, Li Chen, and Xu Yuan. EchoSpot: Spotting your locations via acoustic sensing.

Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3): 113:1–113:21, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478095.

Liu:2021:SEE

[LLF^{+}21] Xin Liu, Yuang Li, Josh Fromm, Yuntao Wang, Ziheng Jiang, Alex Mariakakis, and Shwetak Patel. SplitSR: an end-to-end approach to super-resolution on mobile devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):25:1– 25:20, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3448104.

Liu:2023:TDF

[LLG^{+}23] Jinyi Liu, Wenwei Li, Tao Gu, Ruiyang Gao, Bin Chen, Fusang Zhang, Dan Wu, and Daqing Zhang. Towards a dynamic Fresnel zone model to WiFi-based human activity recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2): 65:1–65:??, June 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3596270.

Lee:2019:QOT

[LLL^{+}19] Lik Hang Lee, Kit Yung Lam, Tong Li, Tristan Braud, Xiang Su, and Pan Hui. Quadmetric optimized thumb-to-finger interaction for force assisted one-handed text entry on mobile headsets. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1–27, September 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3351252.

Liu:2020:PPP

[LLL^{+}20] Bingyan Liu, Yuanchun Li, Yunxin Liu, Yao Guo, and Xiangqun Chen. PMC: a privacy-preserving deep learning model customization framework for edge computing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):139:1– 139:25, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3432208.

Liu:2021:WHW

[LLL21] Yang Liu, Chengdong Lin, and Zhenjiang Li. WR-Hand: Wearable armband can track User’s hand. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):118:1– 118:27, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478112.

Li:2022:LPL

[LLLX22] Dong Li, Jialin Liu, Sunghoon Ivan Lee, and Jie Xiong. LASense: Pushing the limits of finegrained activity sensing using acoustic signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):21:1– 21:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517253.

Lee:2022:DWB

[LLPH22] Jong Taek Lee, Yu Kai Lim, Shijia Pan, and Jun Han. Don’t “Weight” to board: Augmenting vision-based passenger weight prediction via viscoelastic mat. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):168:1– 168:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3569465.

Liu:2022:CHC

[LLW^{+}22] Ziwei Liu, Feng Lin, Chao Wang, Yijie Shen, Zhongjie Ba, Li Lu, Wenyao Xu, and Kui Ren. CamRadar: Hidden camera detection leveraging amplitude-modulated sensor images embedded in electromagnetic emanations. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):173:1– 173:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3569505.

Liu:2021:BLY

[LLWX21] Jialin Liu, Dong Li, Lei Wang, and Jie Xiong. BlinkListener: “listen” to your eye blink using your smartphone. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):73:1– 73:27, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3463521.

Liu:2018:TEM

[LLZ^{+}18] Liang Liu, Wu Liu, Yu Zheng, Huadong Ma, and Cheng Zhang. Third-eye: a mobilephone-enabled crowdsensing system for air quality monitoring. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–26, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191752.

Li:2020:WPW

[LLZ^{+}20] Shengjie Li, Zhaopeng Liu, Yue Zhang, Qin Lv, Xiaopeng Niu, Leye Wang, and Daqing Zhang. Wi-Border: Precise Wi-Fi based boundary sensing via through-wall discrimination. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3): 89:1–89:30, September 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3411834.

Liu:2022:AEA

[LLZ^{+}22] Sicong Liu, Xiaochen Li, Zimu Zhou, Bin Guo, Meng Zhang, Haocheng Shen, and Zhiwen Yu. AdaEnlight: Energy-aware lowlight video stream enhancement on mobile devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4): 172:1–172:??, December 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3569464.

Liu:2023:AAM

[LMS^{+}23] Shinan Liu, Tarun Mangla, Ted Shaowang, Jinjin Zhao, John Paparrizos, Sanjay Krishnan, and Nick Feamster. AMIR: Active multimodal interaction recognition from video and network traffic in connected environments. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1): 21:1–21:??, March 2023. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3580818.

Lachand:2019:TSC

[LMT19] Valentin Lachand, Christine Michel, and Aurélien Tabard. Toccata: Supporting classroom orchestration with activity based computing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–24, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3328924.

Li:2021:HAG

[LNAH21] Tianshi Li, Elijah B. Neundorfer, Yuvraj Agarwal, and Jason I. Hong. Honeysuckle: Annotation-guided code generation of in-app privacy notices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):112:1– 112:27, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478097.

Liu:2019:MLP

[LNT^{+}19] Tony Liu, Jennifer Nicholas, Max M. Theilig, Sharath C. Guntuku, Konrad Kording, David C. Mohr, and Lyle Ungar. Machine learning for phone-based relationship estimation: The need to consider population heterogeneity. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 23, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369820.

Liang:2020:HDF

[LP20] Zilu Liang and Bernd Ploderer. How does Fitbit measure brainwaves: a qualitative study into the credibility of sleep-tracking technologies. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):17:1– 17:29, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3380994.

Lambrichts:2021:STE

[LRH^{+}21] Mannu Lambrichts, Raf Ramakers, Steve Hodges, Sven Coppers, and James Devine. A survey and taxonomy of electronics toolkits for interactive and ubiquitous device prototyping. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):70:1– 70:24, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3463523.

Lambrichts:2023:CSC

[LRH^{+}23] Mannu Lambrichts, Raf Ramakers, Steve Hodges, James Devine, Lorraine Underwood, and Joe Finney. CircuitGIue: a software configurable converter for interconnecting multiple heterogeneous electronic components. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):63:1– 63:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3596265.

Li:2020:EID

[LS20] Boning Li and Akane Sano.

Extraction and interpretation of deep autoencoderbased temporal features from wearables for forecasting personalized mood, health, and stress. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):49:1– 49:26, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397318.

Lee:2021:USS

[LSE21] Jong Ho Lee, Jessica Schroeder, and Daniel A. Epstein. Understanding and supporting self-tracking app selection. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):166:1– 166:25, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3494980.

Lai:2023:ITN

[LSSH23] Hsin-Yu Lai, Charles G.

Sodini, Vivienne Sze, and Thomas Heldt. Individualized tracking of neurocognitivestate-dependent eye-movement features using mobile devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):19:1–19:??, March 2023.

CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3580843.

Liu:2019:RHJ

[LSW^{+}19] Ruibo Liu, Qijia Shao, Siqi Wang, Christina Ru, Devin Balkcom, and Xia Zhou. Reconstructing human joint motion with computational fabrics. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–26, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3314406.

Lu:2018:JMH

[LSY^{+}18] Jin Lu, Chao Shang, Chaoqun Yue, Reynaldo Morillo, Shweta Ware, Jayesh Kamath, Athanasios Bamis, Alexander Russell, Bing Wang, and Jinbo Bi. Joint modeling of heterogeneous sensing data for depression assessment via multitask learning. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–21, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3191753.

Liu:2023:USP

[LSY^{+}23] Xiaoyi Liu, Yingtian Shi, Chun Yu, Cheng Gao, Tianao Yang, Chen Liang, and Yuanchun Shi. Understanding in-situ programming for smart home automation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):66:1– 66:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596254.

Liang:2019:ABA

[LT19] Dawei Liang and Edison Thomaz. Audio-based activities of daily living (ADL) recognition with large-scale acoustic embeddings from online videos. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–18, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314404.

Li:2023:HCB

[LWA23] Youpeng Li, Xuyu Wang, and Lingling An. Hierarchical clustering-based personalized federated learning for robust and fair human activity recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):20:1– 20:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580795.

Lu:2022:SDM

[LWC^{+}22] Wang Lu, Jindong Wang, Yiqiang Chen, Sinno Jialin Pan, Chunyu Hu, and Xin Qin. Semanticdiscriminative mixup for generalizable sensor-based cross-domain activity recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):65:1–65:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3534589.

Liu:2022:MPM

[LWX^{+}22] Xin Liu, Yuntao Wang, Sinan Xie, Xiaoyu Zhang, Zixian Ma, Daniel McDuff, and Shwetak Patel. MobilePhys: Personalized mobile camera-based contactless physiological sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1): 24:1–24:??, March 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3517225.

Lyu:2021:MGV

[LWYZ21] Wenjun Lyu, Guang Wang, Yu Yang, and Desheng Zhang. Mover: Generalizability verification of human mobility models via heterogeneous use cases. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):171:1– 171:21, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3494997.

Liu:2020:RTA

[LWZ^{+}20] Haipeng Liu, Yuheng Wang, Anfu Zhou, Hanyue He, Wei Wang, Kunpeng Wang, Peilin Pan, Yixuan Lu, Liang Liu, and Huadong Ma. Real-time arm gesture recognition in smart home scenarios via millimeter wave sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):140:1– 140:28, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3432235.

Liu:2019:BRC

[LXC^{+}19] Chen Liu, Jie Xiong, Lin Cai, Lin Feng, Xiaojiang Chen, and Dingyi Fang. Beyond respiration: Contactless sleep sound-activity recognition using RF signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 22, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351254.

Lin:2021:DML

[LXW^{+}21] Yuancan Lin, Lei Xie, Chuyu Wang, Yanling Bu, and Sanglu Lu. DropMonitor: Millimeter-level sensing for RFID-based infusion drip rate monitoring. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2): 72:1–72:22, June 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463496.

Lu:2019:ISS

[LYC^{+}19] Li Lu, Jiadi Yu, Yingying Chen, Yanmin Zhu, Minglu Li, and Xiangyu Xu. I3: Sensing scrolling humancomputer interactions for intelligent interest inference on smartphones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1–22, September 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351255.

Lu:2020:VSV

[LYCW20] Li Lu, Jiadi Yu, Yingying Chen, and Yan Wang. VocalLock: Sensing vocal tract for passphraseindependent user authentication leveraging acoustic signals on smartphones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2): 51:1–51:24, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397320.

Liu:2023:TFG

[LYD^{+}23] Jingxiao Liu, Siyuan Yuan, Yiwen Dong, Biondo Biondi, and Hae Young Noh. TelecomTM: a fine-grained and ubiquitous traffic monitoring system using preexisting telecommunication fiber-optic cables as sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2): 64:1–64:??, June 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596262.

Luo:2021:RTR

[LYF^{+}21] Chengwen Luo, Zhongru Yang, Xingyu Feng, Jin Zhang, Hong Jia, Jianqiang Li, Jiawei Wu, and Wen Hu. RFaceID: Towards RFID-based facial recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):170:1– 170:21, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3494985.

Liu:2020:HMS

[LYH^{+}20] Shengzhong Liu, Shuochao Yao, Yifei Huang, Dongxin Liu, Huajie Shao, Yiran Zhao, Jinyang Li, Tianshi Wang, Ruijie Wang, Chaoqi Yang, and Tarek Abdelzaher. Handling missing sensors in topology-aware IoT applications with gated graph neural network. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):90:1– 90:31, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3411818.

Lukoff:2018:WMS

[LYKH18] Kai Lukoff, Cissy Yu, Julie Kientz, and Alexis Hiniker. What makes smartphone use meaningful or meaningless? Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–26, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191754.

Liu:2020:GGA

[LYL^{+}20] Shengzhong Liu, Shuochao Yao, Jinyang Li, Dongxin Liu, Tianshi Wang, Huajie Shao, and Tarek Abdelzaher. GIobalFusion: a global attentional deep learning framework for multisensor information fusion. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):19:1– 19:27, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3380999.

Lian:2021:FDI

[LYLT21] Jie Lian, Xu Yuan, Ming Li, and Nian-Feng Tzeng. Fall detection via inaudible acoustic sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):114:1– 114:21, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478094.

Lee:2022:AUA

[LYP^{+}22] Kyuin Lee, Yucheng Yang, Omkar Prabhune, Aishwarya Lekshmi Chithra, Jack West, Kassem Fawaz, Neil Klingensmith, Suman Banerjee, and Younghyun Kim. AEROKEY: Using ambient electromagnetic radiation for secure and usable wireless device authentication. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):20:1– 20:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517254.

Liang:2021:DES

[LYQ^{+}21] Chen Liang, Chun Yu, Yue Qin, Yuntao Wang, and Yuanchun Shi. DualRing: Enabling subtle and expressive hand interaction with dual IMU rings. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):115:1– 115:27, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478114.

Luo:2022:NIA

[LYWZ22] Zhiqing Luo, Mingxuan Yan, Wei Wang, and Qian Zhang. Non-intrusive anomaly detection of industrial robot operations by exploiting nonlinear effect.

Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4): 175:1–175:??, December 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3569477.

Lu:2017:BEF

[LYY^{+}17] Yiqin Lu, Chun Yu, Xin Yi, Yuanchun Shi, and Shengdong Zhao. Blind-Type: Eyes-free text entry on handheld touchpad by leveraging thumb’s muscle memory. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090083.

Liu:2021:WSB

[LZG^{+}21] Jinyi Liu, Youwei Zeng, Tao Gu, Leye Wang, and Daqing Zhang. WiPhone: Smartphone-based respiration monitoring using ambient reflected WiFi signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1): 23:1–23:19, March 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3448092.

Li:2022:ELP

[LZL^{+}22] Ke Li, Ruidong Zhang, Bo Liang, François Guimbretière, and Cheng Zhang. EarIO: a low-power acoustic sensing earable for continuously tracking detailed facial movements. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):62:1– 62:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3534621.

Liang:2023:AFF

[LZT23] Dawei Liang, Alice Zhang, and Edison Thomaz. Automated face-to-face conversation detection on a commodity smartwatch with acoustic sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3):109:1– 109:??, September 2023. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3610882.

Lee:2020:USL

[LZZ^{+}20] Yeonjoon Lee, Yue Zhao, Jiutian Zeng, Kwangwuk Lee, Nan Zhang, Faysal Hossain Shezan, Yuan Tian, Kai Chen, and XiaoFeng Wang. Using sonar for liveness detection to protect smart speakers against remote attackers. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):16:1– 16:28, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3380991.

Liang:2021:FLC

[LZZ^{+}21] Yumeng Liang, Anfu Zhou, Huanhuan Zhang, Xinzhe Wen, and Huadong Ma. FG-LiquID: a contact-less fine-grained liquid identifier by pushing the limits of millimeter-wave sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3): 116:1–116:27, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478075.

Liang:2023:MDH

[LZZ^{+}23] Kun Liang, Anfu Zhou, Zhan Zhang, Hao Zhou, Huadong Ma, and Chenshu Wu. mmStress: Distilling human stress from daily activities via contactless millimeter-wave sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3): 110:1–110:??, September 2023. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3610926.

Mazankiewicz:2020:IRT

[MBB20] Alan Mazankiewicz, Klemens Böhm, and Mario Berges. Incremental realtime personalization in human activity recognition using domain adaptive batch normalization. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):144:1– 144:20, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3432230.

Mariakakis:2017:BSB

[MBP^{+}17] Alex Mariakakis, Megan A. Banks, Lauren Phillipi, Lei Yu, James Taylor, and Shwetak N. Patel. BiliScreen: Smartphone-based scleral jaundice monitoring for liver and pancreatic disorders. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

                                        //dl.acm.org/citation.
                                        cfm?id=3090085.

Mosenia:2018:PBP

[MBZ^{+}18] Arsalan Mosenia, Jad F. Bechara, Tao Zhang, Prateek Mittal, and Mung Chiang. ProCMotive: Bringing programmability and connectivity into isolated vehicles. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–31, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191758.

Malekzadeh:2021:DDA

[MCCH21] Mohammad Malekzadeh, Richard Clegg, Andrea Cavallaro, and Hamed Haddadi. DANA: Dimension-adaptive neural architecture for multivariate sensor data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):120:1– 120:27, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478074.

Meegahapola:2022:GPM

[MDK^{+}22] Lakmal Meegahapola, William Droz, Peter Kun, Amalia de Götzen, Chaitanya Nutakki, Shyam Diwakar, Salvador Ruiz Correa, Donglei Song, Hao Xu, Miriam Bidoglia, George Gaskell, Altangerel Chagnaa, Amarsanaa Ganbold, Tsolmon Zundui, Carlo Caprini, Daniele Miorandi, Alethia Hume, Jose Luis Zarza, Luca Cernuzzi, Ivano Bison, Marcelo Rodas Britez, Matteo Busso, Ronald Chenu-Abente, Can Günel, Fausto Giunchiglia, Laura Schelenz, and Daniel Gatica-Perez. Generalization and personalization of mobile sensing-based mood inference models: an analysis of College students in eight countries. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):176:1– 176:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3569483.

Meyer:2021:CBH

[MFSK21] Johannes Meyer, Adrian Frank, Thomas Schlebusch, and Enkeljeda Kasneci. A CNN-based human activity recognition system combining a laser feedback interferometry eye movement sensor and an IMU for contextaware smart glasses. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):172:1– 172:24, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3494998.

Mamish:2023:IHD

[MGC^{+}23a] John Mamish, Amy Guo, Thomas Cohen, Julian Richey, Yang Zhang, and Josiah Hester. Interaction harvesting: a design probe of user-powered widgets. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3): 112:1–112:??, September 2023. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3610880.

Murnane:2023:NBV

[MGC^{+}23b] Elizabeth L. Murnane, Yekaterina S. Glazko, Jean Costa, Raymond Yao, Grace Zhao, Paula M. L. Moya, and James A. Landay. Narrative-based visual feedback to encourage sustained physical activity: a field trial of the WhoIsZuki mobile health platform. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):23:1– 23:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580786.

Moore:2021:IME

[MGWM21] Jimmy Moore, Pascal Goffin, Jason Wiese, and Miriah Meyer. An interview method for engaging personal data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(4):173:1– 173:28, December 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3494964.

Miao:2023:TTG

[MGY^{+}23] Yuchen Miao, Chaojie Gu, Zhenyu Yan, Sze Yiu Chau, Rui Tan, Qi Lin, Wen Hu, Shibo He, and Jiming Chen. TouchKey: Touch to generate symmetric keys by skin electric potentials induced by powerline radiation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):70:1– 70:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3596264.

Meier:2023:BBM

[MH23] Manuel Meier and Christian Holz. BMAR: Barometric and motion-based alignment and refinement for offline signal synchronization across devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):69:1– 69:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3596268.

Maiti:2019:LEI

[MJ19] Anindya Maiti and Murtuza Jadliwala. Light Ears: Information leakage via smart lights. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1–27, September 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351256.

Mishra:2021:DRM

[MKK^{+}21] Varun Mishra, Florian Künzler, Jan-Niklas Kramer, Elgar Fleisch, Tobias Kowatsch, and David Kotz. Detecting receptivity for mHealth interventions in the natural environment. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):74:1– 74:24, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463492.

Mahmud:2023:PUB

[MLH^{+}23] Saif Mahmud, Ke Li, Guilin Hu, Hao Chen, Richard Jin, Ruidong Zhang, François Guimbretière, and Cheng Zhang. PoseSonic: 3D upper body pose estimation through egocentric acoustic sensing on smartglasses. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(3): 111:1–111:??, September 2023. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3610895.

Maruri:2018:VSN

[MLMH^{+}18] Héctor A. Cordourier Maruri, Paulo Lopez-Meyer, Jonathan Huang, Willem Marco Beltman, Lama Nachman, and Hong Lu. V-Speech: Noise-robust speech capturing glasses using vibration sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1–23, December 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287058.

Meegahapola:2021:ESC

[MLPGP21] Lakmal Meegahapola, Florian Labhart, Thanh-Trung Phan, and Daniel Gatica-Perez. Examining the social context of alcohol drinking in young adults with smartphone sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):121:1– 121:26, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478126.

Ma:2020:FGA

[MLX^{+}20] Rui Ma, Ning Liu, Xiangxiang Xu, Yue Wang, Hae Young Noh, Pei Zhang, and Lin Zhang. Fine-grained air pollution inference with mobile sensing systems: a weather-related deep autoencoder model. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):52:1– 52:21, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397322.

Mirjafari:2019:DHL

[MMG^{+}19] Shayan Mirjafari, Kizito Masaba, Ted Grover, Weichen Wang, Pino Audia, Andrew T. Campbell, Nitesh V. Chawla, Vedant Das Swain, Munmun De Choudhury, Anind K. Dey, Sidney K. D’Mello, Ge Gao, Julie M. Gregg, Krithika Jagannath, Kaifeng Jiang, Suwen Lin, Qiang Liu, Gloria Mark, Gonzalo J. Martinez, Stephen M. Mattingly, Edward Moskal, Raghu Mulukutla, Subigya Nepal, Kari Nies, Manikanta D. Reddy, Pablo Robles-Granda, Koustuv Saha, Anusha Sirigiri, and Aaron Striegel. Differentiating higher and lower job performers in the workplace using mobile sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–24, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3328908.

Martinez-Maldonado:2020:TTI

[MMMS^{+}20] Roberto Martinez-Maldonado, Katerina Mangaroska, Jurgen Schulte, Doug Elliott, Carmen Axisa, and Simon Buckingham Shum. Teacher tracking with integrity: What indoor positioning can reveal about instructional proxemics. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):22:1– 22:27, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3381017.

Mukherjee:2019:MSM

[MNV^{+}19] Manideepa Mukherjee, Sana Ali Naqvi, Anushika Verma, Debarka Sengupta, and Aman Parnami. MenstruLoss: Sensor for menstrual blood loss monitoring. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–21, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328929.

MontoyaFreire:2021:IFI

[MOD21] Maria L. Montoya Freire, Antti Oulasvirta, and Mario Di Francesco. Inverse foraging: Inferring users’ interest in pervasive displays. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):122:1– 122:18, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478103.

Ma:2023:FAD

[MQZH23] Jiaju Ma, Jing Qian, Tongyu Zhou, and Jeff Huang. FocalPoint: Adaptive direct manipulation for selecting small 3D virtual objects. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):22:1– 22:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580856.

Meegahapola:2021:OMB

[MRCdCRV^{+}21] Lakmal Meegahapola, Salvador Ruiz-Correa, Viridiana del Carmen Robledo-Valero, Emilio Ernesto Hernandez-Huerfano, Leonardo Alvarez-Rivera, Ronald Chenu-Abente, and Daniel Gatica-Perez. One more bite?: Inferring food consumption level of college students using smartphone sensing and self-reports. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):26:1– 26:28, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3448120.

Morshed:2019:PMI

[MSL^{+}19] Mehrab Bin Morshed, Koustuv Saha, Richard Li, Sidney K. D’Mello, Munmun De Choudhury, Gregory D. Abowd, and Thomas Plötz. Prediction of mood instability with passive sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3 (3):1–21, September 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3351233.

Mendel:2022:TPS

[MSTT22] Tamir Mendel, Roei Schuster, Eran Tromer, and Eran Toch. Toward proactive support for older adults: Predicting the right moment for providing mobile safety help. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):25:1– 25:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517249.

Mao:2020:DAR

[MSWQ20] Wenguang Mao, Wei Sun, Mei Wang, and Lili Qiu. DeepRange: Acoustic ranging via deep learning. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):143:1– 143:23, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3432195.

Mendel:2019:MMW

[MT19] Tamir Mendel and Eran Toch. My Mom was getting this popup: Understanding motivations and processes in helping older relatives with mobile security and privacy. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 20, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369821.

Montanari:2018:MIP

[MTF^{+}18] Alessandro Montanari, Zhao Tian, Elena Francu, Benjamin Lucas, Brian Jones, Xia Zhou, and Cecilia Mascolo. Measuring interaction proxemics with wearable light tags. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–30, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191757.

Mai:2019:DMA

[MTH19] Christian Mai, Niklas Thiem, and Heinrich Hussmann. DrawingPresence: a method for assessing temporal fluctuations of presence status in a VR experience. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 21, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369827.

McDonald:2020:KSD

[MVS^{+}20] Denisa Qori McDonald, Richard Vallett, Erin Solovey, Geneviève Dion, and Ali Shokoufandeh. Knitted sensors: Designs and novel approaches for real-time, realworld sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):145:1– 145:25, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3432201.

Motahar:2022:RPI

[MW22] Tamanna Motahar and Jason Wiese. A review of personal informatics research for people with motor disabilities. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):66:1– 66:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534614.

Mohammed:2021:SCF

[MWJ^{+}21] Noor Mohammed, Rui Wang, Robert W. Jackson, Yeonsik Noh, Jeremy Gummeson, and Sunghoon Ivan Lee. ShaZam: Charge-free wearable devices via intrabody power transfer from everyday objects. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):75:1– 75:25, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463505.

Ma:2020:CCO

[MZJ20] Yongsen Ma, Yunze Zeng, and Vivek Jain. CarOSense: Car occupancy sensing with the ultra-wideband keyless infrastructure. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):91:1– 91:28, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3411820.

Maag:2017:SMH

[MZST17] Balz Maag, Zimu Zhou, Olga Saukh, and Lothar Thiele. SCAN: Multihop calibration for mobile sensor arrays. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

                                        //dl.acm.org/citation.
                                        cfm?id=3090084.

Maag:2018:WAE

[MZT18] Balz Maag, Zimu Zhou, and Lothar Thiele. W-Air: Enabling personal air pollution monitoring on wearables. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–25, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3191756.

Ma:2018:SSL

[MZW^{+}18] Yongsen Ma, Gang Zhou, Shuangquan Wang, Hongyang Zhao, and Woosub Jung. SignFi: Sign language recognition using WiFi. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–21, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191755.

Niforatos:2017:CLM

[NCM^{+}17] Evangelos Niforatos, Caterina Cinel, Cathleen Cortis Mack, Marc Langheinrich, and Geoff Ward. Can less be more?: Contrasting limited, unlimited, and automatic picture capture for augmenting memory recall. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090086.

Niu:2020:MMS

[NHL^{+}20] Qun Niu, Tao He, Ning Liu, Suining He, Xiaonan Luo, and Fan Zhou. MAIL: Multi-scale attention-guided indoor localization using geomagnetic sequences. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):54:1– 54:23, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3397335.

Nair:2019:UCT

[NJWFM19] Suraj Nair, Kiran Javkar, Jiahui Wu, and Vanessa Frias-Martinez. Understanding cycling trip purpose and route choice using GPS traces and open data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–26, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3314407.

Nakamura:2023:EEP

[NNMY23] Yugo Nakamura, Rei Nakaoka, Yuki Matsuda, and Keiichi Yasumoto. eat2pic: an eating-painting interactive system to nudge users into making healthier diet choices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):24:1– 24:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3580784.

Nandakumar:2017:CAI

[NTKG17] Rajalakshmi Nandakumar, Alex Takakuwa, Tadayoshi Kohno, and Shyamnath Gollakota. CovertBand: Activity information leakage using music. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(3):87:1– 87:24, ???? 2017. CODEN ???? ISSN ???? URL

http://musicattacks.cs.
washington.edu/activity-
information-leakage.pdf.

Nomiyama:2018:XTP

[NTO^{+}18] Masato Nomiyama, Toshiki Takeuchi, Hiroyuki Onimaru, Tomohiro Tanikawa, Takuji Narumi, and Michitaka Hirose. Xnavi: Travel planning system based on experience flows. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–25, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191759.

Niijima:2021:RMA

[NTT^{+}21] Arinobu Niijima, Toki Takeda, Kentaro Tanaka, Ryosuke Aoki, and Yukio Koike. Reducing muscle activity when playing tremolo by using electrical muscle stimulation to learn efficient motor skills. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):123:1– 123:17, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478110.

Nambi:2018:FSB

[NVP18] Akshay Uttama Nambi, Adtiya Virmani, and Venkata N. Padmanabhan. FarSight: a smartphone-based vehicle ranging system. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1– 22, December 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3287059.

Nguyen:2020:MAS

[NZL^{+}20] Ngoc Thi Nguyen, Agustin Zuniga, Hyowon Lee, Pan Hui, Huber Flores, and Petteri Nurmi. (M)ad to see me?: Intelligent advertisement placement: Balancing user annoyance and advertising effectiveness. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):53:1– 53:26, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3397324.

Ouyang:2018:MFP

[OGG^{+}18] Yi Ouyang, Bin Guo, Tong Guo, Longbing Cao, and Zhiwen Yu. Modeling and forecasting the popularity evolution of mobile apps: a multivariate Hawkes process approach. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(4):1–23, December 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3287060.

Obuchi:2020:PBF

[OHW^{+}20] Mikio Obuchi, Jeremy F.

Huckins, Weichen Wang, Alex daSilva, Courtney Rogers, Eilis Murphy, Elin Hedlund, Paul Holtzheimer, Shayan Mirjafari, and Andrew Campbell. Predicting brain functional connectivity using mobile sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1): 23:1–23:22, March 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3381001.

OHagan:2022:PET

[OSG^{+}22] Joseph O’Hagan, Pejman Saeghe, Jan Gugenheimer, Daniel Medeiros, Karola Marky, Mohamed Khamis, and Mark McGill. Privacyenhancing technology and everyday augmented reality:

Understanding bystanders’ varying needs for awareness and consent. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):177:1– 177:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3569501.

Ott:2020:ODO

[OWH^{+}20] Felix Ott, Mohamad Wehbi, Tim Hamann, Jens Barth, Björn Eskofier, and Christopher Mutschler. The OnHW dataset: Online handwriting recognition from IMUenhanced ballpoint pens with machine learning. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):92:1– 92:20, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3411842.

Parmar:2020:MIP

[PB20] Dhaval Parmar and Timothy Bickmore. Making it personal: Addressing individual audience members in oral presentations using augmented reality. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):55:1– 55:22, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397336.

Pradhan:2018:SBA

[PBM^{+}18] Swadhin Pradhan, Ghufran Baig, Wenguang Mao, Lili Qiu, Guohai Chen, and Bo Yang. Smartphonebased acoustic indoor space mapping. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–26, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214278.

Park:2020:HAL

[PCBK20] Jaeyeon Park, Hyeon Cho, Rajesh Krishna Balan, and JeongGil Ko. HeartQuake: Accurate low-cost noninvasive ECG monitoring using bed-mounted geophones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):93:1– 93:28, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3411843.

Panagiotidou:2023:SSE

[PCF^{+}23] Georgia Panagiotidou, Enrico Costanza, Michael J. Fell, Farhan Samanani, and Hannah Knox. Supporting solar energy coordination among communities. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2): 71:1–71:??, June 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596243.

Peng:2018:ADM

[PCYZ18] Liangying Peng, Ling Chen, Zhenan Ye, and Yi Zhang. AROMA: a deep multitask learning based simple and complex human activity recognition method using wearable sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–16, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214277.

Park:2023:UDJ

[PL23] Joonyoung Park and Uichin Lee. Understanding disengagement in just-in-time mobile health interventions. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2): 72:1–72:??, June 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596240.

Pan:2021:ESE

[PLC^{+}21] Yan Pan, Shining Li, Qianwu Chen, Nan Zhang, Tao Cheng, Zhigang Li, Bin Guo, Qingye Han, and Ting Zhu. Efficient schedule of energy-constrained UAV using crowdsourced buses in last-mile parcel delivery. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1): 28:1–28:23, March 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3448079.

Park:2021:ERT

[PLK21] HyeonJung Park, Youngki Lee, and JeongGil Ko. Enabling real-time sign language translation on mobile platforms with on-board depth cameras. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):77:1– 77:30, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463498.

Ponnada:2022:CBM

[PLW^{+}22] Aditya Ponnada, Jixin Li, Shirlene Wang, Wei-Lin Wang, Bridgette Do, Genevieve F. Dunton, and Stephen S. Intille. Contextual biases in microinteraction ecological momentary assessment (μEMA) non-response. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):26:1– 26:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517259.

Pushp:2018:PMS

[PLX^{+}18] Saumay Pushp, Yunxin Liu, Mengwei Xu, Changyoung Koh, and Junehwa Song. PrivacyShield: a mobile system for supporting subtle just-in-time privacy provisioning through off-screen-based touch gestures. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–38, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214279.

Poyraz:2019:UBS

[PM19] Emirhan Poyraz and Gokhan Memik. Using built-in sensors to predict and utilize user satisfaction for CPU settings on smartphones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–25, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314408.

Phan:2019:DAC

[PMGP19] Thanh-Trung Phan, Skanda Muralidhar, and Daniel Gatica-Perez. Drinks & crowds: Characterizing alcohol consumption through crowdsensing and social media. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3 (2):1–30, June 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3328930.

Park:2021:DEM

[PNL^{+}21] Chunjong Park, Hung Ngo, Libby Rose Lavitt, Vincent Karuri, Shiven Bhatt, Peter Lubell-Doughtie, Anuraj H. Shankar, Leonard Ndwiga, Victor Osoti, Juliana K. Wambua, Philip Bejon, Lynette Isabella Ochola-Oyier, Monique Chilver, Nigel Stocks, Victoria Lyon, Barry R. Lutz, Matthew Thompson, Alex Mariakakis, and Shwetak Patel. The design and evaluation of a mobile system for rapid diagnostic test interpretation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):29:1– 29:26, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3448106.

Paruthi:2018:FSS

[PRC^{+}18] Gaurav Paruthi, Shriti Raj, Natalie Colabianchi, Predrag Klasnja, and Mark W. Newman. Finding the sweet spot(s): Understanding context to support physical activity plans. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–17, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3191761.

Paredes:2021:FEE

[PRC^{+}21] Luis Paredes, Sai Swarup Reddy, Subramanian Chidambaram, Devashri Vagholkar, Yunbo Zhang, Bedrich Benes, and Karthik Ramani. FabHandWear: an end-to-end pipeline from design to fabrication of customized functional hand wearables. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2): 76:1–76:22, June 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3463518.

Prajwal:2023:TEE

[PRS^{+}23] M. Prajwal, Ayush Raj, Sougata Sen, Snehanshu Saha, and Surjya Ghosh. Towards efficient emotion self-report collection using Human–AI collaboration: a case study on smartphone keyboard interaction. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):68:1– 68:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596269.

Price:2017:LYL

[PSC^{+}17] Blaine A. Price, Avelie Stuart, Gul Calikli, Ciaran Mccormick, Vikram Mehta, Luke Hutton, Arosha K. Bandara, Mark Levine, and Bashar Nuseibeh. Logging you, logging me: A replicable study of privacy and sharing behaviour in groups of visual lifeloggers. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090087.

Park:2023:ETS

[PSC^{+}23] Jung Wook Park, Sienna Xin Sun, Tingyu Cheng, Dong Whi Yoo, Jiawei Zhou, Youngwook Do, Gregory D. Abowd, and Rosa I. Arriaga. Exergy: a toolkit to simplify creative applications of wind energy harvesting. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):25:1– 25:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3580814.

Palipana:2021:PMA

[PSLS21] Sameera Palipana, Dariush Salami, Luis A. Leiva, and Stephan Sigg. Pantomime: Mid-air gesture recognition with sparse millimeter-wave radar point clouds. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):27:1– 27:27, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3448110.

Parker:2018:DPS

[PTK18] Callum Parker, Martin Tomitsch, and Judy Kay. Does the public still look at public displays?: a field observation of public displays in the wild. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–24, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214276.

Parilusyan:2022:SNA

[PTMM^{+}22] Brice Parilusyan, Marc Teyssier, Valentin Martinez-Missir, Clément Duhart, and Marcos Serrano. Sensurfaces: a novel approach for embedded touch sensing on everyday surfaces. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2): 67:1–67:??, July 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534616.

Parizi:2019:APE

[PWP19] Farshid Salemi Parizi, Eric Whitmire, and Shwetak Patel. AuraRing: Precise electromagnetic finger tracking. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 28, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369831.

Paredes:2018:JBC

[PZH^{+}18] Pablo E. Paredes, Yijun Zhou, Nur Al-Huda Hamdan, Stephanie Balters, Elizabeth Murnane, Wendy Ju, and James A. Landay. Just breathe: Incar interventions for guided slow breathing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–23, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3191760.

Qian:2018:RMB

[QCP^{+}18] Jing Qian, Arielle Chapin, Alexandra Papoutsaki, Fumeng Yang, Klaas Nelissen, and Jeff Huang. Remotion: a motion-based capture and replay platform of mobile device interaction for remote usability testing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–18, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214280.

Qin:2019:CDA

[QCWY19] Xin Qin, Yiqiang Chen, Jindong Wang, and Chaohui Yu. Cross-dataset activity recognition via adaptive spatial-temporal transfer learning. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1–25, December 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369818.

Qingxin:2019:UFA

[QWK^{+}19] Xia Qingxin, Atsushi Wada, Joseph Korpela, Takuya Maekawa, and Yasuo Namioka. Unsupervised factory activity recognition with wearable sensors using process instruction information. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1– 23, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3328931.

Ruan:2020:DPR

[RBL^{+}20] Sijie Ruan, Jie Bao, Yuxuan Liang, Ruiyuan Li, Tianfu He, Chuishi Meng, Yanhua Li, Yingcai Wu, and Yu Zheng. Dynamic public resource allocation based on human mobility prediction. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1): 25:1–25:22, March 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3380986.

Rashid:2022:TMB

[RDOA22] Nafiul Rashid, Berken Utku Demirel, Mohanad Odema, and Mohammad Abdullah Al Faruque. Template matching based early exit CNN for energy-efficient myocardial infarction detection on low-power wearable devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):68:1– 68:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534580.

Ranganathan:2018:RBF

[RGL^{+}18] Vaishnavi Ranganathan, Sidhant Gupta, Jonathan Lester, Joshua R. Smith, and Desney Tan. RF bandaid: a fully-analog and passive wireless interface for wearable sensors. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–21, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214282.

Radu:2019:VKT

[RH19] Valentin Radu and Maximilian Henne. Vision2Sensor: Knowledge transfer across sensing modalities for human activity recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(3):1– 21, September 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3351242.

Rahman:2023:AGA

[RHI^{+}23] Wasifur Rahman, Masum Hasan, Md Saiful Islam, Titilayo Olubajo, Jeet Thaker, Abdel-Rahman Abdelkader, Phillip Yang, Henry Paulson, Gulin Oz, Alexandra Durr, Thomas Klockgether, Tetsuo Ashizawa, Readisca Investigators, and Ehsan Hoque. Auto-gait: Automatic ataxia risk assessment with computer vision from gait task videos. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1): 26:1–26:??, March 2023. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3580845.

Rinta-Homi:2021:HLC

[RHMZ^{+}21] Mikko Rinta-Homi, Naser Hossein Motlagh, Agustin Zuniga, Huber Flores, and Petteri Nurmi. How low can you go?: Performance trade-offs in low-resolution thermal sensors for occupancy detection: a systematic evaluation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):126:1– 126:22, September 2021.

CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478104.

Ranasinghe:2018:VLU

[RKK18] Champika Ranasinghe, Jakub Krukar, and Christian Kray. Visualizing location uncertainty on mobile devices: Cross-cultural differences in perceptions and preferences. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(1): 1–22, March 2018. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3191762.

Rey:2022:ISP

[RLCI22] Bradley Rey, Bongshin Lee, Eun Kyoung Choe, and Pourang Irani. Investigating in-situ personal health data queries on smartwatches. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4): 179:1–179:??, December 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3569481.

Raj:2019:CDC

[RLGN19] Shriti Raj, Joyce M. Lee, Ashley Garrity, and Mark W. Newman. Clinical data in context: Towards sensemaking tools for interpreting personal health data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–20, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314409.

Rostaminia:2019:WNU

[RLM^{+}19] Soha Rostaminia, Alexander Lamson, Subhransu Maji, Tauhidur Rahman, and Deepak Ganesan. W!NCE: Unobtrusive sensing of upper facial action units with EOG-based eyewear. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–26, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314410.

Rabbi:2019:RCA

[RLY^{+}19] Mashfiqui Rabbi, Katherine Li, H. Yanna Yan, Kelly Hall, Predrag Klasnja, and Susan Murphy. ReVibe: a context-assisted evening recall approach to improve self-report adherence. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 27, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3369806.

Rostaminia:2017:ILP

[RMG^{+}17] Soha Rostaminia, Addison Mayberry, Deepak Ganesan, Benjamin Marlin, and Jeremy Gummeson. iLid: Low-power sensing of fatigue and drowsiness measures on a computational eyeglass. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):1–26, June 2017. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3090088.

Ruiz:2018:IRD

[RPB^{+}18] Carlos Ruiz, Shijia Pan, Adeola Bannis, Xinlei Chen, Carlee Joe-Wong, Hae Young Noh, and Pei Zhang. IDrone: Robust drone identification through motion actuation feedback. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–22, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3214283.

Rabbi:2018:WVR

[RPF^{+}18] Fazlay Rabbi, Taiwoo Park, Biyi Fang, Mi Zhang, and Youngki Lee. When virtual reality meets Internet of Things in the gym: Enabling immersive interactive machine exercises. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–21, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214281.

Rough:2020:EUD

[RQ20] Daniel J. Rough and Aaron Quigley. End-user development of experience sampling smartphone apps — recommendations and requirements. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):56:1– 56:19, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397307.

Rawassizadeh:2022:OFR

[RR22] Reza Rawassizadeh and Yi Rong. ODSearch: Fast and resource efficient ondevice natural language search for fitness trackers’ data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):178:1– 178:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3569488.

Radeta:2023:LDP

[RRS^{+}23] Marko Radeta, Claudio Rodrigues, Francisco Silva, Pedro Abreu, João Pestana, Ngoc Thi Nguyen, Agustin Zuniga, Huber Flores, and Petteri Nurmi. Lost in the deep?: Performance evaluation of dead reckoning techniques in underwater environments. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):73:1– 73:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3596245.

Regmi:2021:SAS

[RSSN21] Hem Regmi, Moh Sabbir Saadat, Sanjib Sur, and Srihari Nelakuditi. SquiggleMilli: Approximating SAR imaging on mobile millimeter-wave devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3): 125:1–125:26, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478113.

Ren:2020:LLS

[RTZ^{+}20] Yili Ren, Sheng Tan, Linghan Zhang, Zi Wang, Zhi Wang, and Jie Yang. Liquid level sensing using commodity WiFi in a smart home environment. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):24:1– 24:30, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3380996.

Ren:2022:GHP

[RWW^{+}22] Yili Ren, Zi Wang, Yichao Wang, Sheng Tan, Yingying Chen, and Jie Yang. Go-Pose: 3D human pose estimation using WiFi. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):69:1– 69:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534605.

Ram:2021:LTE

[RZ21] Ashwin Ram and Shengdong Zhao. LSVP: Towards effective on-the-go video learning using optical headmounted displays. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(1):30:1– 30:27, March 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3448118.

Soha:2017:ILP

[SAD^{+}17] Rostaminia Soha, Mayberry Addison, Ganesan Deepak, Marlin Benjamin, and Gummeson Jeremy. iLid: Low-power sensing of fatigue and drowsiness measures on a computational eyeglass. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

                                        //dl.acm.org/citation.
                                        cfm?id=3090088.

Sangar:2022:PSL

[SBP^{+}22] Yaman Sangar, Yoganand Biradavolu, Kai Pederson, Vaishnavi Ranganathan, and Bhuvana Krishnaswamy. PACT: Scalable, long-range communication for monitoring and tracking systems using battery-less tags. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):180:1– 180:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3569471.

Song:2022:IHF

[SC22] Yunpeng Song and Zhongmin Cai. Integrating handcrafted features with deep representations for smartphone authentication. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):27:1– 27:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517332.

Sun:2020:VRH

[SCZ^{+}20] Wei Sun, Tuochao Chen, Jiayi Zheng, Zhenyu Lei, Lucy Wang, Benjamin Steeper, Peng He, Matthew Dressa, Feng Tian, and Cheng Zhang. VibroSense: Recognizing home activities by deep learning subtle vibrations on an interior surface of a house from a single point using laser Doppler vibrometry. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):96:1– 96:28, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3411828.

Salekin:2018:WSL

[SEG^{+}18] Asif Salekin, Jeremy W.

Eberle, Jeffrey J. Glenn, Bethany A. Teachman, and John A. Stankovic. A weakly supervised learning framework for detecting social anxiety and depression. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1– 26, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3214284.

Swaminathan:2020:OFR

[SFR^{+}20] Saiganesh Swaminathan, Jonathon Fagert, Michael Rivera, Andrew Cao, Gierad Laput, Hae Young Noh, and Scott E. Hudson. OptiStructures: Fabrication of room-scale interactive structures with embedded fiber Bragg grating optical sensors and displays. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):50:1– 50:21, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3397310.

Sun:2023:UWS

[SGO^{+}23] Yanke Sun, Dwaynica A. Greaves, Guido Orgs, Antonia F. de C. Hamilton, Sally Day, and Jamie A. Ward. Using wearable sensors to measure interpersonal synchrony in actors and audience members during a live theatre performance. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):27:1– 27:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580781.

Shao:2023:CBC

[SGZ^{+}23] Shuai Shao, Yu Guan, Bing Zhai, Paolo Missier, and Thomas Plötz. ConvBoost: Boosting ConvNets for sensor-based activity recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):75:1– 75:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596234.

Sheng:2020:WSM

[SH20] Taoran Sheng and Manfred Huber. Weakly supervised multi-task representation learning for human activity analysis using wearables. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):57:1– 57:18, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397330.

Santhalingam:2020:MEI

[SHZ^{+}20] Panneer Selvam Santhalingam, Al Amin Hosain, Ding Zhang, Parth Pathak, Huzefa Rangwala, and Raja Kushalnagar. mmASL: Environmentindependent ASL gesture recognition using 60 GHz millimeter-wave signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):26:1– 26:30, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3381010.

Song:2021:HHM

[SJL^{+}21] Yiwei Song, Dongzhe Jiang, Yunhuai Liu, Zhou Qin, Chang Tan, and Desheng Zhang. HERMAS: a human mobility embedding framework with large-scale cellular signaling data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):128:1– 128:21, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478108.

Schroeder:2019:EOG

[SKM^{+}19] Jessica Schroeder, Ravi Karkar, Natalia Murinova, James Fogarty, and Sean A. Munson. Examining opportunities for goal-directed self-tracking to support chronic condition management. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 26, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369809.

Seo:2021:OSI

[SKS^{+}21] Hyunggoog Seo, Jaedong Kim, Kwanggyoon Seo, Bumki Kim, and Junyong Noh. Overthere: a simple and intuitive object registration method for an absolute mid-air pointing interface. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3): 127:1–127:24, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478128.

Song:2018:MSR

[SLX^{+}18] Chen Song, Zhengxiong Li, Wenyao Xu, Chi Zhou, Zhanpeng Jin, and Kui Ren. My smartphone recognizes genuine QR codes!: Practical unclonable QR code via 3D printing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–20, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214286.

Sharma:2020:ACP

[SNGK20] Kshitij Sharma, Evangelos Niforatos, Michail Giannakos, and Vassilis Kostakos. Assessing cognitive performance using physiological and facial features: Generalizing across contexts. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):95:1– 95:41, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3411811.

Saeed:2019:MTS

[SOL19] Aaqib Saeed, Tanir Ozcelebi, and Johan Lukkien. Multitask self-supervised learning for human activity detection. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–30, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3328932.

Santhalingam:2023:SSI

[SPRK23] Panneer Selvam Santhalingam, Parth Pathak, Huzefa Rangwala, and Jana Kosecka. Synthetic smartwatch IMU data generation from in-the-wild ASL videos. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(2):74:1– 74:??, June 2023. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3596261.

Swaminathan:2019:IOC

[SRK^{+}19] Saiganesh Swaminathan, Michael Rivera, Runchang Kang, Zheng Luo, Kadri Bugra Ozutemiz, and Scott E. Hudson. Input, output and construction methods for custom fabrication of room-scale deployable pneumatic structures. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–17, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328933.

Shahid:2020:LFH

[SRR^{+}20] Farhana Shahid, Wasifur Rahman, M. Saifur Rahman, Sharmin Akther Purabi, Ayesha Seddiqa, Moin Mostakim, Farhan Feroz, Tanjir Rashid Soron, Fahmida Hossain, Nabila Khan, Anika Binte Islam, Nipi Paul, Ehsan Hoque, and A. B. M. Alim Al Islam. Leveraging freehand sketches for potential screening of PTSD. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):94:1– 94:22, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3411835.

Sun:2022:FSV

[SS22] Wei Sun and Kannan Srinivasan. On the feasibility of securing vehiclepavement interaction. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):29:1– 29:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3517230.

Songwa:2021:LLE

[SSB^{+}21] Prince U. C. Songwa, Aaqib Saeed, Sachin Bhardwaj, Thijs W. Kruisselbrink, and Tanir Ozcelebi. LumNet: Learning to estimate vertical visual field luminance for adaptive lighting control. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2): 79:1–79:20, June 2021. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3463500.

Swain:2019:MPC

[SSR^{+}19] Vedant Das Swain, Koustuv Saha, Hemang Rajvanshy, Anusha Sirigiri, Julie M. Gregg, Suwen Lin, Gonzalo J. Martinez, Stephen M. Mattingly, Shayan Mirjafari, Raghu Mulukutla, Subigya Nepal, Kari Nies, Manikanta D. Reddy, Pablo Robles-Granda, Andrew T. Campbell, Nitesh V. Chawla, Sidney D’Mello, Anind K. Dey, Kaifeng Jiang, Qiang Liu, Gloria Mark, Edward Moskal, Aaron Striegel, Louis Tay, Gregory D. Abowd, and Munmun De Choudhury. A multisensor person-centered approach to understand the role of daily activities in job performance with organizational personas. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1–27, December 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369828.

Sapiezynski:2017:IPP

[SSW^{+}17] Piotr Sapiezynski, Arkadiusz Stopczynski, David Kofoed Wind, Jure Leskovec, and Sune Lehmann. Inferring person-to-person proximity using WiFi signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

                                        //dl.acm.org/citation.
                                        cfm?id=3090089.

Shih:2019:BSB

[STL^{+}19] Chen-Hsuan (Iris) Shih, Naofumi Tomita, Yanick X. Lukic, Álvaro Hernández Reguera, Elgar Fleisch, and Tobias Kowatsch. Breeze: Smartphone-based acoustic real-time detection of breathing phases for a gamified biofeedback breathing training. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1–30, December 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3369835.

Sarsenbayeva:2019:MES

[SvBH^{+}19] Zhanna Sarsenbayeva, Niels van Berkel, Danula Hettiachchi, Weiwei Jiang, Tilman Dingler, Eduardo Velloso, Vassilis Kostakos, and Jorge Goncalves. Measuring the effects of stress on mobile interaction. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–18, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3314411.

Sarsenbayeva:2018:EDA

[SvBV^{+}18] Zhanna Sarsenbayeva, Niels van Berkel, Eduardo Velloso, Vassilis Kostakos, and Jorge Goncalves. Effect of distinct ambient noise types on mobile interaction. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–23, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214285.

Su:2022:LDB

[SWLG22] Jie Su, Zhenyu Wen, Tao Lin, and Yu Guan. Learning disentangled behaviour patterns for wearable-based human activity recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1): 28:1–28:??, March 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3517252.

Shen:2021:IFP

[SWLZ21] Meng Shen, Yaqian Wei, Zelin Liao, and Liehuang Zhu. IriTrack: Face presentation attack detection using Iris tracking. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):78:1– 78:21, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3463515.

Sun:2022:EEM

[SXF^{+}22] Xue Sun, Jie Xiong, Chao Feng, Wenwen Deng, Xudong Wei, Dingyi Fang, and Xiaojiang Chen. Earmonitor: In-ear motion-resilient acoustic sensing using commodity earphones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):182:1– 182:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3569472.

Shang:2022:PPL

[SYYL22] Fei Shang, Panlong Yang, Yubo Yan, and Xiang-Yang Li. PackquID: Inpacket liquid identification using RF signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):181:1– 181:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3569469.

Sicong:2017:UBL

[SZJ^{+}17] Liu Sicong, Zhou Zimu, Du Junzhao, Shangguan Longfei, Jun Han, and Xin Wang. UbiEar: Bringing location-independent sound awareness to the hard-of-hearing people with smartphones. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090082.

Shi:2020:RST

[SZZ^{+}20] Yilei Shi, Haimo Zhang, Kaixing Zhao, Jiashuo Cao, Mengmeng Sun, and Suranga Nanayakkara. Ready, steady, touch!: Sensing physical contact with a finger-mounted IMU. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):59:1– 59:25, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397309.

Trinh:2017:RRC

[TAEB17] H. Trinh, R. Asadi, D. Edge, and T. Bickmore. Robo-COP: A robotic coach for oral presentations. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090092.

Terzimehic:2023:TCR

[TAHH23] Nada Terzimehić, Sarah Aragon-Hahner, and Heinrich Hussmann. The tale of a complicated relationship: Insights from users’ Love/ breakup letters to their smartphones before and during the COVID-19 pandemic. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):28:1– 28:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3580792.

Tu:2019:FFC

[TFL^{+}19] Zhen Tu, Yali Fan, Yong Li, Xiang Chen, Li Su, and Depeng Jin. From fingerprint to footprint: Coldstart location recommendation by learning user interest from app data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–22, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3314413.

Tang:2017:HLT

[TK17] Lie Ming Tang and Judy Kay. Harnessing long term physical activity data — how long-term trackers use data and how an adherencebased interface supports new insights. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090091.

Tabassum:2019:IUP

[TKF^{+}19] Madiha Tabassum, Tomasz Kosiński, Alisa Frik, Nathan Malkin, Primal Wijesekera, Serge Egelman, and Heather Richter Lipford. Investigating users’ preferences and expectations for always-listening voice assistants. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1–23, December 2019. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369807.

Talla:2017:BFC

[TKGS17] Vamsi Talla, Bryce Kellogg, Shyamnath Gollakota, and Joshua R. Smith. Batteryfree cellphone. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

                                        //dl.acm.org/citation.
                                        cfm?id=3090090.

Tang:2022:CAC

[TLF22] Xiao Tang, Ruihui Li, and Chi-Wing Fu. CAFI-AR: Contact-aware freehand interaction with AR objects. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4): 183:1–183:??, December 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3569499.

Torkamaan:2023:MMS

[Tor23] Helma Torkamaan. Mood measurement on smartphones: Which measure, which design? Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):29:1– 29:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3580864.

Tapia:2022:SSC

[TXWZ22] Miguel Chávez Tapia, Talia Xu, Zehang Wu, and Marco Zúñiga Zamalloa. SunBox: Screen-to-camera communication with ambient light. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(2):46:1– 46:??, July 2022. CODEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534602.

Tong:2019:BLB

[TZW^{+}19] Xinyu Tong, Fengyuan Zhu, Yang Wan, Xiaohua Tian, and Xinbing Wang. Batch localization based on OFDMA backscatter. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–25, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314412.

Verma:2021:ESF

[VBS^{+}21] Dhruv Verma, Sejal Bhalla, Dhruv Sahnan, Jainendra Shukla, and Aman Parnami. ExpressEar: Sensing fine-grained facial expressions with earables. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):129:1– 129:28, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478085.

Vasisht:2018:DEU

[VJH^{+}18] Deepak Vasisht, Anubhav Jain, Chen-Yu Hsu, Zachary Kabelac, and Dina Katabi.

Duet: Estimating user position and identity in smart homes using intermittent and incomplete RF-Data. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1– 21, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3214287.

Veuskens:2020:RRA

[VLR20] Tom Veuskens, Kris Luyten, and Raf Ramakers. Rataplan: Resilient automation of user interface actions with multi-modal proxies. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):60:1– 60:23, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397329.

Venkatnarayan:2019:EII

[VS19] Raghav H. Venkatnarayan and Muhammad Shahzad. Enhancing indoor inertial odometry with WiFi. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–27, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328918.

Venkatnarayan:2020:LPW

[VSY^{+}20] Raghav H. Venkatnarayan, Muhammad Shahzad, Sangki Yun, Christina Vlachou, and Kyu-Han Kim. Leveraging polarization of WiFi signals to simultaneously track multiple people. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):45:1– 45:24, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397317.

Wu:2019:SSC

[WAL^{+}19] Jason Wu, Karan Ahuja, Richard Li, Victor Chen, and Jeffrey Bigham. ScratchThat: Supporting command-agnostic speech repair in voicedriven assistants. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–17, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328934.

Wan:2019:AHO

[WBG^{+}19] Zhiyuan Wan, Lingfeng Bao, Debin Gao, Eran Toch, Xin Xia, Tamir Mendel, and David Lo. AppMoD: Helping older adults manage mobile security with online social help. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 22, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3369819.

Wang:2021:DVP

[WCW21] Dawei Wang, Kai Chen, and Wei Wang. Demystifying the vetting process of voicecontrolled skills on markets. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3): 130:1–130:28, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478101.

Wang:2019:LTA

[WCY^{+}19] Xiaolei Wang, Andrea Continella, Yuexiang Yang, Yongzhong He, and Sencun Zhu. LeakDoctor: Toward automatically diagnosing privacy leaks in mobile applications. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–25, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3314415.

Wang:2021:CAA

[WGL^{+}21] Hongli Wang, Bin Guo, Jiaqi Liu, Sicong Liu, Yungang Wu, and Zhiwen Yu. Context-aware adaptive surgery: a fast and effective framework for adaptative model partition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):131:1– 131:22, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478073.

Wang:2019:MST

[WLZ^{+}19] Huandong Wang, Yong Li, Sihan Zeng, Gang Wang, Pengyu Zhang, Pan Hui, and Depeng Jin. Modeling spatio-temporal app usage for a large user population. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–23, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314414.

Wu:2021:LWA

[WMS^{+}21] Tong Wu, Nikolas Martelaro, Simon Stent, Jorge Ortiz, and Wendy Ju. Learning when agents can talk to drivers using the IN-AGT dataset and multisensor fusion. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):133:1– 133:28, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478125.

Wang:2020:WOD

[WMW^{+}20] Jiangtao Wang, Junyi Ma, Yasha Wang, Ning Wang, Leye Wang, Daqing Zhang, Feng Wang, and Qin Lv. Will online digital footprints reveal your relationship status?: an empirical study of social applications for sexual-minority men. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):29:1– 29:23, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3380978.

Waghmare:2023:GSB

[WPH^{+}23] Anandghan Waghmare, Farshid Salemi Parizi, Jason Hoffman, Yuntao Wang, Matthew Thompson, and Shwetak Patel. GlucoScreen: a smartphone-based readerless glucose test strip for prediabetes screening. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 7(1):30:1– 30:??, March 2023. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580855.

Wang:2019:RTW

[WPJ^{+}19] Jingxian Wang, Chengfeng Pan, Haojian Jin, Vaibhav Singh, Yash Jain, Jason I. Hong, Carmel Majidi, and Swarun Kumar. RFID Tattoo: a wireless platform for speech recognition. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 24, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3369812.

Wang:2022:ULL

[WRLK22] Jingxian Wang, Vaishnavi Ranganathan, Jonathan Lester, and Swarun Kumar. Ultra low-latency backscatter for fast-moving location tracking. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(1):30:1– 30:??, March 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3517242.

Weigel:2017:DDI

[WS17] Martin Weigel and Jürgen Steimle. DeformWear: Deformation input on tiny wearable devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090093.

Wang:2022:RCD

[WSW^{+}22] Ge Wang, Shouqian Shi, Minmei Wang, Chen Qian, Cong Zhao, Han Ding, Wei Xi, and Jizhong Zhao. RF-Chain: Decentralized, credible, and counterfeit-proof supply chain management with commodity RFIDs. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4): 184:1–184:??, December 2022. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3569493.

Wang:2020:AET

[WTA^{+}20] Xiyue Wang, Kazuki Takashima, Tomoaki Adachi, Patrick Finn, Ehud Sharlin, and Yoshifumi Kitamura. AssessBlocks: Exploring toy block play features for assessing stress in young children after natural disasters. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1): 30:1–30:29, March 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3381016.

Wu:2020:WSE

[WXL^{+}20] Di Wu, Tao Xiao, Xuewen Liao, Jie Luo, Chao Wu, Shigeng Zhang, Yong Li, and Yike Guo. When sharing economy meets IoT: Towards fine-grained urban air quality monitoring through mobile crowdsensing on bike-share system. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2): 61:1–61:26, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397328.

Wang:2018:REH

[WXW^{+}18] Chuyu Wang, Lei Xie, Wei Wang, Yingying Chen, Yanling Bu, and Sanglu Lu. RF-ECG: Heart rate variability assessment based on COTS RFID tag array. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1–26, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3214288.

Wu:2017:GPA

[WXY^{+}17] Chenshu Wu, Jingao Xu, Zheng Yang, Nicholas D.

Lane, and Zuwei Yin. Gain without pain: Accurate WiFi-based localization using fingerprint spatial gradient. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090094.

Wang:2019:RFC

[WXY^{+}19] Zhongqin Wang, Min Xu, Ning Ye, Ruchuan Wang, and Haiping Huang. RF-Focus: Computer visionassisted region-of-interest RFID tag recognition and localization in multipathprevalent environments. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–30, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314416.

Waghmare:2020:USS

[WXZ^{+}20] Anandghan Waghmare, Qiuyue Xue, Dingtian Zhang, Yuhui Zhao, Shivan Mittal, Nivedita Arora, Ceara Byrne, Thad Starner, and Gregory D. Abowd. Ubiqui-Touch: Self sustaining ubiquitous touch interfaces. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):27:1– 27:22, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3380989.

Wang:2021:AKR

[WYL^{+}21] Tianshi Wang, Shuochao Yao, Shengzhong Liu, Jinyang Li, Dongxin Liu, Huajie Shao, Ruijie Wang, and Tarek Abdelzaher. Audio keyword reconstruction from on-device motion sensor signals via neural frequency unfolding. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):132:1– 132:29, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478102.

Wang:2020:FDD

[WZF^{+}20] Guang Wang, Yongfeng Zhang, Zhihan Fang, Shuai Wang, Fan Zhang, and Desheng Zhang. FairCharge: a data-driven fairness-aware charging recommendation system for large-scale electric taxi fleets. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(1):28:1– 28:25, March 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3381003.

Wang:2019:CCF

[WZG^{+}19] Qianru Wang, Junbo Zhang, Bin Guo, Zexia Hao, Yifang Zhou, Junkai Sun, Zhiwen Yu, and Yu Zheng. City-Guard: Citywide fire risk forecasting using a machine learning approach. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(4):1– 21, December 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3369814.

Xie:2020:RRC

[XFW^{+}20] Xiaoyang Xie, Zhihan Fang, Yang Wang, Fan Zhang, and Desheng Zhang. RISC: Resource-constrained urban sensing task scheduling based on commercial fleets. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2): 62:1–62:20, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3397337.

Xia:2020:RUF

[XKNM20] Qingxin Xia, Joseph Korpela, Yasuo Namioka, and Takuya Maekawa. Robust unsupervised factory activity recognition with bodyworn accelerometer using temporal structure of multiple sensor data motifs. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):97:1– 97:30, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3411836.

Xia:2019:RUD

[XL19] Tong Xia and Yong Li. Revealing urban dynamics by learning online and offline behaviours together. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(1): 1–25, March 2019. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3314417.

Xu:2021:CMB

[XLL^{+}21] Chenhan Xu, Huining Li, Zhengxiong Li, Hanbin Zhang, Aditya Singh Rathore, Xingyu Chen, Kun Wang, Ming chun Huang, and Wenyao Xu. CardiacWave: a mmWave-based scheme of non-contact and highdefinition heart activity computing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):135:1– 135:26, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478127.

Xu:2020:SSE

[XLX^{+}20] Fengli Xu, Zongyu Lin, Tong Xia, Diansheng Guo, and Yong Li. SUME: Semantic-enhanced urban mobility network embedding for user demographic inference. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):98:1– 98:25, September 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3411807.

Xu:2020:VPR

[XYG^{+}20] Shuchang Xu, Ciyuan Yang, Wenhao Ge, Chun Yu, and Yuanchun Shi. Virtual paving: Rendering a smooth path for people with visual impairment through vibrotactile and audio feedback. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3): 99:1–99:25, September 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3411814.

Xie:2021:PLL

[XYX21] Binbin Xie, Yuqing Yin, and Jie Xiong. Pushing the limits of long range wireless sensing with LoRa. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):134:1– 134:21, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478080.

Xie:2021:ABU

[XZZ21] Wentao Xie, Qian Zhang, and Jin Zhang. Acousticbased upper facial action recognition for smart eyewear. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(2):41:1–41:28, June 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3448105.

Yang:2020:ECS

[YFX^{+}20] Yu Yang, Zhihan Fang, Xiaoyang Xie, Fan Zhang, Yunhuai Liu, and Desheng Zhang. Extending coverage of stationary sensing systems with mobile sensing systems for human mobility modeling. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):100:1– 100:21, September 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3411827.

Yue:2018:EMP

[YHW^{+}18] Shichao Yue, Hao He, Hao Wang, Hariharan Rahul, and Dina Katabi. Extracting multi-person respiration from entangled RF signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(2):1– 22, July 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3214289.

Yau:2020:HSC

[YLL^{+}20] Yui-Pan Yau, Lik Hang Lee, Zheng Li, Tristan Braud, Yi-Hsuan Ho, and Pan Hui. How subtle can it get?: a trimodal study of ring-sized interfaces for one-handed drone control. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):63:1– 63:29, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3397319.

Young:2020:MNT

[YLMR20] Jacob Young, Tobias Langlotz, Steven Mills, and Holger Regenbrecht. Mobileportation: Nomadic telepresence for mobile devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2): 65:1–65:16, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397331.

Yao:2019:ADI

[YPJdB19] X. Yao, T. Plötz, M. Johnson, and K. de Barbaro. Automated detection of infant holding using wearable sensing: Implications for developmental science and intervention. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 3(2):1–17, June 2019. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3328935.

Yu:2020:SAS

[YXW^{+}20] Yue Yu, Tong Xia, Huandong Wang, Jie Feng, and Yong Li. Semantic-aware spatio-temporal app usage representation via graph convolutional network. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):101:1– 101:24, September 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3411817.

Yue:2020:BMS

[YYW^{+}20] Shichao Yue, Yuzhe Yang, Hao Wang, Hariharan Rahul, and Dina Katabi. Body-Compass: Monitoring sleep posture with wireless signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):66:1–66:25, June 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3397311.

Yang:2021:MBH

[YZ21] Qiang Yang and Yuanqing Zheng. Model-based head orientation estimation for smart devices. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):136:1– 136:24, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478089.

Yin:2020:LRH

[YZS^{+}20] Huanpu Yin, Anfu Zhou, Guangyuan Su, Bo Chen, Liang Liu, and Huadong Ma. Learning to recognize handwriting input with acoustic features. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):64:1– 64:26, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397334.

Zhao:2017:MAM

[ZAP17] Nan Zhao, Asaph Azaria, and Joseph A. Paradiso. Mediated atmospheres: A multimodal mediated work environment. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090096.

Zhao:2021:ULP

[ZCA^{+}21] Minghui Zhao, Tyler Chang, Aditya Arun, Roshan Ayyalasomayajula, Chi Zhang, and Dinesh Bharadia. ULoc: Low-power, scalable and cm-accurate UWB-Tag localization and tracking for indoor applications. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):140:1– 140:31, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478124.

Zheng:2020:VVV

[ZCC^{+}20] Tianyue Zheng, Zhe Chen, Chao Cai, Jun Luo, and Xu Zhang. V2iFi: in-vehicle vital sign monitoring via compact RF sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):70:1– 70:27, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397321.

Zhou:2021:CWL

[ZCL21] Tongqing Zhou, Zhiping Cai, and Fang Liu. The crowd wisdom for location privacy of crowdsensing photos: Spear or shield? Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3): 142:1–142:23, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478106.

Zhang:2020:ELL

[ZCN^{+}20] Fusang Zhang, Zhaoxin Chang, Kai Niu, Jie Xiong, Beihong Jin, Qin Lv, and Daqing Zhang. Exploring LoRa for long-range through-wall sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):68:1– 68:27, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397326.

Zhao:2017:SAC

[ZLB^{+}17] Ru Zhao, Vivian Li, Hugo Barbosa, Gourab Ghoshal, and Mohammed Ehsan Hoque. Semi-automated 8 collaborative online training module for improving communication skills. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090097.

Zhao:2021:ASV

[ZLD^{+}21] Cui Zhao, Zhenjiang Li, Han Ding, Wei Xi, Ge Wang, and Jizhong Zhao. Anti-spoofing voice commands: a generic wireless assisted design. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3): 139:1–139:22, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478116.

Zhang:2020:TRT

[ZLL^{+}20] Shigeng Zhang, Yinggang Li, Xuan Liu, Song Guo, Weiping Wang, Jianxin Wang, Bo Ding, and Di Wu. Towards real-time cooperative deep inference over the cloud and edge end devices.

Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2): 69:1–69:24, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397315.

Zhou:2022:GAI

[ZM22] Heng Zhou and Takuya Maekawa. GPS-assisted indoor pedestrian dead reckoning. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 6(4):166:1– 166:??, December 2022. CO-DEN ???? ISSN 2474-9567 (electronic). URL https://

dl.acm.org/doi/10.1145/
3569467.

Zhai:2020:MSS

[ZPPC^{+}20] Bing Zhai, Ignacio Perez-Pozuelo, Emma A. D. Clifton, Joao Palotti, and Yu Guan. Making sense of sleep: Multimodal sleep stage classification in a large, diverse population using movement and cardiac sensing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):67:1– 67:33, June 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3397325.

Zhang:2020:OTU

[ZPZ^{+}20] Dingtian Zhang, Jung Wook Park, Yang Zhang, Yuhui Zhao, Yiyang Wang, Yunzhi Li, Tanvi Bhagwat, Wen-Fang Chou, Xiaojia Jia, Bernard Kippelen, Canek Fuentes-Hernandez, Thad Starner, and Gregory D. Abowd. OptoSense: Towards ubiquitous self-powered ambient light sensing surfaces. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):103:1– 103:27, September 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3411826.

Zhong:2021:MTP

[ZSG^{+}21] Weida Zhong, Qiuling Suo, Abhishek Gupta, Xiaowei Jia, Chunming Qiao, and Lu Su. MetaTP: Traffic prediction with unevenlydistributed road sensing data via fast adaptation. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3): 141:1–141:28, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478083.

Zeng:2020:MEM

[ZWX^{+}20] Youwei Zeng, Dan Wu, Jie Xiong, Jinyi Liu, Zhaopeng Liu, and Daqing Zhang. MultiSense: Enabling multiperson respiration sensing with commodity WiFi. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(3):102:1– 102:29, September 2020. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3411816.

Zhang:2021:WAS

[ZWZ^{+}21a] Qian Zhang, Dong Wang, Run Zhao, Yinggang Yu, and JiaZhen Jing. Write, attend and spell: Streaming end-to-end free-style handwriting recognition using smartwatches. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):138:1– 138:25, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478100.

Zhang:2021:SHS

[ZWZ^{+}21b] Qian Zhang, Dong Wang, Run Zhao, Yinggang Yu, and Junjie Shen. Sensing to hear: Speech enhancement for mobile devices using acoustic signals. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):137:1– 137:30, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478093.

Zhang:2017:SCT

[ZXW^{+}17] Cheng Zhang, Qiuyue Xue, Anandghan Waghmare, Sumeet Jain, Yiming Pu, Sinan Hersek, Kent Lyons, Kenneth A. Cunefare, Omer T. Inan, and Gregory D. Abowd. SoundTrak: Continuous 3D tracking of a finger using active acoustics. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 1(2):??, June 2017. CODEN ???? ISSN ???? URL http:

//dl.acm.org/citation.
cfm?id=3090095.

Zhou:2021:SVB

[ZXY21] Zhongyi Zhou, Anran Xu, and Koji Yatani. SyncUp: Vision-based practice support for synchronized dancing. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 5(3):143:1– 143:25, September 2021. CODEN ???? ISSN 2474-9567 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478120.

Zhang:2020:NMS

[ZZN^{+}20] Shibo Zhang, Yuqi Zhao, Dzung Tri Nguyen, Runsheng Xu, Sougata Sen, Josiah Hester, and Nabil Alshurafa. NeckSense: a multi-sensor necklace for detecting eating activities in free-living conditions. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(2):72:1– 72:26, June 2020. CODEN ???? ISSN 2474-9567 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3397313.